### v2

# Degree Distribution based Spiking Graph Networks for Domain Adaptation

Abstract
Spiking Graph Networks (SGNs) have garnered significant attraction from both researchers and industry due to their ability to address energy consumption challenges in graph classification. However, SGNs are only effective for in-distribution data and cannot tackle out-of-distribution data. In this paper, we first propose the domain adaptation problem in SGNs, and introduce a novel framework named Degree-aware Spiking Graph Domain Adaptation for Classification (DeSGDA). The proposed DeSGDA addresses the spiking graph domain adaptation problem by three aspects: node degree-aware personalized spiking representation, adversarial feature distribution alignment, and pseudo-label distillation. First, we introduce the personalized spiking representation method for generating degree-dependent spiking signals. Specifically, the threshold of triggering a spike is determined by the node degree, allowing this personalized approach to capture more expressive information for classification. Then, we propose the graph feature distribution alignment module that is adversarially trained using membrane potential against a domain discriminator. Such an alignment module can efficiently maintain high performance and low energy consumption in the case of inconsistent distribution. Additionally, we extract consistent predictions across two spaces to create reliable pseudo-labels, effectively leveraging unlabeled data to enhance graph classification performance. Extensive experiments on benchmark datasets validate the superiority of the proposed DeSGDA compared with competitive baselines.
1 Introduction
Spiking Graph Networks (SGNs) Xu et al. (2021b); Zhu et al. (2022) are a specialized type of artificial neural network engineered to process graph information by mimicking the human brain. SGNs transform static and real-valued graph features into discrete spikes by simulating neurons’ charging and discharging cycles, facilitating spike-based representations for graph node classification. Notably, SGNs excel in capturing semantic spiking representations with low energy consumption, which proves advantageous for event-based processing tasks Yao et al. (2021); Rosenfeld et al. (2021); Chen et al. (2023a, b) such as object recognition Gu et al. (2020); Li et al. (2021b); Liu et al. (2023b); Yi et al. (2024), real-time data analysis Bauer et al. (2019); Zhu et al. (2020); Baronig et al. (2024), and graph classification Li et al. (2023); Yao et al. (2023b); Meng et al. (2024); Ju et al. (2024).
Currently, SGNs are usually tested within the same distribution as the training dataset Li et al. (2023); Yin et al. (2024); Duan et al. (2024). However, in realistic scenarios, the testing set can have different distributions from the training set, and such a distribution shift may lead to a degradation in performance. For instance, Electroencephalography (EEG) data Binnie and Prior (1994); Biasiucci et al. (2019); Peng et al. (2022), typically represented as a graph structure with nodes for neurons and edges for connections, is ideally processed by bio-inspired SGNs that mimic neuronal charging and discharging. Despite the suitability, EEGs often exhibit varying distributions over time or among different groups Zhao et al. (2020, 2021); Wang et al. (2022), leading to suboptimal performance of models trained on specific distributions when applied to others. This significant issue underscores the necessity of exploring domain adaptation for spiking graphs. Traditionally, SNNs transfer learning methods Zhan et al. (2021); Zhang et al. (2021); Zhan et al. (2024); Guo et al. (2024) have been applied in event-based or computer vision scenarios. However, there’s no existing research on spiking graph domain adaptation.
Designing an effective spiking graph domain adaptation framework for classification is non-trivial due to the following major challenges: (1) How to meticulously design an SGN under the circumstance of domain shift? SGNs usually utilize a global threshold for the firing of each node Xu et al. (2021a); Hagenaars et al. (2021); Yin et al. (2024); Zhao et al. (2024). However, we observe that the degree of each node influences the difficulty of triggering spikes. Specifically, nodes with high degrees can integrate more information from neighbors, making it easier for membrane potential to accumulate and trigger a spike. Conversely, nodes with lower degrees are more challenging to reach the firing threshold, denoted as the inflexible architecture challenge. (2) How to design a framework that effectively addresses spiking graph domain adaptation for classification? Current research primarily focuses on graph node classification within the same distribution Li et al. (2023); Yao et al. (2023a); Duan et al. (2024); Cai et al. (2024). However, spike-based graph classification under domain shift remains unexplored. (3) How to guarantee the stability of the proposed framework? Though some works have been proposed to address the spiking transfer learning challenges Zhan et al. (2021); Zhang et al. (2021); Liu et al. (2024a); Zhan et al. (2024), there is still no theoretical research on spiking graphs under domain shift.
To tackle these challenges, we propose a framework named Degree-aware Spiking Graph Domain Adaptation for Classification (DeSGDA), which comprises three components: degree-aware personalized spiking representation, graph feature distribution alignment, and pseudo-label distillation. To address the first challenge, we establish variable node thresholds based on their degrees. By adaptively updating these thresholds, we can achieve a more expressive and personalized spiking representation for each node. Then, we introduce a adversarial feature distribution alignment module that is adversarially trained using membrane potential against a domain discriminator. To further enhance performance, we extract consistent predictions from different spaces to generate reliable pseudo-labels. Additionally, to explore the generalization ability of the proposed DeSGDA, we first propose the error bound for spiking graph domain adaptation and demonstrate that our pseudo-label distillation module effectively reduces this upper bound.
Our contributions can be summarized as follows: (1) Problem Formulation: We first introduce the problem of spiking graph domain adaptation for classification, which is non-trivial due to the challenges of the inflexible architecture of SGNs and theoretical deficiency. (2) Novel Architecture: We propose DeSGDA, a framework that efficiently learns personalized spiking representations for nodes using degree-aware thresholds and aligns domain distributions through adversarial training on membrane potential. Furthermore, we utilize pseudo-label distillation to improve the performance further. (3) Theoretical Analysis: To guarantee the stability of DeSGDA, we provide theoretical proof of the error bound for spiking graph domain adaptation. Furthermore, we demonstrate that DeSGDA maintains a lower theoretical bound than standard spiking graph domain adaptation through the effective use of the pseudo-label distillation module. (4) Extensive Experiments. We evaluate the proposed DeSGDA on extensive spiking graph domain adaptation learning datasets, which shows that our proposed DeSGDA outperforms the variety of state-of-the-art methods.
2 Related work
Spiking Graph Networks (SGNs). SGNs are a specialized type of neural network that combines Spiking Neural Networks (SNNs) with Graph Neural Networks (GNNs), preserving energy efficiency while achieving competitive performance in various graph tasks Li et al. (2023); Yao et al. (2023a); Qiu et al. (2024); Duan et al. (2024). Existing research on SGNs focuses on capturing the dynamic temporal information contained within graphs and enhancing model scalability. For instance, Xu et al. (2021a) utilizes spatial-temporal feature normalization within SNNs to effectively process dynamic graph data, ensuring robust learning and improved predictive performance. Zhao et al. (2024) propose a method that dynamically adapts to evolving graph structures and relationships through a novel architecture that updates node representations in real time. Additionally, Yin et al. (2024) adapts SNNs to dynamic graph settings and employs implicit differentiation for the node classification task. However, existing methods still suffer from data distribution shift issues when training and testing data come from different domains, resulting in degraded model performance and generalization. To address this, we propose a novel domain adaptation method based on SGNs to tackle these challenges.
Spiking Transfer Learning. Spiking transfer learning focuses on adjusting SNNs to handle data distribution shifts across various domains effectively. Recent advances in spiking transfer learning have been extensively applied in vision tasks, enhancing model performance while maintaining energy efficiency Zhan et al. (2021); Zhang et al. (2021); Zhan et al. (2024). For instance, Guo et al. (2024) leverages a Jaccard attention mechanism within SNNs to effectively adapt to target domains without requiring source domain data. Similarly, He et al. (2024) facilitates the transfer of learned representations from static to dynamic event-based domains by adapting SNNs to process temporal information. Additionally, Zhan et al. (2024) converts RGB images into spike-based neuromorphic data, enabling SNNs to process visual information across various domains efficiently. However, the difficulty of graph topologies makes it infeasible to apply spiking transfer learning to SGNs directly. To this end, we introduce a specialized domain adaptation method tailored for SGNs.
3 Preliminaries
Bound for Graph Domain Adaptation (GDA). Applying GDA with optimal transport (OT), if the covariate shift holds on representations that $\mathbb{P}_{S}(Y|Z)=\mathbb{P}_{T}(Y|Z)$, the target risk $\epsilon_{T}(h,\hat{h})$ is bounded with the theorem:
Theorem 1
(You et al., 2023) Assuming that the learned discriminator is $C_{g}$-Lipschitz continuous as described in (Redko et al., 2017), and the graph feature extractor $f$ (also referred to as GNN) is $C_{f}$-Lipschitz that $||f||_{Lip}=\max_{G_{1},G_{2}}\frac{||f(G_{1})-f(G_{2})||_{2}}{\eta(G_{1},G_{2})}=C_{f}$ for some graph distance measure $\eta$. Let :=$\mathcal{H}:=\{h:\mathcal{G}\rightarrow\mathcal{Y}\}$ be the set of bounded real-valued functions with the pseudo-dimension $Pdim(\mathcal{H})=d$ that $h=g\circ f\in\mathcal{H}$, with probability at least $1-\delta$ the following inequality holds:
$\epsilon_{T}(h,\hat{h})\leq$ $\hat{\epsilon}_{S}(h,\hat{h})+\sqrt{\frac{4d}{N_{S}}\log(\frac{eN_{S}}{d})+\frac{1}{N_{S}}\log(\frac{1}{\delta})}+2C_{f}C_{g}W_{1}(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G))+\omega,$
where the (empirical) source and target risks are $\hat{\epsilon}_{S}(h,\hat{h})=\frac{1}{N_{S}}\sum_{n=1}^{N_{S}}|h(G_{n})-\hat{h}(G_{n})|$ and $\epsilon_{T}(h,\hat{h})=\mathbb{E}_{\mathbb{P}_{T}(G}\{|h(G)-\hat{h}(G)|\}$, respectively, where $\hat{h}:\mathcal{G}\rightarrow\mathcal{Y}$ is the labeling function for graphs and $\omega=\min_{||g||_{Lip}\leq C_{g},||f||_{Lip}\leq C_{f}}\{\epsilon_{S}(h,\hat{h})+\epsilon_{T}(h,\hat{h})\}$. The first Wasserstein distance is defined as (Villani et al., 2009): $W_{1}(\mathbb{P},\mathbb{Q})=\sup_{||g||_{Lip}\leq 1}\left\{\mathbb{E}_{\mathbb{P}_{S}(Z)}g(Z)-\mathbb{E}_{\mathbb{P}_{T}(Z)}g(Z)\right\}$.
The comprehensive justification of the OT-based graph domain adaptation bound demonstrates that the generalization gap relies on both the domain divergence $2C_{f}C_{g}W_{1}(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G))$ and model discriminability $\omega$.
Spiking Graph Networks. In contrast to traditional artificial neural networks, SGNs (Xu et al., 2021a; Zhu et al., 2022) convert input data into binary spikes over time, with each neuron in the SGNs maintaining a membrane potential that accumulates input spikes. A spike is produced as an output when the membrane potential exceeds a threshold, which is formulated as:
$u_{\tau+1,i}=\lambda(u_{\tau,i}-V_{th}s_{\tau,i})+\sum_{j}w_{ij}\mathcal{A}\left(A,s_{\tau,j}\right)+b,\quad s_{\tau+1,i}=\mathbb{H}(u_{\tau+1,i}-V_{th}),$ (1)
where $\mathbb{H}(x)$ is the Heaviside function, which is the non-differentiable spiking function. $\mathcal{A}$ is the graph aggregation operation, and $A$ is the adjacency matrix of graph. $s_{\tau,i}$ denotes the binary spike train of neuron $i$, and $\lambda$ is the constant. $w_{ij}$ and $b$ are the weights and bias of each neuron.
4 Methodology
This work studies the spiking graph domain adaptation problem and proposes a new approach DeSGDA. DeSGDA consists of three parts: Degree-aware personalized spiking representation utilizes different thresholds for different degrees, effectively addressing the inflexible architecture challenge; Adversarial distribution alignment uses the adversarial training on membrane potential against a domain discriminatory to align distribution between different domains, and Pseudo-label distillation further applies the pseudo-label to enhance model performance. We provide the theoretical guarantee of DeSGDA to ensure the effectiveness. The overview of DeSGDA is shown in Figure 1.
Figure 1: An overview of the proposed DeSGDA. DeSGDA first utilizes different thresholds for different degree nodes to obtain personalized spiking representation. To align domain distribution, we use adversarial training on membrane potential against domain discrimination. Besides, DeSGDA applies the pseudo-label to enhance model performance.
Problem Setup. Given a graph $G=(V,E,\mathbf{X})$ with the node set $V$, the edge set $E$, and the node attribute matrix $\mathbf{X}$. Denote $S$ as the binary input sampled from Bernoulli distribution with probability of $\mathbf{X}$. The labeled source domain is denoted as $\mathcal{D}^{s}=\{(G_{i}^{s},y_{i}^{s})\}_{i=1}^{N_{s}}$, where $y_{i}^{s}$ denotes the labels of $G_{i}^{s}$; the unlabeled target domain is $\mathcal{D}^{t}=\{G_{j}^{t}\}_{j=1}^{N_{t}}$, where $N^{s}$ and $N^{t}$ denote the number of source graphs and target graphs. Both domains share the same label space $\mathcal{Y}$ but have different distributions in the graph space. We aim to train a spiking graph model using labeled source graphs and unlabeled target graphs to achieve superior performance in the target domain.
4.1 Degree-aware Personalized Spiking Representation
In this part, we first study the disadvantages of directly applying SNNs to graphs and then propose the degree-aware personalized spiking representation. Existing SGNs Li et al. (2023); Yao et al. (2023a); Duan et al. (2024) usually employ a global threshold for membrane potential firing. However, the global threshold can lead to the inflexible architecture issue since nodes with higher degrees are more likely to trigger spikes than those with lower degrees. As shown in Eq. 1, nodes with higher degrees have more neighbors, and the aggregation operation allows for more significant feature accumulation, making it easier for these nodes to trigger spikes compared to those with fewer neighbors. To alleviate this issue, we propose the degree-aware thresholds and iteratively update their values.
Specifically, we first $set$ all the degrees of nodes in the source domain graphs, i.e., $D^{s}=set(D_{1}^{s}\cup\cdots\cup D_{N_{s}}^{s})$, where $D_{i}$ denotes the degree set of graph $G_{i}^{s}$, and $set(\cdot)$ operation is an unordered sequence of non-repeating elements. Considering that low-degree nodes are more challenging to trigger while high-degree nodes trigger more easily, we propose setting higher thresholds for high-degree nodes and lower thresholds for low-degree nodes, which is formulated as:
$s_{\tau}^{d_{i}^{s}}=\mathbb{H}(u_{\tau}-V_{th}^{d_{i}^{s}}),\quad S^{d_{i}^{s}}=avg(s_{\tau}^{d_{i}^{s}}),\quad V_{th}^{d_{i}^{s}}=(1-\alpha)V_{th}^{d_{i}^{s}}+\alpha S^{d_{i}^{s}},$ (2)
where $V_{th}^{d_{i}^{s}}$ is the threshold of degree $d_{i}^{s}\in D^{s}$, initially set to $V_{th}$, and $\alpha$ is a hyper-parameter. The $avg(\cdot)$ operation takes the average of spiking representation with degree $d_{i}^{s}$. Consequently, high-degree nodes tend to achieve high $S^{d_{i}^{s}}$, which leads to an iterative increase in the threshold corresponding to degree $d_{i}^{s}$ and conversely for lower-degree nodes.
With different thresholds for different node degrees, we can obtain the personalized node spiking representation $\mathbf{s}_{v\in G_{i}^{s}}^{d_{j}^{s}}$. Then, we summarize all node representations with a readout function into the graph-level representation and output the prediction with a multi-layer perception (MLP) classifier:
$\mathbf{s}_{i}=\operatorname{READOUT}\left(\left\{\mathbf{s}_{v}^{d_{j}^{s}}\right\}_{v\in G_{i}^{s}}\right),\quad\hat{y}_{i}^{s}=H(\mathbf{s}_{i}),$ (3)
where $\hat{y}_{i}$ is the predicted result and $H(\cdot)$ is the classifier. After that, the source classification loss is:
$\mathcal{L}_{S}=\mathbb{E}_{G_{i}^{s}\in\mathcal{D}^{s}}{l}(y_{i}^{s},\hat{y}_{i}^{s}),$ (4)
where $l(\cdot)$ is the loss function and $y_{i}^{s}$ is the ground truth of the $i$-th graph $G_{i}^{s}$ in the source domain.
However, in the scenario of domain adaptation, two significant issues remain unresolved. The first issue is that degrees in the target domain may be unseen in the source, rendering the thresholds ineffective for these degrees. The second is that the thresholds in the target may differ from those in the source, simply applying the source domain’s thresholds could lead to performance degradation.
To alleviate the first issue, we initialize the threshold $V_{th}^{d_{i}^{t}}$ with the same value, where $d_{i}^{t}\notin D^{s}$. Then, with the training process of adversarial alignment, we iteratively update the threshold for degree $d_{i}^{t}$ with Eq. 2. To address the second issue, we incorporate the pseudo-label distillation module in Section 4.3 to guide the update of source degree thresholds on the target domain.
4.2 Adversarial Distribution Alignment
To eliminate the discrepancy between the source and target domains, we propose the adversarial distribution alignment module. Specifically, for each source graph $G_{i}^{s}$ and target graph $G_{i}^{t}$, we use the degree-aware personalized spiking GNNs-based encoder $F(\cdot)$ and semantic classifier $H(\cdot)$ to produce predicted labels. Then, a domain discriminator $Q(\cdot)$ is trained to distinguish features from the source and target domains. The encoder and classifier are adversarial trained to align the feature spaces of the source and target domains.
$\mathcal{L}_{AD}=\mathbb{E}_{G_{i}^{s}\in\mathcal{D}^{s}}\log Q\left(F(G_{i}^{s}),H(G_{i}^{s})|V_{th}^{D^{s}}\right)+\mathbb{E}_{G_{j}^{t}\in\mathcal{D}^{t}}\log\left(1-Q\left(F\left(G_{j}^{t}\right),H(G_{j}^{t})|V_{th}^{D^{s}}\right)\right).$
However, the degree in the target domain may be unseen by the source. Thus, we further initialize the threshold with $V_{th}^{d_{j}^{t}}$ and $d_{j}^{t}\notin D^{s}$, which is formulated as:
$\mathcal{L}_{AD}=$ $\mathbb{E}_{\begin{subarray}{c}G_{j}^{t}\in\mathcal{D}^{t}\\ D_{j}^{t}\subset{D^{s}}\end{subarray}}\log\left(1-Q\left(F\left(G_{j}^{t}\right),H(G_{j}^{t})|V_{th}^{D^{s}}\right)\right)+\mathbb{E}_{G_{i}^{s}\in\mathcal{D}^{s}}\log Q\left(F(G_{i}^{s}),H(G_{i}^{s})|V_{th}^{D^{s}}\right)$ (5)
$+\mathbb{E}_{\begin{subarray}{c}G_{j}^{t}\in\mathcal{D}^{t}\\ \exists d_{j}^{t}\notin{D^{s}}\end{subarray}}\log\left(1-Q\left(F\left(G_{j}^{t}\right),H(G_{j}^{t})|V_{th}^{D^{s}},V_{th}^{D^{t}}\right)\right),$
where $D^{t}=\{d_{i}^{t}|d_{i}^{t}\in D^{t},d_{i}^{t}\notin D^{s}\}$. Then, we iteratively update $V_{th}^{D^{t}}$ with Eq. 2 on each latency. Furthermore, we present an upper bound on the adversarial distribution alignment.
Theorem 2
Assuming that the learned discriminator is $C_{g}$-Lipschitz continuous as described in Theorem 1, the graph feature extractor $f$ (also referred to as GNN) is $C_{f}$-Lipschitz that $||f||_{Lip}=\max_{G_{1},G_{2}}\frac{||f(G_{1})-f(G_{2})||_{2}}{\eta(G_{1},G_{2})}=C_{f}$ for some graph distance measure $\eta$ and the loss function bounded by $C>0$. Let :=$\mathcal{H}:=\{h:\mathcal{G}\rightarrow\mathcal{Y}\}$ be the set of bounded real-valued functions with the pseudo-dimension $Pdim(\mathcal{H})=d$ that $h=g\circ f\in\mathcal{H}$, and provided the spike training data set $S_{n}=\{(\mathbf{X}_{i},y_{i})\in\mathcal{X}\times\mathcal{Y}\}_{i\in[n]}$ drawn from $\mathcal{D}^{s}$, with probability at least $1-\delta$ the following inequality :
$\epsilon_{T}(h,\hat{h}_{T}(\mathbf{X}))\leq$ $\hat{\epsilon}_{S}(h,\hat{h}_{S}(\mathbf{S}))+2\mathbb{E}\left[\sup\sup\frac{1}{N_{S}}\sum_{i=1}^{N_{S}}\epsilon_{i}h(\mathbf{X}_{i},y_{i},p_{i})\right]+C\sqrt{\frac{ln(2/\delta)}{N_{S}}}$ (6)
$+\min\left(|\epsilon_{S}(h,\hat{h}_{S}(\mathbf{X}))-\epsilon_{S}(h,\hat{h}_{T}(\mathbf{X}))|,|\epsilon_{T}(h,\hat{h}_{S}(\mathbf{X}))-\epsilon_{T}(h,\hat{h}_{T}(\mathbf{X}))|\right)$
$+2C_{f}C_{g}W_{1}\left(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G)\right),$
where the (empirical) source and target risks are $\hat{\epsilon}_{S}(h,\hat{h}(\mathbf{S}))=\frac{1}{N_{S}}\sum_{n=1}^{N_{S}}|h(\mathbf{S}_{n})-\hat{h}(\mathbf{S}_{n})|$ and $\epsilon_{T}(h,\hat{h}(\mathbf{X}))=\mathbb{E}_{\mathbb{P}_{T}(G}\{|h(G)-\hat{h}(G)|\}$, respectively, where $\hat{h}:\mathcal{G}\rightarrow\mathcal{Y}$ is the labeling function for graphs and $\omega=\min_{||g||_{Lip}\leq C_{g},||f||_{Lip}\leq C_{f}}\{\epsilon_{S}(h,\hat{h}(\mathbf{X}))+\epsilon_{T}(h,\hat{h}(\mathbf{X}))\}$, $\epsilon_{i}$ is the Rademacher variable and $p_{i}$ is the $i^{th}$ row of $\mathbf{P}$, which is the probability matrix with:
$\mathbf{P}_{kt}=\left\{\begin{array}[]{lr}\exp\left(\frac{u_{k}(t)-V_{th}}{\sigma(u_{k}(t)-u_{reset})}\right),&if\quad u_{\theta}\leq u(t)\leq V_{th},\\ 0,&if\quad u_{reset}\leq u_{k}(t)\leq u_{\theta}.\end{array}\right.$ (7)
Theorem 2 proves the generalization bound of spiking graph domain adaptation. More details can be found in Appendix A.
4.3 Pseudo-label Distillation for Discrimination Learning
To further address the variance in thresholds between the target and source domains, we incorporate the pseudo-label distillation module into the DeSGDA framework. With reliable pseudo-labels, we can effectively update the source degree thresholds in the target domain.
The goal of the pseudo-label distilling procedure is to keep those examples and their corresponding pseudo-labels from the deep feature space that aligns with the shallow feature space. Specifically, we denote $\mathbf{s^{\prime}}_{i}^{t}$ as the shallow spiking graph representation on the $L\textquoteright$ ’-th layer, where $L\textquoteright<L$, and $\hat{y}_{i}^{t}$ as the prediction of graph $G_{i}^{t}$ on the $L$-th layer.
Then, to enhance alignment between the shallow and deep feature spaces and facilitate the generation of more accurate pseudo-labels, we cluster the shallow features $\mathbf{s^{\prime}}^{t}$ into $C$ clusters and each cluster $\mathcal{E}_{j}$ includes graphs $\{G_{j}^{t}\}$.
After that, we find the dominating labels $e_{r}$ in the cluster, i.e., $\max_{r}|\{\mathcal{E}_{r}:e_{r}=\hat{y}_{j}^{t}\}|$, and remove other instances with the same pseudo-label but in different clusters. Formally, the pseudo-labels are signed with:
$\mathcal{P}=\left\{\left({G}_{j}^{t},\hat{y}_{j}^{t}\right):e_{j}=\max_{r}\left|\left\{\mathcal{E}_{r}:e_{r}=\hat{y}_{j}^{t}\right\}\right|\right\}.$ (8)
Finally, we utilize the distilled pseudo-labels to guide the update of source degree thresholds on the target domain with Eq. 2, and to direct classification in the target domain:
$\mathcal{L}_{T}=\mathbb{E}_{{G}_{j}^{t}\in\mathcal{P}}l\left({H(\mathbf{s}}_{j}^{t}),\hat{y}_{j}^{t}\right),$ (9)
where $H(\cdot)$ and $\mathbf{s}_{j}^{t}$ are the classifier and spiking graph representation, respectively, which are defined in Eq. 3. $l(\cdot)$ is the loss function, and we implement it with cross-entropy loss.
Theorem 3
Under the assumption of Theorem 1, we further assume that there exists a small amount of i.i.d. samples with pseudo labels $\{(G_{n},Y_{n})\}_{n=1}^{N^{\prime}_{T}}$ from the target distribution $\mathbb{P}_{T}(G,Y)$ $(N^{\prime}_{T}\ll N_{S})$ and bring in the conditional shift assumption that domains have different labeling function $\hat{h}_{S}\neq\hat{h}_{T}$ and $\max_{G_{1},G_{2}}\frac{|\hat{h}_{D}(G_{1})-\hat{h}_{D}(G_{2})|}{\eta(G_{1},G_{2})}=C_{h}\leq C_{f}C_{g}(D\in\{S,T\})$ for some constant $C_{h}$ and distance measure $\eta$, and the loss function bounded by $C>0$. Let :=$\mathcal{H}:=\{h:\mathcal{G}\rightarrow\mathcal{Y}\}$ be the set of bounded real-valued functions with the pseudo-dimension $Pdim(\mathcal{H})=d$, and provided the spike training data set $S_{n}=\{(\mathbf{X}_{i}^{s},y_{i}^{s})\}_{i\in[n]}$, with probability at least $1-\delta$ the following inequality holds:
$\epsilon_{T}(h,\hat{h}_{T}(\mathbf{X}))\leq$ $\frac{N^{\prime}_{T}}{N_{S}+N^{\prime}_{T}}\hat{\epsilon}_{T}(h,\hat{h}_{T}(S))+\frac{N_{S}}{N_{S}+N^{\prime}_{T}}\Bigg{(}\hat{\epsilon}_{S}(h,\hat{h}_{S}(S))+2C_{f}C_{g}W_{1}\left(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G)\right)$ (10)
$+2\mathbb{E}\left[\sup\frac{1}{N_{S}}\sum_{i=1}^{N_{S}}\epsilon_{i}h(\mathbf{X}_{i},y_{i},p_{i})\right]+C\sqrt{\frac{ln(2/\delta)}{N_{S}}}$
$+\min\left(|\epsilon_{S}(h,\hat{h}_{S}(\mathbf{X})))-\epsilon_{S}(h,\hat{h}_{T}(\mathbf{X})))|,|\epsilon_{T}(h,\hat{h}_{S}(\mathbf{X})))-\epsilon_{T}(h,\hat{h}_{T}(\mathbf{X})))|\right)\Bigg{)}$
$\leq$ $\hat{\epsilon}_{S}(h,\hat{h}_{S}(S))+2\mathbb{E}\left[\sup\frac{1}{N_{S}}\sum_{i=1}^{N_{S}}\epsilon_{i}h(\mathbf{X}_{i},y_{i},p_{i})\right]+C\sqrt{\frac{ln(2/\delta)}{N_{S}}}$
$+2C_{f}C_{g}W_{1}\left(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G)\right)+\omega^{\prime},$
where the (empirical) source and target risks are $\hat{\epsilon}_{S}(h,\hat{h})=\frac{1}{N_{S}}\sum_{n=1}^{N_{S}}|h(G_{n})-\hat{h}(G_{n})|$ and $\epsilon_{T}(h,\hat{h})=\mathbb{E}_{\mathbb{P}_{T}(G}\{|h(G)-\hat{h}(G)|\}$, respectively, where $\hat{h}:\mathcal{G}\rightarrow\mathcal{Y}$ is the labeling function for graphs and $\omega^{\prime}=\min_{||g||_{Lip}\leq C_{g},||f||_{Lip}\leq C_{f}}\{\epsilon_{S}(h,\hat{h})+\epsilon_{T}(h,\hat{h})\}$, $\epsilon_{i}$ is the Rademacher variable and $p_{i}$ is the $i^{th}$ row of $\mathbf{P}$, which is defined in Eq. 7.
The proof is detailed in Appendix B. From Theorem 3, we observe that the bound of DeSGDA is lower than simply aligning the distributions by incorporating the highly reliable pseudo-labels, demonstrating the effectiveness of pseudo labels for spiking graph domain adaptation.
4.4 Learning Framework
Finally, the overall training objective of DeSGDA integrates classification loss $\mathcal{L}_{S}$, adversarial training loss $\mathcal{L}_{AD}$, and pseudo-label distillation loss $\mathcal{L}_{T}$, which is formulated as:
$\mathcal{L}=\mathcal{L}_{S}+\mathcal{L}_{T}-\lambda\mathcal{L}_{AD},$ (11)
where $\lambda$ is a hyper-parameter to balance the adversarial training loss and classification loss. The learning procedure is illustrated in Algorithm C, and the complexity is shown in Appendix D .
[…]
WL subtree 68.7 82.3 50.7 82.3 58.1 83.8 64.0 74.1 43.7 70.5 71.3 60.1 67.5
GCN 73.4$\pm$0.2 83.5$\pm$0.3 57.6$\pm$0.2 84.2$\pm$1.8 24.0$\pm$0.1 16.6$\pm$0.4 57.6$\pm$0.2 73.7$\pm$0.4 24.0$\pm$0.1 26.6$\pm$0.2 39.9$\pm$0.9 42.5$\pm$0.1 50.3
GIN 62.5$\pm$4.7 74.9$\pm$3.7 53.0$\pm$4.6 59.6$\pm$4.2 73.7$\pm$0.8 64.7$\pm$3.4 60.6$\pm$2.7 69.8$\pm$0.6 31.1$\pm$2.8 63.1$\pm$3.4 72.3$\pm$2.7 64.6$\pm$1.4 62.5
GMT 73.4$\pm$0.3 83.5$\pm$0.2 57.6$\pm$0.1 83.5$\pm$0.3 24.0$\pm$0.1 83.5$\pm$0.1 57.4$\pm$0.2 73.4$\pm$0.2 24.1$\pm$0.1 73.4$\pm$0.3 24.0$\pm$0.1 57.6$\pm$0.2 59.6
CIN 74.5$\pm$0.2 84.1$\pm$0.5 57.8$\pm$0.2 82.7$\pm$0.9 75.6$\pm$0.6 79.2$\pm$2.2 61.5$\pm$2.7 74.0$\pm$1.0 75.5$\pm$0.8 72.5$\pm$2.1 76.0$\pm$0.3 60.9$\pm$1.2 72.9
SpikeGCN 71.8$\pm$0.8 79.5$\pm$1.3 63.8$\pm$1.0 78.9$\pm$1.4 68.6$\pm$1.1 76.5$\pm$1.8 62.3$\pm$2.2 72.1$\pm$1.5 68.1$\pm$2.1 67.2$\pm$1.9 69.2$\pm$2.1 64.2$\pm$1.8 70.2
DRSGNN 72.6$\pm$0.6 80.1$\pm$1.6 63.1$\pm$1.4 79.5$\pm$1.8 70.4$\pm$1.9 78.6$\pm$2.1 64.1$\pm$1.7 70.7$\pm$2.3 67.8$\pm$1.6 65.6$\pm$1.4 71.3$\pm$1.3 62.1$\pm$1.0 70.5
CDAN 72.2$\pm$1.8 82.4$\pm$1.6 59.8$\pm$2.1 76.8$\pm$2.4 69.3$\pm$4.1 71.8$\pm$3.7 64.4$\pm$2.5 74.3$\pm$0.4 46.3$\pm$2.0 69.8$\pm$1.8 74.4$\pm$1.7 62.6$\pm$2.3 68.7
ToAlign 73.4$\pm$0.1 83.5$\pm$0.2 57.6$\pm$0.1 83.5$\pm$0.2 24.0$\pm$0.3 83.5$\pm$0.4 57.6$\pm$0.1 73.4$\pm$0.1 24.0$\pm$0.2 73.4$\pm$0.2 24.0$\pm$0.1 57.6$\pm$0.3 59.6
MetaAlign 75.5$\pm$0.9 84.9$\pm$0.6 64.8$\pm$1.6 85.9$\pm$1.1 69.3$\pm$2.7 83.3$\pm$0.6 68.7$\pm$1.2 74.2$\pm$0.7 73.3$\pm$3.3 72.2$\pm$0.9 69.9$\pm$1.8 63.6$\pm$2.3 73.8
DEAL 76.5$\pm$0.4 83.1$\pm$0.4 67.5$\pm$1.3 77.6$\pm$1.8 76.0$\pm$0.2 80.1$\pm$2.7 66.1$\pm$1.3 75.4$\pm$1.5 42.3$\pm$4.1 68.1$\pm$3.7 73.1$\pm$2.2 67.8$\pm$1.2 71.1
CoCo 75.5$\pm$0.2 84.2$\pm$0.4 59.8$\pm$0.5 83.4$\pm$0.2 73.6$\pm$2.3 81.6$\pm$2.4 65.8$\pm$0.3 76.2$\pm$0.2 75.8$\pm$0.2 71.1$\pm$2.1 76.1$\pm$0.2 67.1$\pm$0.6 74.2
SGDA 63.8$\pm$0.6 65.2$\pm$1.3 66.7$\pm$1.0 59.1$\pm$1.5 60.1$\pm$0.8 64.4$\pm$1.2 65.2$\pm$0.7 63.9$\pm$0.9 64.5$\pm$0.6 61.1$\pm$1.3 58.9$\pm$1.4 64.9$\pm$1.2 63.2
DGDA 58.7$\pm$0.8 59.9$\pm$1.2 57.1$\pm$0.6 57.9$\pm$0.8 59.2$\pm$1.3 58.9$\pm$0.4 61.1$\pm$1.2 60.3$\pm$1.6 58.6$\pm$0.9 57.5$\pm$1.2 58.4$\pm$0.5 62.3$\pm$1.5 59.2
A2GNN 65.4$\pm$1.3 66.3$\pm$1.1 68.2$\pm$1.4 66.3$\pm$1.2 65.4$\pm$0.7 65.9$\pm$0.9 66.9$\pm$1.3 65.4$\pm$1.2 65.6$\pm$0.9 65.5$\pm$1.2 66.1$\pm$2.0 66.0$\pm$1.8 66.1
PA-BOTH 63.1$\pm$0.7 67.2$\pm$1.1 64.3$\pm$0.5 72.1$\pm$1.8 66.3$\pm$0.7 64.1$\pm$1.2 69.7$\pm$2.1 67.5$\pm$1.8 61.2$\pm$1.4 67.7$\pm$2.3 61.2$\pm$1.6 65.5$\pm$0.6 65.9
DeSGDA 76.7$\pm$0.8 84.6$\pm$0.9 69.4$\pm$0.6 85.2$\pm$1.5 76.2$\pm$1.1 83.9$\pm$1.2 69.9$\pm$0.6 76.3$\pm$1.4 75.9$\pm$1.0 73.5$\pm$1.3 76.3$\pm$1.6 68.3$\pm$0.7 76.4
5 Experiment
5.1 Experimental Settings
Dataset. To demonstrate the effectiveness of DeSGDA, we conduct extensive experiments on four widely-used graph classification datasets from TUDataset  [^1] https://chrsmrrs.github.io/datasets/, including PROTEINS Dobson and Doig (2003), NCI1 Wale et al. (2008), FRANKENSTEIN Orsini et al. (2015), and MUTAGENICITY Kazius et al. (2005). To better address the variation in domain distributions within each dataset, we divided them into source and target domains based on the edge density, node density, and graph flux (i.e., the ratio of the number of nodes to the number of edges). The specific statistics, distribution visualization, and details introduction of experimental datasets are presented in Appendix E.
Baselines. We compare DeSGDA with competitive baselines on the aforementioned datasets, including one graph kernel method: WL subtree Shervashidze et al. (2011); four general graph neural networks: GCN Kipf and Welling (2017), GIN Xu et al. (2018), CIN Bodnar et al. (2021) and GMT Baek et al. (2021); two spiking graph neural networks: SpikeGCN Zhu et al. (2022) and DRSGNN Zhao et al. (2024); three recent domain adaptation methods: CDAN Long et al. (2018), ToAlign Wei et al. (2021b), and MetaAlign Wei et al. (2021a); and six graph domain adaptation methods: DEAL Yin et al. (2022), CoCo Yin et al. (2023), SGDA Qiao et al. (2023), DGDA Cai et al. (2024), A2GNN Liu et al. (2024b) and PA-BOTH Liu et al. (2024c). More details about the compared baselines can be found in Appendix F.
Implementation Details. DeSGDA and all baseline models are implemented using PyTorch [^2] https://pytorch.org/ and PyTorch Geometric [^3] https://www.pyg.org/. For DeSGDA, we deploy the GIN Xu et al. (2018) as the backbone of the degree-aware personalized spiking graph encoder, incorporating a mean-pooling layer for the readout function. We conduct experiments for DeSGDA and all baselines on NVIDIA A100 GPUs for a fair comparison, where the learning rate of Adam optimizer set to ${10^{-4}}$, hidden embedding dimension 256, weight decay ${10^{-12}}$, and GNN layers 4. Additionally, DeSGDA and all baseline models are trained using all labeled source samples and evaluated on unlabeled target samples Wu et al. (2020); Pang et al. (2023); Liu et al. (2023a); You et al. (2023). The performances of all models are measured and averaged on all samples for five runs.
[…]
WL subtree 75.9 70.4 64.3 63.9 60.6 64.7 73.2 78.9 66.8 69.2 74.2 72.9 69.6
GCN 49.2$\pm$1.7 55.8$\pm$1.5 46.8$\pm$0.5 54.6$\pm$2.2 43.4$\pm$0.6 46.7$\pm$0.2 50.0$\pm$1.8 57.2$\pm$2.2 44.2$\pm$0.4 51.6$\pm$0.8 62.7$\pm$2.1 56.8$\pm$1.3 51.6
GIN 68.8$\pm$2.5 70.6$\pm$1.0 64.2$\pm$1.1 67.2$\pm$2.4 62.2$\pm$1.8 62.5$\pm$1.5 68.7$\pm$2.4 72.5$\pm$0.6 63.3$\pm$1.6 65.2$\pm$0.6 62.4$\pm$0.3 70.9$\pm$0.5 66.6
GMT 66.7$\pm$0.3 58.2$\pm$0.5 63.9$\pm$0.3 58.4$\pm$0.3 63.8$\pm$0.4 56.7$\pm$0.5 63.9$\pm$0.7 66.3$\pm$1.0 63.8$\pm$1.1 66.6$\pm$0.4 63.8$\pm$0.2 62.6$\pm$0.7 62.9
CIN 58.7$\pm$2.4 54.9$\pm$0.2 52.0$\pm$0.3 54.8$\pm$0.1 56.6$\pm$0.2 54.9$\pm$0.1 52.9$\pm$1.4 52.8$\pm$0.5 56.5$\pm$0.6 52.8$\pm$2.1 58.5$\pm$0.8 56.6$\pm$1.4 55.1
SpikeGCN 58.9$\pm$0.9 65.2$\pm$1.2 60.8$\pm$1.3 62.0$\pm$1.5 62.3$\pm$0.8 58.7$\pm$1.6 64.1$\pm$0.6 66.7$\pm$1.3 60.5$\pm$1.7 63.8$\pm$1.4 62.2$\pm$2.1 61.1$\pm$1.5 62.1
DRSGNN 58.0$\pm$0.6 64.3$\pm$1.1 61.2$\pm$0.8 62.2$\pm$1.0 62.9$\pm$1.5 64.0$\pm$1.3 60.6$\pm$1.6 64.0$\pm$1.4 67.6$\pm$2.1 62.4$\pm$1.9 71.3$\pm$2.3 68.8$\pm$2.0 63.9
CDAN 64.0$\pm$1.1 68.1$\pm$0.3 60.1$\pm$0.5 64.0$\pm$1.3 60.9$\pm$0.2 57.8$\pm$1.0 64.3$\pm$1.6 61.2$\pm$0.2 66.3$\pm$0.7 59.0$\pm$0.5 68.9$\pm$0.3 63.7$\pm$0.6 63.2
ToAlign 52.8$\pm$0.5 54.8$\pm$0.2 48.2$\pm$1.1 54.8$\pm$1.5 44.0$\pm$0.8 54.8$\pm$2.0 48.2$\pm$1.7 52.8$\pm$0.6 44.0$\pm$0.2 52.8$\pm$0.3 44.0$\pm$1.0 48.2$\pm$1.2 50.0
MetaAlign 63.1$\pm$0.3 63.8$\pm$1.3 58.9$\pm$2.4 58.5$\pm$0.4 59.1$\pm$2.1 59.2$\pm$1.6 70.1$\pm$0.8 63.3$\pm$1.4 66.5$\pm$2.7 60.9$\pm$1.1 71.4$\pm$0.2 67.5$\pm$0.8 63.5
DEAL 70.7$\pm$0.9 72.3$\pm$0.2 69.9$\pm$0.8 68.9$\pm$0.7 64.1$\pm$0.6 65.6$\pm$0.9 71.9$\pm$0.4 69.9$\pm$1.7 70.6$\pm$0.4 66.5$\pm$0.3 71.6$\pm$0.7 69.9$\pm$0.5 69.3
CoCo 64.0$\pm$1.3 63.9$\pm$0.6 65.8$\pm$1.8 59.9$\pm$1.7 62.2$\pm$2.1 60.6$\pm$1.6 65.0$\pm$2.1 64.8$\pm$1.4 60.0$\pm$0.8 61.3$\pm$0.5 68.5$\pm$0.4 67.1$\pm$0.6 63.6
SGDA OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
DGDA OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
A2GNN 58.9$\pm$0.9 60.1$\pm$0.7 59.8$\pm$1.2 59.4$\pm$1.0 62.3$\pm$1.5 60.9$\pm$1.6 61.6$\pm$1.3 59.9$\pm$1.9 64.9$\pm$1.6 62.9$\pm$2.1 65.4$\pm$1.5 63.3$\pm$2.3 61.7
PA-BOTH 61.1$\pm$0.5 60.9$\pm$0.4 61.6$\pm$0.6 61.2$\pm$0.8 60.8$\pm$0.6 61.5$\pm$0.5 62.2$\pm$1.0 61.9$\pm$0.7 61.8$\pm$1.1 61.1$\pm$0.9 60.9$\pm$0.8 61.3$\pm$1.2 61.6
DeSGDA 68.5$\pm$1.2 71.4$\pm$1.3 70.1$\pm$0.7 69.0$\pm$1.1 68.9$\pm$1.0 66.3$\pm$1.4 69.6$\pm$1.3 70.2$\pm$1.7 71.1$\pm$1.6 69.3$\pm$1.8 74.4$\pm$1.6 70.0$\pm$1.9 70.1
[…]
WL subtree 65.7 71.8 57.9 71.1 47.4 43.4 65.5 75.1 45.3 34.9 52.7 49.8 56.7
GCN 70.6$\pm$2.1 60.3$\pm$1.5 60.5$\pm$3.4 62.3$\pm$1.1 58.4$\pm$0.5 43.2$\pm$0.2 63.8$\pm$1.0 70.3$\pm$0.3 50.6$\pm$1.0 32.8$\pm$0.3 50.1$\pm$0.4 42.2$\pm$0.2 55.4
GIN 66.7$\pm$2.1 73.7$\pm$2.4 57.3$\pm$3.1 69.4$\pm$2.3 58.6$\pm$0.4 43.1$\pm$0.3 66.4$\pm$2.7 74.8$\pm$1.8 42.2$\pm$1.6 33.5$\pm$1.0 57.4$\pm$0.8 43.9$\pm$2.3 57.2
GMT 67.3$\pm$0.3 56.8$\pm$0.4 58.0$\pm$0.2 56.8$\pm$0.2 60.6$\pm$0.3 56.8$\pm$0.5 57.8$\pm$0.1 67.3$\pm$0.1 39.5$\pm$0.3 67.3$\pm$0.2 39.5$\pm$0.5 57.8$\pm$0.4 57.1
CIN 67.6$\pm$0.4 63.7$\pm$2.1 58.9$\pm$1.0 56.8$\pm$0.4 63.6$\pm$0.4 59.5$\pm$2.7 58.7$\pm$1.2 67.0$\pm$0.5 61.7$\pm$1.6 67.8$\pm$0.7 62.2$\pm$2.1 56.0$\pm$1.3 61.9
SpikeGCN 67.2$\pm$0.5 57.2$\pm$1.2 57.9$\pm$0.8 57.1$\pm$0.6 61.1$\pm$1.3 58.9$\pm$1.6 60.0$\pm$1.2 67.2$\pm$0.9 53.9$\pm$2.1 64.4$\pm$0.8 57.8$\pm$1.0 59.9$\pm$1.2 60.2
DRSGNN 67.4$\pm$0.4 58.4$\pm$1.0 59.0$\pm$1.2 57.4$\pm$0.5 62.3$\pm$1.1 60.4$\pm$1.3 61.1$\pm$1.6 67.9$\pm$1.5 56.2$\pm$1.8 66.2$\pm$2.1 60.9$\pm$1.4 58.6$\pm$2.5 61.3
CDAN 72.9$\pm$0.4 72.7$\pm$0.4 65.4$\pm$0.3 72.9$\pm$0.1 61.2$\pm$0.3 70.3$\pm$0.2 65.7$\pm$0.4 72.7$\pm$0.1 61.0$\pm$0.1 72.1$\pm$1.2 60.7$\pm$0.2 65.3$\pm$0.6 67.7
ToAlign 32.7$\pm$2.0 43.2$\pm$0.1 42.2$\pm$1.3 43.2$\pm$0.9 60.5$\pm$0.7 43.2$\pm$1.2 42.2$\pm$0.4 32.7$\pm$1.2 60.5$\pm$0.9 32.7$\pm$0.3 60.5$\pm$0.7 42.2$\pm$0.4 44.7
MetaAlign 67.3$\pm$0.7 56.8$\pm$0.2 57.8$\pm$0.6 56.8$\pm$0.4 60.5$\pm$1.3 56.8$\pm$0.8 57.8$\pm$1.1 67.3$\pm$1.2 60.5$\pm$0.4 67.3$\pm$0.6 60.5$\pm$0.7 57.8$\pm$0.6 60.6
DEAL 75.0$\pm$0.9 76.3$\pm$2.4 65.9$\pm$1.8 77.5$\pm$2.7 60.3$\pm$4.5 69.7$\pm$3.2 67.2$\pm$1.5 75.3$\pm$1.7 57.4$\pm$4.1 71.1$\pm$2.2 65.7$\pm$2.7 66.4$\pm$1.6 69.0
CoCo 74.2$\pm$1.7 74.3$\pm$0.6 65.9$\pm$1.2 72.7$\pm$2.1 61.1$\pm$0.2 71.0$\pm$1.7 68.6$\pm$0.3 75.9$\pm$0.2 60.7$\pm$0.2 73.9$\pm$0.4 59.7$\pm$1.1 67.3$\pm$0.8 68.8
SGDA 55.9$\pm$0.6 57.1$\pm$0.5 56.1$\pm$0.4 54.6$\pm$0.8 55.8$\pm$1.1 57.7$\pm$0.6 54.3$\pm$0.7 53.6$\pm$1.3 59.1$\pm$0.8 56.7$\pm$0.6 55.4$\pm$1.2 53.8$\pm$0.5 55.9
DGDA OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
A2GNN 55.9$\pm$0.7 55.7$\pm$0.4 56.6$\pm$0.6 57.1$\pm$1.0 56.1$\pm$1.2 55.8$\pm$0.5 56.5$\pm$0.7 55.5$\pm$0.4 55.9$\pm$0.8 56.2$\pm$0.6 56.5$\pm$1.5 56.0$\pm$0.5 56.2
PA-BOTH 56.4$\pm$0.5 55.9$\pm$0.6 56.0$\pm$0.5 56.4$\pm$0.4 56.3$\pm$0.6 57.7$\pm$0.7 56.6$\pm$0.2 58.8$\pm$0.9 56.9$\pm$0.7 57.2$\pm$0.3 56.5$\pm$0.5 58.3$\pm$0.8 56.9
DeSGDA 75.2$\pm$0.8 74.3$\pm$1.2 68.0$\pm$1.5 73.6$\pm$2.0 62.3$\pm$1.7 71.2$\pm$2.5 68.8$\pm$1.3 76.0$\pm$1.2 61.9$\pm$2.2 71.4$\pm$1.6 62.8$\pm$1.2 65.7$\pm$0.9 69.3
5.2 Performance Comparision
We present the results of the proposed DeSGDA with all baseline models under the setting of graph domain adaptation on different datasets in Table 1, 2, 3. From these tables, we observe that: (1) The performance of graph domain adaptation methods surpasses that of graph and spike-based graph methods. We attribute this improvement to the fact that domain distribution shifts degrade the performance of traditional graph methods. (2) The graph domain adaptation methods (DEAL and CoCo) outperform the spike-based graph methods (SpikeGCN and DRSGNN), underscoring the necessity of the research in spiking graph domain adaptation. (3) Our DeSGDA outperforms all baselines for most cases, which demonstrates its superiority over other methods. The remarkable performance of DeSGDA lies in two main reasons: (i) The degree-aware personalized spiking representations can capture more expressive information for graph classification by dynamically adjusting the thresholds of nodes in SNNs. (ii) The adversarial distribution alignment effectively addresses domain discrepancies by adversarially training the encoder and domain discriminator to align feature spaces. Moreover, the pseudo-label distillation aids in updating the source degree thresholds in the target domain, thereby ensuring optimal performance. More results evaluated on other datasets can be found in Appendix G.1.
Figure 2: The performance with different GNN architectures on PROTEINS.
Additionally, we conduct experiments to explore the flexibility of the proposed DeSGDA. Specifically, we replace the backbone of the degree-aware personalized spiking graph encoder (GIN) with various GNNs methods (i.e., GCN and GAT), and the results are shown in Figure 2. From the results, we observe that GIN consistently outperforms other GNNs architectures in most cases, demonstrating its powerful representation capability. This phenomenon also justifies our choice of using GIN to enhance the performance of the proposed DeSGDA. More results are reported in Appendix G.1.
5.3 Energy Efficiency Analysis
Figure 3: Energy consumption of DeSGDA on different datasets.
To assess the energy efficiency of DeSGDA, we use the metric from Zhu et al. (2022) and quantify the energy consumption for graph classification in the inference stage. Specifically, the graph domain adaption methods are evaluated on GPUs (NVIDIA A100), and the spiking-based methods are evaluated on neuromorphic chips (ROLLS Indiveri et al. (2015)) following Zhu et al. (2022). The results are shown in Figure 3, from the results, we find that compared with traditional graph domain adaptation methods, the spike-based methods (DeSGDA and DRSGNN) have significantly lower energy consumption, demonstrating the superior energy efficiency of SGNs. Moreover, although the energy consumption of DeSGDA is slightly higher than DRSGNN due to additional computations required for domain adaptation, the performance improvement justifies the deployment of DeSGDA in low-power devices.
Table 4: The results of ablation studies on PROTEINS (source → target). Bold results indicate the best performance. Bold results indicate the best performance.
[…]
DeSGDA w/o CA 72.3 72.6 71.9 75.1 74.1 71.3 71.5 70.4 71.6 70.5 78.3 71.5 72.6
DeSGDA w/o PL 72.1 68.7 67.8 69.4 63.7 55.7 68.3 69.5 70.2 69.6 76.7 66.9 68.3
DeSGDA w/o CF 56.7 54.2 60.7 62.1 73.4 68.3 55.3 69.5 76.7 65.5 73.7 63.3 65.7
DeSGDA w/o TL 61.7 43.8 56.1 71.2 49.5 69.0 42.4 68.3 73.7 64.0 70.5 43.5 59.5
DeSGDA w/ PT 71.5 68.3 66.1 72.4 66.7 70.1 67.9 70.8 66.3 71.8 74.1 72.9 70.0
DeSGDA w/ CL 74.9 73.3 73.7 75.1 77.0 71.3 73.9 70.0 78.1 77.6 78.8 74.9 75.1
DeSGDA 78.7 78.4 74.8 77.6 79.5 76.7 74.9 71.2 79.5 72.8 81.0 75.1 76.4
5.4 Ablation Study
We conduct ablation studies to examine the contributions of each component in the proposed DeSGDA: (1) DeSGDA w/o CA: It removes the adversarial distribution alignment module; (2) DeSGDA w/o PL: It removes the pseudo-label distilling module; (3) DeSGDA w/o CF: It removes the classification loss $\mathcal{L}_{S}$; (4) DeSGDA w/o TL: It utilizes the global thresholds on all nodes; (5) DeSGDA w/ PT: It deploys the adaptive perturbations Yin et al. (2022) on source data for alignment; (6) DeSGDA w/ CL: It replaces the adversarial learning with the cross-domain contrastive learning Yin et al. (2023).
Experimental results are shown in Table 4. From the table, we find that: (1) DeSGDA outperforms DeSGDA w/o CA, DeSGDA w/o PL, and DeSGDA w/o CF, demonstrating that the adversarial distribution alignment module can effectively reduce domain discrepancies, ensuring well-aligned feature spaces between source and target domains. Additionally, the pseudo-label distillation module can address the variance in thresholds across domains, while the classification loss $\mathcal{L}_{S}$ enables DeSGDA to effectively learn from labeled source data and generalize to the target domain. (2) DeSGDA w/o TL shows lower performance compared to DeSGDA, showing that the degree-aware thresholds, which are iteratively updated during model training, can resolve the issue of the inflexible architecture in SGNs. By using these thresholds, DeSGDA can effectively learn meaningful representations for nodes with various degrees. (3) DeSGDA outperforms DeSGDA w/ PT and DeSGDA w/ CL. We attribute that the adaptive perturbations method (DeSGDA w/ PT) can not effectively help DeSGDA overcome the domain discrepancy. Additionally, the cross-domain contrastive learning method (DeSGDA w/ CL) is less effective at aligning the source and target distribution compared to adversarial learning. More ablation results on other datasets are reported in Appendix G.2.
(a) Time Latency $\tau$
(b) Initial Threshold $V_{th}^{degree}$
(c) Balance Ratio $\alpha$
Figure 4: Hyperparameter sensitivity analysis of time latency $\tau$, initial threshold $V_{th}^{degree}$ in SNNs, and balance ratio $\alpha$ on PROTEINS.
5.5 Sensitivity Analysis
We study the sensitivity analysis of DeSGDA with respect to the impact of its hyperparameters: time latency $\tau$, degree threshold value $V_{th}^{degree}$ in SNNs, and balance ratio $\alpha$, which plays a crucial role in the performance of DeSGDA. In particular, $\tau$ controls the number of SNNs propagation steps; $V_{th}^{degree}$ determines when a neuron fires; $\alpha$ governs the changing ratio of degree-aware thresholds.
Figure 4 illustrates how $\tau$, $V_{th}^{degree}$, and $\alpha$ affects the performance of DeSGDA on the PROTEINS dataset. More results on other datasets are shown in Appendix G.3. We vary $\tau$ within the range of $\{5,6,7,8,9,10\}$, $V_{th}^{degree}$ in $\{0.05,0.1,0.2,0.5,1.0,2.0,5.0\}$, and $\alpha$ in $\{0.1,0.3,0.5,0.7,0.9\}$. From the results, we observe that: (1) The performance of DeSGDA in Figure 4(a) generally exhibits an increasing trend at the beginning and then stabilizes when $\tau$ is greater than 8. We attribute this to smaller values of $\tau$ potentially losing important information for representation, while larger values significantly increase model complexity. To balance performance and complexity, we set $\tau$ to 8 as default. (2) Figure 4(b) indicates an initial increase followed by a decreasing trend in performance as $V_{th}^{degree}$ increases. This trend occurs because a lower threshold may trigger more spikes for high-degree nodes, leading to a drastic change in the threshold, which can degrade performance. Conversely, a higher threshold for low-degree nodes could result in fewer spikes, affecting the model’s ability to process information effectively. Therefore, we set $V_{th}^{degree}$ to 0.5 as default. (3) From Figure 4(c), we find that the performance of DeSGDA initially increases and then decreases as $\alpha$ increases. The potential reason is that the smaller $\alpha$ may delay the updating of the threshold, leading to performance degradation. Contrarily, a larger $\alpha$ tends to introduce more spikes that change dynamically at each step, resulting in instability in the model’s performance. Therefore, we set $\alpha$ to 0.5 as default.
6 Conclusion
In this paper, we first propose the problem of spiking graph domain adaptation and introduce a novel framework DeSGDA for graph classification. This framework enhances the adaptability and performance of SGNs through three key aspects: node degree-aware personalized spiking representation, adversarial feature distribution alignment, and pseudo-label distillation. Our approach enables more expressive information capture through degree-dependent spiking thresholds, aligns feature distributions via adversarial training, and utilizes pseudo-labels to leverage unlabeled data effectively. The extensive experimental validation across benchmark datasets has demonstrated that DeSGDA not only surpasses existing methods in accuracy but also maintains efficient energy consumption, making it a promising solution for advancing the domain adaptation capabilities of spiking graph networks. In the future, we will apply SGNs in the scenarios of source-free domain adaptation and domain generalization.
Appendix A Proof of Theorem 2
Theorem 2 Assuming that the learned discriminator is $C_{g}$-Lipschitz continuous as described in Theorem 1, the graph feature extractor $f$ (also referred to as GNN) is $C_{f}$-Lipschitz that $||f||_{Lip}=\max_{G_{1},G_{2}}\frac{||f(G_{1})-f(G_{2})||_{2}}{\eta(G_{1},G_{2})}=C_{f}$ for some graph distance measure $\eta$ and the loss function bounded by $C>0$. Let :=$\mathcal{H}:=\{h:\mathcal{G}\rightarrow\mathcal{Y}\}$ be the set of bounded real-valued functions with the pseudo-dimension $Pdim(\mathcal{H})=d$ that $h=g\circ f\in\mathcal{H}$, and provided the spike training data set $S_{n}=\{(\mathbf{X}_{i},y_{i})\in\mathcal{X}\times\mathcal{Y}\}_{i\in[n]}$ drawn from $\mathcal{D}^{s}$, with probability at least $1-\delta$ the following inequality:
$\epsilon_{T}(h,\hat{h}_{T}(\mathbf{X}))\leq$ $\hat{\epsilon}_{S}(h,\hat{h}_{S}(S_{n}))+2\mathbb{E}\left[\sup\frac{1}{N_{S}}\sum_{i=1}^{N_{S}}\epsilon_{i}h(\mathbf{X}_{i},y_{i},p_{i})\right]+C\sqrt{\frac{ln(2/\delta)}{N_{S}}}$ (12)
$+\min\left(|\epsilon_{S}(h,\hat{h}_{S}(\mathbf{X}))-\epsilon_{S}(h,\hat{h}_{T}(\mathbf{X}))|,|\epsilon_{T}(h,\hat{h}_{S}(\mathbf{X}))-\epsilon_{T}(h,\hat{h}_{T}(\mathbf{X}))|\right)$
$+2C_{f}C_{g}W_{1}\left(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G)\right),$
where the (empirical) source and target risks are $\hat{\epsilon}_{S}(h,\hat{h})=\frac{1}{N_{S}}\sum_{n=1}^{N_{S}}|h(G_{n})-\hat{h}(G_{n})|$ and $\epsilon_{T}(h,\hat{h})=\mathbb{E}_{\mathbb{P}_{T}(G}\{|h(G)-\hat{h}(G)|\}$, respectively, where $\hat{h}:\mathcal{G}\rightarrow\mathcal{Y}$ is the labeling function for graphs and $\omega=\min_{||g||_{Lip}\leq C_{g},||f||_{Lip}\leq C_{f}}\{\epsilon_{S}(h,\hat{h})+\epsilon_{T}(h,\hat{h})\}$, $\epsilon_{i}$ is the Rademacher variable and $p_{i}$ is the $i^{th}$ row of $\mathbf{P}$, which is the probability matrix with:
$\mathbf{P}_{kt}=\left\{\begin{array}[]{lr}\exp\left(\frac{u_{k}(t)-V_{th}}{\sigma(u_{k}(t)-u_{reset})}\right),&if\quad u_{\theta}\leq u(t)\leq V_{th},\\ 0,&if\quad u_{reset}\leq u_{k}(t)\leq u_{\theta}.\end{array}\right.$ (13)
Proof. Before showing the designated lemma, we first introduce the following inequality to be used that:
$|\epsilon_{S}(h,\hat{h}_{S})-\epsilon_{T}(h,\hat{h}_{T})|$ $=|\epsilon_{S}(h,\hat{h}_{S})-\epsilon_{S}(h,\hat{h}_{T})+\epsilon_{S}(h,\hat{h}_{T})-\epsilon_{T}(h,\hat{h}_{T})|$ (14)
$\leq|\epsilon_{S}(h,\hat{h}_{S})-\epsilon_{S}(h,\hat{h}_{T})|+|\epsilon_{S}(h,\hat{h}_{T})-\epsilon_{T}(h,\hat{h}_{T})|$
$\overset{(a)}{\leq}|\epsilon_{S}(h,\hat{h}_{S})-\epsilon_{S}(h,\hat{h}_{T})|+2C_{f}C_{g}W_{1}\left(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G)\right),$
where $(a)$ results from Shen et al. [2018] Theorem 1 with the assumption $\max(||h||_{Lip},\max_{G_{1},G_{2}}\frac{|\hat{h}_{D}(G_{1})-\hat{h}_{D}(G_{2})|}{\eta(G_{1},G_{2})})\leq C_{f}C_{g}$, $D\in\{S,T\}$. Similarly, we obtain:
$|\epsilon_{S}(h,\hat{h}_{S})-\epsilon_{T}(h,\hat{h}_{T})|\leq|\epsilon_{T}(h,\hat{h}_{S})-\epsilon_{T}(h,\hat{h}_{T})|+2C_{f}C_{g}W_{1}(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G)).$ (15)
We therefore combine them into:
$|\epsilon_{S}(h,\hat{h}_{S})-\epsilon_{T}(h,\hat{h}_{T})|\leq$ $2C_{f}C_{g}W_{1}(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G))$ (16)
$+\min\left(|\epsilon_{S}(h,\hat{h}_{S})-\epsilon_{S}(h,\hat{h}_{T})|,|\epsilon_{T}(h,\hat{h}_{S})-\epsilon_{T}(h,\hat{h}_{T})|\right),$
i.e. the following holds to bound the target risk $\epsilon_{T}(h,\hat{h}_{T})$:
$\epsilon_{T}(h,\hat{h}_{T})\leq$ $\epsilon_{S}(h,\hat{h}_{S})+2C_{f}C_{g}W_{1}\left(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G)\right)$ (17)
$+\min\left(|\epsilon_{S}(h,\hat{h}_{S})-\epsilon_{S}(h,\hat{h}_{T})|,|\epsilon_{T}(h,\hat{h}_{S})-\epsilon_{T}(h,\hat{h}_{T})|\right).$
We next link the bound with the empirical risk and labeled sample size by showing, with probability at least $1-\delta$ that:
$\epsilon_{T}(h,\hat{h}_{T})\leq$ $\epsilon_{S}(h,\hat{h}_{S})+2C_{f}C_{g}W_{1}\left(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G)\right)$ (18)
$+\min\left(|\epsilon_{S}(h,\hat{h}_{S})-\epsilon_{S}(h,\hat{h}_{T})|,|\epsilon_{T}(h,\hat{h}_{S})-\epsilon_{T}(h,\hat{h}_{T})|\right)$
The $\hat{h}$ above is the abbreviation of $\hat{h}(x)$, which means the input is the continuous feature. Provided the spike training data set $S_{n}=\{(\mathbf{X}_{i},y_{i})\in\mathcal{X}\times\mathcal{Y}\}_{i\in[n]}$ drawn from $\mathcal{D}$, and motivated by Yin et al. [2024], we have:
$\lim_{\tau\to\infty}P\left(\hat{h}(S_{n})_{\tau,i}>\hat{h}(\mathbf{X}_{\tau,i})+\epsilon\right)\leq e^{-\epsilon^{2}/2(\sigma+\hat{w}_{i}\epsilon/3)},$ (19)
where $\hat{w}_{i}=max\{w_{i1},\cdots,w_{id}\}$ and $h(\mathbf{x}_{ij})=\sum_{j=1}^{d}w_{ij}\bm{x}_{ij}$. From Equation 2, we observe that as $\tau\to\infty$, the difference between spike and real-valued features will be with the probability of $p=e^{-\epsilon^{2}/2(\sigma+\hat{w}_{i}\epsilon/3)}$ to exceed the upper and lower bounds.
Furthermore, motivated by the techniques given by Bartlett and Mendelson [2002], we have:
$\epsilon_{S}(h,\hat{h}_{S}(S_{n}))\leq\hat{\epsilon}_{S}(h,\hat{h}_{S}(S_{n}))+\underbrace{\sup[\epsilon_{S}(h,\hat{h}_{S}(S_{n}))-\hat{\epsilon}_{S}(h,\hat{h}_{S}(S_{n}))]}_{R(S_{n},\mathbf{P})},$ (20)
where $\mathbf{P}$ is the probability matrix with:
$\mathbf{P}_{kt}=\left\{\begin{array}[]{lr}\exp\left(\frac{u_{k}(t)-V_{th}}{\sigma(u_{k}(t)-u_{reset})}\right),&if\quad u_{\theta}\leq u(t)\leq V_{th},\\ 0,&if\quad u_{reset}\leq u_{k}(t)\leq u_{\theta},\end{array}\right.$ (21)
where $k$ indicates the $k-th$ spiking neuron and the membrane threshold $u_{theta}$ is relative to the excitation probability threshold $p_{\theta}\in(0,1]$. Let $p_{k}$ is the $k-th$ row vector of $\mathbf{P}$. Thus, we have the probability at least $1-e^{-\epsilon^{2}/2(\sigma+\hat{w}_{i}\epsilon/3)}$ to hold:
$\epsilon_{S}(h,\hat{h}_{S}(\mathbf{X}_{n}))\leq\hat{\epsilon}_{S}(h,\hat{h}_{S}(S_{n}))+\underbrace{\sup[\epsilon_{S}(h,\hat{h}_{S}(S_{n}))-\hat{\epsilon}_{S}(h,\hat{h}_{S}(S_{n}))]}_{R(S_{n},\mathbf{P})},$ (22)
Let $S^{\prime}_{n}$ denote the sample set that the $i^{th}$ sample $(\mathbf{X}_{i},y_{i})$ is replaced by $(\mathbf{X}^{\prime}_{i},y^{\prime}_{i})$, and correspondingly $\mathbf{P}^{\prime}$ is the possibility matrix that the $i^{th}$ row vector $p_{i}$ is replaced by $p^{\prime}_{i}$, for $i\in[n]$. For the loss function bounded by $C>0$, we have:
$\left\{\begin{array}[]{lr}|R(S_{n},\mathbf{P})-R(S^{\prime}_{n},\mathbf{P})|\leq C/n,&\\ |R(S_{n},\mathbf{P})-R(S_{n},\mathbf{P}^{\prime})|\leq C/n.&\end{array}\right.$ (23)
From McDiarmid’s inequality McDiarmid et al. [1989], with probability at least $1-\delta$, we have:
$R(S_{n},\mathbf{P})\leq\mathbb{E}_{S_{n}\in\mathcal{D},\mathbf{P}}[R(S_{n},\mathbf{P})]+C\sqrt{\frac{ln(2/\delta)}{N_{S}}}.$ (24)
It is observed that:
$R(S_{n},\mathbf{P})=\sup\mathbb{E}_{\tilde{S}_{n}\in\mathcal{D},\tilde{\mathbf{P}}}[\hat{\epsilon}(\hat{h}(S_{n});\tilde{S}_{n},\tilde{\mathbf{P}})-\tilde{\mathbf{P}}[\hat{\epsilon}(\hat{h}(S_{n});{S}_{n},{\mathbf{P}})],$ (25)
where $\tilde{S}_{n}$ is another collection drawn from $\mathcal{D}$ as well as $\tilde{\mathbf{P}}$. Thus, we have
$\mathbb{E}_{S_{n}\in\mathcal{D},\mathbf{P}}[R(S_{n},\mathbf{P})]$ $\leq\mathbb{E}\left[\sup\left[\hat{\epsilon}(\hat{h}(S_{n});\tilde{S}_{n},\tilde{\mathbf{P}})-\tilde{\mathbf{P}}[\hat{\epsilon}(\hat{h}(S_{n});{S}_{n},{\mathbf{P}})\right]\right]$ (26)
$=\mathbb{E}\left[\sup\frac{1}{n}\sum_{i=1}^{n}[\hat{h}(\tilde{\mathbf{X}}_{i},\tilde{y}_{i},\tilde{p}_{i})-\hat{h}(\mathbf{X}_{i},y_{i},p_{i})]\right]$
$\leq 2\mathbb{E}\left[\sup\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}\hat{h}(\mathbf{X}_{i},y_{i},p_{i})\right],$
where $\epsilon_{i}$ is the Rademacher variable. Combining Eq. 23 24 26, we have:
$\epsilon_{S}(h,\hat{h}_{S}(\mathbf{X}_{n}))\leq\hat{\epsilon}_{S}(h,\hat{h}_{S}(S_{n}))+2\mathbb{E}\left[\sup\frac{1}{N_{S}}\sum_{i=1}^{N_{S}}\epsilon_{i}h(\mathbf{X}_{i},y_{i},p_{i})\right]+C\sqrt{\frac{ln(2/\delta)}{N_{S}}}.$ (27)
Finally, we have:
$\epsilon_{T}(h,\hat{h}_{T}(\mathbf{X}))\leq$ $\epsilon_{S}(h,\hat{h}_{S}(\mathbf{X}))+2C_{f}C_{g}W_{1}\left(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G)\right)$ (28)
$+\min\left(|\epsilon_{S}(h,\hat{h}_{S}(\mathbf{X}))-\epsilon_{S}(h,\hat{h}_{T}(\mathbf{X}))|,|\epsilon_{T}(h,\hat{h}_{S}(\mathbf{X}))-\epsilon_{T}(h,\hat{h}_{T}(\mathbf{X}))|\right)$
$\leq$ $\hat{\epsilon}_{S}(h,\hat{h}_{S}(S_{n}))+2\mathbb{E}\left[\sup\frac{1}{N_{S}}\sum_{i=1}^{N_{S}}\epsilon_{i}h(\mathbf{X}_{i},y_{i},p_{i})\right]+C\sqrt{\frac{ln(2/\delta)}{N_{S}}}$
$+\min\left(|\epsilon_{S}(h,\hat{h}_{S}(\mathbf{X}))-\epsilon_{S}(h,\hat{h}_{T}(\mathbf{X}))|,|\epsilon_{T}(h,\hat{h}_{S}(\mathbf{X}))-\epsilon_{T}(h,\hat{h}_{T}(\mathbf{X}))|\right)$
$+2C_{f}C_{g}W_{1}\left(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G)\right).$
Appendix B Proof of Theorem 3
Theorem 3 Under the assumption of Theorem 1, we further assume that there exists a small amount of i.i.d. samples with pseudo labels $\{(G_{n},Y_{n})\}_{n=1}^{N^{\prime}_{T}}$ from the target distribution $\mathbb{P}_{T}(G,Y)$ $(N^{\prime}_{T}\ll N_{S})$ and bring in the conditional shift assumption that domains have different labeling function $\hat{h}_{S}\neq\hat{h}_{T}$ and $\max_{G_{1},G_{2}}\frac{|\hat{h}_{D}(G_{1})-\hat{h}_{D}(G_{2})|}{\eta(G_{1},G_{2})}=C_{h}\leq C_{f}C_{g}(D\in\{S,T\})$ for some constant $C_{h}$ and distance measure $\eta$, and the loss function bounded by $C>0$. Let :=$\mathcal{H}:=\{h:\mathcal{G}\rightarrow\mathcal{Y}\}$ be the set of bounded real-valued functions with the pseudo-dimension $Pdim(\mathcal{H})=d$, and provided the spike training data set $S_{n}=\{(\mathbf{X}_{i},y_{i})\in\mathcal{X}\times\mathcal{Y}\}_{i\in[n]}$ drawn from $\mathcal{D}^{s}$, with probability at least $1-\delta$ the following inequality holds:
$\epsilon_{T}(h,\hat{h}_{T}(\mathbf{X}))\leq$ $\frac{N^{\prime}_{T}}{N_{S}+N^{\prime}_{T}}\hat{\epsilon}_{T}(h,\hat{h}_{T}(S))+\frac{N_{S}}{N_{S}+N^{\prime}_{T}}\Bigg{(}\hat{\epsilon}_{S}(h,\hat{h}_{S}(S))+2C_{f}C_{g}W_{1}\left(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G)\right)$ (29)
$+2\mathbb{E}\left[\sup\frac{1}{N_{S}}\sum_{i=1}^{N_{S}}\epsilon_{i}h(\mathbf{X}_{i},y_{i},p_{i})\right]+C\sqrt{\frac{ln(2/\delta)}{N_{S}}}$
$+\min\left(|\epsilon_{S}(h,\hat{h}_{S}(\mathbf{X})))-\epsilon_{S}(h,\hat{h}_{T}(\mathbf{X})))|,|\epsilon_{T}(h,\hat{h}_{S}(\mathbf{X})))-\epsilon_{T}(h,\hat{h}_{T}(\mathbf{X})))|\right)\Bigg{)}$
$\leq$ $\hat{\epsilon}_{S}(h,\hat{h}_{S}(S))+2\mathbb{E}\left[\sup\frac{1}{N_{S}}\sum_{i=1}^{N_{S}}\epsilon_{i}h(\mathbf{X}_{i},y_{i},p_{i})\right]+C\sqrt{\frac{ln(2/\delta)}{N_{S}}}$
$+2C_{f}C_{g}W_{1}\left(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G)\right)+\omega^{\prime}.$
where the (empirical) source and target risks are $\hat{\epsilon}_{S}(h,\hat{h})=\frac{1}{N_{S}}\sum_{n=1}^{N_{S}}|h(G_{n})-\hat{h}(G_{n})|$ and $\epsilon_{T}(h,\hat{h})=\mathbb{E}_{\mathbb{P}_{T}(G}\{|h(G)-\hat{h}(G)|\}$, respectively, where $\hat{h}:\mathcal{G}\rightarrow\mathcal{Y}$ is the labeling function for graphs and $\omega^{\prime}=\min_{||g||_{Lip}\leq C_{g},||f||_{Lip}\leq C_{f}}\{\epsilon_{S}(h,\hat{h})+\epsilon_{T}(h,\hat{h})\}$, $\epsilon_{i}$ is the Rademacher variable and $p_{i}$ is the $i^{th}$ row of $\mathbf{P}$, which is the probability matrix with:
$\mathbf{P}_{kt}=\left\{\begin{array}[]{lr}\exp\left(\frac{u_{k}(t)-V_{th}}{\sigma(u_{k}(t)-u_{reset})}\right),&if\quad u_{\theta}\leq u(t)\leq V_{th},\\ 0,&if\quad u_{reset}\leq u_{k}(t)\leq u_{\theta}.\end{array}\right.$ (30)
Proof. As proved in Theorem 2, we have:
$\epsilon_{T}(h,\hat{h}_{T}(\mathbf{X}))\leq$ $\hat{\epsilon}_{S}(h,\hat{h}_{S}(S_{n}))+2\mathbb{E}\left[\sup\frac{1}{N_{S}}\sum_{i=1}^{N_{S}}\epsilon_{i}h(\mathbf{X}_{i},y_{i},p_{i})\right]+C\sqrt{\frac{ln(2/\delta)}{N_{S}}}$ (31)
$+\min\left(|\epsilon_{S}(h,\hat{h}_{S}(\mathbf{X}))-\epsilon_{S}(h,\hat{h}_{T}(\mathbf{X}))|,|\epsilon_{T}(h,\hat{h}_{S}(\mathbf{X}))-\epsilon_{T}(h,\hat{h}_{T}(\mathbf{X}))|\right)$
$+2C_{f}C_{g}W_{1}\left(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G)\right).$
Similar with Eq. 27, there exists:
$\epsilon_{T}(h,\hat{h}_{T}(\mathbf{X}_{n}))\leq\hat{\epsilon}_{T}(h,\hat{h}_{T}(S_{n}))+2\mathbb{E}\left[\sup\frac{1}{N^{\prime}_{T}}\sum_{i=1}^{N^{\prime}_{T}}\epsilon_{i}h(\mathbf{X}_{i},y_{i},p_{i})\right]+C\sqrt{\frac{ln(2/\delta)}{N^{\prime}_{T}}}.$ (32)
Combining Eq. 31 and 32, we have:
$\epsilon_{T}(h,\hat{h}_{T}(\mathbf{X}))\overset{(a)}{\leq}$ $\frac{N^{\prime}_{T}}{N_{S}+N^{\prime}_{T}}\left(\hat{\epsilon}_{T}(h,\hat{h}_{T}(S))+2\mathbb{E}\left[\sup\frac{1}{N^{\prime}_{T}}\sum_{i=1}^{N^{\prime}_{T}}\epsilon_{i}h(\mathbf{X}_{i},y_{i},p_{i})\right]+C\sqrt{\frac{ln(2/\delta)}{N^{\prime}_{T}}}\right)$
$+\frac{N_{S}}{N_{S}+N^{\prime}_{T}}\left(\hat{\epsilon}_{S}(h,\hat{h}_{S}(S))+2\mathbb{E}\left[\sup\frac{1}{N_{S}}\sum_{i=1}^{N_{S}}\epsilon_{i}h(\mathbf{X}_{i},y_{i},p_{i})\right]+C\sqrt{\frac{ln(2/\delta)}{N_{S}}}\right)$
$+\frac{N_{S}}{N_{S}+N^{\prime}_{T}}\Bigg{(}2C_{f}C_{g}W_{1}\left(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G)\right)$
$+\min\left(|\epsilon_{S}(h,\hat{h}_{S}(\mathbf{X}))-\epsilon_{S}(h,\hat{h}_{T}(\mathbf{X}))|,|\epsilon_{T}(h,\hat{h}_{S}(\mathbf{X}))-\epsilon_{T}(h,\hat{h}_{T}(\mathbf{X}))|\right)\Bigg{)}$
[…]
$+\frac{N_{S}}{N_{S}+N^{\prime}_{T}}\Bigg{(}2C_{f}C_{g}W_{1}\left(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G)\right)$
$+\min\left(|\epsilon_{S}(h,\hat{h}_{S}(\mathbf{X})-\epsilon_{S}(h,\hat{h}_{T}((\mathbf{X}))|,|\epsilon_{T}(h,\hat{h}_{S}((\mathbf{X}))-\epsilon_{T}(h,\hat{h}_{T}((\mathbf{X}))|\right)\Bigg{)}$
$+\frac{N^{\prime}_{T}}{N_{S}+N^{\prime}_{T}}\left(2\mathbb{E}\left[\sup\frac{1}{N^{\prime}_{T}}\sum_{i=1}^{N^{\prime}_{T}}\epsilon_{i}h(\mathbf{X}_{i},y_{i},p_{i})\right]+C\sqrt{\frac{ln(2/\delta)}{N^{\prime}_{T}}}\right)$
$+\frac{N_{S}}{N_{S}+N^{\prime}_{T}}\left(2\mathbb{E}\left[\sup\frac{1}{N_{S}}\sum_{i=1}^{N_{S}}\epsilon_{i}h(\mathbf{X}_{i},y_{i},p_{i})\right]+C\sqrt{\frac{ln(2/\delta)}{N_{S}}}\right)$
$\overset{(b)}{\doteq}$ $\frac{N^{\prime}_{T}}{N_{S}+N^{\prime}_{T}}\hat{\epsilon}_{T}(h,\hat{h}_{T}(S))+\frac{N_{S}}{N_{S}+N^{\prime}_{T}}\hat{\epsilon}_{S}(h,\hat{h}_{S}(S))$
$+\frac{N_{S}}{N_{S}+N^{\prime}_{T}}\left(2\mathbb{E}\left[\sup\frac{1}{N_{S}}\sum_{i=1}^{N_{S}}\epsilon_{i}h(\mathbf{X}_{i},y_{i},p_{i})\right]+C\sqrt{\frac{ln(2/\delta)}{N_{S}}}\right)$
$+\frac{N_{S}}{N_{S}+N^{\prime}_{T}}\Bigg{(}2C_{f}C_{g}W_{1}\left(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G)\right)$
$+\min\left(|\epsilon_{S}(h,\hat{h}_{S}(\mathbf{X}))-\epsilon_{S}(h,\hat{h}_{T}(\mathbf{X}))|,|\epsilon_{T}(h,\hat{h}_{S}(\mathbf{X}))-\epsilon_{T}(h,\hat{h}_{T}(\mathbf{X}))|\right)\Bigg{)}$
$=$ $\frac{N^{\prime}_{T}}{N_{S}+N^{\prime}_{T}}\hat{\epsilon}_{T}(h,\hat{h}_{T}(S))+\frac{N_{S}}{N_{S}+N^{\prime}_{T}}\Bigg{(}\hat{\epsilon}_{S}(h,\hat{h}_{S}(S))+2C_{f}C_{g}W_{1}\left(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G)\right)$
$+2\mathbb{E}\left[\sup\frac{1}{N_{S}}\sum_{i=1}^{N_{S}}\epsilon_{i}h(\mathbf{X}_{i},y_{i},p_{i})\right]+C\sqrt{\frac{ln(2/\delta)}{N_{S}}}$
$+\min\left(|\epsilon_{S}(h,\hat{h}_{S}(\mathbf{X})))-\epsilon_{S}(h,\hat{h}_{T}(\mathbf{X})))|,|\epsilon_{T}(h,\hat{h}_{S}(\mathbf{X})))-\epsilon_{T}(h,\hat{h}_{T}(\mathbf{X})))|\right)\Bigg{)}$
where (a) is the outcome of applying the union bound with coefficient $\frac{N^{\prime}_{T}}{N_{S}+N^{\prime}_{T}}$, $\frac{N_{S}}{N_{S}+N^{\prime}_{T}}$ respectively; (b) additionally adopt the assumption $N^{\prime}_{T}\ll N_{S}$, following the sleight-of-hand in Li et al. [2021a] Theorem 3.2.
Due to the sampels are selected with high confidence, thus, we have the following assumption:
$\hat{\epsilon}_{T}\leq\epsilon_{T}\leq$ $\hat{\epsilon}_{S}(h,\hat{h}(\mathbf{X})))+2\mathbb{E}\left[\sup\frac{1}{N_{S}}\sum_{i=1}^{N_{S}}\epsilon_{i}h(\mathbf{X}_{i},y_{i},p_{i})\right]$ (33)
$+C\sqrt{\frac{ln(2/\delta)}{N_{S}}}+2C_{f}C_{g}W_{1}(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G))+\omega^{\prime},$
where $\omega^{\prime}=\min_{||g||_{Lip}\leq C_{g},||f||_{Lip}\leq C_{f}}\{\epsilon_{S}(h,\hat{h})+\epsilon_{T}(h,\hat{h})\}$, $\hat{\epsilon}_{T}$ is the empirical risk on the high confidence samples, $\epsilon_{T}$ is the empirical risk on the target domain. Besides, we have:
$\min(|\epsilon_{S}(h,\hat{h}_{S}(\mathbf{X})))-$ $\epsilon_{S}(h,\hat{h}_{T}(\mathbf{X})))|,|\epsilon_{T}(h,\hat{h}_{S}(\mathbf{X})))-\epsilon_{T}(h,\hat{h}_{T})|(\mathbf{X})))\leq$ (34)
$\min\left(\epsilon_{S}(h,\hat{h}_{S}(\mathbf{X})))+\epsilon_{T}(h,\hat{h}_{S}(\mathbf{X})))\right)$
Then,
$\epsilon_{T}(h,\hat{h}_{T}(\mathbf{X}))\leq$ $\frac{N^{\prime}_{T}}{N_{S}+N^{\prime}_{T}}\hat{\epsilon}_{T}(h,\hat{h}_{T}(S))+\frac{N_{S}}{N_{S}+N^{\prime}_{T}}\Bigg{(}\hat{\epsilon}_{S}(h,\hat{h}_{S}(S))+2C_{f}C_{g}W_{1}\left(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G)\right)$ (35)
$+2\mathbb{E}\left[\sup\frac{1}{N_{S}}\sum_{i=1}^{N_{S}}\epsilon_{i}h(\mathbf{X}_{i},y_{i},p_{i})\right]+C\sqrt{\frac{ln(2/\delta)}{N_{S}}}$
$+\min\left(|\epsilon_{S}(h,\hat{h}_{S}(\mathbf{X})))-\epsilon_{S}(h,\hat{h}_{T}(\mathbf{X})))|,|\epsilon_{T}(h,\hat{h}_{S}(\mathbf{X})))-\epsilon_{T}(h,\hat{h}_{T}(\mathbf{X})))|\right)\Bigg{)}$
$\leq$ $\hat{\epsilon}_{S}(h,\hat{h}_{S}(S))+2\mathbb{E}\left[\sup\frac{1}{N_{S}}\sum_{i=1}^{N_{S}}\epsilon_{i}h(\mathbf{X}_{i},y_{i},p_{i})\right]+C\sqrt{\frac{ln(2/\delta)}{N_{S}}}$
$+2C_{f}C_{g}W_{1}\left(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G)\right)+\omega^{\prime}.$
Appendix C Algorithm
Algorithm 1 Learning Algorithm of DeSGDA
Input: Source data $\mathcal{D}^{s}$; Target data $\mathcal{D}^{t}$.
Output: Node degree-aware personalized spiking graph encoder parameters $\theta$.
1:Initialize model parameters.
2:while not convergence do
3:     Sample mini-batches $\mathcal{B}^{s}$ and $\mathcal{B}^{t}$ from source and target data, respectively;
4:     Forward propagation $\mathcal{B}^{s}$ and $\mathcal{B}^{t}$ through node degree-aware personalized spiking graph encoder;
5:     Pseudo-label Distilling;
6:     Calculate the loss function by Eq. 11;
7:     Update model parameters through back propagation;
8:end while
Appendix D Complexity Analysis
Here we analyze the computational complexity of the proposed DeSGDA. The computational complexity primarily relies on node degree-aware personalized spiking representations. For a given graph $G,\|A\|_{0}$ denotes the number of nonzeros in the adjacency matrix. $d$ is the feature dimension. $L$ denote the layer number of GIN. $|V|$ is the number of nodes. $T$ denotes the number of time latency. The spiking graph encoder takes $\mathcal{O}\left(T\cdot L\cdot\left(\|A\|_{0}\cdot d+|V|\cdot d^{2}\right)\right)$ computational time for each graph. As a result, the complexity of our DeSGDA is proportional to both $|V|$ and $\|A\|_{0}$.
Figure 5: Visualization of different distributions on PROTEINS.
Figure 6: Visualization of different distributions on NCI1.
Figure 7: Visualization of different distributions on FRANKENSTEIN.
Figure 8: Visualization of different distributions on MUTAGNENICITY.
Appendix E Dataset
Table 5: Statistics of the experimental datasets.
Datasets Graphs Avg. Nodes Avg. Edges Classes
PROTEINS 1,113 39.1 72.8 2
NCI1 4,110 29.87 32.30 2
MUTAGENICITY 4,337 30.32 30.77 2
FRANKENSTEIN 4,337 16.9 17.88 2
E.1 Dataset Description
We conduct extensive experiments on four public benchmark graph datasets from TUDataset. The dataset statistics can be found in Table 5, and their details are shown as follows:
  • PROTEINS. The PROTEINS dataset Dobson and Doig [2003] consists of protein graphs. Each label indicates whether the protein graph is an enzyme. Nodes within these graphs represent amino acids connected by edges if neighbors along the amino acid sequence are spaced less than 6 Angstroms apart. We divide the PROTEINS dataset into four parts, i.e., P0, P1, P2, and P3, based on edge density, node density, and graph flux; the sub-datasets exhibit substantial domain disparities among them.
  • NCI1. The NCI1 Wale et al. [2008] dataset contains 4,100 chemical compounds with atoms as nodes and bonds as edges. Each label indicates the characteristics that hinder the growth of cancer cells. Like the PROTEINS dataset, we divide the NCI1 dataset into four parts, i.e., N0, N1, N2, and N3, based on edge density, node density, and graph flux.
  • FRANKENSTEIN. The FRANKENSTEIN Orsini et al. [2015] dataset consists of 4,337 molecular graphs, where nodes represent atoms and edges depict chemical bonds. Each graph is labeled to classify molecules based on their biological activity. Like the PROTEINS dataset, the entire FRANKENSTEIN dataset is divided into four segments (F0, F1, F2, and F3) based on edge density, node density, and graph flux.
  • MUTAGENICITY. The MUTAGENICITY Kazius et al. [2005] dataset contains 4,337 chemical compounds, each represented as a graph where nodes represent atoms and edges indicate bonds. Each graph can be used to identify mutagenic compounds, aiding studies in toxicology and chemical safety. Like the PROTEINS dataset, the entire MUTAGENICITY dataset is divided into four segments (M0, M1, M2, and M3) based on edge density, node density, and graph flux.
Table 6: The results of ablation studies on NCI1 (source → target). Bold results indicate the best performance per column.
[…]
DeSGDA w/o CA 58.9 62.6 61.7 60.1 58.2 62.0 63.4 61.7 63.5 60.7 68.3 64.0 62.1
DeSGDA w/o PL 62.8 65.1 55.0 56.9 55.3 54.4 58.6 62.9 61.2 62.2 59.7 64.3 59.7
DeSGDA w/o CF 52.7 62.6 58.5 56.5 56.3 63.8 63.1 52.7 65.4 65.1 58.2 63.7 60.0
DeSGDA w/o TL 64.6 62.8 60.6 54.9 44.9 65.4 60.8 64.8 56.1 64.1 57.8 54.0 59.3
DeSGDA w PT 64.4 59.2 64.6 62.2 64.5 60.1 61.1 60.8 62.0 64.2 62.7 65.5 62.7
DeSGDA w CL 65.3 70.1 68.9 65.2 65.6 63.0 67.6 64.6 68.9 66.8 68.6 70.7 67.1
DeSGDA 68.5 72.4 70.1 69.0 68.9 66.3 69.6 70.2 71.1 69.3 74.4 70.0 70.1
Table 7: The results of ablation studies on FRANKENSTEIN (source → target). Bold results indicate the best performance. Bold results indicate the best performance.
[…]
DeSGDA w/o CA 59.6 60.7 60.2 60.7 77.9 58.5 61.3 60.5 76.8 60.0 80.1 63.9 65.0
DeSGDA w/o PL 58.9 56.6 55.9 56.6 69.8 53.6 55.0 57.7 70.4 53.2 68.8 61.7 60.0
DeSGDA w/o CF 55.9 59.2 59.5 54.5 74.9 52.9 59.9 54.6 78.8 56.1 74.5 58.5 61.7
DeSGDA w/o TL 56.2 57.2 60.7 48.2 53.3 45.8 59.5 53.0 48.8 54.8 51.1 54.7 53.6
DeSGDA w PT 60.9 60.2 61.0 61.5 76.2 60.9 60.3 62.2 75.3 57.2 74.8 63.3 64.5
DeSGDA w CL 62.2 61.3 61.7 60.7 79.3 59.7 63.6 64.1 79.5 61.1 78.6 62.7 66.2
DeSGDA 63.5 64.1 62.6 63.0 81.1 62.4 65.5 65.4 81.9 60.8 82.0 65.9 68.1
E.2 Data processing
In our implementation, we process these four TUDatasets by adding a self-loop connection for each node. Additionally, we utilize one-hot embeddings to represent node attributes for datasets (FRANKENSTEIN and MUTAGENICITY) where node features are unavailable.
Appendix F Baselines
In this part, we introduce the details of the compared baselines as follows:
Graph kernel method. We compare DeSGDA with one graph kernel method:
  • WL subtree: The Weisfeiler-Lehman (WL) subtree Shervashidze et al. [2011] method is a graph kernel method, which calculates the graph similarity by a kernel function, where It encodes local neighborhood structures into subtree patterns, efficiently capturing the topology information contained in graphs.
Graph neural networks. We compare DeSGDA with four widely used general graph neural networks:
  • GCN: GCN Kipf and Welling [2017] is a spectral-based neural network that iteratively updates node representations by aggregating information from neighboring nodes, effectively capturing both local graph structure and node features.
  • GIN: GIN Xu et al. [2018] is a message-passing neural network designed to distinguish graph structures using an injective aggregation function, theoretically achieving the expressive power of the Weisfeiler-Lehman test.
  • CIN: CIN Bodnar et al. [2021] extends the Weisfeiler-Lehman framework by integrating cellular complexes into graph neural networks, allowing for the capture of higher-dimensional topological features.
  • GMT: GMT Baek et al. [2021] utilizes self-attention mechanisms to dynamically adjust the importance of nodes based on their structural dependencies, thereby enhancing both adaptability and performance.
Spiking graph neural networks. We compare DeSGDA with two spiking graph neural networks:
  • SpikeGCN: SpikeGCN Zhu et al. [2022] introduces an end-to-end framework designed to integrate the fidelity characteristics of SNNs with graph node representations.
  • DRSGNN: DRSGNN Zhao et al. [2024] dynamically adapts to evolving graph structures and relationships through a novel architecture that updates node representations in real-time..
Table 8: The results of ablation studies on MUTAGENICITY (source → target). Bold results indicate the best performance. Bold results indicate the best performance.
[…]
DeSGDA w/o CA 58.3 63.3 62.1 60.1 81.4 62.7 63.8 59.1 80.3 57.1 78.9 63.0 65.8
DeSGDA w/o PL 58.8 57.6 60.2 61.4 73.3 54.5 63.7 61.0 61.2 61.9 71.0 60.2 62.1
DeSGDA w/o CF 54.7 58.3 59.6 58.7 72.1 60.7 57.9 54.9 78.0 57.4 63.2 60.4 61.3
DeSGDA w/o TL 56.5 60.7 59.1 60.7 43.4 56.3 51.0 56.7 53.3 56.9 63.1 52.5 55.9
DeSGDA w PT 62.1 64.3 63.9 61.8 77.3 62.4 66.6 63.9 73.5 66.1 82.1 64.9 67.4
DeSGDA w CL 66.1 62.6 63.3 64.7 81.5 60.7 69.8 67.0 82.2 67.8 82.4 63.6 69.6
DeSGDA 65.4 65.9 65.5 65.6 82.8 63.6 70.7 68.2 82.9 67.6 83.9 66.5 70.1
Domain adaption methods. We compare DeSGDA with two recent domain adaption methods:
  • CDAN: CDAN Long et al. [2018] employs a conditional adversarial learning strategy to reduce domain discrepancy by conditioning adversarial adaptation on discriminative information from multiple domains.
  • ToAlign: ToAlign Wei et al. [2021b] uses token-level alignment strategies within Transformer architectures to enhance cross-lingual transfer, optimizing the alignment of semantic representations across languages.
  • MetaAlign: MetaAlign Wei et al. [2021a] is a meta-learning framework for domain adaptation that dynamically aligns feature distributions across domains by learning domain-invariant representations.
Graph domain adaptation methods. We compare DeSGDA with six SOTA graph domain adaption methods:
  • DEAL: DEAL Yin et al. [2022] uses domain adversarial learning to align graph representations across different domains without labeled data, overcoming discrepancies between the source and target domains.
  • CoCo: CoCo Yin et al. [2023] leverages contrastive learning to align graph representations between source and target domains, enhancing domain adaptation by promoting intra-domain cohesion and inter-domain separation in an unsupervised manner.
  • SGDA: SGDA Qiao et al. [2023] utilizes labeled data from the source domain along with a limited amount of labeled data from the target domain to learn domain-invariant graph representations.
  • DGDA: DGDA Cai et al. [2024] employs generative models to capture the underlying distribution of graph data across domains, facilitating the transfer of graph structures and features by learning shared latent spaces.
  • A2GNN: A2GNN Liu et al. [2024b] introduces a novel propagation mechanism to enhance feature transferability across domains, improving the alignment of graph structures and node features in an unsupervised setting.
  • PA-BOTH: PA-BOTH Liu et al. [2024c] aligns node pairs between source and target graphs, optimizing feature correspondence at a granular level to improve the transferability of structural and feature information across domains.
For GCN and GIN, we use the Pytorch Geometric  [^4] https://www.pyg.org/ to implement the model. For other baseline methods, we use the source codes provided by the corresponding paper. For all baseline methods, we vary the dropout rate in the range of {0.1,0.3,0.5,0.7} and then choose the best one. The hidden dimension in these methods is set to 256 for a fair comparison.
Appendix G More experimental results
G.1 More performance comparison
In this part, we provide additional results for our proposed method DeSGDA compared with all baseline models across various datasets, as illustrated in Table 9 to Table 17. These results consistently show that DeSGDA outperforms the baseline models in most cases, validating the superiority of our proposed method.
Additionally, the performance of DeSGDA with different GNN architectures on the NCI1, FRANKENSTEIN, and MUTAGENICITY datasets is shown in Figure 12. It is evident that GIN consistently outperforms other GNN architectures in most cases.
G.2 More Ablation study
To validate the effectiveness of the different components in DeSGDA, we conduct more experiments with sive variants on NCI1, FRANKENSTEIN and MUTAGENICITY datasets, i.e., DeSGDA w/o CA, DeSGDA w/o PL, DeSGDA w/o CF, DeSGDA w/o TL, DeSGDA w PT and DeSGDA w CL. The results are shown in Table 6 , 7 and 8. From the results, we have similar observations as summarized in Section 5.4.
G.3 More Sensitivity Analysis
In this part, we provide additional sensitivity analysis of the proposed DeSGDA with respect to the impact of its hyperparameters: the time latency $\tau$, degree threshold value $V_{th}^{degree}$ in SNNs, and balance ratio $\alpha$ on NCI1, FRANKENSTEIN, and MUTAGENICITY datasets. The results are illustrated in Figure 9, 10 and 11, where we observe trends similar to those discussed in Section 5.5.
(a) Time Latency
(b) Initial Threshold
(c) Balance Tatio
Figure 9: Hyperparameter sensitivity analysis of time latency $\tau$, initial threshold $V_{th}^{degree}$ in SNNs, and balance ratio $\alpha$ on NCI1.
(a) Time Latency
(b) Initial Threshold
(c) Balance Ratio
Figure 10: Hyperparameter sensitivity analysis of time latency $\tau$, initial threshold $V_{th}^{degree}$ in SNNs, and balance ratio $\alpha$ on FRANKENSTEIN.
(a) Time Latency
(b) Initial Threshold
(c) Balance Ratio
Figure 11: Hyperparameter sensitivity analysis of time latency $\tau$, initial threshold $V_{th}^{degree}$ in SNNs, and balance ratio $\alpha$ on MUTAGENICITY.
(a) NCI1
(b) FRANKENSTEIN
(c) MUTAGENICITY
Figure 12: The performance with different GNN architectures on different datasets.
[…]
WL subtree 73.4 72.7 70.5 73.0 72.8 59.0 66.5 71.6 60.6 58.3 76.3 64.0 68.2
GCN 57.2$\pm$2.7 62.8$\pm$1.7 67.6$\pm$0.5 58.5$\pm$1.3 67.7$\pm$0.4 61.0$\pm$0.3 65.0$\pm$0.8 51.1$\pm$1.3 65.6$\pm$2.2 55.4$\pm$0.4 68.5$\pm$3.1 67.7$\pm$0.5 62.3
GIN 69.3$\pm$2.3 65.8$\pm$0.8 69.3$\pm$1.7 69.8$\pm$1.6 71.4$\pm$2.1 52.4$\pm$1.8 64.0$\pm$2.4 65.7$\pm$3.2 53.4$\pm$3.7 58.1$\pm$0.8 72.6$\pm$0.3 64.6$\pm$2.3 64.7
GMT 67.8$\pm$1.3 69.6$\pm$0.7 74.5$\pm$0.5 67.6$\pm$2.5 69.9$\pm$2.1 55.8$\pm$0.7 74.8$\pm$1.4 60.1$\pm$2.4 71.4$\pm$3.3 51.5$\pm$0.5 69.0$\pm$0.5 63.3$\pm$1.3 66.3
CIN 62.6$\pm$0.5 59.4$\pm$0.5 64.0$\pm$0.9 58.5$\pm$1.8 71.9$\pm$1.7 60.6$\pm$2.1 63.7$\pm$0.5 61.2$\pm$2.1 73.2$\pm$0.5 57.7$\pm$3.0 68.1$\pm$0.4 58.5$\pm$2.7 63.3
SpikeGCN 66.1$\pm$0.4 67.8$\pm$0.7 68.5$\pm$0.9 67.2$\pm$0.3 70.1$\pm$1.1 68.6$\pm$1.3 66.1$\pm$0.7 65.8$\pm$1.2 74.5$\pm$0.9 65.4$\pm$0.6 73.8$\pm$1.1 70.1$\pm$1.6 68.6
DRSGNN 68.3$\pm$0.5 71.9$\pm$0.9 71.5$\pm$1.1 75.1$\pm$1.7 76.7$\pm$1.3 74.4$\pm$0.8 70.5$\pm$0.7 67.8$\pm$0.8 77.0$\pm$1.5 65.9$\pm$1.6 75.2$\pm$1.9 69.2$\pm$2.1 71.8
CDAN 75.6$\pm$0.5 70.5$\pm$0.6 71.6$\pm$0.5 69.8$\pm$0.5 76.6$\pm$0.8 71.4$\pm$0.3 71.4$\pm$0.3 72.1$\pm$0.3 75.5$\pm$0.7 74.3$\pm$0.8 78.2$\pm$1.1 74.0$\pm$0.8 73.4
ToAlign 51.1$\pm$0.6 55.8$\pm$0.1 63.3$\pm$0.2 55.8$\pm$0.4 68.1$\pm$0.7 55.8$\pm$0.3 63.3$\pm$0.5 51.1$\pm$0.2 68.1$\pm$1.0 51.1$\pm$0.4 68.1$\pm$0.6 63.3$\pm$0.2 59.6
MetaAlign 59.4$\pm$1.1 62.2$\pm$1.0 68.9$\pm$0.3 65.3$\pm$0.8 75.1$\pm$0.7 67.5$\pm$2.1 70.9$\pm$1.4 60.6$\pm$2.3 72.4$\pm$1.4 59.4$\pm$0.6 74.6$\pm$0.7 67.8$\pm$1.3 67.0
DEAL 76.6$\pm$0.4 62.8$\pm$0.8 72.8$\pm$1.3 67.3$\pm$2.2 77.2$\pm$2.3 67.6$\pm$1.9 71.2$\pm$1.6 56.0$\pm$2.5 73.9$\pm$2.1 66.0$\pm$0.3 76.4$\pm$1.1 65.5$\pm$2.1 69.4
CoCo 73.4$\pm$0.5 73.6$\pm$0.8 73.4$\pm$1.0 71.6$\pm$0.5 75.2$\pm$1.6 74.6$\pm$0.3 70.7$\pm$0.8 68.4$\pm$1.5 75.0$\pm$0.2 72.7$\pm$0.4 76.3$\pm$1.1 75.0$\pm$1.8 73.3
SGDA 63.8$\pm$0.8 65.2$\pm$0.5 66.7$\pm$0.3 59.1$\pm$1.2 62.3$\pm$0.7 60.6$\pm$0.4 65.2$\pm$0.9 61.8$\pm$1.0 64.5$\pm$1.3 60.9$\pm$0.8 59.4$\pm$1.2 64.9$\pm$1.1 62.9
DGDA 59.4$\pm$0.7 62.3$\pm$1.1 63.1$\pm$0.5 61.2$\pm$0.9 60.4$\pm$0.6 58.8$\pm$1.0 60.3$\pm$0.8 63.5$\pm$1.2 61.9$\pm$0.8 60.4$\pm$1.6 64.2$\pm$1.3 62.6$\pm$1.4 61.5
A2GNN 65.4$\pm$0.7 66.4$\pm$1.1 65.7$\pm$1.3 66.0$\pm$0.6 64.9$\pm$1.2 65.8$\pm$1.6 65.5$\pm$1.8 66.0$\pm$1.4 65.8$\pm$2.1 65.6$\pm$1.9 66.1$\pm$1.7 66.0$\pm$2.0 65.8
PA-BOTH 66.9$\pm$0.5 67.1$\pm$0.8 67.3$\pm$1.1 65.8$\pm$0.7 69.1$\pm$1.0 66.1$\pm$1.4 66.7$\pm$1.3 67.4$\pm$1.4 66.3$\pm$1.8 66.0$\pm$1.2 66.8$\pm$0.8 66.3$\pm$1.5 66.8
DeSGDA 78.7$\pm$1.3 78.4$\pm$1.1 74.8$\pm$0.6 77.6$\pm$0.9 79.5$\pm$1.2 76.7$\pm$0.8 74.9$\pm$0.7 71.2$\pm$1.7 79.5$\pm$1.4 72.8$\pm$0.8 81.0$\pm$1.5 75.1$\pm$1.0 76.7
[…]
WL subtree 58.4 51.8 58.7 51.3 64.3 48.9 64.9 58.9 78.5 54.6 57.1 61.3 59.1
GCN 56.2$\pm$0.2 59.0$\pm$1.3 41.4$\pm$0.4 45.8$\pm$0.5 21.2$\pm$0.7 41.4$\pm$1.7 42.5$\pm$1.6 49.0$\pm$0.4 24.1$\pm$1.6 44.8$\pm$0.7 81.4$\pm$0.3 58.8$\pm$0.2 47.1
GIN 60.7$\pm$0.6 58.0$\pm$1.0 61.0$\pm$2.3 58.9$\pm$2.3 77.5$\pm$2.2 45.3$\pm$2.5 62.5$\pm$0.2 59.2$\pm$3.0 71.4$\pm$2.8 49.8$\pm$1.7 77.9$\pm$1.4 59.9$\pm$0.5 61.8
GMT 56.2$\pm$0.4 59.8$\pm$0.2 41.4$\pm$0.3 59.8$\pm$0.7 21.2$\pm$1.1 59.8$\pm$0.5 41.4$\pm$0.2 56.2$\pm$0.2 21.1$\pm$1.1 56.2$\pm$1.4 78.8$\pm$0.6 58.6$\pm$0.8 50.9
CIN 57.8$\pm$1.1 60.1$\pm$0.7 58.6$\pm$0.2 59.8$\pm$0.2 78.9$\pm$0.1 59.9$\pm$0.4 58.8$\pm$0.3 57.4$\pm$0.5 78.8$\pm$0.6 57.7$\pm$1.2 78.8$\pm$0.7 60.1$\pm$1.1 63.9
SpikeGCN 56.1$\pm$0.7 59.7$\pm$1.0 58.8$\pm$0.6 57.8$\pm$0.2 77.1$\pm$1.3 53.2$\pm$1.6 41.4$\pm$1.9 56.1$\pm$1.5 70.1$\pm$0.9 59.9$\pm$1.5 76.8$\pm$1.8 58.5$\pm$1.4 60.4
DRSGNN 60.2$\pm$0.9 59.9$\pm$0.8 57.3$\pm$1.2 59.0$\pm$1.0 74.2$\pm$1.9 54.6$\pm$1.7 58.5$\pm$1.5 58.9$\pm$1.8 77.7$\pm$2.3 56.9$\pm$2.0 78.9$\pm$2.4 58.8$\pm$1.6 62.9
CDAN 60.9$\pm$0.7 59.8$\pm$0.5 61.1$\pm$1.3 61.0$\pm$0.2 80.5$\pm$1.2 59.8$\pm$0.3 64.0$\pm$0.4 61.4$\pm$0.1 81.8$\pm$0.1 58.0$\pm$1.2 81.8$\pm$0.3 63.8$\pm$0.7 66.1
ToAlign 56.2$\pm$0.2 59.8$\pm$0.2 41.4$\pm$0.1 59.8$\pm$0.2 21.1$\pm$0.3 59.8$\pm$0.7 41.4$\pm$1.1 56.2$\pm$1.2 21.1$\pm$0.4 56.2$\pm$0.6 21.1$\pm$1.3 41.4$\pm$0.5 44.6
MetaAlign 57.3$\pm$2.4 59.1$\pm$1.1 60.9$\pm$1.5 60.2$\pm$0.4 80.3$\pm$2.1 60.4$\pm$0.6 64.0$\pm$1.1 64.9$\pm$0.6 81.4$\pm$1.2 58.5$\pm$2.3 80.8$\pm$0.5 63.4$\pm$1.8 65.9
DEAL 65.3$\pm$0.6 64.0$\pm$0.2 61.3$\pm$0.6 61.0$\pm$0.9 78.3$\pm$2.1 55.5$\pm$1.8 64.9$\pm$1.2 64.8$\pm$1.1 80.1$\pm$1.3 60.1$\pm$2.1 81.8$\pm$0.4 65.7$\pm$0.7 66.9
CoCo 63.5$\pm$2.4 61.5$\pm$1.0 64.4$\pm$1.0 61.2$\pm$0.7 81.7$\pm$0.4 55.0$\pm$1.6 64.5$\pm$0.6 64.6$\pm$1.1 80.4$\pm$1.5 60.6$\pm$1.5 81.5$\pm$0.6 62.2$\pm$1.7 66.8
SGDA 55.7$\pm$0.5 55.4$\pm$0.9 54.8$\pm$0.3 55.3$\pm$0.7 56.1$\pm$0.5 55.4$\pm$0.8 53.2$\pm$1.1 55.1$\pm$0.6 58.4$\pm$0.4 55.3$\pm$0.5 57.7$\pm$1.0 54.9$\pm$0.6 55.7
DGDA OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
A2GNN 56.0$\pm$0.3 56.3$\pm$0.6 55.6$\pm$0.4 57.3$\pm$0.7 58.6$\pm$0.6 55.9$\pm$0.9 55.5$\pm$0.5 55.3$\pm$0.2 61.2$\pm$1.3 56.6$\pm$0.9 65.5$\pm$0.8 56.0$\pm$1.0 57.5
PA-BOTH 62.2$\pm$0.5 60.7$\pm$0.7 61.5$\pm$0.6 61.2$\pm$1.0 61.9$\pm$1.3 61.1$\pm$0.8 62.3$\pm$0.4 61.7$\pm$0.8 62.0$\pm$0.9 61.1$\pm$1.2 61.2$\pm$1.5 60.9$\pm$0.6 61.5
DeSGDA 63.5$\pm$1.1 64.1$\pm$0.9 62.6$\pm$1.3 63.0$\pm$0.8 81.1$\pm$1.2 62.4$\pm$1.5 65.5$\pm$0.6 65.4$\pm$1.7 81.9$\pm$1.0 60.8$\pm$1.4 82.0$\pm$2.1 65.9$\pm$1.8 68.1
[…]
WL subtree 74.4 72.9 64.9 68.9 49.1 59.8 70.0 70.5 76.9 60.7 82.6 70.5 68.5
GCN 63.1$\pm$1.0 68.1$\pm$0.3 48.8$\pm$0.4 62.6$\pm$0.3 29.1$\pm$2.1 38.8$\pm$0.3 54.3$\pm$0.1 61.8$\pm$0.5 30.4$\pm$0.2 43.6$\pm$0.3 67.8$\pm$0.1 57.9$\pm$1.3 52.2
GIN 68.1$\pm$1.6 74.2$\pm$0.6 59.6$\pm$2.3 65.2$\pm$1.4 40.3$\pm$2.7 54.6$\pm$1.8 61.3$\pm$1.1 63.1$\pm$3.2 71.6$\pm$3.0 60.0$\pm$1.4 79.7$\pm$1.3 69.2$\pm$0.7 63.9
GMT 56.5$\pm$0.3 60.7$\pm$0.4 57.9$\pm$0.2 40.2$\pm$1.2 80.6$\pm$0.4 39.3$\pm$0.6 57.9$\pm$1.1 45.0$\pm$2.1 80.6$\pm$0.5 43.5$\pm$1.1 80.6$\pm$1.4 57.9$\pm$2.2 58.4
CIN 64.1$\pm$3.0 61.3$\pm$0.5 63.5$\pm$2.3 63.6$\pm$1.5 78.2$\pm$0.5 63.9$\pm$2.7 60.6$\pm$1.5 57.0$\pm$0.4 73.7$\pm$3.2 61.4$\pm$1.0 79.1$\pm$2.1 61.1$\pm$1.9 65.6
SpikeGCN 56.4$\pm$1.2 60.7$\pm$0.9 59.5$\pm$1.3 57.7$\pm$1.6 59.0$\pm$1.1 60.1$\pm$1.8 54.2$\pm$1.0 59.9$\pm$2.1 50.2$\pm$2.6 55.1$\pm$1.7 80.1$\pm$2.5 57.9$\pm$2.2 59.3
DRSGNN 56.7$\pm$0.7 61.0$\pm$1.1 57.2$\pm$1.0 57.7$\pm$1.6 52.1$\pm$1.4 55.2$\pm$1.5 59.4$\pm$1.8 56.3$\pm$1.7 75.9$\pm$2.3 60.7$\pm$1.9 80.6$\pm$0.8 58.0$\pm$1.3 60.9
CDAN 62.8$\pm$0.3 68.2$\pm$0.6 63.6$\pm$0.6 66.9$\pm$1.7 81.2$\pm$0.5 65.0$\pm$2.1 65.8$\pm$0.2 64.7$\pm$1.2 80.7$\pm$0.1 62.5$\pm$2.3 82.4$\pm$0.4 66.0$\pm$0.5 69.1
ToAlign 43.5$\pm$0.4 39.3$\pm$0.7 57.9$\pm$1.0 39.3$\pm$1.4 80.6$\pm$1.1 39.3$\pm$0.7 57.9$\pm$0.3 43.5$\pm$2.1 80.6$\pm$1.8 43.5$\pm$0.4 80.6$\pm$0.9 57.9$\pm$1.0 55.3
MetaAlign 63.1$\pm$2.5 68.8$\pm$2.6 63.3$\pm$0.6 65.2$\pm$2.2 81.9$\pm$0.1 64.5$\pm$1.4 65.0$\pm$0.6 68.3$\pm$0.6 81.0$\pm$0.3 65.2$\pm$0.2 82.5$\pm$0.4 68.3$\pm$0.6 69.7
DEAL 64.6$\pm$0.5 65.5$\pm$0.8 64.2$\pm$1.0 63.1$\pm$2.1 82.7$\pm$0.8 62.8$\pm$0.7 70.2$\pm$0.4 67.3$\pm$0.4 79.6$\pm$0.1 63.9$\pm$1.4 75.7$\pm$0.3 67.0$\pm$0.2 68.9
CoCo 65.7$\pm$1.8 74.1$\pm$0.7 65.1$\pm$0.2 67.6$\pm$0.9 80.5$\pm$1.3 56.5$\pm$1.7 68.4$\pm$1.3 70.7$\pm$0.4 78.9$\pm$1.2 67.3$\pm$0.3 83.7$\pm$0.1 71.5$\pm$0.9 70.8
SGDA OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
DGDA OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
A2GNN 55.4$\pm$0.3 55.7$\pm$0.7 55.6$\pm$0.5 54.7$\pm$0.8 63.3$\pm$1.0 56.6$\pm$0.9 55.3$\pm$0.6 55.7$\pm$0.5 65.5$\pm$0.8 56.6$\pm$1.2 69.9$\pm$1.4 55.0$\pm$0.5 58.3
PA-BOTH 61.2$\pm$0.9 62.0$\pm$0.4 60.7$\pm$0.8 61.7$\pm$0.5 60.9$\pm$1.2 61.1$\pm$0.7 61.5$\pm$0.9 60.2$\pm$1.2 61.3$\pm$1.5 61.8$\pm$0.8 62.2$\pm$0.9 62.0$\pm$1.0 61.4
DeSGDA 65.4$\pm$1.3 65.9$\pm$0.9 65.5$\pm$1.4 64.8$\pm$1.1 82.8$\pm$1.6 63.6$\pm$1.5 70.7$\pm$1.8 68.2$\pm$0.7 82.9$\pm$1.2 67.6$\pm$0.8 83.9$\pm$1.0 66.5$\pm$1.2 70.9
[…]
WL subtree 69.1 59.7 61.2 75.9 41.6 83.5 61.5 72.7 24.7 72.7 63.1 62.9 62.4
GCN 73.7$\pm$0.3 82.7$\pm$0.4 57.6$\pm$0.2 84.0$\pm$1.3 24.4$\pm$0.4 17.3$\pm$0.2 57.6$\pm$0.1 70.9$\pm$0.7 24.4$\pm$0.5 26.3$\pm$0.1 37.5$\pm$0.2 42.5$\pm$0.8 49.9
GIN 71.8$\pm$2.7 70.2$\pm$4.7 58.5$\pm$4.3 56.9$\pm$4.9 74.2$\pm$1.7 78.2$\pm$3.3 63.3$\pm$2.7 67.1$\pm$3.8 35.9$\pm$4.2 61.0$\pm$2.4 71.9$\pm$2.1 65.1$\pm$1.0 64.5
GMT 73.7$\pm$0.2 82.7$\pm$0.1 57.6$\pm$0.3 83.1$\pm$0.5 75.6$\pm$1.4 17.3$\pm$0.6 57.6$\pm$1.5 73.7$\pm$0.6 75.6$\pm$0.4 26.3$\pm$1.2 75.6$\pm$0.7 42.4$\pm$0.5 61.8
CIN 74.1$\pm$0.6 83.8$\pm$1.0 60.1$\pm$2.1 78.6$\pm$3.1 75.6$\pm$0.2 74.8$\pm$3.7 63.9$\pm$2.7 74.1$\pm$0.6 57.0$\pm$4.3 58.9$\pm$3.3 75.6$\pm$0.7 63.6$\pm$1.0 70.0
SpikeGCN 71.8$\pm$0.9 80.9$\pm$1.2 64.9$\pm$1.4 79.1$\pm$2.2 71.1$\pm$1.9 73.8$\pm$1.6 62.4$\pm$2.0 71.8$\pm$2.3 70.1$\pm$2.4 66.9$\pm$1.9 72.1$\pm$1.9 64.5$\pm$1.7 70.9
DRSGNN 73.6$\pm$1.1 81.3$\pm$1.5 64.6$\pm$1.2 80.6$\pm$1.4 70.2$\pm$1.7 76.1$\pm$2.3 64.1$\pm$1.5 71.9$\pm$1.9 70.4$\pm$2.0 64.1$\pm$3.1 74.7$\pm$1.4 64.3$\pm$1.1 71.3
CDAN 75.9$\pm$1.0 83.1$\pm$0.6 60.8$\pm$0.6 82.6$\pm$0.2 75.8$\pm$0.3 70.9$\pm$2.4 64.7$\pm$0.3 77.7$\pm$0.6 73.3$\pm$1.8 75.4$\pm$0.7 75.8$\pm$0.4 67.1$\pm$0.8 73.6
ToAlign 73.7$\pm$0.4 82.7$\pm$0.3 57.6$\pm$0.6 82.7$\pm$0.8 24.4$\pm$0.1 82.7$\pm$0.3 57.6$\pm$0.4 73.7$\pm$0.2 24.4$\pm$0.7 73.7$\pm$0.3 24.4$\pm$0.5 57.6$\pm$0.4 59.6
MetaAlign 74.3$\pm$0.8 83.3$\pm$2.2 60.6$\pm$1.7 71.2$\pm$2.1 76.3$\pm$0.3 77.3$\pm$2.4 64.6$\pm$1.2 72.0$\pm$1.0 76.0$\pm$0.5 73.3$\pm$1.8 74.4$\pm$1.7 56.9$\pm$1.4 71.7
DEAL 75.4$\pm$1.2 78.0$\pm$2.4 68.1$\pm$1.9 80.8$\pm$2.1 73.8$\pm$1.4 80.6$\pm$2.3 65.7$\pm$1.7 74.7$\pm$2.4 74.7$\pm$1.6 71.0$\pm$2.1 68.1$\pm$2.6 70.3$\pm$0.4 73.4
CoCo 74.8$\pm$0.6 84.1$\pm$1.1 65.5$\pm$0.4 83.6$\pm$1.1 72.4$\pm$2.9 83.1$\pm$0.4 69.7$\pm$0.5 75.8$\pm$0.7 71.4$\pm$2.3 73.4$\pm$1.3 72.5$\pm$2.7 66.4$\pm$1.7 74.4
SGDA 64.2$\pm$0.5 61.0$\pm$0.7 66.9$\pm$1.2 61.9$\pm$0.9 65.4$\pm$1.6 66.5$\pm$1.0 64.6$\pm$1.1 60.1$\pm$0.5 66.3$\pm$1.3 59.3$\pm$0.8 66.0$\pm$1.6 66.2$\pm$1.3 64.1
DGDA 58.1$\pm$0.4 58.6$\pm$0.6 58.9$\pm$1.0 61.0$\pm$0.9 59.6$\pm$0.7 60.2$\pm$1.5 56.7$\pm$0.6 56.8$\pm$0.8 58.1$\pm$0.4 58.8$\pm$1.1 57.0$\pm$1.2 62.2$\pm$1.6 58.9
A2GNN 65.7$\pm$0.6 65.9$\pm$0.8 66.3$\pm$0.9 65.6$\pm$1.1 65.2$\pm$1.4 65.6$\pm$1.3 65.9$\pm$1.7 65.8$\pm$1.6 65.0$\pm$1.5 66.1$\pm$1.2 65.2$\pm$1.9 65.9$\pm$1.8 65.7
PA-BOTH 61.0$\pm$0.8 61.2$\pm$1.3 60.3$\pm$0.6 66.7$\pm$2.1 63.7$\pm$1.5 61.9$\pm$2.0 66.2$\pm$1.4 69.9$\pm$2.3 68.0$\pm$0.7 69.4$\pm$1.8 61.5$\pm$0.4 67.6$\pm$1.0 64.9
DeSGDA 77.6$\pm$0.9 84.3$\pm$1.1 70.5$\pm$0.6 84.8$\pm$1.4 76.6$\pm$0.7 83.9$\pm$0.9 71.9$\pm$0.6 76.9$\pm$1.1 76.1$\pm$0.8 73.7$\pm$1.0 76.0$\pm$1.2 70.4$\pm$0.7 76.8
[…]
WL subtree 73.5 79.5 64.8 75.9 58.9 68.4 72.5 72.0 69.7 63.6 76.1 74.0 70.7
GCN 51.2$\pm$0.1 71.1$\pm$0.4 42.7$\pm$0.4 27.8$\pm$0.3 32.1$\pm$1.1 27.0$\pm$0.2 55.2$\pm$0.6 50.5$\pm$0.7 50.9$\pm$1.1 49.1$\pm$0.3 67.1$\pm$0.6 57.3$\pm$0.6 48.5
GIN 66.9$\pm$2.2 78.9$\pm$2.3 60.3$\pm$3.1 72.8$\pm$0.3 51.1$\pm$0.6 68.6$\pm$1.8 63.5$\pm$2.1 67.8$\pm$3.7 65.9$\pm$1.7 60.3$\pm$1.8 71.1$\pm$1.1 67.2$\pm$1.3 66.2
GMT 50.9$\pm$0.5 73.0$\pm$0.1 57.3$\pm$0.3 73.0$\pm$0.4 66.5$\pm$0.2 73.0$\pm$0.3 72.4$\pm$0.6 50.9$\pm$0.1 66.5$\pm$0.4 58.3$\pm$0.2 66.5$\pm$0.5 72.8$\pm$0.3 65.1
CIN 60.1$\pm$0.7 73.1$\pm$1.1 57.5$\pm$0.2 73.0$\pm$0.4 66.5$\pm$1.1 73.1$\pm$0.7 58.5$\pm$2.1 52.9$\pm$1.4 66.5$\pm$1.3 56.1$\pm$0.1 66.5$\pm$0.4 57.4$\pm$0.7 63.4
SpikeGCN 63.3$\pm$0.4 72.6$\pm$0.7 60.6$\pm$0.2 73.1$\pm$0.8 65.4$\pm$0.6 66.6$\pm$1.2 64.3$\pm$1.4 64.7$\pm$1.0 65.6$\pm$1.2 59.8$\pm$0.8 70.1$\pm$1.7 60.9$\pm$1.4 65.6
DRSGNN 63.9$\pm$0.8 73.3$\pm$1.2 56.9$\pm$1.5 72.5$\pm$1.3 66.7$\pm$2.1 64.4$\pm$1.8 65.6$\pm$1.3 64.8$\pm$1.9 66.7$\pm$1.4 58.4$\pm$1.6 68.6$\pm$2.0 61.4$\pm$1.3 65.3
CDAN 57.1$\pm$0.4 75.0$\pm$0.7 61.2$\pm$0.4 73.7$\pm$0.1 68.2$\pm$0.4 73.3$\pm$0.3 60.2$\pm$0.1 56.5$\pm$1.4 68.2$\pm$0.2 53.9$\pm$1.4 68.4$\pm$0.2 59.6$\pm$0.5 64.6
ToAlign 49.1$\pm$0.3 27.0$\pm$0.6 57.3$\pm$0.5 27.0$\pm$0.4 66.5$\pm$0.5 27.0$\pm$0.2 57.3$\pm$0.3 49.1$\pm$0.4 66.5$\pm$0.2 49.1$\pm$0.3 66.5$\pm$0.1 57.3$\pm$0.4 50.0
MetaAlign 65.6$\pm$1.8 77.7$\pm$0.2 63.5$\pm$1.4 75.7$\pm$0.7 66.4$\pm$0.3 74.0$\pm$0.3 66.3$\pm$1.1 64.6$\pm$1.2 66.7$\pm$0.2 59.5$\pm$2.6 66.7$\pm$0.3 66.7$\pm$2.7 67.8
DEAL 64.0$\pm$0.9 71.9$\pm$1.2 61.4$\pm$0.3 73.3$\pm$0.3 64.9$\pm$1.4 71.9$\pm$1.9 62.5$\pm$2.1 66.2$\pm$0.5 54.2$\pm$1.4 55.6$\pm$0.8 64.6$\pm$0.4 58.8$\pm$0.4 64.1
CoCo 69.7$\pm$0.1 80.4$\pm$0.4 64.7$\pm$1.2 76.5$\pm$0.4 65.0$\pm$1.7 73.9$\pm$0.3 68.9$\pm$1.3 70.7$\pm$0.9 68.2$\pm$1.2 61.4$\pm$1.7 73.0$\pm$0.1 65.2$\pm$0.9 69.8
SGDA OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
DGDA OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
A2GNN 59.0$\pm$0.6 58.3$\pm$1.1 58.5$\pm$0.8 58.6$\pm$1.3 58.7$\pm$1.0 59.0$\pm$0.7 58.5$\pm$1.1 58.7$\pm$1.5 59.1$\pm$0.6 58.3$\pm$1.2 58.6$\pm$0.7 59.0$\pm$0.5 58.7
PA-BOTH 57.7$\pm$0.4 58.0$\pm$0.6 57.9$\pm$0.5 56.9$\pm$0.8 57.4$\pm$0.6 58.3$\pm$0.5 57.1$\pm$1.2 58.8$\pm$0.9 58.1$\pm$0.7 58.0$\pm$0.9 57.9$\pm$0.5 58.3$\pm$0.8 57.9
DeSGDA 64.4$\pm$1.2 76.9$\pm$1.5 64.9$\pm$0.9 76.6$\pm$1.2 68.6$\pm$1.8 74.1$\pm$1.3 66.9$\pm$0.8 65.1$\pm$1.2 69.9$\pm$1.5 63.9$\pm$2.0 70.9$\pm$1.6 64.2$\pm$1.4 69.0
[…]
WL subtree 78.0 68.7 70.1 70.5 59.0 61.2 71.7 78.0 49.9 56.3 69.4 71.9 67.1
GCN 74.5$\pm$0.2 60.8$\pm$2.1 69.7$\pm$0.4 68.5$\pm$1.7 54.1$\pm$0.9 55.2$\pm$0.9 68.6$\pm$1.6 75.5$\pm$0.5 51.5$\pm$1.3 46.4$\pm$1.7 58.6$\pm$0.4 60.2$\pm$0.2 61.9
GIN 77.9$\pm$3.1 70.7$\pm$2.4 70.9$\pm$0.8 69.2$\pm$1.2 64.1$\pm$1.0 61.9$\pm$2.4 78.5$\pm$0.2 79.8$\pm$3.3 65.5$\pm$2.7 71.5$\pm$0.9 69.5$\pm$1.8 73.5$\pm$2.6 71.1
GMT 67.3$\pm$0.2 52.5$\pm$0.1 59.9$\pm$0.3 47.5$\pm$0.2 53.5$\pm$0.2 52.5$\pm$0.4 59.9$\pm$0.1 67.3$\pm$0.2 46.7$\pm$0.5 67.3$\pm$0.3 53.3$\pm$0.1 59.9$\pm$0.4 57.1
CIN 70.8$\pm$1.1 66.9$\pm$3.4 61.7$\pm$0.6 62.6$\pm$2.4 56.3$\pm$3.1 62.9$\pm$1.3 65.1$\pm$1.0 68.8$\pm$1.7 56.6$\pm$1.4 66.9$\pm$1.0 58.1$\pm$1.3 62.5$\pm$0.9 63.3
SpikeGCN 63.2$\pm$0.6 55.9$\pm$1.7 59.8$\pm$1.2 56.5$\pm$1.6 54.1$\pm$2.2 60.3$\pm$1.2 60.1$\pm$0.8 68.8$\pm$1.4 54.4$\pm$2.0 64.2$\pm$1.6 56.6$\pm$0.8 58.5$\pm$1.5 59.4
DRSGNN 56.9$\pm$0.9 53.4$\pm$1.2 58.9$\pm$1.0 57.7$\pm$1.4 53.2$\pm$1.6 61.2$\pm$1.5 60.6$\pm$2.1 67.9$\pm$1.2 57.1$\pm$1.4 67.3$\pm$2.0 59.9$\pm$1.9 57.1$\pm$0.8 59.3
CDAN 75.5$\pm$0.1 71.3$\pm$0.4 70.7$\pm$0.3 70.3$\pm$0.1 58.7$\pm$0.6 58.4$\pm$0.6 70.2$\pm$0.5 76.1$\pm$0.5 58.5$\pm$0.6 69.4$\pm$1.5 59.0$\pm$0.1 63.7$\pm$1.4 66.8
ToAlign 67.3$\pm$0.2 47.5$\pm$0.4 59.9$\pm$0.6 47.5$\pm$0.5 46.7$\pm$0.4 47.5$\pm$0.2 59.9$\pm$0.7 67.3$\pm$0.3 46.7$\pm$0.1 67.3$\pm$0.4 46.7$\pm$0.5 59.9$\pm$0.3 55.4
MetaAlign 76.5$\pm$0.4 71.8$\pm$1.1 71.8$\pm$0.8 71.4$\pm$0.9 59.3$\pm$0.8 63.0$\pm$1.0 74.2$\pm$1.6 78.0$\pm$0.2 61.7$\pm$1.2 69.9$\pm$1.6 62.2$\pm$0.4 68.3$\pm$1.5 69.0
DEAL 76.6$\pm$0.8 68.8$\pm$1.0 69.9$\pm$0.4 66.4$\pm$0.8 59.3$\pm$2.1 64.2$\pm$2.2 79.1$\pm$0.1 81.9$\pm$0.6 64.5$\pm$1.1 75.3$\pm$0.6 69.8$\pm$1.6 76.5$\pm$0.2 71.0
CoCo 75.5$\pm$0.4 71.7$\pm$0.7 68.7$\pm$1.1 69.2$\pm$2.0 60.8$\pm$1.1 65.7$\pm$0.3 79.2$\pm$1.2 76.8$\pm$0.6 63.8$\pm$0.5 73.8$\pm$0.4 64.6$\pm$0.8 70.1$\pm$1.1 70.0
SGDA OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
DGDA OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
A2GNN 55.4$\pm$0.6 56.3$\pm$0.2 55.6$\pm$0.8 55.1$\pm$0.5 55.3$\pm$1.1 55.9$\pm$0.4 56.1$\pm$0.7 55.7$\pm$0.6 57.1$\pm$0.3 56.6$\pm$1.2 55.2$\pm$0.7 56.8$\pm$1.0 55.9
PA-BOTH 55.9$\pm$1.0 56.0$\pm$0.5 56.1$\pm$0.7 56.6$\pm$1.2 55.9$\pm$0.6 56.0$\pm$0.7 57.3$\pm$0.8 56.8$\pm$1.3 55.9$\pm$1.2 56.3$\pm$1.0 56.4$\pm$0.9 57.1$\pm$1.3 56.4
DeSGDA 76.7$\pm$1.0 72.5$\pm$0.6 72.0$\pm$2.3 71.5$\pm$1.5 61.0$\pm$1.2 67.6$\pm$0.8 75.5$\pm$1.0 79.8$\pm$1.2 62.8$\pm$1.6 75.9$\pm$1.3 65.5$\pm$2.1 72.3$\pm$1.7 71.2
[…]
WL subtree 72.6 80.3 62.7 75.5 52.0 63.6 69.1 69.8 70.7 59.4 80.0 70.6 68.9
GCN 49.5$\pm$0.4 71.1$\pm$0.4 46.8$\pm$0.5 33.7$\pm$2.8 32.7$\pm$0.4 27.4$\pm$0.1 56.2$\pm$1.5 55.3$\pm$0.4 58.2$\pm$1.7 51.0$\pm$0.2 60.7$\pm$3.7 53.2$\pm$0.2 49.6
GIN 67.3$\pm$2.7 67.9$\pm$4.8 61.5$\pm$4.2 65.4$\pm$3.7 58.9$\pm$4.1 61.0$\pm$3.4 62.5$\pm$3.2 66.2$\pm$2.1 69.7$\pm$0.9 56.8$\pm$0.7 72.4$\pm$2.8 64.0$\pm$1.6 64.5
GMT 50.3$\pm$1.2 42.5$\pm$3.4 51.1$\pm$3.7 42.5$\pm$4.5 56.1$\pm$4.7 42.5$\pm$4.1 53.2$\pm$4.9 51.0$\pm$0.2 68.2$\pm$0.4 51.0$\pm$0.3 68.2$\pm$0.5 53.2$\pm$0.4 52.5
CIN 51.1$\pm$0.2 72.6$\pm$0.1 54.0$\pm$0.9 72.6$\pm$0.2 68.2$\pm$0.3 71.5$\pm$1.3 55.0$\pm$2.1 53.5$\pm$1.8 68.2$\pm$0.3 52.0$\pm$0.3 68.3$\pm$0.1 53.6$\pm$0.6 61.7
SpikeGCN 62.8$\pm$0.8 73.1$\pm$1.2 61.4$\pm$0.8 70.9$\pm$0.6 57.7$\pm$1.6 66.2$\pm$1.1 61.2$\pm$1.6 64.5$\pm$1.0 62.3$\pm$1.4 57.3$\pm$0.9 68.9$\pm$1.2 60.1$\pm$1.0 63.6
DRSGNN 64.3$\pm$0.6 76.3$\pm$0.9 56.7$\pm$1.1 73.2$\pm$0.8 58.6$\pm$1.4 63.9$\pm$1.8 63.0$\pm$2.1 65.1$\pm$1.6 64.1$\pm$1.9 59.2$\pm$2.2 70.8$\pm$2.5 56.6$\pm$1.4 64.3
CDAN 59.6$\pm$0.3 73.8$\pm$0.5 56.7$\pm$1.4 73.7$\pm$0.3 71.2$\pm$0.4 73.2$\pm$0.3 55.5$\pm$0.2 57.3$\pm$1.1 69.9$\pm$0.2 54.6$\pm$2.0 69.8$\pm$1.4 56.6$\pm$0.3 64.3
ToAlign 51.0$\pm$0.2 27.4$\pm$0.1 53.2$\pm$0.4 27.4$\pm$0.2 68.2$\pm$0.3 27.4$\pm$0.3 53.2$\pm$0.1 51.0$\pm$0.2 68.2$\pm$0.2 51.0$\pm$0.4 68.2$\pm$0.3 53.2$\pm$0.2 50.0
MetaAlign 65.0$\pm$0.7 77.6$\pm$1.6 62.0$\pm$0.6 77.1$\pm$0.9 68.2$\pm$0.8 74.5$\pm$2.0 64.2$\pm$0.9 65.4$\pm$0.3 68.0$\pm$0.3 56.1$\pm$2.3 68.2$\pm$0.1 66.2$\pm$1.1 67.7
DEAL 65.6$\pm$0.6 73.0$\pm$0.9 58.0$\pm$0.3 71.6$\pm$1.6 60.1$\pm$2.8 73.1$\pm$0.5 62.8$\pm$1.0 65.0$\pm$2.4 65.8$\pm$0.8 53.9$\pm$2.6 57.6$\pm$2.8 56.7$\pm$3.1 63.6
CoCo 70.4$\pm$0.7 80.4$\pm$0.9 62.4$\pm$0.8 75.8$\pm$1.2 65.7$\pm$2.0 73.7$\pm$0.3 67.0$\pm$0.8 70.4$\pm$0.7 69.7$\pm$0.4 62.7$\pm$0.9 74.4$\pm$0.5 63.7$\pm$0.9 69.7
SGDA OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
DGDA OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
A2GNN 59.2$\pm$0.8 58.7$\pm$0.5 59.0$\pm$1.1 58.7$\pm$0.8 58.9$\pm$0.6 59.2$\pm$1.2 58.7$\pm$0.6 58.6$\pm$1.2 59.0$\pm$1.0 59.5$\pm$0.6 58.7$\pm$0.5 58.5$\pm$1.1 58.9
PA-BOTH 57.6$\pm$0.5 58.4$\pm$0.4 58.9$\pm$0.6 57.4$\pm$0.6 57.1$\pm$1.0 58.4$\pm$0.5 58.0$\pm$1.0 58.1$\pm$0.5 58.4$\pm$0.6 57.7$\pm$1.1 57.5$\pm$0.6 58.0$\pm$0.4 58.0
DeSGDA 66.0$\pm$0.5 76.2$\pm$0.4 62.8$\pm$0.6 77.6$\pm$0.6 68.5$\pm$1.0 74.6$\pm$0.5 65.1$\pm$1.0 65.4$\pm$0.5 70.9$\pm$0.6 64.8$\pm$1.1 72.6$\pm$0.6 66.3$\pm$0.4 69.3
[…]
WL subtree 71.6 72.1 62.1 71.2 57.8 67.7 64.0 75.3 41.1 59.2 55.9 55.4 62.8
GCN 66.5$\pm$0.4 60.0$\pm$0.8 55.4$\pm$0.3 60.0$\pm$0.1 39.6$\pm$0.3 40.0$\pm$0.4 55.4$\pm$0.2 66.5$\pm$0.1 39.6$\pm$0.6 33.5$\pm$0.3 39.6$\pm$0.1 44.7$\pm$0.2 50.1
GIN 71.4$\pm$4.7 73.4$\pm$3.4 60.8$\pm$2.7 66.0$\pm$3.4 50.5$\pm$3.7 51.6$\pm$1.8 64.8$\pm$1.0 71.3$\pm$3.5 48.3$\pm$4.2 57.4$\pm$3.8 55.1$\pm$3.4 52.6$\pm$4.3 60.3
GMT 67.4$\pm$1.0 61.7$\pm$2.1 55.8$\pm$0.7 57.0$\pm$2.4 60.2$\pm$0.5 58.2$\pm$2.0 57.8$\pm$2.1 65.7$\pm$1.3 60.2$\pm$0.3 57.3$\pm$2.3 60.7$\pm$0.6 57.1$\pm$1.2 59.9
CIN 70.4$\pm$2.8 66.5$\pm$4.3 58.5$\pm$2.6 64.2$\pm$2.7 60.6$\pm$3.0 64.2$\pm$3.2 58.7$\pm$2.4 69.1$\pm$2.7 57.5$\pm$3.4 67.7$\pm$2.1 59.5$\pm$2.3 56.1$\pm$1.2 62.7
SpikeGCN 66.5$\pm$0.9 60.1$\pm$1.5 55.9$\pm$0.8 60.2$\pm$0.6 54.8$\pm$2.1 59.9$\pm$1.6 55.8$\pm$0.8 62.9$\pm$1.6 58.4$\pm$1.2 61.1$\pm$1.3 58.8$\pm$1.9 62.1$\pm$1.5 59.8
DRSGNN 67.5$\pm$1.2 61.2$\pm$0.7 55.6$\pm$1.4 61.1$\pm$0.9 52.4$\pm$2.3 61.0$\pm$1.2 56.9$\pm$0.7 66.7$\pm$1.5 60.3$\pm$0.5 62.0$\pm$2.0 59.8$\pm$1.4 59.2$\pm$1.1 60.3
CDAN 72.9$\pm$0.2 74.0$\pm$0.3 62.7$\pm$0.3 73.8$\pm$0.5 61.2$\pm$1.0 70.0$\pm$1.2 62.8$\pm$0.1 73.0$\pm$0.3 60.6$\pm$0.2 71.6$\pm$1.5 60.5$\pm$0.2 61.1$\pm$1.4 67.0
ToAlign 68.0$\pm$3.8 73.4$\pm$2.7 64.5$\pm$1.1 63.7$\pm$2.4 60.6$\pm$1.2 61.9$\pm$1.3 64.8$\pm$1.3 74.0$\pm$1.3 60.0$\pm$0.6 65.7$\pm$3.1 61.0$\pm$1.4 56.2$\pm$2.3 64.5
MetaAlign 73.6$\pm$0.2 72.7$\pm$1.9 63.9$\pm$1.0 67.9$\pm$4.3 60.4$\pm$0.7 65.4$\pm$1.8 65.2$\pm$0.8 73.2$\pm$2.3 60.0$\pm$0.6 66.7$\pm$2.4 61.2$\pm$1.1 56.8$\pm$2.1 65.6
DEAL 75.4$\pm$0.3 74.6$\pm$1.1 66.1$\pm$0.6 74.6$\pm$0.8 53.8$\pm$1.0 69.6$\pm$1.8 66.4$\pm$0.3 73.9$\pm$0.6 61.6$\pm$1.4 69.8$\pm$0.2 60.7$\pm$1.0 58.3$\pm$0.9 67.1
CoCo 74.6$\pm$0.9 77.2$\pm$0.6 64.1$\pm$3.4 73.8$\pm$1.1 60.5$\pm$0.2 71.5$\pm$0.7 65.9$\pm$0.5 76.0$\pm$0.5 61.4$\pm$0.4 72.6$\pm$0.6 59.6$\pm$1.0 64.7$\pm$1.0 68.5
SGDA 56.6$\pm$0.6 56.9$\pm$0.8 55.3$\pm$1.2 54.6$\pm$0.5 57.9$\pm$1.3 58.3$\pm$0.4 56.1$\pm$0.9 55.9$\pm$0.6 54.6$\pm$1.3 56.7$\pm$0.5 53.3$\pm$0.7 56.8$\pm$1.1 56.1
DGDA OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
A2GNN 55.4$\pm$0.8 56.1$\pm$0.6 56.7$\pm$1.0 55.3$\pm$0.5 54.9$\pm$0.7 57.2$\pm$0.9 55.7$\pm$0.5 56.5$\pm$1.3 54.5$\pm$0.6 56.8$\pm$0.5 56.2$\pm$1.0 58.8$\pm$0.8 56.1
PA-BOTH 56.1$\pm$0.5 56.0$\pm$0.4 56.3$\pm$0.7 56.4$\pm$0.4 56.0$\pm$0.6 57.1$\pm$0.7 56.2$\pm$1.1 58.3$\pm$0.9 56.5$\pm$0.6 57.2$\pm$0.9 56.9$\pm$0.4 57.7$\pm$0.8 56.8
DeSGDA 75.9$\pm$1.0 74.7$\pm$1.2 66.8$\pm$1.5 71.3$\pm$1.9 63.2$\pm$1.5 69.4$\pm$1.7 66.5$\pm$1.3 74.3$\pm$2.4 62.6$\pm$1.5 73.0$\pm$1.8 61.6$\pm$2.2 63.6$\pm$1.5 68.6
[…]
WL subtree 74.9 74.8 67.3 69.9 57.8 57.9 73.7 80.2 60.0 57.9 70.2 73.1 68.1
GCN 73.0$\pm$1.7 68.7$\pm$1.5 66.8$\pm$3.5 69.2$\pm$0.9 53.9$\pm$3.4 53.4$\pm$2.7 69.3$\pm$0.8 74.0$\pm$1.1 55.1$\pm$1.3 42.6$\pm$1.9 55.5$\pm$3.5 57.9$\pm$2.9 61.6
GIN 74.1$\pm$1.8 73.4$\pm$3.4 65.4$\pm$1.5 70.4$\pm$2.9 58.9$\pm$2.7 61.2$\pm$1.1 73.2$\pm$3.8 77.7$\pm$3.0 63.1$\pm$3.7 63.9$\pm$2.4 67.4$\pm$2.3 73.2$\pm$1.9 68.5
GMT 69.0$\pm$4.0 67.4$\pm$3.8 60.3$\pm$4.2 66.5$\pm$3.8 54.9$\pm$1.6 54.8$\pm$3.6 65.6$\pm$4.2 70.4$\pm$3.2 64.0$\pm$2.3 56.8$\pm$4.3 64.7$\pm$1.5 61.1$\pm$3.5 63.0
CIN 68.5$\pm$2.1 65.1$\pm$2.6 65.4$\pm$1.3 63.6$\pm$2.8 57.3$\pm$3.4 59.0$\pm$3.1 59.3$\pm$1.5 68.3$\pm$1.3 58.1$\pm$2.4 71.1$\pm$3.1 60.7$\pm$1.7 61.7$\pm$2.4 63.2
SpikeGCN 66.7$\pm$1.5 65.5$\pm$2.0 57.9$\pm$0.4 60.2$\pm$1.6 53.2$\pm$1.4 60.1$\pm$1.5 57.7$\pm$1.2 67.3$\pm$1.5 57.7$\pm$2.1 60.1$\pm$1.9 59.9$\pm$2.4 63.3$\pm$1.8 60.1
DRSGNN 66.9$\pm$1.2 62.1$\pm$0.7 57.1$\pm$1.2 63.3$\pm$2.1 56.6$\pm$0.9 62.1$\pm$1.3 56.9$\pm$1.0 67.2$\pm$1.8 58.1$\pm$0.6 61.3$\pm$2.5 58.8$\pm$1.0 64.7$\pm$1.7 61.3
CDAN 74.2$\pm$0.3 73.7$\pm$0.5 68.8$\pm$0.2 71.8$\pm$0.4 59.9$\pm$2.0 58.6$\pm$1.9 70.7$\pm$1.4 74.3$\pm$0.3 59.2$\pm$1.2 69.0$\pm$1.6 60.0$\pm$1.2 62.7$\pm$1.3 66.9
ToAlign 75.5$\pm$1.9 67.1$\pm$3.8 68.1$\pm$1.5 63.3$\pm$2.7 55.6$\pm$1.2 67.3$\pm$4.3 69.4$\pm$3.3 77.0$\pm$1.2 57.6$\pm$1.6 74.9$\pm$2.4 59.0$\pm$3.3 64.6$\pm$3.4 66.6
MetaAlign 74.5$\pm$0.9 73.8$\pm$0.6 69.4$\pm$1.2 72.6$\pm$1.3 59.8$\pm$1.8 70.7$\pm$2.7 72.0$\pm$0.5 75.6$\pm$0.6 62.4$\pm$2.1 72.3$\pm$1.9 62.2$\pm$1.1 72.0$\pm$1.2 69.7
DEAL 76.3$\pm$0.2 72.4$\pm$0.7 68.8$\pm$1.0 72.5$\pm$0.7 57.6$\pm$0.6 67.6$\pm$1.9 77.4$\pm$0.6 80.0$\pm$0.7 64.9$\pm$0.7 72.8$\pm$1.4 70.3$\pm$0.3 76.2$\pm$1.3 71.4
CoCo 77.5$\pm$0.4 75.7$\pm$1.3 68.3$\pm$3.7 74.9$\pm$0.5 65.1$\pm$2.1 74.0$\pm$0.4 76.9$\pm$0.6 77.4$\pm$3.4 66.4$\pm$1.5 71.2$\pm$2.7 62.8$\pm$4.2 77.1$\pm$0.6 72.2
SGDA OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
DGDA OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
A2GNN 55.3$\pm$0.3 54.9$\pm$0.6 55.8$\pm$0.4 55.1$\pm$0.8 54.2$\pm$1.0 57.1$\pm$1.2 56.1$\pm$0.5 55.2$\pm$0.7 57.9$\pm$1.5 56.3$\pm$0.6 54.4$\pm$0.5 58.1$\pm$1.5 55.8
PA-BOTH 56.3$\pm$0.5 57.7$\pm$0.9 56.9$\pm$0.6 56.2$\pm$1.0 55.7$\pm$0.8 56.5$\pm$0.9 57.8$\pm$1.2 56.9$\pm$2.1 56.5$\pm$1.5 56.2$\pm$1.8 56.8$\pm$1.4 57.4$\pm$0.7 56.8
DeSGDA 75.8$\pm$1.4 74.5$\pm$1.7 69.5$\pm$1.3 75.0$\pm$2.0 61.0$\pm$1.5 69.2$\pm$1.3 69.5$\pm$1.6 76.1$\pm$1.5 65.0$\pm$1.4 75.5$\pm$2.2 63.4$\pm$1.8 68.3$\pm$1.3 70.3

### v3

@@
 Abstract
-Spiking Graph Networks (SGNs) have garnered significant attraction from both researchers and industry due to their ability to address energy consumption challenges in graph classification.
+Spiking Graph Networks (SGNs) have garnered significant attraction from researchers and industry due to their ability to address energy consumption challenges in graph classification.
 However, SGNs are only effective for in-distribution data and cannot tackle out-of-distribution data.
 In this paper, we first propose the domain adaptation problem in SGNs, and introduce a novel framework named Degree-aware Spiking Graph Domain Adaptation for Classification (DeSGDA).
-The proposed DeSGDA addresses the spiking graph domain adaptation problem by three aspects: node degree-aware personalized spiking representation, adversarial feature distribution alignment, and pseudo-label distillation.
+The proposed DeSGDA addresses the spiking graph domain adaptation problem in three aspects: node degree-aware personalized spiking representation, adversarial feature distribution alignment, and pseudo-label distillation.
 First, we introduce the personalized spiking representation method for generating degree-dependent spiking signals.
-Specifically, the threshold of triggering a spike is determined by the node degree, allowing this personalized approach to capture more expressive information for classification.
+Specifically, the node degree determines the threshold of triggering a spike, allowing this personalized approach to capture more expressive information for classification.
 Then, we propose the graph feature distribution alignment module that is adversarially trained using membrane potential against a domain discriminator.
 Such an alignment module can efficiently maintain high performance and low energy consumption in the case of inconsistent distribution.
-Additionally, we extract consistent predictions across two spaces to create reliable pseudo-labels, effectively leveraging unlabeled data to enhance graph classification performance.
-Extensive experiments on benchmark datasets validate the superiority of the proposed DeSGDA compared with competitive baselines.
+Additionally, we extract consistent predictions across two spaces to create reliable pseudo-labels, effectively leveraging unlabeled data to enhance classification performance.
+Extensive experiments on benchmark datasets validate the superiority of the DeSGDA compared with competitive baselines.
 1 Introduction
-Spiking Graph Networks (SGNs) Xu et al.
-(2021b); Zhu et al.
-(2022) are a specialized type of artificial neural network engineered to process graph information by mimicking the human brain.
+Spiking Graph Networks (SGNs) Zhu et al.
+(2022); Xu et al.
+(2021b) are a specialized artificial neural network engineered to process graph information by mimicking the human brain.
 SGNs transform static and real-valued graph features into discrete spikes by simulating neurons’ charging and discharging cycles, facilitating spike-based representations for graph node classification.
 Notably, SGNs excel in capturing semantic spiking representations with low energy consumption, which proves advantageous for event-based processing tasks Yao et al.
-(2021); Rosenfeld et al.
-(2021); Chen et al.
-(2023a, b) such as object recognition Gu et al.
+(2021); Wang et al.
+(2024) such as object recognition Gu et al.
 (2020); Li et al.
-(2021b); Liu et al.
-(2023b); Yi et al.
-(2024), real-time data analysis Bauer et al.
+(2021b); Chen et al.
+(2024); Yin et al.
+(2024a), real-time data analysis Zhu et al.
+(2020); Bauer et al.
+(2019); Yin et al.
+(2023b), and graph classification Li et al.
 (2023); Zhu et al.
-(2020); Baronig et al.
-(2024), and graph classification Li et al.
-(2023); Yao et al.
-(2023b); Meng et al.
-(2024); Ju et al.
-(2024).
+(2022); Liu et al.
+(2024c); Yin et al.
+(2022a).
 Currently, SGNs are usually tested within the same distribution as the training dataset Li et al.
 (2023); Yin et al.
-(2024); Duan et al.
+(2024a); Duan et al.
 (2024).
-However, in realistic scenarios, the testing set can have different distributions from the training set, and such a distribution shift may lead to a degradation in performance.
+However, in realistic scenarios, the testing set can have different distributions from the training set, and such a distribution shift may degrade performance.
 For instance, Electroencephalography (EEG) data Binnie and Prior (1994); Biasiucci et al.
-(2019); Peng et al.
 (2019), typically represented as a graph structure with nodes for neurons and edges for connections, is ideally processed by bio-inspired SGNs that mimic neuronal charging and discharging.
@@
 Traditionally, SNNs transfer learning methods Zhan et al.
-(2021); Zhang et al.
-(2021); Zhan et al.
-(2024); Guo et al.
+(2021); Yin et al.
+(2023c, 2024b); Pang et al.
 (2023) have been applied in event-based or computer vision scenarios.
 However, there’s no existing research on spiking graph domain adaptation.
-Designing an effective spiking graph domain adaptation framework for classification is non-trivial due to the following major challenges: (1) How to meticulously design an SGN under the circumstance of domain shift?
+In this paper, we address the development of energy-efficient SGNs tailored for scenarios involving distribution shifts.
+Both domain adaptation and SGNs are particularly well-suited for real-world applications where data distributions vary across environments, and efficient processing of graph-structured, dynamic data under resource constraints is crucial.
+These challenges are common across numerous fields that require solutions capable of handling distribution shifts while minimizing energy consumption.
+However, designing an effective spiking graph domain adaptation framework is non-trivial due to the following major challenges: (1) How to meticulously design an SGN under the circumstance of domain shift?
 SGNs usually utilize a global threshold for the firing of each node Xu et al.
-(2021a); Hagenaars et al.
-(2021); Yin et al.
-(2024); Zhao et al.
+(2021a); Yin et al.
+(2024a); Zhao et al.
 (2024).
@@
 (2023a); Duan et al.
-(2024); Cai et al.
 (2024).
@@
 (2021); Zhang et al.
-(2021); Liu et al.
-(2024a); Zhan et al.
+(2021); Zhan et al.
 (2024), there is still no theoretical research on spiking graphs under domain shift.
@@
 To address the first challenge, we establish variable node thresholds based on their degrees.
-By adaptively updating these thresholds, we can achieve a more expressive and personalized spiking representation for each node.
-Then, we introduce a adversarial feature distribution alignment module that is adversarially trained using membrane potential against a domain discriminator.
+By adaptively updating these thresholds, we achieve a more expressive and personalized spiking representation for each node.
+Then, we introduce an adversarial feature distribution alignment module trained using membrane potential against a domain discriminator.
 To further enhance performance, we extract consistent predictions from different spaces to generate reliable pseudo-labels.
-Additionally, to explore the generalization ability of the proposed DeSGDA, we first propose the error bound for spiking graph domain adaptation and demonstrate that our pseudo-label distillation module effectively reduces this upper bound.
-Our contributions can be summarized as follows: (1) Problem Formulation: We first introduce the problem of spiking graph domain adaptation for classification, which is non-trivial due to the challenges of the inflexible architecture of SGNs and theoretical deficiency.
-(2) Novel Architecture: We propose DeSGDA, a framework that efficiently learns personalized spiking representations for nodes using degree-aware thresholds and aligns domain distributions through adversarial training on membrane potential.
+Additionally, to explore the generalization ability of the proposed DeSGDA, we propose the error bound for spiking graph domain adaptation and demonstrate that our pseudo-label distillation module effectively reduces this upper bound.
+In summary, we utilize simple yet effective techniques to address a novel problem while providing an insightful analysis of the background mechanisms and capabilities of our proposed method.
+Our contributions can be summarized as follows:
+  • Problem Formulation: We first introduce the problem of spiking graph domain adaptation for classification, which is non-trivial due to the challenges of the inflexible architecture of SGNs and theoretical deficiency.
+  • Novel Architecture: We propose DeSGDA, a framework that efficiently learns personalized spiking representations for nodes using degree-aware thresholds and aligns domain distributions through adversarial training on membrane potential.
 Furthermore, we utilize pseudo-label distillation to improve the performance further.
-(3) Theoretical Analysis: To guarantee the stability of DeSGDA, we provide theoretical proof of the error bound for spiking graph domain adaptation.
-Furthermore, we demonstrate that DeSGDA maintains a lower theoretical bound than standard spiking graph domain adaptation through the effective use of the pseudo-label distillation module.
-(4) Extensive Experiments.
+  • Theoretical Analysis: To guarantee the stability of DeSGDA, we provide theoretical proof of the error bound for spiking graph domain adaptation.
+Furthermore, we demonstrate that DeSGDA maintains a lower theoretical bound than standard spiking graph domain adaptation through the use of the pseudo-label distillation module.
+  • Extensive Experiments.
 We evaluate the proposed DeSGDA on extensive spiking graph domain adaptation learning datasets, which shows that our proposed DeSGDA outperforms the variety of state-of-the-art methods.
 2 Related work
+Figure 1: An overview of the proposed DeSGDA.
+To achieve personalized spiking representations, DeSGDA employs adaptive thresholds based on node degrees, enabling the generation of tailored spiking representations.
+To align domain distributions, DeSGDA leverages adversarial training on membrane potentials to counter domain discrimination.
+Furthermore, DeSGDA utilizes pseudo-labeling to identify and select reliable samples, thereby enhancing overall model performance.
 Spiking Graph Networks (SGNs).
@@
 (2023); Yao et al.
-(2023a); Qiu et al.
-(2024); Duan et al.
+(2023a, b); Duan et al.
+(2024); Liang et al.
 (2023).
-Existing research on SGNs focuses on capturing the dynamic temporal information contained within graphs and enhancing model scalability.
+Existing research on SGNs focuses on capturing the dynamic temporal information within graphs and enhancing scalability.
 For instance, Xu et al.
@@
 Additionally, Yin et al.
-(2024) adapts SNNs to dynamic graph settings and employs implicit differentiation for the node classification task.
-However, existing methods still suffer from data distribution shift issues when training and testing data come from different domains, resulting in degraded model performance and generalization.
+(2024a) adapts SNNs to dynamic graph settings and employs implicit differentiation for the node classification task.
+However, existing methods still suffer from data distribution shift issues when training and testing data come from different domains, resulting in degraded performance and generalization.
 To address this, we propose a novel domain adaptation method based on SGNs to tackle these challenges.
 Spiking Transfer Learning.
-Spiking transfer learning focuses on adjusting SNNs to handle data distribution shifts across various domains effectively.
-Recent advances in spiking transfer learning have been extensively applied in vision tasks, enhancing model performance while maintaining energy efficiency Zhan et al.
+Spiking transfer learning focuses on adjusting SNNs to handle data distribution shifts across various domains more effectively.
+Recent advances in spiking transfer learning have been extensively applied in vision tasks, significantly enhancing model performance while maintaining energy efficiency Zhan et al.
 (2021); Zhang et al.
@@
 For instance, Guo et al.
-(2024) leverages a Jaccard attention mechanism within SNNs to effectively adapt to target domains without requiring source domain data.
+(2024) leverages a Jaccard attention mechanism within SNNs to adapt to target domains without requiring source domain data.
 Similarly, He et al.
@@
 (2024) converts RGB images into spike-based neuromorphic data, enabling SNNs to process visual information across various domains efficiently.
-However, the difficulty of graph topologies makes it infeasible to apply spiking transfer learning to SGNs directly.
-To this end, we introduce a specialized domain adaptation method tailored for SGNs.
+However, the inherent difficulty of complex graph topologies makes it infeasible to apply spiking transfer learning to SGNs directly.
+To this end, we introduce a specialized domain adaptation method specifically tailored for SGNs.
 3 Preliminaries
@@
 Theorem 1
-(You et al., 2023) Assuming that the learned discriminator is $C_{g}$-Lipschitz continuous as described in (Redko et al., 2017), and the graph feature extractor $f$ (also referred to as GNN) is $C_{f}$-Lipschitz that $||f||_{Lip}=\max_{G_{1},G_{2}}\frac{||f(G_{1})-f(G_{2})||_{2}}{\eta(G_{1},G_{2})}=C_{f}$ for some graph distance measure $\eta$.
+You et al.
+(2023) Assuming that the learned discriminator is $C_{g}$-Lipschitz continuous as described in Redko et al.
+(2017), and the graph feature extractor $f$ (also referred to as GNN) is $C_{f}$-Lipschitz that $||f||_{Lip}=\max_{G_{1},G_{2}}\frac{||f(G_{1})-f(G_{2})||_{2}}{\eta(G_{1},G_{2})}=C_{f}$ for some graph distance measure $\eta$.
 Let :=$\mathcal{H}:=\{h:\mathcal{G}\rightarrow\mathcal{Y}\}$ be the set of bounded real-valued functions with the pseudo-dimension $Pdim(\mathcal{H})=d$ that $h=g\circ f\in\mathcal{H}$, with probability at least $1-\delta$ the following inequality holds:
-$\epsilon_{T}(h,\hat{h})\leq$ $\hat{\epsilon}_{S}(h,\hat{h})+\sqrt{\frac{4d}{N_{S}}\log(\frac{eN_{S}}{d})+\frac{1}{N_{S}}\log(\frac{1}{\delta})}+2C_{f}C_{g}W_{1}(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G))+\omega,$
+$\epsilon_{T}(h,\hat{h})\leq$ $\ \hat{\epsilon}_{S}(h,\hat{h})+\sqrt{\frac{4d}{N_{S}}\log\left(\frac{eN_{S}}{d}\right)+\frac{1}{N_{S}}\log\left(\frac{1}{\delta}\right)}$
+$+2C_{f}C_{g}W_{1}(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G))+\omega,$
 where the (empirical) source and target risks are $\hat{\epsilon}_{S}(h,\hat{h})=\frac{1}{N_{S}}\sum_{n=1}^{N_{S}}|h(G_{n})-\hat{h}(G_{n})|$ and $\epsilon_{T}(h,\hat{h})=\mathbb{E}_{\mathbb{P}_{T}(G}\{|h(G)-\hat{h}(G)|\}$, respectively, where $\hat{h}:\mathcal{G}\rightarrow\mathcal{Y}$ is the labeling function for graphs and $\omega=\min_{||g||_{Lip}\leq C_{g},||f||_{Lip}\leq C_{f}}\{\epsilon_{S}(h,\hat{h})+\epsilon_{T}(h,\hat{h})\}$.
-The first Wasserstein distance is defined as (Villani et al., 2009): $W_{1}(\mathbb{P},\mathbb{Q})=\sup_{||g||_{Lip}\leq 1}\left\{\mathbb{E}_{\mathbb{P}_{S}(Z)}g(Z)-\mathbb{E}_{\mathbb{P}_{T}(Z)}g(Z)\right\}$.
+The first Wasserstein distance is defined as Villani et al.
+(2009): $W_{1}(\mathbb{P},\mathbb{Q})=\sup_{||g||_{Lip}\leq 1}\left\{\mathbb{E}_{\mathbb{P}_{S}(Z)}g(Z)-\mathbb{E}_{\mathbb{P}_{T}(Z)}g(Z)\right\}$.
 The comprehensive justification of the OT-based graph domain adaptation bound demonstrates that the generalization gap relies on both the domain divergence $2C_{f}C_{g}W_{1}(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G))$ and model discriminability $\omega$.
 Spiking Graph Networks.
-In contrast to traditional artificial neural networks, SGNs (Xu et al., 2021a; Zhu et al., 2022) convert input data into binary spikes over time, with each neuron in the SGNs maintaining a membrane potential that accumulates input spikes.
+In contrast to traditional artificial neural networks, SGNs Xu et al.
+(2021a); Zhu et al.
+(2022) convert input data into binary spikes over time, with each neuron in the SGNs maintaining a membrane potential that accumulates input spikes.
 A spike is produced as an output when the membrane potential exceeds a threshold, which is formulated as:
-$u_{\tau+1,i}=\lambda(u_{\tau,i}-V_{th}s_{\tau,i})+\sum_{j}w_{ij}\mathcal{A}\left(A,s_{\tau,j}\right)+b,\quad s_{\tau+1,i}=\mathbb{H}(u_{\tau+1,i}-V_{th}),$ (1)
-where $\mathbb{H}(x)$ is the Heaviside function, which is the non-differentiable spiking function.
-$\mathcal{A}$ is the graph aggregation operation, and $A$ is the adjacency matrix of graph.
+$u_{\tau+1,i}=\lambda(u_{\tau,i}-V_{th}s_{\tau,i})+\sum_{j}w_{ij}\mathcal{A}\left(A,s_{\tau,j}\right)+b,\vspace{-4pt}$
+$s_{\tau+1,i}=\mathbb{H}(u_{\tau+1,i}-V_{th}),$ (1)
+where $\mathbb{H}(x)$ is the Heaviside function, which is the non-differentiable spiking function, $\mathcal{A}$ is the graph aggregation operation, and $A$ is the adjacency matrix of a graph.
 $s_{\tau,i}$ denotes the binary spike train of neuron $i$, and $\lambda$ is the constant.
@@
 The overview of DeSGDA is shown in Figure 1.
-Figure 1: An overview of the proposed DeSGDA.
-DeSGDA first utilizes different thresholds for different degree nodes to obtain personalized spiking representation.
-To align domain distribution, we use adversarial training on membrane potential against domain discrimination.
-Besides, DeSGDA applies the pseudo-label to enhance model performance.
 Problem Setup.
@@
 Specifically, we first $set$ all the degrees of nodes in the source domain graphs, i.e., $D^{s}=set(D_{1}^{s}\cup\cdots\cup D_{N_{s}}^{s})$, where $D_{i}$ denotes the degree set of graph $G_{i}^{s}$, and $set(\cdot)$ operation is an unordered sequence of non-repeating elements.
-Considering that low-degree nodes are more challenging to trigger while high-degree nodes trigger more easily, we propose setting higher thresholds for high-degree nodes and lower thresholds for low-degree nodes, which is formulated as:
-$s_{\tau}^{d_{i}^{s}}=\mathbb{H}(u_{\tau}-V_{th}^{d_{i}^{s}}),\quad S^{d_{i}^{s}}=avg(s_{\tau}^{d_{i}^{s}}),\quad V_{th}^{d_{i}^{s}}=(1-\alpha)V_{th}^{d_{i}^{s}}+\alpha S^{d_{i}^{s}},$ (2)
+Considering that low-degree nodes are more challenging to trigger while high-degree nodes trigger more easily, we propose setting higher thresholds for high-degree nodes and lower for low-degree nodes, which is formulated as:
+$s_{\tau}^{d_{i}^{s}}=\mathbb{H}(u_{\tau}$ $-V_{th}^{d_{i}^{s}}),\quad S^{d_{i}^{s}}=avg(s_{\tau}^{d_{i}^{s}}),$ (2)
+$V_{th}^{d_{i}^{s}}$ $=(1-\alpha)V_{th}^{d_{i}^{s}}+\alpha S^{d_{i}^{s}},$
 where $V_{th}^{d_{i}^{s}}$ is the threshold of degree $d_{i}^{s}\in D^{s}$, initially set to $V_{th}$, and $\alpha$ is a hyper-parameter.
@@
 Consequently, high-degree nodes tend to achieve high $S^{d_{i}^{s}}$, which leads to an iterative increase in the threshold corresponding to degree $d_{i}^{s}$ and conversely for lower-degree nodes.
+To further explore the background mechanism of the hypothesis, we have the following analysis.
+Hypothesis 1
+In graph spiking networks, nodes with low-degree are more challenging to trigger while high-degree nodes trigger more easily.
+The details analysis are introduced in Appendix A.
 With different thresholds for different node degrees, we can obtain the personalized node spiking representation $\mathbf{s}_{v\in G_{i}^{s}}^{d_{j}^{s}}$.
 Then, we summarize all node representations with a readout function into the graph-level representation and output the prediction with a multi-layer perception (MLP) classifier:
-$\mathbf{s}_{i}=\operatorname{READOUT}\left(\left\{\mathbf{s}_{v}^{d_{j}^{s}}\right\}_{v\in G_{i}^{s}}\right),\quad\hat{y}_{i}^{s}=H(\mathbf{s}_{i}),$ (3)
+$\mathbf{s}_{i}=$ $\operatorname{READOUT}\left(\left\{\mathbf{s}_{v}^{d_{j}^{s}}\right\}_{v\in G_{i}^{s}}\right),$ (3)
+$\hat{y}_{i}^{s}=H(\mathbf{s}_{i}),$
 where $\hat{y}_{i}$ is the predicted result and $H(\cdot)$ is the classifier. After that, the source classification loss is:
@@
 The encoder and classifier are adversarial trained to align the feature spaces of the source and target domains.
-$\mathcal{L}_{AD}=\mathbb{E}_{G_{i}^{s}\in\mathcal{D}^{s}}\log Q\left(F(G_{i}^{s}),H(G_{i}^{s})|V_{th}^{D^{s}}\right)+\mathbb{E}_{G_{j}^{t}\in\mathcal{D}^{t}}\log\left(1-Q\left(F\left(G_{j}^{t}\right),H(G_{j}^{t})|V_{th}^{D^{s}}\right)\right).$
+$\mathcal{L}_{AD}$ $=\mathbb{E}_{G_{i}^{s}\in\mathcal{D}^{s}}\log Q\left(F(G_{i}^{s}),H(G_{i}^{s})|V_{th}^{D^{s}}\right)$
+$+\mathbb{E}_{G_{j}^{t}\in\mathcal{D}^{t}}\log\left(1-Q\left(F\left(G_{j}^{t}\right),H(G_{j}^{t})|V_{th}^{D^{s}}\right)\right).$
 However, the degree in the target domain may be unseen by the source. Thus, we further initialize the threshold with $V_{th}^{d_{j}^{t}}$ and $d_{j}^{t}\notin D^{s}$, which is formulated as:
-$\mathcal{L}_{AD}=$ $\mathbb{E}_{\begin{subarray}{c}G_{j}^{t}\in\mathcal{D}^{t}\\ D_{j}^{t}\subset{D^{s}}\end{subarray}}\log\left(1-Q\left(F\left(G_{j}^{t}\right),H(G_{j}^{t})|V_{th}^{D^{s}}\right)\right)+\mathbb{E}_{G_{i}^{s}\in\mathcal{D}^{s}}\log Q\left(F(G_{i}^{s}),H(G_{i}^{s})|V_{th}^{D^{s}}\right)$ (5)
+$\mathcal{L}_{AD}$ $=\mathbb{E}_{\begin{subarray}{c}G_{j}^{t}\in\mathcal{D}^{t}\\ D_{j}^{t}\subset{D^{s}}\end{subarray}}\log\left(1-Q\left(F\left(G_{j}^{t}\right),H(G_{j}^{t})|V_{th}^{D^{s}}\right)\right)$
+$+\mathbb{E}_{G_{i}^{s}\in\mathcal{D}^{s}}\log Q\left(F(G_{i}^{s}),H(G_{i}^{s})|V_{th}^{D^{s}}\right)$
 $+\mathbb{E}_{\begin{subarray}{c}G_{j}^{t}\in\mathcal{D}^{t}\\ \exists d_{j}^{t}\notin{D^{s}}\end{subarray}}\log\left(1-Q\left(F\left(G_{j}^{t}\right),H(G_{j}^{t})|V_{th}^{D^{s}},V_{th}^{D^{t}}\right)\right),$
 where $D^{t}=\{d_{i}^{t}|d_{i}^{t}\in D^{t},d_{i}^{t}\notin D^{s}\}$.
-Then, we iteratively update $V_{th}^{D^{t}}$ with Eq.
+We iteratively update $V_{th}^{D^{t}}$ with Eq.
 2 on each latency.
-Furthermore, we present an upper bound on the adversarial distribution alignment.
+Furthermore, we present an upper bound for adversarial distribution alignment.
 Theorem 2
@@
 Let :=$\mathcal{H}:=\{h:\mathcal{G}\rightarrow\mathcal{Y}\}$ be the set of bounded real-valued functions with the pseudo-dimension $Pdim(\mathcal{H})=d$ that $h=g\circ f\in\mathcal{H}$, and provided the spike training data set $S_{n}=\{(\mathbf{X}_{i},y_{i})\in\mathcal{X}\times\mathcal{Y}\}_{i\in[n]}$ drawn from $\mathcal{D}^{s}$, with probability at least $1-\delta$ the following inequality :
-$\epsilon_{T}(h,\hat{h}_{T}(\mathbf{X}))\leq$ $\hat{\epsilon}_{S}(h,\hat{h}_{S}(\mathbf{S}))+2\mathbb{E}\left[\sup\sup\frac{1}{N_{S}}\sum_{i=1}^{N_{S}}\epsilon_{i}h(\mathbf{X}_{i},y_{i},p_{i})\right]+C\sqrt{\frac{ln(2/\delta)}{N_{S}}}$ (6)
-$+\min\left(|\epsilon_{S}(h,\hat{h}_{S}(\mathbf{X}))-\epsilon_{S}(h,\hat{h}_{T}(\mathbf{X}))|,|\epsilon_{T}(h,\hat{h}_{S}(\mathbf{X}))-\epsilon_{T}(h,\hat{h}_{T}(\mathbf{X}))|\right)$
-$+2C_{f}C_{g}W_{1}\left(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G)\right),$
+$\epsilon_{T}(h,$ $\hat{h}_{T}(\mathbf{X}))\leq\hat{\epsilon}_{S}(h,\hat{h}_{S}(\mathbf{S}))+2\mathbb{E}\left[\sup\frac{1}{N_{S}}\sum_{i=1}^{N_{S}}\epsilon_{i}h(\mathbf{X}_{i},y_{i},p_{i})\right]$
+$+C\sqrt{\frac{ln(2/\delta)}{N_{S}}}+\min\Big{(}|\epsilon_{S}(h,\hat{h}_{S}(\mathbf{X}))-\epsilon_{S}(h,\hat{h}_{T}(\mathbf{X}))|,$
+$|\epsilon_{T}(h,\hat{h}_{S}(\mathbf{X}))-\epsilon_{T}(h,\hat{h}_{T}(\mathbf{X}))|\Big{)}+2C_{f}C_{g}W_{1}\left(\mathbb{P}_{S},\mathbb{P}_{T}\right),$
 where the (empirical) source and target risks are $\hat{\epsilon}_{S}(h,\hat{h}(\mathbf{S}))=\frac{1}{N_{S}}\sum_{n=1}^{N_{S}}|h(\mathbf{S}_{n})-\hat{h}(\mathbf{S}_{n})|$ and $\epsilon_{T}(h,\hat{h}(\mathbf{X}))=\mathbb{E}_{\mathbb{P}_{T}(G}\{|h(G)-\hat{h}(G)|\}$, respectively, where $\hat{h}:\mathcal{G}\rightarrow\mathcal{Y}$ is the labeling function for graphs and $\omega=\min_{||g||_{Lip}\leq C_{g},||f||_{Lip}\leq C_{f}}\{\epsilon_{S}(h,\hat{h}(\mathbf{X}))+\epsilon_{T}(h,\hat{h}(\mathbf{X}))\}$, $\epsilon_{i}$ is the Rademacher variable and $p_{i}$ is the $i^{th}$ row of $\mathbf{P}$, which is the probability matrix with:
-$\mathbf{P}_{kt}=\left\{\begin{array}[]{lr}\exp\left(\frac{u_{k}(t)-V_{th}}{\sigma(u_{k}(t)-u_{reset})}\right),&if\quad u_{\theta}\leq u(t)\leq V_{th},\\ 0,&if\quad u_{reset}\leq u_{k}(t)\leq u_{\theta}.\end{array}\right.$ (7)
-Theorem 2 proves the generalization bound of spiking graph domain adaptation. More details can be found in Appendix A.
+$\mathbf{P}_{kt}=\begin{array}[]{lr}\exp\left(\frac{u_{k}(t)-V_{th}}{\sigma(u_{k}(t)-u_{reset})}\right),&if\quad u_{\theta}\leq u(t)\leq V_{th},\\ 0,&if\quad u_{reset}\leq u_{k}(t)\leq u_{\theta}.\end{array}$ (5)
+Theorem 2 proves the generalization bound of spiking graph domain adaptation. More details can be found in Appendix B.
 4.3 Pseudo-label Distillation for Discrimination Learning
@@
 The goal of the pseudo-label distilling procedure is to keep those examples and their corresponding pseudo-labels from the deep feature space that aligns with the shallow feature space.
-Specifically, we denote $\mathbf{s^{\prime}}_{i}^{t}$ as the shallow spiking graph representation on the $L\textquoteright$ ’-th layer, where $L\textquoteright<L$, and $\hat{y}_{i}^{t}$ as the prediction of graph $G_{i}^{t}$ on the $L$-th layer.
+Specifically, we denote $\mathbf{s^{\prime}}_{i}^{t}$ as the shallow spiking graph representation on the $L^{\prime}$-th layer, where $L^{\prime}<L$, and $\hat{y}_{i}^{t}$ as the prediction of graph $G_{i}^{t}$ on the $L$-th layer.
 Then, to enhance alignment between the shallow and deep feature spaces and facilitate the generation of more accurate pseudo-labels, we cluster the shallow features $\mathbf{s^{\prime}}^{t}$ into $C$ clusters and each cluster $\mathcal{E}_{j}$ includes graphs $\{G_{j}^{t}\}$.
@@
 $\mathcal{L}_{T}=\mathbb{E}_{{G}_{j}^{t}\in\mathcal{P}}l\left({H(\mathbf{s}}_{j}^{t}),\hat{y}_{j}^{t}\right),$ (7)
-where $H(\cdot)$ and $\mathbf{s}_{j}^{t}$ are the classifier and spiking graph representation, respectively, which are defined in Eq.
-3.
-$l(\cdot)$ is the loss function, and we implement it with cross-entropy loss.
+where $H(\cdot)$ and $\mathbf{s}_{j}^{t}$ are the classifier and spiking graph representation, which are defined in Eq. 3. $l(\cdot)$ is the loss function, and we implement it with cross-entropy loss.
 Theorem 3
@@
 Let :=$\mathcal{H}:=\{h:\mathcal{G}\rightarrow\mathcal{Y}\}$ be the set of bounded real-valued functions with the pseudo-dimension $Pdim(\mathcal{H})=d$, and provided the spike training data set $S_{n}=\{(\mathbf{X}_{i}^{s},y_{i}^{s})\}_{i\in[n]}$, with probability at least $1-\delta$ the following inequality holds:
-$\epsilon_{T}(h,\hat{h}_{T}(\mathbf{X}))\leq$ $\frac{N^{\prime}_{T}}{N_{S}+N^{\prime}_{T}}\hat{\epsilon}_{T}(h,\hat{h}_{T}(S))+\frac{N_{S}}{N_{S}+N^{\prime}_{T}}\Bigg{(}\hat{\epsilon}_{S}(h,\hat{h}_{S}(S))+2C_{f}C_{g}W_{1}\left(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G)\right)$ (10)
+$\epsilon_{T}(h,\hat{h}_{T}($ $\mathbf{X}))\leq\frac{N^{\prime}_{T}}{N_{S}+N^{\prime}_{T}}\hat{\epsilon}_{T}(h,\hat{h}_{T}(S))+\frac{N_{S}}{N_{S}+N^{\prime}_{T}}\cdot\Bigg{(}\hat{\epsilon}_{S}(h,\hat{h}_{S}(S))+2C_{f}C_{g}W_{1}\left(\mathbb{P}_{S},\mathbb{P}_{T}\right)$
 $+2\mathbb{E}\left[\sup\frac{1}{N_{S}}\sum_{i=1}^{N_{S}}\epsilon_{i}h(\mathbf{X}_{i},y_{i},p_{i})\right]+C\sqrt{\frac{ln(2/\delta)}{N_{S}}}$
-$+\min\left(|\epsilon_{S}(h,\hat{h}_{S}(\mathbf{X})))-\epsilon_{S}(h,\hat{h}_{T}(\mathbf{X})))|,|\epsilon_{T}(h,\hat{h}_{S}(\mathbf{X})))-\epsilon_{T}(h,\hat{h}_{T}(\mathbf{X})))|\right)\Bigg{)}$
-$\leq$ $\hat{\epsilon}_{S}(h,\hat{h}_{S}(S))+2\mathbb{E}\left[\sup\frac{1}{N_{S}}\sum_{i=1}^{N_{S}}\epsilon_{i}h(\mathbf{X}_{i},y_{i},p_{i})\right]+C\sqrt{\frac{ln(2/\delta)}{N_{S}}}$
-$+2C_{f}C_{g}W_{1}\left(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G)\right)+\omega^{\prime},$
-where the (empirical) source and target risks are $\hat{\epsilon}_{S}(h,\hat{h})=\frac{1}{N_{S}}\sum_{n=1}^{N_{S}}|h(G_{n})-\hat{h}(G_{n})|$ and $\epsilon_{T}(h,\hat{h})=\mathbb{E}_{\mathbb{P}_{T}(G}\{|h(G)-\hat{h}(G)|\}$, respectively, where $\hat{h}:\mathcal{G}\rightarrow\mathcal{Y}$ is the labeling function for graphs and $\omega^{\prime}=\min_{||g||_{Lip}\leq C_{g},||f||_{Lip}\leq C_{f}}\{\epsilon_{S}(h,\hat{h})+\epsilon_{T}(h,\hat{h})\}$, $\epsilon_{i}$ is the Rademacher variable and $p_{i}$ is the $i^{th}$ row of $\mathbf{P}$, which is defined in Eq.
-7.
-The proof is detailed in Appendix B.
+$+\min\Big{(}|\epsilon_{S}(h,\hat{h}_{S}(\mathbf{X})))-\epsilon_{S}(h,\hat{h}_{T}(\mathbf{X})))|,$
+$|\epsilon_{T}(h,\hat{h}_{S}(\mathbf{X})))-\epsilon_{T}(h,\hat{h}_{T}(\mathbf{X})))|\Big{)}\Bigg{)}\hat{\epsilon}_{S}(h,\hat{h}_{S}(S))+2\mathbb{E}\left[\sup\frac{1}{N_{S}}\sum_{i=1}^{N_{S}}\epsilon_{i}h(\mathbf{X}_{i},y_{i},p_{i})\right]$
+$+C\sqrt{\frac{ln(2/\delta)}{N_{S}}}+2C_{f}C_{g}W_{1}\left(\mathbb{P}_{S},\mathbb{P}_{T}\right)+\omega^{\prime},$
+where the (empirical) source and target risks are $\hat{\epsilon}_{S}(h,\hat{h})=\frac{1}{N_{S}}\sum_{n=1}^{N_{S}}|h(G_{n})-\hat{h}(G_{n})|$ and $\epsilon_{T}(h,\hat{h})=\mathbb{E}_{\mathbb{P}_{T}(G}\{|h(G)-\hat{h}(G)|\}$, respectively, where $\hat{h}:\mathcal{G}\rightarrow\mathcal{Y}$ is the labeling function for graphs and $\omega^{\prime}=\min_{||g||_{Lip}\leq C_{g},||f||_{Lip}\leq C_{f}}\{\epsilon_{S}(h,\hat{h})+\epsilon_{T}(h,\hat{h})\}$, $\epsilon_{i}$ is the Rademacher variable and $p_{i}$ is the $i^{th}$ row of $\mathbf{P}$, which is defined in Theorem 2.
+The proof is detailed in Appendix C.
 From Theorem 3, we observe that the bound of DeSGDA is lower than simply aligning the distributions by incorporating the highly reliable pseudo-labels, demonstrating the effectiveness of pseudo labels for spiking graph domain adaptation.
@@
 $\mathcal{L}=\mathcal{L}_{S}+\mathcal{L}_{T}-\lambda\mathcal{L}_{AD},$ (8)
-where $\lambda$ is a hyper-parameter to balance the adversarial training loss and classification loss. The learning procedure is illustrated in Algorithm C, and the complexity is shown in Appendix D .
-[…]
-[…]
-WL subtree 68.7 82.3 50.7 82.3 58.1 83.8 64.0 74.1 43.7 70.5 71.3 60.1 67.5
-GCN 73.4$\pm$0.2 83.5$\pm$0.3 57.6$\pm$0.2 84.2$\pm$1.8 24.0$\pm$0.1 16.6$\pm$0.4 57.6$\pm$0.2 73.7$\pm$0.4 24.0$\pm$0.1 26.6$\pm$0.2 39.9$\pm$0.9 42.5$\pm$0.1 50.3
-GIN 62.5$\pm$4.7 74.9$\pm$3.7 53.0$\pm$4.6 59.6$\pm$4.2 73.7$\pm$0.8 64.7$\pm$3.4 60.6$\pm$2.7 69.8$\pm$0.6 31.1$\pm$2.8 63.1$\pm$3.4 72.3$\pm$2.7 64.6$\pm$1.4 62.5
-GMT 73.4$\pm$0.3 83.5$\pm$0.2 57.6$\pm$0.1 83.5$\pm$0.3 24.0$\pm$0.1 83.5$\pm$0.1 57.4$\pm$0.2 73.4$\pm$0.2 24.1$\pm$0.1 73.4$\pm$0.3 24.0$\pm$0.1 57.6$\pm$0.2 59.6
-CIN 74.5$\pm$0.2 84.1$\pm$0.5 57.8$\pm$0.2 82.7$\pm$0.9 75.6$\pm$0.6 79.2$\pm$2.2 61.5$\pm$2.7 74.0$\pm$1.0 75.5$\pm$0.8 72.5$\pm$2.1 76.0$\pm$0.3 60.9$\pm$1.2 72.9
-SpikeGCN 71.8$\pm$0.8 79.5$\pm$1.3 63.8$\pm$1.0 78.9$\pm$1.4 68.6$\pm$1.1 76.5$\pm$1.8 62.3$\pm$2.2 72.1$\pm$1.5 68.1$\pm$2.1 67.2$\pm$1.9 69.2$\pm$2.1 64.2$\pm$1.8 70.2
-DRSGNN 72.6$\pm$0.6 80.1$\pm$1.6 63.1$\pm$1.4 79.5$\pm$1.8 70.4$\pm$1.9 78.6$\pm$2.1 64.1$\pm$1.7 70.7$\pm$2.3 67.8$\pm$1.6 65.6$\pm$1.4 71.3$\pm$1.3 62.1$\pm$1.0 70.5
-CDAN 72.2$\pm$1.8 82.4$\pm$1.6 59.8$\pm$2.1 76.8$\pm$2.4 69.3$\pm$4.1 71.8$\pm$3.7 64.4$\pm$2.5 74.3$\pm$0.4 46.3$\pm$2.0 69.8$\pm$1.8 74.4$\pm$1.7 62.6$\pm$2.3 68.7
-ToAlign 73.4$\pm$0.1 83.5$\pm$0.2 57.6$\pm$0.1 83.5$\pm$0.2 24.0$\pm$0.3 83.5$\pm$0.4 57.6$\pm$0.1 73.4$\pm$0.1 24.0$\pm$0.2 73.4$\pm$0.2 24.0$\pm$0.1 57.6$\pm$0.3 59.6
-MetaAlign 75.5$\pm$0.9 84.9$\pm$0.6 64.8$\pm$1.6 85.9$\pm$1.1 69.3$\pm$2.7 83.3$\pm$0.6 68.7$\pm$1.2 74.2$\pm$0.7 73.3$\pm$3.3 72.2$\pm$0.9 69.9$\pm$1.8 63.6$\pm$2.3 73.8
-DEAL 76.5$\pm$0.4 83.1$\pm$0.4 67.5$\pm$1.3 77.6$\pm$1.8 76.0$\pm$0.2 80.1$\pm$2.7 66.1$\pm$1.3 75.4$\pm$1.5 42.3$\pm$4.1 68.1$\pm$3.7 73.1$\pm$2.2 67.8$\pm$1.2 71.1
-CoCo 75.5$\pm$0.2 84.2$\pm$0.4 59.8$\pm$0.5 83.4$\pm$0.2 73.6$\pm$2.3 81.6$\pm$2.4 65.8$\pm$0.3 76.2$\pm$0.2 75.8$\pm$0.2 71.1$\pm$2.1 76.1$\pm$0.2 67.1$\pm$0.6 74.2
-SGDA 63.8$\pm$0.6 65.2$\pm$1.3 66.7$\pm$1.0 59.1$\pm$1.5 60.1$\pm$0.8 64.4$\pm$1.2 65.2$\pm$0.7 63.9$\pm$0.9 64.5$\pm$0.6 61.1$\pm$1.3 58.9$\pm$1.4 64.9$\pm$1.2 63.2
-DGDA 58.7$\pm$0.8 59.9$\pm$1.2 57.1$\pm$0.6 57.9$\pm$0.8 59.2$\pm$1.3 58.9$\pm$0.4 61.1$\pm$1.2 60.3$\pm$1.6 58.6$\pm$0.9 57.5$\pm$1.2 58.4$\pm$0.5 62.3$\pm$1.5 59.2
-A2GNN 65.4$\pm$1.3 66.3$\pm$1.1 68.2$\pm$1.4 66.3$\pm$1.2 65.4$\pm$0.7 65.9$\pm$0.9 66.9$\pm$1.3 65.4$\pm$1.2 65.6$\pm$0.9 65.5$\pm$1.2 66.1$\pm$2.0 66.0$\pm$1.8 66.1
-PA-BOTH 63.1$\pm$0.7 67.2$\pm$1.1 64.3$\pm$0.5 72.1$\pm$1.8 66.3$\pm$0.7 64.1$\pm$1.2 69.7$\pm$2.1 67.5$\pm$1.8 61.2$\pm$1.4 67.7$\pm$2.3 61.2$\pm$1.6 65.5$\pm$0.6 65.9
-DeSGDA 76.7$\pm$0.8 84.6$\pm$0.9 69.4$\pm$0.6 85.2$\pm$1.5 76.2$\pm$1.1 83.9$\pm$1.2 69.9$\pm$0.6 76.3$\pm$1.4 75.9$\pm$1.0 73.5$\pm$1.3 76.3$\pm$1.6 68.3$\pm$0.7 76.4
+where $\lambda$ is a hyper-parameter to balance the adversarial training loss and classification loss. The learning procedure is illustrated in Algorithm D, and the complexity is shown in Appendix E .
 5 Experiment
@@
 To better address the variation in domain distributions within each dataset, we divided them into source and target domains based on the edge density, node density, and graph flux (i.e., the ratio of the number of nodes to the number of edges).
-The specific statistics, distribution visualization, and details introduction of experimental datasets are presented in Appendix E.
+The specific statistics, distribution visualization, and details introduction of experimental datasets are presented in Appendix F.
 Baselines.
-We compare DeSGDA with competitive baselines on the aforementioned datasets, including one graph kernel method: WL subtree Shervashidze et al.
+We compare the proposed DeSGDA with competitive baselines on the aforementioned datasets, including one graph kernel method: WL subtree Shervashidze et al.
 (2011); four general graph neural networks: GCN Kipf and Welling (2017), GIN Xu et al.
@@
 (2022) and DRSGNN Zhao et al.
-(2024); three recent domain adaptation methods: CDAN Long et al.
+(2024); three domain adaptation methods: CDAN Long et al.
 (2018), ToAlign Wei et al.
@@
 (2021a); and six graph domain adaptation methods: DEAL Yin et al.
-(2022), CoCo Yin et al.
-(2023), SGDA Qiao et al.
+(2022b), CoCo Yin et al.
+(2023a), SGDA Qiao et al.
 (2023), DGDA Cai et al.
 (2024), A2GNN Liu et al.
-(2024b) and PA-BOTH Liu et al.
-(2024c).
-More details about the compared baselines can be found in Appendix F.
+(2024a) and PA-BOTH Liu et al.
+(2024b).
+More detailed settings about the compared baselines are introduced in Appendix G.
+[…]
+[…]
+WL subtree 73.4 72.7 70.5 73.0 72.8 59.0 66.5 71.6 60.6 58.3 76.3 64.0 68.2
+GCN 57.2$\pm$2.7 62.8$\pm$1.7 67.6$\pm$0.5 58.5$\pm$1.3 67.7$\pm$0.4 61.0$\pm$0.3 65.0$\pm$0.8 51.1$\pm$1.3 65.6$\pm$2.2 55.4$\pm$0.4 68.5$\pm$3.1 67.7$\pm$0.5 62.3
+GIN 69.3$\pm$2.3 65.8$\pm$0.8 69.3$\pm$1.7 69.8$\pm$1.6 71.4$\pm$2.1 52.4$\pm$1.8 64.0$\pm$2.4 65.7$\pm$3.2 53.4$\pm$3.7 58.1$\pm$0.8 72.6$\pm$0.3 64.6$\pm$2.3 64.7
+GMT 67.8$\pm$1.3 69.6$\pm$0.7 74.5$\pm$0.5 67.6$\pm$2.5 69.9$\pm$2.1 55.8$\pm$0.7 74.8$\pm$1.4 60.1$\pm$2.4 71.4$\pm$3.3 51.5$\pm$0.5 69.0$\pm$0.5 63.3$\pm$1.3 66.3
+CIN 62.6$\pm$0.5 59.4$\pm$0.5 64.0$\pm$0.9 58.5$\pm$1.8 71.9$\pm$1.7 60.6$\pm$2.1 63.7$\pm$0.5 61.2$\pm$2.1 73.2$\pm$0.5 57.7$\pm$3.0 68.1$\pm$0.4 58.5$\pm$2.7 63.3
+SpikeGCN 66.1$\pm$0.4 67.8$\pm$0.7 68.5$\pm$0.9 67.2$\pm$0.3 70.1$\pm$1.1 68.6$\pm$1.3 66.1$\pm$0.7 65.8$\pm$1.2 74.5$\pm$0.9 65.4$\pm$0.6 73.8$\pm$1.1 70.1$\pm$1.6 68.6
+DRSGNN 68.3$\pm$0.5 71.9$\pm$0.9 71.5$\pm$1.1 75.1$\pm$1.7 76.7$\pm$1.3 74.4$\pm$0.8 70.5$\pm$0.7 67.8$\pm$0.8 77.0$\pm$1.5 65.9$\pm$1.6 75.2$\pm$1.9 69.2$\pm$2.1 71.8
+CDAN 75.6$\pm$0.5 70.5$\pm$0.6 71.6$\pm$0.5 69.8$\pm$0.5 76.6$\pm$0.8 71.4$\pm$0.3 71.4$\pm$0.3 72.1$\pm$0.3 75.5$\pm$0.7 74.3$\pm$0.8 78.2$\pm$1.1 74.0$\pm$0.8 73.4
+ToAlign 51.1$\pm$0.6 55.8$\pm$0.1 63.3$\pm$0.2 55.8$\pm$0.4 68.1$\pm$0.7 55.8$\pm$0.3 63.3$\pm$0.5 51.1$\pm$0.2 68.1$\pm$1.0 51.1$\pm$0.4 68.1$\pm$0.6 63.3$\pm$0.2 59.6
+MetaAlign 59.4$\pm$1.1 62.2$\pm$1.0 68.9$\pm$0.3 65.3$\pm$0.8 75.1$\pm$0.7 67.5$\pm$2.1 70.9$\pm$1.4 60.6$\pm$2.3 72.4$\pm$1.4 59.4$\pm$0.6 74.6$\pm$0.7 67.8$\pm$1.3 67.0
+DEAL 76.6$\pm$0.4 62.8$\pm$0.8 72.8$\pm$1.3 67.3$\pm$2.2 77.2$\pm$2.3 67.6$\pm$1.9 71.2$\pm$1.6 56.0$\pm$2.5 73.9$\pm$2.1 66.0$\pm$0.3 76.4$\pm$1.1 65.5$\pm$2.1 69.4
+CoCo 73.4$\pm$0.5 73.6$\pm$0.8 73.4$\pm$1.0 71.6$\pm$0.5 75.2$\pm$1.6 74.6$\pm$0.3 70.7$\pm$0.8 68.4$\pm$1.5 75.0$\pm$0.2 72.7$\pm$0.4 76.3$\pm$1.1 75.0$\pm$1.8 73.3
+SGDA 63.8$\pm$0.8 65.2$\pm$0.5 66.7$\pm$0.3 59.1$\pm$1.2 62.3$\pm$0.7 60.6$\pm$0.4 65.2$\pm$0.9 61.8$\pm$1.0 64.5$\pm$1.3 60.9$\pm$0.8 59.4$\pm$1.2 64.9$\pm$1.1 62.9
+DGDA 59.4$\pm$0.7 62.3$\pm$1.1 63.1$\pm$0.5 61.2$\pm$0.9 60.4$\pm$0.6 58.8$\pm$1.0 60.3$\pm$0.8 63.5$\pm$1.2 61.9$\pm$0.8 60.4$\pm$1.6 64.2$\pm$1.3 62.6$\pm$1.4 61.5
+A2GNN 65.4$\pm$0.7 66.4$\pm$1.1 65.7$\pm$1.3 66.0$\pm$0.6 64.9$\pm$1.2 65.8$\pm$1.6 65.5$\pm$1.8 66.0$\pm$1.4 65.8$\pm$2.1 65.6$\pm$1.9 66.1$\pm$1.7 66.0$\pm$2.0 65.8
+PA-BOTH 66.9$\pm$0.5 67.1$\pm$0.8 67.3$\pm$1.1 65.8$\pm$0.7 69.1$\pm$1.0 66.1$\pm$1.4 66.7$\pm$1.3 67.4$\pm$1.4 66.3$\pm$1.8 66.0$\pm$1.2 66.8$\pm$0.8 66.3$\pm$1.5 66.8
+DeSGDA 78.7$\pm$1.3 78.4$\pm$1.1 74.8$\pm$0.6 77.6$\pm$0.9 79.5$\pm$1.2 76.7$\pm$0.8 74.9$\pm$0.7 71.2$\pm$1.7 79.5$\pm$1.4 72.8$\pm$0.8 81.0$\pm$1.5 75.1$\pm$1.0 76.7
 Implementation Details.
@@
 Additionally, DeSGDA and all baseline models are trained using all labeled source samples and evaluated on unlabeled target samples Wu et al.
-(2020); Pang et al.
-(2023); Liu et al.
-(2023a); You et al.
 (2020).
-The performances of all models are measured and averaged on all samples for five runs.
+The performances of all models are measured and averaged on all samples for five different runs.
+Table 2: The results of ablation studies on PROTEINS (source → target). Bold results indicate the best performance.
 […]
-[…]
-WL subtree 75.9 70.4 64.3 63.9 60.6 64.7 73.2 78.9 66.8 69.2 74.2 72.9 69.6
-GCN 49.2$\pm$1.7 55.8$\pm$1.5 46.8$\pm$0.5 54.6$\pm$2.2 43.4$\pm$0.6 46.7$\pm$0.2 50.0$\pm$1.8 57.2$\pm$2.2 44.2$\pm$0.4 51.6$\pm$0.8 62.7$\pm$2.1 56.8$\pm$1.3 51.6
-GIN 68.8$\pm$2.5 70.6$\pm$1.0 64.2$\pm$1.1 67.2$\pm$2.4 62.2$\pm$1.8 62.5$\pm$1.5 68.7$\pm$2.4 72.5$\pm$0.6 63.3$\pm$1.6 65.2$\pm$0.6 62.4$\pm$0.3 70.9$\pm$0.5 66.6
-GMT 66.7$\pm$0.3 58.2$\pm$0.5 63.9$\pm$0.3 58.4$\pm$0.3 63.8$\pm$0.4 56.7$\pm$0.5 63.9$\pm$0.7 66.3$\pm$1.0 63.8$\pm$1.1 66.6$\pm$0.4 63.8$\pm$0.2 62.6$\pm$0.7 62.9
-CIN 58.7$\pm$2.4 54.9$\pm$0.2 52.0$\pm$0.3 54.8$\pm$0.1 56.6$\pm$0.2 54.9$\pm$0.1 52.9$\pm$1.4 52.8$\pm$0.5 56.5$\pm$0.6 52.8$\pm$2.1 58.5$\pm$0.8 56.6$\pm$1.4 55.1
-SpikeGCN 58.9$\pm$0.9 65.2$\pm$1.2 60.8$\pm$1.3 62.0$\pm$1.5 62.3$\pm$0.8 58.7$\pm$1.6 64.1$\pm$0.6 66.7$\pm$1.3 60.5$\pm$1.7 63.8$\pm$1.4 62.2$\pm$2.1 61.1$\pm$1.5 62.1
-DRSGNN 58.0$\pm$0.6 64.3$\pm$1.1 61.2$\pm$0.8 62.2$\pm$1.0 62.9$\pm$1.5 64.0$\pm$1.3 60.6$\pm$1.6 64.0$\pm$1.4 67.6$\pm$2.1 62.4$\pm$1.9 71.3$\pm$2.3 68.8$\pm$2.0 63.9
-CDAN 64.0$\pm$1.1 68.1$\pm$0.3 60.1$\pm$0.5 64.0$\pm$1.3 60.9$\pm$0.2 57.8$\pm$1.0 64.3$\pm$1.6 61.2$\pm$0.2 66.3$\pm$0.7 59.0$\pm$0.5 68.9$\pm$0.3 63.7$\pm$0.6 63.2
-ToAlign 52.8$\pm$0.5 54.8$\pm$0.2 48.2$\pm$1.1 54.8$\pm$1.5 44.0$\pm$0.8 54.8$\pm$2.0 48.2$\pm$1.7 52.8$\pm$0.6 44.0$\pm$0.2 52.8$\pm$0.3 44.0$\pm$1.0 48.2$\pm$1.2 50.0
-MetaAlign 63.1$\pm$0.3 63.8$\pm$1.3 58.9$\pm$2.4 58.5$\pm$0.4 59.1$\pm$2.1 59.2$\pm$1.6 70.1$\pm$0.8 63.3$\pm$1.4 66.5$\pm$2.7 60.9$\pm$1.1 71.4$\pm$0.2 67.5$\pm$0.8 63.5
-DEAL 70.7$\pm$0.9 72.3$\pm$0.2 69.9$\pm$0.8 68.9$\pm$0.7 64.1$\pm$0.6 65.6$\pm$0.9 71.9$\pm$0.4 69.9$\pm$1.7 70.6$\pm$0.4 66.5$\pm$0.3 71.6$\pm$0.7 69.9$\pm$0.5 69.3
-CoCo 64.0$\pm$1.3 63.9$\pm$0.6 65.8$\pm$1.8 59.9$\pm$1.7 62.2$\pm$2.1 60.6$\pm$1.6 65.0$\pm$2.1 64.8$\pm$1.4 60.0$\pm$0.8 61.3$\pm$0.5 68.5$\pm$0.4 67.1$\pm$0.6 63.6
-SGDA OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
-DGDA OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
-A2GNN 58.9$\pm$0.9 60.1$\pm$0.7 59.8$\pm$1.2 59.4$\pm$1.0 62.3$\pm$1.5 60.9$\pm$1.6 61.6$\pm$1.3 59.9$\pm$1.9 64.9$\pm$1.6 62.9$\pm$2.1 65.4$\pm$1.5 63.3$\pm$2.3 61.7
-PA-BOTH 61.1$\pm$0.5 60.9$\pm$0.4 61.6$\pm$0.6 61.2$\pm$0.8 60.8$\pm$0.6 61.5$\pm$0.5 62.2$\pm$1.0 61.9$\pm$0.7 61.8$\pm$1.1 61.1$\pm$0.9 60.9$\pm$0.8 61.3$\pm$1.2 61.6
-DeSGDA 68.5$\pm$1.2 71.4$\pm$1.3 70.1$\pm$0.7 69.0$\pm$1.1 68.9$\pm$1.0 66.3$\pm$1.4 69.6$\pm$1.3 70.2$\pm$1.7 71.1$\pm$1.6 69.3$\pm$1.8 74.4$\pm$1.6 70.0$\pm$1.9 70.1
-[…]
-[…]
-WL subtree 65.7 71.8 57.9 71.1 47.4 43.4 65.5 75.1 45.3 34.9 52.7 49.8 56.7
-GCN 70.6$\pm$2.1 60.3$\pm$1.5 60.5$\pm$3.4 62.3$\pm$1.1 58.4$\pm$0.5 43.2$\pm$0.2 63.8$\pm$1.0 70.3$\pm$0.3 50.6$\pm$1.0 32.8$\pm$0.3 50.1$\pm$0.4 42.2$\pm$0.2 55.4
-GIN 66.7$\pm$2.1 73.7$\pm$2.4 57.3$\pm$3.1 69.4$\pm$2.3 58.6$\pm$0.4 43.1$\pm$0.3 66.4$\pm$2.7 74.8$\pm$1.8 42.2$\pm$1.6 33.5$\pm$1.0 57.4$\pm$0.8 43.9$\pm$2.3 57.2
-GMT 67.3$\pm$0.3 56.8$\pm$0.4 58.0$\pm$0.2 56.8$\pm$0.2 60.6$\pm$0.3 56.8$\pm$0.5 57.8$\pm$0.1 67.3$\pm$0.1 39.5$\pm$0.3 67.3$\pm$0.2 39.5$\pm$0.5 57.8$\pm$0.4 57.1
-CIN 67.6$\pm$0.4 63.7$\pm$2.1 58.9$\pm$1.0 56.8$\pm$0.4 63.6$\pm$0.4 59.5$\pm$2.7 58.7$\pm$1.2 67.0$\pm$0.5 61.7$\pm$1.6 67.8$\pm$0.7 62.2$\pm$2.1 56.0$\pm$1.3 61.9
-SpikeGCN 67.2$\pm$0.5 57.2$\pm$1.2 57.9$\pm$0.8 57.1$\pm$0.6 61.1$\pm$1.3 58.9$\pm$1.6 60.0$\pm$1.2 67.2$\pm$0.9 53.9$\pm$2.1 64.4$\pm$0.8 57.8$\pm$1.0 59.9$\pm$1.2 60.2
-DRSGNN 67.4$\pm$0.4 58.4$\pm$1.0 59.0$\pm$1.2 57.4$\pm$0.5 62.3$\pm$1.1 60.4$\pm$1.3 61.1$\pm$1.6 67.9$\pm$1.5 56.2$\pm$1.8 66.2$\pm$2.1 60.9$\pm$1.4 58.6$\pm$2.5 61.3
-CDAN 72.9$\pm$0.4 72.7$\pm$0.4 65.4$\pm$0.3 72.9$\pm$0.1 61.2$\pm$0.3 70.3$\pm$0.2 65.7$\pm$0.4 72.7$\pm$0.1 61.0$\pm$0.1 72.1$\pm$1.2 60.7$\pm$0.2 65.3$\pm$0.6 67.7
-ToAlign 32.7$\pm$2.0 43.2$\pm$0.1 42.2$\pm$1.3 43.2$\pm$0.9 60.5$\pm$0.7 43.2$\pm$1.2 42.2$\pm$0.4 32.7$\pm$1.2 60.5$\pm$0.9 32.7$\pm$0.3 60.5$\pm$0.7 42.2$\pm$0.4 44.7
-MetaAlign 67.3$\pm$0.7 56.8$\pm$0.2 57.8$\pm$0.6 56.8$\pm$0.4 60.5$\pm$1.3 56.8$\pm$0.8 57.8$\pm$1.1 67.3$\pm$1.2 60.5$\pm$0.4 67.3$\pm$0.6 60.5$\pm$0.7 57.8$\pm$0.6 60.6
-DEAL 75.0$\pm$0.9 76.3$\pm$2.4 65.9$\pm$1.8 77.5$\pm$2.7 60.3$\pm$4.5 69.7$\pm$3.2 67.2$\pm$1.5 75.3$\pm$1.7 57.4$\pm$4.1 71.1$\pm$2.2 65.7$\pm$2.7 66.4$\pm$1.6 69.0
-CoCo 74.2$\pm$1.7 74.3$\pm$0.6 65.9$\pm$1.2 72.7$\pm$2.1 61.1$\pm$0.2 71.0$\pm$1.7 68.6$\pm$0.3 75.9$\pm$0.2 60.7$\pm$0.2 73.9$\pm$0.4 59.7$\pm$1.1 67.3$\pm$0.8 68.8
-SGDA 55.9$\pm$0.6 57.1$\pm$0.5 56.1$\pm$0.4 54.6$\pm$0.8 55.8$\pm$1.1 57.7$\pm$0.6 54.3$\pm$0.7 53.6$\pm$1.3 59.1$\pm$0.8 56.7$\pm$0.6 55.4$\pm$1.2 53.8$\pm$0.5 55.9
-DGDA OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
-A2GNN 55.9$\pm$0.7 55.7$\pm$0.4 56.6$\pm$0.6 57.1$\pm$1.0 56.1$\pm$1.2 55.8$\pm$0.5 56.5$\pm$0.7 55.5$\pm$0.4 55.9$\pm$0.8 56.2$\pm$0.6 56.5$\pm$1.5 56.0$\pm$0.5 56.2
-PA-BOTH 56.4$\pm$0.5 55.9$\pm$0.6 56.0$\pm$0.5 56.4$\pm$0.4 56.3$\pm$0.6 57.7$\pm$0.7 56.6$\pm$0.2 58.8$\pm$0.9 56.9$\pm$0.7 57.2$\pm$0.3 56.5$\pm$0.5 58.3$\pm$0.8 56.9
-DeSGDA 75.2$\pm$0.8 74.3$\pm$1.2 68.0$\pm$1.5 73.6$\pm$2.0 62.3$\pm$1.7 71.2$\pm$2.5 68.8$\pm$1.3 76.0$\pm$1.2 61.9$\pm$2.2 71.4$\pm$1.6 62.8$\pm$1.2 65.7$\pm$0.9 69.3
+DeSGDA w/o CA 72.3 72.6 71.9 75.1 74.1 71.3 71.5 70.4 71.6 70.5 78.3 71.5 72.6
+DeSGDA w/o PL 72.1 68.7 67.8 69.4 63.7 55.7 68.3 69.5 70.2 69.6 76.7 66.9 68.3
+DeSGDA w/o CF 56.7 54.2 60.7 62.1 73.4 68.3 55.3 69.5 76.7 65.5 73.7 63.3 65.0
+DeSGDA w/o TL 61.7 43.8 56.1 71.2 49.5 69.0 42.4 68.3 73.7 64.0 70.5 43.5 59.5
+DeSGDA w/ PT 71.5 68.3 66.1 72.4 66.7 70.1 67.9 70.8 66.3 71.8 74.1 72.9 70.0
+DeSGDA w/ CL 74.9 73.3 73.7 75.1 77.0 71.3 73.9 70.0 78.1 77.6 78.8 74.9 74.9
+DeSGDA 78.7 78.4 74.8 77.6 79.5 76.7 74.9 71.2 79.5 72.8 81.0 75.1 76.7
 5.2 Performance Comparision
-We present the results of the proposed DeSGDA with all baseline models under the setting of graph domain adaptation on different datasets in Table 1, 2, 3.
-From these tables, we observe that: (1) The performance of graph domain adaptation methods surpasses that of graph and spike-based graph methods.
+We present the results of our proposed DeSGDA with all baseline models under the setting of graph domain adaptation on different datasets in Tables 1 and 13-23.
+From these tables, we observe that: (1) The performance of graph domain adaptation methods surpasses graph-based and spike-based graph methods.
 We attribute this improvement to the fact that domain distribution shifts degrade the performance of traditional graph methods.
 (2) The graph domain adaptation methods (DEAL and CoCo) outperform the spike-based graph methods (SpikeGCN and DRSGNN), underscoring the necessity of the research in spiking graph domain adaptation.
+(3) The WL subtree method outperforms SGDA, DGDA, A2GNN, and PA-BOTH but falls short compared to DEAL and CoCO.
+We attribute this to the relatively limited research specifically addressing the graph domain adaptation problem (e.g., DEAL and CoCo).
+To bridge this gap, we adapted node classification methods for graph classification tasks (e.g., SGDA, DGDA, A2GNN, and PA-BOTH).
+While the WL subtree method demonstrates superior performance over these adapted node classification-based methods, it remains less effective than dedicated graph domain adaptation methods tailored for graph classification tasks.
 (4) Our DeSGDA outperforms all baselines for most cases, which demonstrates its superiority over other methods.
@@
 Moreover, the pseudo-label distillation aids in updating the source degree thresholds in the target domain, thereby ensuring optimal performance.
-More results evaluated on other datasets can be found in Appendix G.1.
+More results on other datasets can be found in Appendix H.1.
 Figure 2: The performance with different GNN architectures on PROTEINS.
-Additionally, we conduct experiments to explore the flexibility of the proposed DeSGDA.
-Specifically, we replace the backbone of the degree-aware personalized spiking graph encoder (GIN) with various GNNs methods (i.e., GCN and GAT), and the results are shown in Figure 2.
-From the results, we observe that GIN consistently outperforms other GNNs architectures in most cases, demonstrating its powerful representation capability.
+Additionally, we conduct experiments to explore the flexibility of DeSGDA.
+Specifically, we replace the backbone of the degree-aware personalized spiking graph encoder (GIN) with various GNN methods (i.e., GCN and GAT), and the results are shown in Figure 2.
+From those results, we observe that GIN consistently outperforms other GNN architectures in most cases, demonstrating its powerful representation capability.
 This phenomenon also justifies our choice of using GIN to enhance the performance of the proposed DeSGDA.
-More results are reported in Appendix G.1.
+More results are reported in Appendix H.1.
 5.3 Energy Efficiency Analysis
-Figure 3: Energy consumption of DeSGDA on different datasets.
+Figure 3: Energy consumption of DeSGDA and baselines on different datasets.
 To assess the energy efficiency of DeSGDA, we use the metric from Zhu et al.
@@
 (2022).
-The results are shown in Figure 3, from the results, we find that compared with traditional graph domain adaptation methods, the spike-based methods (DeSGDA and DRSGNN) have significantly lower energy consumption, demonstrating the superior energy efficiency of SGNs.
+The results are shown in Figure 3, where we find that compared with traditional graph domain adaptation methods, the spike-based methods (DeSGDA and DRSGNN) have significantly lower energy consumption, demonstrating the superior energy efficiency of SGNs.
 Moreover, although the energy consumption of DeSGDA is slightly higher than DRSGNN due to additional computations required for domain adaptation, the performance improvement justifies the deployment of DeSGDA in low-power devices.
-Table 4: The results of ablation studies on PROTEINS (source → target). Bold results indicate the best performance. Bold results indicate the best performance.
-[…]
-DeSGDA w/o CA 72.3 72.6 71.9 75.1 74.1 71.3 71.5 70.4 71.6 70.5 78.3 71.5 72.6
-DeSGDA w/o PL 72.1 68.7 67.8 69.4 63.7 55.7 68.3 69.5 70.2 69.6 76.7 66.9 68.3
-DeSGDA w/o CF 56.7 54.2 60.7 62.1 73.4 68.3 55.3 69.5 76.7 65.5 73.7 63.3 65.7
-DeSGDA w/o TL 61.7 43.8 56.1 71.2 49.5 69.0 42.4 68.3 73.7 64.0 70.5 43.5 59.5
-DeSGDA w/ PT 71.5 68.3 66.1 72.4 66.7 70.1 67.9 70.8 66.3 71.8 74.1 72.9 70.0
-DeSGDA w/ CL 74.9 73.3 73.7 75.1 77.0 71.3 73.9 70.0 78.1 77.6 78.8 74.9 75.1
-DeSGDA 78.7 78.4 74.8 77.6 79.5 76.7 74.9 71.2 79.5 72.8 81.0 75.1 76.4
+Additionally, we present a comparison of training time and memory usage between DeSGDA and other graph domain adaptation methods.
+The results are detailed in Tables 9 and  10.
 5.4 Ablation Study
 We conduct ablation studies to examine the contributions of each component in the proposed DeSGDA: (1) DeSGDA w/o CA: It removes the adversarial distribution alignment module; (2) DeSGDA w/o PL: It removes the pseudo-label distilling module; (3) DeSGDA w/o CF: It removes the classification loss $\mathcal{L}_{S}$; (4) DeSGDA w/o TL: It utilizes the global thresholds on all nodes; (5) DeSGDA w/ PT: It deploys the adaptive perturbations Yin et al.
-(2022) on source data for alignment; (6) DeSGDA w/ CL: It replaces the adversarial learning with the cross-domain contrastive learning Yin et al.
-(2023).
+(2022b) on source data for alignment; (6) DeSGDA w/ CL: It replaces the adversarial learning with the cross-domain contrastive learning Yin et al.
+(2023a).
 Experimental results are shown in Table 2.
-From the table, we find that: (1) DeSGDA outperforms DeSGDA w/o CA, DeSGDA w/o PL, and DeSGDA w/o CF, demonstrating that the adversarial distribution alignment module can effectively reduce domain discrepancies, ensuring well-aligned feature spaces between source and target domains.
-Additionally, the pseudo-label distillation module can address the variance in thresholds across domains, while the classification loss $\mathcal{L}_{S}$ enables DeSGDA to effectively learn from labeled source data and generalize to the target domain.
-(2) DeSGDA w/o TL shows lower performance compared to DeSGDA, showing that the degree-aware thresholds, which are iteratively updated during model training, can resolve the issue of the inflexible architecture in SGNs.
-By using these thresholds, DeSGDA can effectively learn meaningful representations for nodes with various degrees.
+From the table, we find that: (1) DeSGDA outperforms DeSGDA w/o CA, DeSGDA w/o PL, and DeSGDA w/o CF, demonstrating that the adversarial distribution alignment module effectively reduces domain discrepancies, ensuring well-aligned feature spaces between source and target domains.
+Additionally, the pseudo-label distillation module addresses variance in thresholds across domains, while the classification loss $\mathcal{L}_{S}$ enables DeSGDA to effectively learn from labeled source data and generalize to the target domain.
+(2) DeSGDA w/o TL shows lower performance compared to DeSGDA, showing that degree-aware thresholds, iteratively updated during training, resolve the issue of inflexible architecture in SGNs.
+By using these thresholds, DeSGDA effectively learns meaningful representations for nodes with various degrees.
 (3) DeSGDA outperforms DeSGDA w/ PT and DeSGDA w/ CL.
-We attribute that the adaptive perturbations method (DeSGDA w/ PT) can not effectively help DeSGDA overcome the domain discrepancy.
+We attribute this to the adaptive perturbations method (DeSGDA w/ PT) fail to help DeSGDA overcome domain discrepancies effectively.
 Additionally, the cross-domain contrastive learning method (DeSGDA w/ CL) is less effective at aligning the source and target distribution compared to adversarial learning.
-More ablation results on other datasets are reported in Appendix G.2.
+Furthermore, we provide ablation studies to examine the effect of directly replacing SGNs with commonly used Graph Neural Networks (GNNs) for generating representations for DeSGDA, and the results are shown in Tables 11, 12.
+More details about ablation results on other datasets are reported in Appendix H.3.
 (a) Time Latency $\tau$
@@
 5.5 Sensitivity Analysis
-We study the sensitivity analysis of DeSGDA with respect to the impact of its hyperparameters: time latency $\tau$, degree threshold value $V_{th}^{degree}$ in SNNs, and balance ratio $\alpha$, which plays a crucial role in the performance of DeSGDA.
+We conduct the sensitivity analysis of DeSGDA with respect to the impact of its hyperparameters: time latency $\tau$, degree threshold value $V_{th}^{degree}$ in SNNs, and balance ratio $\alpha$, which plays a crucial role in the performance of DeSGDA.
 In particular, $\tau$ controls the number of SNNs propagation steps; $V_{th}^{degree}$ determines when a neuron fires; $\alpha$ governs the changing ratio of degree-aware thresholds.
 Figure 4 illustrates how $\tau$, $V_{th}^{degree}$, and $\alpha$ affects the performance of DeSGDA on the PROTEINS dataset.
-More results on other datasets are shown in Appendix G.3.
+More results on other datasets are shown in Appendix H.4.
 We vary $\tau$ within the range of $\{5,6,7,8,9,10\}$, $V_{th}^{degree}$ in $\{0.05,0.1,0.2,0.5,1.0,2.0,5.0\}$, and $\alpha$ in $\{0.1,0.3,0.5,0.7,0.9\}$.
@@
 (2) Figure 4(b) indicates an initial increase followed by a decreasing trend in performance as $V_{th}^{degree}$ increases.
-This trend occurs because a lower threshold may trigger more spikes for high-degree nodes, leading to a drastic change in the threshold, which can degrade performance.
-Conversely, a higher threshold for low-degree nodes could result in fewer spikes, affecting the model’s ability to process information effectively.
-Therefore, we set $V_{th}^{degree}$ to 0.5 as default.
+This trend occurs because a lower threshold may trigger more spikes for high-degree nodes, leading to a drastic change in threshold, which degrades performance.
+Conversely, a higher threshold for low-degree nodes could result in fewer spikes, affecting model’s ability to process information.
+Thus, we set $V_{th}^{degree}$ to 0.5 as default.
 (3) From Figure 4(c), we find that the performance of DeSGDA initially increases and then decreases as $\alpha$ increases.
-The potential reason is that the smaller $\alpha$ may delay the updating of the threshold, leading to performance degradation.
+The potential reason is that smaller $\alpha$ may delay the updating of the threshold, leading to performance degradation.
 Contrarily, a larger $\alpha$ tends to introduce more spikes that change dynamically at each step, resulting in instability in the model’s performance.
@@
 6 Conclusion
-In this paper, we first propose the problem of spiking graph domain adaptation and introduce a novel framework DeSGDA for graph classification.
+In this paper, we propose the problem of spiking graph domain adaptation and introduce a novel framework, DeSGDA, for graph classification.
 This framework enhances the adaptability and performance of SGNs through three key aspects: node degree-aware personalized spiking representation, adversarial feature distribution alignment, and pseudo-label distillation.
-Our approach enables more expressive information capture through degree-dependent spiking thresholds, aligns feature distributions via adversarial training, and utilizes pseudo-labels to leverage unlabeled data effectively.
-The extensive experimental validation across benchmark datasets has demonstrated that DeSGDA not only surpasses existing methods in accuracy but also maintains efficient energy consumption, making it a promising solution for advancing the domain adaptation capabilities of spiking graph networks.
-In the future, we will apply SGNs in the scenarios of source-free domain adaptation and domain generalization.
-Appendix A Proof of Theorem 2
+Our approach enables expressive information capture through degree-dependent spiking thresholds, aligns feature distributions via adversarial training, and utilizes pseudo-labels to leverage unlabeled data effectively.
+Extensive experimental validation across benchmark datasets demonstrates that DeSGDA surpasses existing methods in accuracy while maintaining efficient energy consumption, making it a promising solution for advancing the domain adaptation capabilities of spiking graph networks.
+In the future, we will apply SGNs in source-free domain adaptation and domain generalization scenarios.
+Appendix A Proof of Hypothesis 1
+Assuming that the node feature $h_{i}$ follows a normal distribution $\mathcal{N}(\mu,\sigma^{2})$, then for each node in the graph, we follow the message-passing mechanism and have the information aggregation as:
+$h_{i}=h_{i}+\sum_{j\in N(i)}w_{ij}h_{j}.$ (9)
+Therefore, we have the expectation:
+$\mathbb{E}(h_{i})=\mathbb{E}(h_{i})+\sum_{j\in N(i)}w_{ij}\mathbb{E}(h_{j}).$ (10)
+Since $\mathbb{E}(h_{j})\sim\mathcal{N}(\mu,\sigma^{2})$, we have:
+$\mathbb{E}(h_{i})\sim\mathcal{N}\left((1+\sum_{j\in N(i)}w_{ij})\mu,(1+\sum_{j\in N(i)}w_{ij})\sigma^{2}\right).$ (11)
+From the results, we observe that node $i$ follows a normal distribution with a mean of $(1+\sum_{j\in N(i)}w_{ij})\mu$, determined by the aggregated weights of its neighboring nodes.
+To provide a more intuitive understanding, we visualize the aggregated neighbor weights of GCN Kipf and Welling [2017] and GIN Xu et al.
+[2018] in Figure 5.
+The results show that as the degree increases, the aggregated weights also increase progressively.
+Consequently, high-degree nodes tend to follow a normal distribution with a higher mean and variance.
+In other words, nodes with higher degrees accumulate greater signals, making them more likely to trigger spiking.
+Based on this, we propose assigning higher thresholds to high-degree nodes and lower thresholds to low-degree nodes.
+Another observation is that methods that normalize neighbor weights to 1 (e.g., GAT Veličković et al.
+[2017], GraphSAGE Hamilton et al.
+[2017]) still result in aggregated features following the same normal distribution.
+This normalization diminishes the ability to distinguish between nodes with varying degrees, ultimately degrading performance.
+This explains why, when using GAT as the backbone of DeSGDA, the performance is the weakest.
+Figure 5: Visualization of degree-aware thresholds and aggregation weights on PROTEINS and NCI1.
+Appendix B Proof of Theorem 2
 Theorem 2 Assuming that the learned discriminator is $C_{g}$-Lipschitz continuous as described in Theorem 1, the graph feature extractor $f$ (also referred to as GNN) is $C_{f}$-Lipschitz that $||f||_{Lip}=\max_{G_{1},G_{2}}\frac{||f(G_{1})-f(G_{2})||_{2}}{\eta(G_{1},G_{2})}=C_{f}$ for some graph distance measure $\eta$ and the loss function bounded by $C>0$.
@@
 $\mathbf{P}_{kt}=\left\{\begin{array}[]{lr}\exp\left(\frac{u_{k}(t)-V_{th}}{\sigma(u_{k}(t)-u_{reset})}\right),&if\quad u_{\theta}\leq u(t)\leq V_{th},\\ 0,&if\quad u_{reset}\leq u_{k}(t)\leq u_{\theta}.\end{array}\right.$ (13)
-Proof. Before showing the designated lemma, we first introduce the following inequality to be used that:
+Proof.
+Before showing the designated lemma, we first introduce the following inequality to be used that:
 $|\epsilon_{S}(h,\hat{h}_{S})-\epsilon_{T}(h,\hat{h}_{T})|$ $=|\epsilon_{S}(h,\hat{h}_{S})-\epsilon_{S}(h,\hat{h}_{T})+\epsilon_{S}(h,\hat{h}_{T})-\epsilon_{T}(h,\hat{h}_{T})|$ (14)
@@
 Provided the spike training data set $S_{n}=\{(\mathbf{X}_{i},y_{i})\in\mathcal{X}\times\mathcal{Y}\}_{i\in[n]}$ drawn from $\mathcal{D}$, and motivated by Yin et al.
-[2024], we have:
+[2024a], we have:
 $\lim_{\tau\to\infty}P\left(\hat{h}(S_{n})_{\tau,i}>\hat{h}(\mathbf{X}_{\tau,i})+\epsilon\right)\leq e^{-\epsilon^{2}/2(\sigma+\hat{w}_{i}\epsilon/3)},$ (19)
@@
 $+2C_{f}C_{g}W_{1}\left(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G)\right).$
-Appendix B Proof of Theorem 3
+Appendix C Proof of Theorem 3
 Theorem 3 Under the assumption of Theorem 1, we further assume that there exists a small amount of i.i.d.
@@
 $\mathbf{P}_{kt}=\left\{\begin{array}[]{lr}\exp\left(\frac{u_{k}(t)-V_{th}}{\sigma(u_{k}(t)-u_{reset})}\right),&if\quad u_{\theta}\leq u(t)\leq V_{th},\\ 0,&if\quad u_{reset}\leq u_{k}(t)\leq u_{\theta}.\end{array}\right.$ (30)
-Proof. As proved in Theorem 2, we have:
+Proof.
+As proved in Theorem 2, we have:
 $\epsilon_{T}(h,\hat{h}_{T}(\mathbf{X}))\leq$ $\hat{\epsilon}_{S}(h,\hat{h}_{S}(S_{n}))+2\mathbb{E}\left[\sup\frac{1}{N_{S}}\sum_{i=1}^{N_{S}}\epsilon_{i}h(\mathbf{X}_{i},y_{i},p_{i})\right]+C\sqrt{\frac{ln(2/\delta)}{N_{S}}}$ (31)
@@
 $+2C_{f}C_{g}W_{1}\left(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G)\right)+\omega^{\prime}.$
-Appendix C Algorithm
+Appendix D Algorithm
 Algorithm 1 Learning Algorithm of DeSGDA
-Input: Source data $\mathcal{D}^{s}$; Target data $\mathcal{D}^{t}$.
-Output: Node degree-aware personalized spiking graph encoder parameters $\theta$.
-1:Initialize model parameters.
-2:while not convergence do
-3:     Sample mini-batches $\mathcal{B}^{s}$ and $\mathcal{B}^{t}$ from source and target data, respectively;
-4:     Forward propagation $\mathcal{B}^{s}$ and $\mathcal{B}^{t}$ through node degree-aware personalized spiking graph encoder;
-5:     Pseudo-label Distilling;
-6:     Calculate the loss function by Eq. 11;
-7:     Update model parameters through back propagation;
+Input: Source data $\mathcal{D}^{s}$, Target data $\mathcal{D}^{t}$
+Output: Node degree-aware personalized spiking graph encoder parameters $\theta$, domain discriminator $\gamma$
+1:Initialize model parameters $\theta$ and $\gamma$
+2:while not converged do
+3:     Sample mini-batches $\mathcal{B}^{s}$ and $\mathcal{B}^{t}$ from $\mathcal{D}^{s}$ and $\mathcal{D}^{t}$
+4:     Forward propagate $\mathcal{B}^{s}$ and $\mathcal{B}^{t}$ through the node degree-aware personalized spiking graph encoder
+5:     Perform pseudo-label distillation
+6:     Compute the loss function using Eq. 8
+7:     Update model parameters $\theta$ and $\gamma$ via backpropagation
 8:end while
-Appendix D Complexity Analysis
+Appendix E Complexity Analysis
 Here we analyze the computational complexity of the proposed DeSGDA.
 The computational complexity primarily relies on node degree-aware personalized spiking representations.
-For a given graph $G,\|A\|_{0}$ denotes the number of nonzeros in the adjacency matrix.
+For a given graph $G$, $\|A\|_{0}$ denotes the number of nonzeros in the adjacency matrix of $G$.
 $d$ is the feature dimension.
-$L$ denote the layer number of GIN.
+$L$ denote the layer number of spiking graph encoder.
 $|V|$ is the number of nodes.
@@
 Figure 9: Visualization of different distributions on MUTAGNENICITY.
-Appendix E Dataset
+Appendix F Dataset
 Table 3: Statistics of the experimental datasets.
@@
 FRANKENSTEIN 4,337 16.9 17.88 2
-E.1 Dataset Description
+F.1 Dataset Description
 We conduct extensive experiments on four public benchmark graph datasets from TUDataset. The dataset statistics can be found in Table 3, and their details are shown as follows:
@@
 DeSGDA w/o CA 58.9 62.6 61.7 60.1 58.2 62.0 63.4 61.7 63.5 60.7 68.3 64.0 62.1
-DeSGDA w/o PL 62.8 65.1 55.0 56.9 55.3 54.4 58.6 62.9 61.2 62.2 59.7 64.3 59.7
-DeSGDA w/o CF 52.7 62.6 58.5 56.5 56.3 63.8 63.1 52.7 65.4 65.1 58.2 63.7 60.0
+DeSGDA w/o PL 62.8 65.1 55.0 56.9 55.3 54.4 58.6 62.9 61.2 62.2 59.7 64.3 59.8
+DeSGDA w/o CF 52.7 62.6 58.5 56.5 56.3 63.8 63.1 52.7 65.4 65.1 58.2 63.7 59.9
 DeSGDA w/o TL 64.6 62.8 60.6 54.9 44.9 65.4 60.8 64.8 56.1 64.1 57.8 54.0 59.3
-DeSGDA w PT 64.4 59.2 64.6 62.2 64.5 60.1 61.1 60.8 62.0 64.2 62.7 65.5 62.7
+DeSGDA w PT 64.4 59.2 64.6 62.2 64.5 60.1 61.1 60.8 62.0 64.2 62.7 65.5 62.6
 DeSGDA w CL 65.3 70.1 68.9 65.2 65.6 63.0 67.6 64.6 68.9 66.8 68.6 70.7 67.1
-DeSGDA 68.5 72.4 70.1 69.0 68.9 66.3 69.6 70.2 71.1 69.3 74.4 70.0 70.1
-Table 7: The results of ablation studies on FRANKENSTEIN (source → target). Bold results indicate the best performance. Bold results indicate the best performance.
+DeSGDA 68.5 71.4 70.1 69.0 68.9 66.3 69.6 70.2 71.1 69.3 74.4 70.0 69.9
+Table 5: The results of ablation studies on FRANKENSTEIN (source → target). Bold results indicate the best performance.
 […]
 DeSGDA w/o CA 59.6 60.7 60.2 60.7 77.9 58.5 61.3 60.5 76.8 60.0 80.1 63.9 65.0
-DeSGDA w/o PL 58.9 56.6 55.9 56.6 69.8 53.6 55.0 57.7 70.4 53.2 68.8 61.7 60.0
-DeSGDA w/o CF 55.9 59.2 59.5 54.5 74.9 52.9 59.9 54.6 78.8 56.1 74.5 58.5 61.7
+DeSGDA w/o PL 58.9 56.6 55.9 56.6 69.8 53.6 55.0 57.7 70.4 53.2 68.8 61.7 59.9
+DeSGDA w/o CF 55.9 59.2 59.5 54.5 74.9 52.9 59.9 54.6 78.8 56.1 74.5 58.5 61.6
 DeSGDA w/o TL 56.2 57.2 60.7 48.2 53.3 45.8 59.5 53.0 48.8 54.8 51.1 54.7 53.6
@@
 DeSGDA w CL 62.2 61.3 61.7 60.7 79.3 59.7 63.6 64.1 79.5 61.1 78.6 62.7 66.2
-DeSGDA 63.5 64.1 62.6 63.0 81.1 62.4 65.5 65.4 81.9 60.8 82.0 65.9 68.1
-E.2 Data processing
+DeSGDA 63.5 64.1 62.6 63.0 81.1 62.4 65.5 65.4 81.9 60.8 82.0 65.9 68.2
+Table 6: The results of ablation studies on MUTAGENICITY (source → target). Bold results indicate the best performance.
+[…]
+DeSGDA w/o CA 58.3 63.3 62.1 60.1 81.4 62.7 63.8 59.1 80.3 57.1 78.9 63.0 65.8
+DeSGDA w/o PL 58.8 57.6 60.2 61.4 73.3 54.5 63.7 61.0 61.2 61.9 71.0 60.2 62.1
+DeSGDA w/o CF 54.7 58.3 59.6 58.7 72.1 60.7 57.9 54.9 78.0 57.4 63.2 60.4 61.3
+DeSGDA w/o TL 56.5 60.7 59.1 60.7 43.4 56.3 51.0 56.7 53.3 56.9 63.1 52.5 55.9
+DeSGDA w PT 62.1 64.3 63.9 61.8 77.3 62.4 66.6 63.9 73.5 66.1 82.1 64.9 67.4
+DeSGDA w CL 66.1 62.6 63.3 64.7 81.5 60.7 69.8 67.0 82.2 67.8 82.4 63.6 69.3
+DeSGDA 65.4 65.9 65.5 65.6 82.8 63.6 70.7 68.2 82.9 67.6 83.9 66.5 70.8
+F.2 Data processing
 In our implementation, we process these four TUDatasets by adding a self-loop connection for each node.
 Additionally, we utilize one-hot embeddings to represent node attributes for datasets (FRANKENSTEIN and MUTAGENICITY) where node features are unavailable.
-Appendix F Baselines
+Appendix G Baselines
 In this part, we introduce the details of the compared baselines as follows:
 Graph kernel method. We compare DeSGDA with one graph kernel method:
-  • WL subtree: The Weisfeiler-Lehman (WL) subtree Shervashidze et al.
-[2011] method is a graph kernel method, which calculates the graph similarity by a kernel function, where It encodes local neighborhood structures into subtree patterns, efficiently capturing the topology information contained in graphs.
+  • WL subtree: Weisfeiler-Lehman (WL) subtree Shervashidze et al.
+[2011] is a graph kernel method, which calculates the graph similarity by a kernel function, where It encodes local neighborhood structures into subtree patterns, efficiently capturing the topology information contained in graphs.
 Graph neural networks. We compare DeSGDA with four widely used general graph neural networks:
@@
   • SpikeGCN: SpikeGCN Zhu et al. [2022] introduces an end-to-end framework designed to integrate the fidelity characteristics of SNNs with graph node representations.
-  • DRSGNN: DRSGNN Zhao et al. [2024] dynamically adapts to evolving graph structures and relationships through a novel architecture that updates node representations in real-time..
-Table 8: The results of ablation studies on MUTAGENICITY (source → target). Bold results indicate the best performance. Bold results indicate the best performance.
-[…]
-DeSGDA w/o CA 58.3 63.3 62.1 60.1 81.4 62.7 63.8 59.1 80.3 57.1 78.9 63.0 65.8
-DeSGDA w/o PL 58.8 57.6 60.2 61.4 73.3 54.5 63.7 61.0 61.2 61.9 71.0 60.2 62.1
-DeSGDA w/o CF 54.7 58.3 59.6 58.7 72.1 60.7 57.9 54.9 78.0 57.4 63.2 60.4 61.3
-DeSGDA w/o TL 56.5 60.7 59.1 60.7 43.4 56.3 51.0 56.7 53.3 56.9 63.1 52.5 55.9
-DeSGDA w PT 62.1 64.3 63.9 61.8 77.3 62.4 66.6 63.9 73.5 66.1 82.1 64.9 67.4
-DeSGDA w CL 66.1 62.6 63.3 64.7 81.5 60.7 69.8 67.0 82.2 67.8 82.4 63.6 69.6
-DeSGDA 65.4 65.9 65.5 65.6 82.8 63.6 70.7 68.2 82.9 67.6 83.9 66.5 70.1
+  • DRSGNN: DRSGNN Zhao et al. [2024] dynamically adapts to evolving graph structures and relationships through a novel architecture that updates node representations in real-time.
 Domain adaption methods. We compare DeSGDA with two recent domain adaption methods:
@@
   • DEAL: DEAL Yin et al.
-[2022] uses domain adversarial learning to align graph representations across different domains without labeled data, overcoming discrepancies between the source and target domains.
+[2022b] uses domain adversarial learning to align graph representations across different domains without labeled data, overcoming discrepancies between the source and target domains.
   • CoCo: CoCo Yin et al.
-[2023] leverages contrastive learning to align graph representations between source and target domains, enhancing domain adaptation by promoting intra-domain cohesion and inter-domain separation in an unsupervised manner.
+[2023a] leverages contrastive learning to align graph representations between source and target domains, enhancing domain adaptation by promoting intra-domain cohesion and inter-domain separation in an unsupervised manner.
   • SGDA: SGDA Qiao et al. [2023] utilizes labeled data from the source domain along with a limited amount of labeled data from the target domain to learn domain-invariant graph representations.
@@
   • A2GNN: A2GNN Liu et al.
-[2024b] introduces a novel propagation mechanism to enhance feature transferability across domains, improving the alignment of graph structures and node features in an unsupervised setting.
+[2024a] introduces a novel propagation mechanism to enhance feature transferability across domains, improving the alignment of graph structures and node features in an unsupervised setting.
   • PA-BOTH: PA-BOTH Liu et al.
-[2024c] aligns node pairs between source and target graphs, optimizing feature correspondence at a granular level to improve the transferability of structural and feature information across domains.
+[2024b] aligns node pairs between source and target graphs, optimizing feature correspondence at a granular level to improve the transferability of structural and feature information across domains.
 For GCN and GIN, we use the Pytorch Geometric  [^4] https://www.pyg.org/ to implement the model.
@@
 The hidden dimension in these methods is set to 256 for a fair comparison.
-Appendix G More experimental results
-G.1 More performance comparison
-In this part, we provide additional results for our proposed method DeSGDA compared with all baseline models across various datasets, as illustrated in Table 9 to Table 17.
-These results consistently show that DeSGDA outperforms the baseline models in most cases, validating the superiority of our proposed method.
+Table 7: Statistics of the SEED dataset.
+Datasets Graphs Avg. Nodes Avg. Edges Classes
+SEED 7,636 62.0 168.2 3
+[…]
+[…]
+GCN 46.0 $\pm$ 0.9 47.8 $\pm$ 1.0 47.7 $\pm$ 1.4 49.7 $\pm$ 0.7 51.3 $\pm$ 0.8 49.8 $\pm$ 1.0
+GIN 48.9 $\pm$ 0.5 50.3 $\pm$ 0.6 49.5 $\pm$ 0.7 50.7 $\pm$ 1.0 52.7 $\pm$ 1.0 52.1 $\pm$ 0.9
+DEAL 53.5 $\pm$ 0.4 56.2 $\pm$ 0.7 53.2 $\pm$ 0.8 53.7 $\pm$ 1.1 55.1 $\pm$ 0.8 56.0 $\pm$ 0.7
+CoCo 53.9 $\pm$ 0.5 53.0 $\pm$ 0.6 54.1 $\pm$ 0.7 54.3 $\pm$ 0.7 55.3 $\pm$ 1.0 55.9 $\pm$ 0.6
+DeSGDA 54.5 $\pm$ 0.6 55.6 $\pm$ 0.7 54.6 $\pm$ 0.5 54.5 $\pm$ 0.7 55.8 $\pm$ 1.1 56.6 $\pm$ 0.8
+Table 9: GPU memory consumption of different graph domain methods in training stage for each training epoch (in GB).
+DeSGDA DEAL CoCo A2GNN PA-BOTH
+PROTEINS 2.3 1.4 1.2 33.1 5.6
+NCI1 5.7 2.9 3.3 34.8 11.9
+MUTAGENICITY 5.0 2.8 3.2 35.0 12.3
+FRANKENSTEIN 2.9 2.4 1.9 33.6 7.3
+Table 10: Time consumption of different graph domain methods in training stage for each training epoch (in seconds).
+DeSGDA DEAL CoCo A2GNN PA-BOTH
+PROTEINS 0.587 0.126 22.123 0.869 0.283
+NCI1 0.855 0.518 58.564 1.483 0.597
+MUTAGENICITY 0.887 0.612 52.740 1.553 0.511
+FRANKENSTEIN 0.663 0.375 26.837 1.016 0.275
+Appendix H More experimental results
+H.1 More performance comparison
+In this part, we provide additional results for our proposed method DeSGDA compared with all baseline models across various datasets, as illustrated in Table 1 and Table 13-23.
+These results consistently show that DeSGDA outperforms the baselines in most cases, validating the superiority of our proposed method.
 Additionally, the performance of DeSGDA with different GNN architectures on the NCI1, FRANKENSTEIN, and MUTAGENICITY datasets is shown in Figure 10.
 It is evident that GIN consistently outperforms other GNN architectures in most cases.
-G.2 More Ablation study
+To verify the efffectiveness of DeSGDA in EEG data and multi-class scenarios, we conducted additional experiments using the SEED dataset Duan et al.
+[2013], Zheng and Lu [2015], which is a well-known EEG dataset for emotion classification.
+For EEG data processing, we utilized the torcheeg  [^5] https://github.com/torcheeg/torcheeg library to convert standard EEG data into graph structures.
+During graph construction, we remove edges from each graph and partitioning the dataset into source and target domains based on edge density Klepl et al.
+[2022].
+The statistics of SEED dataset is shown in Table 7.
+We compared the DeSGDA with two general graph neural networks (GCN Kipf and Welling [2017] and GIN Xu et al.
+[2018]) and two graph domain adaptation methods (DEAL Yin et al.
+[2022b] and CoCo Yin et al.
+[2023a]).
+The results in Table 8 show that DeSGDA still outperforms the other methods in most cases.
+H.2 Training time and memory comparison
+We provide detailed comparisons of GPU memory consumption and training time per epoch for DeSGDA and other graph domain adaptation methods under identical experimental settings in this part, as shown in Tables 9 and 10.
+It is worth noting that the training phase is typically conducted on more powerful hardware to achieve optimal performance within a reasonable time frame.
+Table 11: The results of DeSGDA with different widely used graph neural networks (GIN, GCN and SAGE) on PROTEINS and NCI1 dataset. Bold results indicate the best performance.
+[…]
+DeSGDA w GCN 76.6 70.5 71.8 74.1 66.3 68.3 68.5 67.1
+DeSGDA w SAGE 75.8 73.3 72.4 75.2 67.2 69.5 66.6 68.4
+DeSGDA w GIN 77.3 75.8 73.8 77.1 69.0 69.8 68.8 68.8
+DeSGDA 78.7 78.4 74.8 77.6 68.5 71.4 70.1 69.0
+Table 12: The results of DeSGDA with different widely used graph neural networks (GIN, GCN and SAGE) on MUTAGENICITY and FRANKENSTEIN dataset. Bold results indicate the best performance.
+[…]
+DeSGDA w GCN 60.9 60.7 63.6 60.6 61.3 62.2 60.7 61.1
+DeSGDA w SAGE 61.1 62.3 64.2 61.1 61.9 62.6 61.3 61.6
+DeSGDA w GIN 64.5 65.4 65.0 63.9 62.8 63.6 61.8 62.8
+DeSGDA 65.4 65.9 65.5 65.6 63.5 64.1 62.6 63.0
+H.3 More Ablation study
 To validate the effectiveness of the different components in DeSGDA, we conduct more experiments with sive variants on NCI1, FRANKENSTEIN and MUTAGENICITY datasets, i.e., DeSGDA w/o CA, DeSGDA w/o PL, DeSGDA w/o CF, DeSGDA w/o TL, DeSGDA w PT and DeSGDA w CL.
-The results are shown in Table 6 , 7 and 8.
+The results are shown in Table 4 , 5 and 6.
 From the results, we have similar observations as summarized in Section 5.4.
-G.3 More Sensitivity Analysis
+Additionally, we conduct ablation studies to examine the effect of directly replacing the SGNs with commonly used Graph Neural Networks (GNNs) for generating representations for DeSGDA: (1) DeSGDA w GCN: It replaces SGNs with GCN Kipf and Welling [2017]; (2) DeSGDA w GIN: It replaces SGNs with GIN Xu et al.
+[2018]; (3) DeSGDA w SAGE: It replaces SGNs with GraphSAGE Hamilton et al.
+[2017].
+The experimental results across the PROTEINS, NCI1, MUTAGENICITY, and FRANKENSTEIN datasets are shown in Table 11 and 12.
+However, the critical aspect of our work lies in the specific problem we set up, i.e., low-power and distribution shift environments.
+In this context, directly replacing SGNs with commonly used GNNs like GIN or GCN is not feasible, as these models are unsuitable for deployment on low-energy devices.
+As demonstrated in Section 5.3, GNN based methods have much higher energy consumption than the spike based methods.
+H.4 More Sensitivity Analysis
 In this part, we provide additional sensitivity analysis of the proposed DeSGDA with respect to the impact of its hyperparameters: the time latency $\tau$, degree threshold value $V_{th}^{degree}$ in SNNs, and balance ratio $\alpha$ on NCI1, FRANKENSTEIN, and MUTAGENICITY datasets.
-The results are illustrated in Figure 9, 10 and 11, where we observe trends similar to those discussed in Section 5.5.
+The results are illustrated in Figure 11, 12 and 13, where we observe trends similar to those discussed in Section 5.5.
+(a) NCI1
+(b) FRANKENSTEIN
+(c) MUTAGENICITY
+Figure 10: The performance with different GNN architectures on different datasets.
 (a) Time Latency
@@
 Figure 13: Hyperparameter sensitivity analysis of time latency $\tau$, initial threshold $V_{th}^{degree}$ in SNNs, and balance ratio $\alpha$ on MUTAGENICITY.
-(a) NCI1
-(b) FRANKENSTEIN
-(c) MUTAGENICITY
-Figure 12: The performance with different GNN architectures on different datasets.
 […]
 […]
-WL subtree 73.4 72.7 70.5 73.0 72.8 59.0 66.5 71.6 60.6 58.3 76.3 64.0 68.2
-GCN 57.2$\pm$2.7 62.8$\pm$1.7 67.6$\pm$0.5 58.5$\pm$1.3 67.7$\pm$0.4 61.0$\pm$0.3 65.0$\pm$0.8 51.1$\pm$1.3 65.6$\pm$2.2 55.4$\pm$0.4 68.5$\pm$3.1 67.7$\pm$0.5 62.3
-GIN 69.3$\pm$2.3 65.8$\pm$0.8 69.3$\pm$1.7 69.8$\pm$1.6 71.4$\pm$2.1 52.4$\pm$1.8 64.0$\pm$2.4 65.7$\pm$3.2 53.4$\pm$3.7 58.1$\pm$0.8 72.6$\pm$0.3 64.6$\pm$2.3 64.7
-GMT 67.8$\pm$1.3 69.6$\pm$0.7 74.5$\pm$0.5 67.6$\pm$2.5 69.9$\pm$2.1 55.8$\pm$0.7 74.8$\pm$1.4 60.1$\pm$2.4 71.4$\pm$3.3 51.5$\pm$0.5 69.0$\pm$0.5 63.3$\pm$1.3 66.3
-CIN 62.6$\pm$0.5 59.4$\pm$0.5 64.0$\pm$0.9 58.5$\pm$1.8 71.9$\pm$1.7 60.6$\pm$2.1 63.7$\pm$0.5 61.2$\pm$2.1 73.2$\pm$0.5 57.7$\pm$3.0 68.1$\pm$0.4 58.5$\pm$2.7 63.3
-SpikeGCN 66.1$\pm$0.4 67.8$\pm$0.7 68.5$\pm$0.9 67.2$\pm$0.3 70.1$\pm$1.1 68.6$\pm$1.3 66.1$\pm$0.7 65.8$\pm$1.2 74.5$\pm$0.9 65.4$\pm$0.6 73.8$\pm$1.1 70.1$\pm$1.6 68.6
-DRSGNN 68.3$\pm$0.5 71.9$\pm$0.9 71.5$\pm$1.1 75.1$\pm$1.7 76.7$\pm$1.3 74.4$\pm$0.8 70.5$\pm$0.7 67.8$\pm$0.8 77.0$\pm$1.5 65.9$\pm$1.6 75.2$\pm$1.9 69.2$\pm$2.1 71.8
-CDAN 75.6$\pm$0.5 70.5$\pm$0.6 71.6$\pm$0.5 69.8$\pm$0.5 76.6$\pm$0.8 71.4$\pm$0.3 71.4$\pm$0.3 72.1$\pm$0.3 75.5$\pm$0.7 74.3$\pm$0.8 78.2$\pm$1.1 74.0$\pm$0.8 73.4
-ToAlign 51.1$\pm$0.6 55.8$\pm$0.1 63.3$\pm$0.2 55.8$\pm$0.4 68.1$\pm$0.7 55.8$\pm$0.3 63.3$\pm$0.5 51.1$\pm$0.2 68.1$\pm$1.0 51.1$\pm$0.4 68.1$\pm$0.6 63.3$\pm$0.2 59.6
-MetaAlign 59.4$\pm$1.1 62.2$\pm$1.0 68.9$\pm$0.3 65.3$\pm$0.8 75.1$\pm$0.7 67.5$\pm$2.1 70.9$\pm$1.4 60.6$\pm$2.3 72.4$\pm$1.4 59.4$\pm$0.6 74.6$\pm$0.7 67.8$\pm$1.3 67.0
-DEAL 76.6$\pm$0.4 62.8$\pm$0.8 72.8$\pm$1.3 67.3$\pm$2.2 77.2$\pm$2.3 67.6$\pm$1.9 71.2$\pm$1.6 56.0$\pm$2.5 73.9$\pm$2.1 66.0$\pm$0.3 76.4$\pm$1.1 65.5$\pm$2.1 69.4
-CoCo 73.4$\pm$0.5 73.6$\pm$0.8 73.4$\pm$1.0 71.6$\pm$0.5 75.2$\pm$1.6 74.6$\pm$0.3 70.7$\pm$0.8 68.4$\pm$1.5 75.0$\pm$0.2 72.7$\pm$0.4 76.3$\pm$1.1 75.0$\pm$1.8 73.3
-SGDA 63.8$\pm$0.8 65.2$\pm$0.5 66.7$\pm$0.3 59.1$\pm$1.2 62.3$\pm$0.7 60.6$\pm$0.4 65.2$\pm$0.9 61.8$\pm$1.0 64.5$\pm$1.3 60.9$\pm$0.8 59.4$\pm$1.2 64.9$\pm$1.1 62.9
-DGDA 59.4$\pm$0.7 62.3$\pm$1.1 63.1$\pm$0.5 61.2$\pm$0.9 60.4$\pm$0.6 58.8$\pm$1.0 60.3$\pm$0.8 63.5$\pm$1.2 61.9$\pm$0.8 60.4$\pm$1.6 64.2$\pm$1.3 62.6$\pm$1.4 61.5
-A2GNN 65.4$\pm$0.7 66.4$\pm$1.1 65.7$\pm$1.3 66.0$\pm$0.6 64.9$\pm$1.2 65.8$\pm$1.6 65.5$\pm$1.8 66.0$\pm$1.4 65.8$\pm$2.1 65.6$\pm$1.9 66.1$\pm$1.7 66.0$\pm$2.0 65.8
-PA-BOTH 66.9$\pm$0.5 67.1$\pm$0.8 67.3$\pm$1.1 65.8$\pm$0.7 69.1$\pm$1.0 66.1$\pm$1.4 66.7$\pm$1.3 67.4$\pm$1.4 66.3$\pm$1.8 66.0$\pm$1.2 66.8$\pm$0.8 66.3$\pm$1.5 66.8
-DeSGDA 78.7$\pm$1.3 78.4$\pm$1.1 74.8$\pm$0.6 77.6$\pm$0.9 79.5$\pm$1.2 76.7$\pm$0.8 74.9$\pm$0.7 71.2$\pm$1.7 79.5$\pm$1.4 72.8$\pm$0.8 81.0$\pm$1.5 75.1$\pm$1.0 76.7
+WL subtree 75.9 70.4 64.3 63.9 60.6 64.7 73.2 78.9 66.8 69.2 74.2 72.9 69.6
+GCN 49.2$\pm$1.7 55.8$\pm$1.5 46.8$\pm$0.5 54.6$\pm$2.2 43.4$\pm$0.6 46.7$\pm$0.2 50.0$\pm$1.8 57.2$\pm$2.2 44.2$\pm$0.4 51.6$\pm$0.8 62.7$\pm$2.1 56.8$\pm$1.3 51.6
+GIN 68.8$\pm$2.5 70.6$\pm$1.0 64.2$\pm$1.1 67.2$\pm$2.4 62.2$\pm$1.8 62.5$\pm$1.5 68.7$\pm$2.4 72.5$\pm$0.6 63.3$\pm$1.6 65.2$\pm$0.6 62.4$\pm$0.3 70.9$\pm$0.5 66.6
+GMT 66.7$\pm$0.3 58.2$\pm$0.5 63.9$\pm$0.3 58.4$\pm$0.3 63.8$\pm$0.4 56.7$\pm$0.5 63.9$\pm$0.7 66.3$\pm$1.0 63.8$\pm$1.1 66.6$\pm$0.4 63.8$\pm$0.2 62.6$\pm$0.7 62.9
+CIN 58.7$\pm$2.4 54.9$\pm$0.2 52.0$\pm$0.3 54.8$\pm$0.1 56.6$\pm$0.2 54.9$\pm$0.1 52.9$\pm$1.4 52.8$\pm$0.5 56.5$\pm$0.6 52.8$\pm$2.1 58.5$\pm$0.8 56.6$\pm$1.4 55.1
+SpikeGCN 58.9$\pm$0.9 65.2$\pm$1.2 60.8$\pm$1.3 62.0$\pm$1.5 62.3$\pm$0.8 58.7$\pm$1.6 64.1$\pm$0.6 66.7$\pm$1.3 60.5$\pm$1.7 63.8$\pm$1.4 62.2$\pm$2.1 61.1$\pm$1.5 62.1
+DRSGNN 58.0$\pm$0.6 64.3$\pm$1.1 61.2$\pm$0.8 62.2$\pm$1.0 62.9$\pm$1.5 64.0$\pm$1.3 60.6$\pm$1.6 64.0$\pm$1.4 67.6$\pm$2.1 62.4$\pm$1.9 71.3$\pm$2.3 68.8$\pm$2.0 63.9
+CDAN 64.0$\pm$1.1 68.1$\pm$0.3 60.1$\pm$0.5 64.0$\pm$1.3 60.9$\pm$0.2 57.8$\pm$1.0 64.3$\pm$1.6 61.2$\pm$0.2 66.3$\pm$0.7 59.0$\pm$0.5 68.9$\pm$0.3 63.7$\pm$0.6 63.2
+ToAlign 52.8$\pm$0.5 54.8$\pm$0.2 48.2$\pm$1.1 54.8$\pm$1.5 44.0$\pm$0.8 54.8$\pm$2.0 48.2$\pm$1.7 52.8$\pm$0.6 44.0$\pm$0.2 52.8$\pm$0.3 44.0$\pm$1.0 48.2$\pm$1.2 50.0
+MetaAlign 63.1$\pm$0.3 63.8$\pm$1.3 58.9$\pm$2.4 58.5$\pm$0.4 59.1$\pm$2.1 59.2$\pm$1.6 70.1$\pm$0.8 63.3$\pm$1.4 66.5$\pm$2.7 60.9$\pm$1.1 71.4$\pm$0.2 67.5$\pm$0.8 63.5
+DEAL 70.7$\pm$0.9 72.3$\pm$0.2 69.9$\pm$0.8 68.9$\pm$0.7 64.1$\pm$0.6 65.6$\pm$0.9 71.9$\pm$0.4 69.9$\pm$1.7 70.6$\pm$0.4 66.5$\pm$0.3 71.6$\pm$0.7 69.9$\pm$0.5 69.3
+CoCo 64.0$\pm$1.3 63.9$\pm$0.6 65.8$\pm$1.8 59.9$\pm$1.7 62.2$\pm$2.1 60.6$\pm$1.6 65.0$\pm$2.1 64.8$\pm$1.4 60.0$\pm$0.8 61.3$\pm$0.5 68.5$\pm$0.4 67.1$\pm$0.6 63.6
+SGDA OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
+DGDA OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
+A2GNN 58.9$\pm$0.9 60.1$\pm$0.7 59.8$\pm$1.2 59.4$\pm$1.0 62.3$\pm$1.5 60.9$\pm$1.6 61.6$\pm$1.3 59.9$\pm$1.9 64.9$\pm$1.6 62.9$\pm$2.1 65.4$\pm$1.5 63.3$\pm$2.3 61.7
+PA-BOTH 61.1$\pm$0.5 60.9$\pm$0.4 61.6$\pm$0.6 61.2$\pm$0.8 60.8$\pm$0.6 61.5$\pm$0.5 62.2$\pm$1.0 61.9$\pm$0.7 61.8$\pm$1.1 61.1$\pm$0.9 60.9$\pm$0.8 61.3$\pm$1.2 61.6
+DeSGDA 68.5$\pm$1.2 71.4$\pm$1.3 70.1$\pm$0.7 69.0$\pm$1.1 68.9$\pm$1.0 66.3$\pm$1.4 69.6$\pm$1.3 70.2$\pm$1.7 71.1$\pm$1.6 69.3$\pm$1.8 74.4$\pm$1.6 70.0$\pm$1.9 69.9
 […]
@@
 PA-BOTH 62.2$\pm$0.5 60.7$\pm$0.7 61.5$\pm$0.6 61.2$\pm$1.0 61.9$\pm$1.3 61.1$\pm$0.8 62.3$\pm$0.4 61.7$\pm$0.8 62.0$\pm$0.9 61.1$\pm$1.2 61.2$\pm$1.5 60.9$\pm$0.6 61.5
-DeSGDA 63.5$\pm$1.1 64.1$\pm$0.9 62.6$\pm$1.3 63.0$\pm$0.8 81.1$\pm$1.2 62.4$\pm$1.5 65.5$\pm$0.6 65.4$\pm$1.7 81.9$\pm$1.0 60.8$\pm$1.4 82.0$\pm$2.1 65.9$\pm$1.8 68.1
+DeSGDA 63.5$\pm$1.1 64.1$\pm$0.9 62.6$\pm$1.3 63.0$\pm$0.8 81.1$\pm$1.2 62.4$\pm$1.5 65.5$\pm$0.6 65.4$\pm$1.7 81.9$\pm$1.0 60.8$\pm$1.4 82.0$\pm$2.1 65.9$\pm$1.8 68.2
 […]
@@
 PA-BOTH 61.2$\pm$0.9 62.0$\pm$0.4 60.7$\pm$0.8 61.7$\pm$0.5 60.9$\pm$1.2 61.1$\pm$0.7 61.5$\pm$0.9 60.2$\pm$1.2 61.3$\pm$1.5 61.8$\pm$0.8 62.2$\pm$0.9 62.0$\pm$1.0 61.4
-DeSGDA 65.4$\pm$1.3 65.9$\pm$0.9 65.5$\pm$1.4 64.8$\pm$1.1 82.8$\pm$1.6 63.6$\pm$1.5 70.7$\pm$1.8 68.2$\pm$0.7 82.9$\pm$1.2 67.6$\pm$0.8 83.9$\pm$1.0 66.5$\pm$1.2 70.9
+DeSGDA 65.4$\pm$1.3 65.9$\pm$0.9 65.5$\pm$1.4 65.6$\pm$1.1 82.8$\pm$1.6 63.6$\pm$1.5 70.7$\pm$1.8 68.2$\pm$0.7 82.9$\pm$1.2 67.6$\pm$0.8 83.9$\pm$1.0 66.5$\pm$1.2 70.8
 […]
@@
 PA-BOTH 61.0$\pm$0.8 61.2$\pm$1.3 60.3$\pm$0.6 66.7$\pm$2.1 63.7$\pm$1.5 61.9$\pm$2.0 66.2$\pm$1.4 69.9$\pm$2.3 68.0$\pm$0.7 69.4$\pm$1.8 61.5$\pm$0.4 67.6$\pm$1.0 64.9
-DeSGDA 77.6$\pm$0.9 84.3$\pm$1.1 70.5$\pm$0.6 84.8$\pm$1.4 76.6$\pm$0.7 83.9$\pm$0.9 71.9$\pm$0.6 76.9$\pm$1.1 76.1$\pm$0.8 73.7$\pm$1.0 76.0$\pm$1.2 70.4$\pm$0.7 76.8
+DeSGDA 77.6$\pm$0.9 84.3$\pm$1.1 70.5$\pm$0.6 84.8$\pm$1.4 76.6$\pm$0.7 83.9$\pm$0.9 71.9$\pm$0.6 76.9$\pm$1.1 76.1$\pm$0.8 73.7$\pm$1.0 76.0$\pm$1.2 70.4$\pm$0.7 76.9
 […]
@@
 PA-BOTH 57.7$\pm$0.4 58.0$\pm$0.6 57.9$\pm$0.5 56.9$\pm$0.8 57.4$\pm$0.6 58.3$\pm$0.5 57.1$\pm$1.2 58.8$\pm$0.9 58.1$\pm$0.7 58.0$\pm$0.9 57.9$\pm$0.5 58.3$\pm$0.8 57.9
-DeSGDA 64.4$\pm$1.2 76.9$\pm$1.5 64.9$\pm$0.9 76.6$\pm$1.2 68.6$\pm$1.8 74.1$\pm$1.3 66.9$\pm$0.8 65.1$\pm$1.2 69.9$\pm$1.5 63.9$\pm$2.0 70.9$\pm$1.6 64.2$\pm$1.4 69.0
+DeSGDA 64.4$\pm$1.2 76.9$\pm$1.5 64.9$\pm$0.9 76.6$\pm$1.2 68.6$\pm$1.8 74.1$\pm$1.3 66.9$\pm$0.8 65.1$\pm$1.2 69.9$\pm$1.5 63.9$\pm$2.0 70.9$\pm$1.6 64.2$\pm$1.4 68.9
+[…]
+[…]
+WL subtree 65.7 71.8 57.9 71.1 47.4 43.4 65.5 75.1 45.3 34.9 52.7 49.8 56.7
+GCN 70.6$\pm$2.1 60.3$\pm$1.5 60.5$\pm$3.4 62.3$\pm$1.1 58.4$\pm$0.5 43.2$\pm$0.2 63.8$\pm$1.0 70.3$\pm$0.3 50.6$\pm$1.0 32.8$\pm$0.3 50.1$\pm$0.4 42.2$\pm$0.2 55.4
+GIN 66.7$\pm$2.1 73.7$\pm$2.4 57.3$\pm$3.1 69.4$\pm$2.3 58.6$\pm$0.4 43.1$\pm$0.3 66.4$\pm$2.7 74.8$\pm$1.8 42.2$\pm$1.6 33.5$\pm$1.0 57.4$\pm$0.8 43.9$\pm$2.3 57.2
+GMT 67.3$\pm$0.3 56.8$\pm$0.4 58.0$\pm$0.2 56.8$\pm$0.2 60.6$\pm$0.3 56.8$\pm$0.5 57.8$\pm$0.1 67.3$\pm$0.1 39.5$\pm$0.3 67.3$\pm$0.2 39.5$\pm$0.5 57.8$\pm$0.4 57.1
+CIN 67.6$\pm$0.4 63.7$\pm$2.1 58.9$\pm$1.0 56.8$\pm$0.4 63.6$\pm$0.4 59.5$\pm$2.7 58.7$\pm$1.2 67.0$\pm$0.5 61.7$\pm$1.6 67.8$\pm$0.7 62.2$\pm$2.1 56.0$\pm$1.3 61.9
+SpikeGCN 67.2$\pm$0.5 57.2$\pm$1.2 57.9$\pm$0.8 57.1$\pm$0.6 61.1$\pm$1.3 58.9$\pm$1.6 60.0$\pm$1.2 67.2$\pm$0.9 53.9$\pm$2.1 64.4$\pm$0.8 57.8$\pm$1.0 59.9$\pm$1.2 60.2
+DRSGNN 67.4$\pm$0.4 58.4$\pm$1.0 59.0$\pm$1.2 57.4$\pm$0.5 62.3$\pm$1.1 60.4$\pm$1.3 61.1$\pm$1.6 67.9$\pm$1.5 56.2$\pm$1.8 66.2$\pm$2.1 60.9$\pm$1.4 58.6$\pm$2.5 61.3
+CDAN 72.9$\pm$0.4 72.7$\pm$0.4 65.4$\pm$0.3 72.9$\pm$0.1 61.2$\pm$0.3 70.3$\pm$0.2 65.7$\pm$0.4 72.7$\pm$0.1 61.0$\pm$0.1 72.1$\pm$1.2 60.7$\pm$0.2 65.3$\pm$0.6 67.7
+ToAlign 32.7$\pm$2.0 43.2$\pm$0.1 42.2$\pm$1.3 43.2$\pm$0.9 60.5$\pm$0.7 43.2$\pm$1.2 42.2$\pm$0.4 32.7$\pm$1.2 60.5$\pm$0.9 32.7$\pm$0.3 60.5$\pm$0.7 42.2$\pm$0.4 44.7
+MetaAlign 67.3$\pm$0.7 56.8$\pm$0.2 57.8$\pm$0.6 56.8$\pm$0.4 60.5$\pm$1.3 56.8$\pm$0.8 57.8$\pm$1.1 67.3$\pm$1.2 60.5$\pm$0.4 67.3$\pm$0.6 60.5$\pm$0.7 57.8$\pm$0.6 60.6
+DEAL 75.0$\pm$0.9 76.3$\pm$2.4 65.9$\pm$1.8 77.5$\pm$2.7 60.3$\pm$4.5 69.7$\pm$3.2 67.2$\pm$1.5 75.3$\pm$1.7 57.4$\pm$4.1 71.1$\pm$2.2 65.7$\pm$2.7 66.4$\pm$1.6 69.0
+CoCo 74.2$\pm$1.7 74.3$\pm$0.6 65.9$\pm$1.2 72.7$\pm$2.1 61.1$\pm$0.2 71.0$\pm$1.7 68.6$\pm$0.3 75.9$\pm$0.2 60.7$\pm$0.2 73.9$\pm$0.4 59.7$\pm$1.1 67.3$\pm$0.8 68.8
+SGDA 55.9$\pm$0.6 57.1$\pm$0.5 56.1$\pm$0.4 54.6$\pm$0.8 55.8$\pm$1.1 57.7$\pm$0.6 54.3$\pm$0.7 53.6$\pm$1.3 59.1$\pm$0.8 56.7$\pm$0.6 55.4$\pm$1.2 53.8$\pm$0.5 55.9
+DGDA OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
+A2GNN 55.9$\pm$0.7 55.7$\pm$0.4 56.6$\pm$0.6 57.1$\pm$1.0 56.1$\pm$1.2 55.8$\pm$0.5 56.5$\pm$0.7 55.5$\pm$0.4 55.9$\pm$0.8 56.2$\pm$0.6 56.5$\pm$1.5 56.0$\pm$0.5 56.2
+PA-BOTH 56.4$\pm$0.5 55.9$\pm$0.6 56.0$\pm$0.5 56.4$\pm$0.4 56.3$\pm$0.6 57.7$\pm$0.7 56.6$\pm$0.2 58.8$\pm$0.9 56.9$\pm$0.7 57.2$\pm$0.3 56.5$\pm$0.5 58.3$\pm$0.8 56.9
+DeSGDA 75.2$\pm$0.8 74.3$\pm$1.2 68.0$\pm$1.5 73.6$\pm$2.0 62.3$\pm$1.7 71.2$\pm$2.5 68.8$\pm$1.3 76.0$\pm$1.2 61.9$\pm$2.2 71.4$\pm$1.6 62.8$\pm$1.2 65.7$\pm$0.9 69.3
 […]
@@
 PA-BOTH 55.9$\pm$1.0 56.0$\pm$0.5 56.1$\pm$0.7 56.6$\pm$1.2 55.9$\pm$0.6 56.0$\pm$0.7 57.3$\pm$0.8 56.8$\pm$1.3 55.9$\pm$1.2 56.3$\pm$1.0 56.4$\pm$0.9 57.1$\pm$1.3 56.4
-DeSGDA 76.7$\pm$1.0 72.5$\pm$0.6 72.0$\pm$2.3 71.5$\pm$1.5 61.0$\pm$1.2 67.6$\pm$0.8 75.5$\pm$1.0 79.8$\pm$1.2 62.8$\pm$1.6 75.9$\pm$1.3 65.5$\pm$2.1 72.3$\pm$1.7 71.2
+DeSGDA 76.7$\pm$1.0 72.5$\pm$0.6 72.0$\pm$2.3 71.5$\pm$1.5 61.0$\pm$1.2 67.6$\pm$0.8 75.5$\pm$1.0 79.8$\pm$1.2 62.8$\pm$1.6 75.9$\pm$1.3 65.5$\pm$2.1 72.3$\pm$1.7 71.1
+[…]
+[…]
+WL subtree 68.7 82.3 50.7 82.3 58.1 83.8 64.0 74.1 43.7 70.5 71.3 60.1 67.5
+GCN 73.4$\pm$0.2 83.5$\pm$0.3 57.6$\pm$0.2 84.2$\pm$1.8 24.0$\pm$0.1 16.6$\pm$0.4 57.6$\pm$0.2 73.7$\pm$0.4 24.0$\pm$0.1 26.6$\pm$0.2 39.9$\pm$0.9 42.5$\pm$0.1 50.3
+GIN 62.5$\pm$4.7 74.9$\pm$3.7 53.0$\pm$4.6 59.6$\pm$4.2 73.7$\pm$0.8 64.7$\pm$3.4 60.6$\pm$2.7 69.8$\pm$0.6 31.1$\pm$2.8 63.1$\pm$3.4 72.3$\pm$2.7 64.6$\pm$1.4 62.5
+GMT 73.4$\pm$0.3 83.5$\pm$0.2 57.6$\pm$0.1 83.5$\pm$0.3 24.0$\pm$0.1 83.5$\pm$0.1 57.4$\pm$0.2 73.4$\pm$0.2 24.1$\pm$0.1 73.4$\pm$0.3 24.0$\pm$0.1 57.6$\pm$0.2 59.6
+CIN 74.5$\pm$0.2 84.1$\pm$0.5 57.8$\pm$0.2 82.7$\pm$0.9 75.6$\pm$0.6 79.2$\pm$2.2 61.5$\pm$2.7 74.0$\pm$1.0 75.5$\pm$0.8 72.5$\pm$2.1 76.0$\pm$0.3 60.9$\pm$1.2 72.9
+SpikeGCN 71.8$\pm$0.8 79.5$\pm$1.3 63.8$\pm$1.0 78.9$\pm$1.4 68.6$\pm$1.1 76.5$\pm$1.8 62.3$\pm$2.2 72.1$\pm$1.5 68.1$\pm$2.1 67.2$\pm$1.9 69.2$\pm$2.1 64.2$\pm$1.8 70.2
+DRSGNN 72.6$\pm$0.6 80.1$\pm$1.6 63.1$\pm$1.4 79.5$\pm$1.8 70.4$\pm$1.9 78.6$\pm$2.1 64.1$\pm$1.7 70.7$\pm$2.3 67.8$\pm$1.6 65.6$\pm$1.4 71.3$\pm$1.3 62.1$\pm$1.0 70.5
+CDAN 72.2$\pm$1.8 82.4$\pm$1.6 59.8$\pm$2.1 76.8$\pm$2.4 69.3$\pm$4.1 71.8$\pm$3.7 64.4$\pm$2.5 74.3$\pm$0.4 46.3$\pm$2.0 69.8$\pm$1.8 74.4$\pm$1.7 62.6$\pm$2.3 68.7
+ToAlign 73.4$\pm$0.1 83.5$\pm$0.2 57.6$\pm$0.1 83.5$\pm$0.2 24.0$\pm$0.3 83.5$\pm$0.4 57.6$\pm$0.1 73.4$\pm$0.1 24.0$\pm$0.2 73.4$\pm$0.2 24.0$\pm$0.1 57.6$\pm$0.3 59.6
+MetaAlign 75.5$\pm$0.9 84.9$\pm$0.6 64.8$\pm$1.6 85.9$\pm$1.1 69.3$\pm$2.7 83.3$\pm$0.6 68.7$\pm$1.2 74.2$\pm$0.7 73.3$\pm$3.3 72.2$\pm$0.9 69.9$\pm$1.8 63.6$\pm$2.3 73.8
+DEAL 76.5$\pm$0.4 83.1$\pm$0.4 67.5$\pm$1.3 77.6$\pm$1.8 76.0$\pm$0.2 80.1$\pm$2.7 66.1$\pm$1.3 75.4$\pm$1.5 42.3$\pm$4.1 68.1$\pm$3.7 73.1$\pm$2.2 67.8$\pm$1.2 71.1
+CoCo 75.5$\pm$0.2 84.2$\pm$0.4 59.8$\pm$0.5 83.4$\pm$0.2 73.6$\pm$2.3 81.6$\pm$2.4 65.8$\pm$0.3 76.2$\pm$0.2 75.8$\pm$0.2 71.1$\pm$2.1 76.1$\pm$0.2 67.1$\pm$0.6 74.2
+SGDA 63.8$\pm$0.6 65.2$\pm$1.3 66.7$\pm$1.0 59.1$\pm$1.5 60.1$\pm$0.8 64.4$\pm$1.2 65.2$\pm$0.7 63.9$\pm$0.9 64.5$\pm$0.6 61.1$\pm$1.3 58.9$\pm$1.4 64.9$\pm$1.2 63.2
+DGDA 58.7$\pm$0.8 59.9$\pm$1.2 57.1$\pm$0.6 57.9$\pm$0.8 59.2$\pm$1.3 58.9$\pm$0.4 61.1$\pm$1.2 60.3$\pm$1.6 58.6$\pm$0.9 57.5$\pm$1.2 58.4$\pm$0.5 62.3$\pm$1.5 59.2
+A2GNN 65.4$\pm$1.3 66.3$\pm$1.1 68.2$\pm$1.4 66.3$\pm$1.2 65.4$\pm$0.7 65.9$\pm$0.9 66.9$\pm$1.3 65.4$\pm$1.2 65.6$\pm$0.9 65.5$\pm$1.2 66.1$\pm$2.0 66.0$\pm$1.8 66.1
+PA-BOTH 63.1$\pm$0.7 67.2$\pm$1.1 64.3$\pm$0.5 72.1$\pm$1.8 66.3$\pm$0.7 64.1$\pm$1.2 69.7$\pm$2.1 67.5$\pm$1.8 61.2$\pm$1.4 67.7$\pm$2.3 61.2$\pm$1.6 65.5$\pm$0.6 65.9
+DeSGDA 76.7$\pm$0.8 84.6$\pm$0.9 69.4$\pm$0.6 85.2$\pm$1.5 76.2$\pm$1.1 83.9$\pm$1.2 69.9$\pm$0.6 76.3$\pm$1.4 75.9$\pm$1.0 73.5$\pm$1.3 76.3$\pm$1.6 68.3$\pm$0.7 76.4
 […]

### v4

@@
-Degree Distribution based Spiking Graph Networks for Domain Adaptation
+Degree-Conscious Spiking Graph for Cross-Domain Adaptation
 Abstract
-Spiking Graph Networks (SGNs) have garnered significant attraction from researchers and industry due to their ability to address energy consumption challenges in graph classification.
-However, SGNs are only effective for in-distribution data and cannot tackle out-of-distribution data.
-In this paper, we first propose the domain adaptation problem in SGNs, and introduce a novel framework named Degree-aware Spiking Graph Domain Adaptation for Classification (DeSGDA).
-The proposed DeSGDA addresses the spiking graph domain adaptation problem in three aspects: node degree-aware personalized spiking representation, adversarial feature distribution alignment, and pseudo-label distillation.
-First, we introduce the personalized spiking representation method for generating degree-dependent spiking signals.
-Specifically, the node degree determines the threshold of triggering a spike, allowing this personalized approach to capture more expressive information for classification.
-Then, we propose the graph feature distribution alignment module that is adversarially trained using membrane potential against a domain discriminator.
-Such an alignment module can efficiently maintain high performance and low energy consumption in the case of inconsistent distribution.
-Additionally, we extract consistent predictions across two spaces to create reliable pseudo-labels, effectively leveraging unlabeled data to enhance classification performance.
-Extensive experiments on benchmark datasets validate the superiority of the DeSGDA compared with competitive baselines.
+Spiking Graph Networks (SGNs) have demonstrated significant potential in graph classification by emulating brain-inspired neural dynamics to achieve energy-efficient computation.
+However, existing SGNs are generally constrained to in-distribution scenarios and struggle with distribution shifts.
+In this paper, we first propose the domain adaptation problem in SGNs, and introduce a novel framework named Degree-Consicious Spiking Graph for Cross-Domain Adaptation (DeSGraDA).
+DeSGraDA enhances generalization across domains with three key components.
+First, we introduce the degree-conscious spiking representation module by adapting spike thresholds based on node degrees, enabling more expressive and structure-aware signal encoding.
+Then, we perform temporal distribution alignment by adversarially matching membrane potentials between domains, ensuring effective performance under domain shift while preserving energy efficiency.
+Additionally, we extract consistent predictions across two spaces to create reliable pseudo-labels, effectively leveraging unlabeled data to enhance graph classification performance.
+Furthermore, we establish the first generalization bound for SGDA, providing theoretical insights into its adaptation performance.
+Extensive experiments on benchmark datasets validate that DeSGraDA consistently outperforms state-of-the-art methods in both classification accuracy and energy efficiency.
 1 Introduction
-Spiking Graph Networks (SGNs) Zhu et al.
-(2022); Xu et al.
-(2021b) are a specialized artificial neural network engineered to process graph information by mimicking the human brain.
-SGNs transform static and real-valued graph features into discrete spikes by simulating neurons’ charging and discharging cycles, facilitating spike-based representations for graph node classification.
-Notably, SGNs excel in capturing semantic spiking representations with low energy consumption, which proves advantageous for event-based processing tasks Yao et al.
-(2021); Wang et al.
-(2024) such as object recognition Gu et al.
-(2020); Li et al.
-(2021b); Chen et al.
-(2024); Yin et al.
-(2024a), real-time data analysis Zhu et al.
-(2020); Bauer et al.
-(2019); Yin et al.
-(2023b), and graph classification Li et al.
-(2023); Zhu et al.
-(2022); Liu et al.
-(2024c); Yin et al.
-(2022a).
-Currently, SGNs are usually tested within the same distribution as the training dataset Li et al.
-(2023); Yin et al.
-(2024a); Duan et al.
-(2024).
-However, in realistic scenarios, the testing set can have different distributions from the training set, and such a distribution shift may degrade performance.
-For instance, Electroencephalography (EEG) data Binnie and Prior (1994); Biasiucci et al.
-(2019), typically represented as a graph structure with nodes for neurons and edges for connections, is ideally processed by bio-inspired SGNs that mimic neuronal charging and discharging.
-Despite the suitability, EEGs often exhibit varying distributions over time or among different groups Zhao et al.
-(2020, 2021); Wang et al.
-(2022), leading to suboptimal performance of models trained on specific distributions when applied to others.
-This significant issue underscores the necessity of exploring domain adaptation for spiking graphs.
-Traditionally, SNNs transfer learning methods Zhan et al.
-(2021); Yin et al.
-(2023c, 2024b); Pang et al.
-(2023) have been applied in event-based or computer vision scenarios.
-However, there’s no existing research on spiking graph domain adaptation.
-In this paper, we address the development of energy-efficient SGNs tailored for scenarios involving distribution shifts.
-Both domain adaptation and SGNs are particularly well-suited for real-world applications where data distributions vary across environments, and efficient processing of graph-structured, dynamic data under resource constraints is crucial.
-These challenges are common across numerous fields that require solutions capable of handling distribution shifts while minimizing energy consumption.
-However, designing an effective spiking graph domain adaptation framework is non-trivial due to the following major challenges: (1) How to meticulously design an SGN under the circumstance of domain shift?
-SGNs usually utilize a global threshold for the firing of each node Xu et al.
-(2021a); Yin et al.
-(2024a); Zhao et al.
-(2024).
-However, we observe that the degree of each node influences the difficulty of triggering spikes.
-Specifically, nodes with high degrees can integrate more information from neighbors, making it easier for membrane potential to accumulate and trigger a spike.
-Conversely, nodes with lower degrees are more challenging to reach the firing threshold, denoted as the inflexible architecture challenge.
-(2) How to design a framework that effectively addresses spiking graph domain adaptation for classification?
-Current research primarily focuses on graph node classification within the same distribution Li et al.
-(2023); Yao et al.
-(2023a); Duan et al.
-(2024).
-However, spike-based graph classification under domain shift remains unexplored.
-(3) How to guarantee the stability of the proposed framework?
-Though some works have been proposed to address the spiking transfer learning challenges Zhan et al.
-(2021); Zhang et al.
-(2021); Zhan et al.
-(2024), there is still no theoretical research on spiking graphs under domain shift.
-To tackle these challenges, we propose a framework named Degree-aware Spiking Graph Domain Adaptation for Classification (DeSGDA), which comprises three components: degree-aware personalized spiking representation, graph feature distribution alignment, and pseudo-label distillation.
-To address the first challenge, we establish variable node thresholds based on their degrees.
-By adaptively updating these thresholds, we achieve a more expressive and personalized spiking representation for each node.
-Then, we introduce an adversarial feature distribution alignment module trained using membrane potential against a domain discriminator.
-To further enhance performance, we extract consistent predictions from different spaces to generate reliable pseudo-labels.
-Additionally, to explore the generalization ability of the proposed DeSGDA, we propose the error bound for spiking graph domain adaptation and demonstrate that our pseudo-label distillation module effectively reduces this upper bound.
-In summary, we utilize simple yet effective techniques to address a novel problem while providing an insightful analysis of the background mechanisms and capabilities of our proposed method.
-Our contributions can be summarized as follows:
-  • Problem Formulation: We first introduce the problem of spiking graph domain adaptation for classification, which is non-trivial due to the challenges of the inflexible architecture of SGNs and theoretical deficiency.
-  • Novel Architecture: We propose DeSGDA, a framework that efficiently learns personalized spiking representations for nodes using degree-aware thresholds and aligns domain distributions through adversarial training on membrane potential.
-Furthermore, we utilize pseudo-label distillation to improve the performance further.
-  • Theoretical Analysis: To guarantee the stability of DeSGDA, we provide theoretical proof of the error bound for spiking graph domain adaptation.
-Furthermore, we demonstrate that DeSGDA maintains a lower theoretical bound than standard spiking graph domain adaptation through the use of the pseudo-label distillation module.
-  • Extensive Experiments.
-We evaluate the proposed DeSGDA on extensive spiking graph domain adaptation learning datasets, which shows that our proposed DeSGDA outperforms the variety of state-of-the-art methods.
+Spiking Graph Networks (SGNs) (Zhu et al., 2022; Xu et al., 2021b) as a specialized neural network combining Spiking Neural Networks (SNNs) (Gerstner and Kistler, 2002; Maass, 1997) with Graph Neural Networks (GNNs) (Kipf and Welling, 2017; Scarselli et al., 2009; Chen et al., 2024), have emerged as a groundbreaking paradigm in artificial neural networks, uniquely designed to process graph-structured data by mimicking the bio-inspired mechanisms of the human brain.
+SGNs convert graph features into binary spiking signals, replacing matrix multiplications with simple additions to achieve high energy efficiency.
+They further exploit temporal spiking representations, encoding information in spike timing to enable asynchronous, event-driven processing.
+It is particularly critical for applications where energy consumption is a bottleneck, such as real-time brain-computer interfaces (Kumar et al., 2022; Nason et al., 2020), large-scale sensor networks (Yao et al., 2021; Wilson et al., 2024), and temporal analysis (Yin et al., 2024; Zhu et al., 2024; Zhou et al., 2021).
+Despite their potential for energy-efficient graph representation, existing SGNs are primarily studied under closed-world assumptions, where source and target data share identical distributions (Li et al., 2023; Yin et al., 2024; Duan et al., 2024).
+This assumption is inadequate for many real-world scenarios, such as cross-subject brain–computer interfaces (BCIs) (Binnie and Prior, 1994; Biasiucci et al., 2019) where distribution shifts can significantly degrade performance(Zhao et al., 2020, 2021; Wang et al., 2022).
+Although recent advancements in transfer learning for SNNs have shown promise in vision tasks by leveraging neuromorphic adaptations (Zhang et al., 2021; Zhan et al., 2024; Guo et al., 2024), they are primarily designed for grid-like inputs and fail to generalize to graph data.
+The non-Euclidean nature and inherent irregularity of graphs introduce fundamental challenges (Bronstein et al., 2017; Wang et al., 2024b; Yao et al., 2023b), as existing methods neglect the topological dependencies and message-passing mechanisms crucial for effective graph learning.
+Consequently, directly applying these methods to SGNs results in suboptimal adaptation under domain shifts.
+In this paper, we investigate the development of energy-efficient SGNs for scenarios involving distribution shifts.
+However, designing an effective domain adaptation framework for SGNs poses several fundamental challenges: (1) How to adapt spiking representations to account for the structural diversity of graph-structured data?
+Traditional SGNs typically assign a fixed firing threshold to all nodes, ignoring the structural diversity in graphs (Xu et al., 2021a; Yin et al., 2024).
+This uniform treatment leads to under-activation in nodes with fewer connections, where important features are missed, and over-activation in highly connected nodes, where excessive signals distort the learned representation.
+Both cases reduce the model’s representational effectiveness.
+(2) How to design domain adaptation strategies that account for the temporal-based representations?
+Unlike static neural models, SGNs encode information through temporal spike sequences, making them more sensitive to domain shifts (Zhou et al., 2023; Zhan et al., 2021).
+Existing methods fail to address how these shifts impact the spike sequences, resulting in suboptimal alignment across domains.
+(3) How to theoretically characterize and bound the generalization error of SGNs under domain shift?
+Despite empirical advances (Zhang et al., 2021; Zhan et al., 2024), the theoretical understanding of SGN domain adaptation remains limited.
+Without a principled framework to quantify generalization under distribution shift, it is difficult to design adaptation methods with guaranteed performance.
+To tackle these challenges, we propose a framework called Degree-Conscious Spiking Graph for Cross-Domain Adaptation (DeSGraDA).
+This has three components: (i) degree-conscious spiking representation, which assigns variable firing thresholds to nodes based on their degrees, enabling adaptive control over spiking sensitivity.
+This degree-conscious mechanism balances the firing frequencies of high- and low-degree nodes, preventing information loss in sparsely connected nodes and avoiding distortion from excessive signals in highly connected ones, thus enhancing the expressiveness of the spiking representations; (ii) temporal distribution alignment, which explicitly aligns the time-evolving spiking representations between the source and target domains.
+By leveraging membrane potential dynamics as evolving signals, the model captures domain-specific patterns, improving robustness to temporal shifts; and (iii) pseudo-label distillation assigns reliable pseudo-labels by aligning consistent predictions from shallow and deep network layers.
+We also demonstrate that this pseudo-label distillation module can effectively reduce a generalization bound tailored for spiking graph domain adaptation.
+In summary, DeSGraDA provides a simple yet effective solution to a novel and underexplored problem, and offers deep insights into the spiking graph domain adaptation.
+Our contributions can be summarized as follows: (1) Problem Formulation: We first introduce the spiking graph domain adaptation for graph classification, highlighting the challenges posed by the inflexible threshold mechanism of SGNs and theoretical limitations that hinder effective adaptation.
+(2) Novel Architecture and Theoretical Analysis: We propose DeSGraDA, a framework combining degree-conscious spiking representation and temporal distribution alignment.
+Moreover, we provide a generalization bound for spiking graph domain adaptation.
+(3) Extensive Experiments.
+We evaluate the proposed DeSGraDA on extensive spiking graph domain adaptation learning datasets, demonstrating that it can outperform various state-of-the-art methods.
 2 Related Work
-Figure 1: An overview of the proposed DeSGDA.
-To achieve personalized spiking representations, DeSGDA employs adaptive thresholds based on node degrees, enabling the generation of tailored spiking representations.
-To align domain distributions, DeSGDA leverages adversarial training on membrane potentials to counter domain discrimination.
-Furthermore, DeSGDA utilizes pseudo-labeling to identify and select reliable samples, thereby enhancing overall model performance.
+Domain Adaptation (DA).
+DA transfers knowledge from a labeled source domain to an unlabeled target domain by mitigating the distributional shift between the two domains (Redko et al., 2017; Long et al., 2018; Shen et al., 2018).
+It has been widely applied to vision and language tasks (Wei et al., 2021b, a; Shi et al., 2024).
+Recently, DA has been extended to graph data to address the unique challenges posed by complex relationships, leading to the emergence of Graph Domain Adaptation (GDA) (Yin et al., 2023; Liu et al., 2024a; Cai et al., 2024).
+Most existing GDA approaches first leverage GNNs to generate node and graph representations (Zhu et al., 2021; Yin et al., 2022; Liu et al., 2024b), followed by adversarial learning to implicitly align feature distributions and reduce domain discrepancies.
+They also apply structure-aware strategies to explicitly align graph-level semantics and topological structures, thereby enhancing generalization across diverse graph domains (You et al., 2023; Liu et al., 2024b; Luo et al., 2024).
+However, GDA remains under-explored in the context of spiking graphs, where energy efficiency becomes a critical requirement for real-world applications.
+To bridge this gap, we introduce the novel problem of Spiking Graph Domain Adaptation (SGDA), extending GDA to spiking graphs for energy-efficient domain adaptation.
 Spiking Graph Networks (SGNs).
-SGNs are a specialized type of neural network that combines Spiking Neural Networks (SNNs) with Graph Neural Networks (GNNs), preserving energy efficiency while achieving competitive performance in various graph tasks Li et al.
-(2023); Yao et al.
-(2023a, b); Duan et al.
-(2024); Liang et al.
-(2023).
-Existing research on SGNs focuses on capturing the dynamic temporal information within graphs and enhancing scalability.
+SGNs are a specialized neural network combining SNNs (Gerstner and Kistler, 2002; Maass, 1997) with GNNs, preserving energy efficiency while achieving competitive performance on various tasks (Li et al., 2023; Yao et al., 2023a; Duan et al., 2024).
+Existing research on SGNs focuses on capturing temporal information within graphs and enhancing scalability.
 For instance, Xu et al.
-(2021a) utilizes spatial-temporal feature normalization within SNNs to effectively process dynamic graph data, ensuring robust learning and improved predictive performance.
+(2021a) utilizes spatial-temporal feature normalization within SNNs to effectively process dynamic graph data, ensuring robust learning and improved performance.
 Zhao et al.
-(2024) propose a method that dynamically adapts to evolving graph structures and relationships through a novel architecture that updates node representations in real time.
+(2024) proposes a method that adapts to evolving graph structures through a novel architecture that updates node representations in real time.
 Additionally, Yin et al.
-(2024a) adapts SNNs to dynamic graph settings and employs implicit differentiation for the node classification task.
-However, existing methods still suffer from data distribution shift issues when training and testing data come from different domains, resulting in degraded performance and generalization.
-To address this, we propose a novel domain adaptation method based on SGNs to tackle these challenges.
-Spiking Transfer Learning.
-Spiking transfer learning focuses on adjusting SNNs to handle data distribution shifts across various domains more effectively.
-Recent advances in spiking transfer learning have been extensively applied in vision tasks, significantly enhancing model performance while maintaining energy efficiency Zhan et al.
-(2021); Zhang et al.
-(2021); Zhan et al.
-(2024).
-For instance, Guo et al.
-(2024) leverages a Jaccard attention mechanism within SNNs to adapt to target domains without requiring source domain data.
-Similarly, He et al.
-(2024) facilitates the transfer of learned representations from static to dynamic event-based domains by adapting SNNs to process temporal information.
-Additionally, Zhan et al.
-(2024) converts RGB images into spike-based neuromorphic data, enabling SNNs to process visual information across various domains efficiently.
-However, the inherent difficulty of complex graph topologies makes it infeasible to apply spiking transfer learning to SGNs directly.
-To this end, we introduce a specialized domain adaptation method specifically tailored for SGNs.
+(2024) adapts SNNs to dynamic graph settings and employs implicit differentiation for the node classification task.
+However, existing methods still suffer from data distribution shift issues when the training and test data come from different domains, resulting in degraded performance and generalization.
+To tackle these challenges, we propose a novel domain adaptation method for spiking graph networks.
 3 Preliminaries
-Bound for Graph Domain Adaptation (GDA).
-Applying GDA with optimal transport (OT), if the covariate shift holds on representations that $\mathbb{P}_{S}(Y|Z)=\mathbb{P}_{T}(Y|Z)$, the target risk $\epsilon_{T}(h,\hat{h})$ is bounded with the theorem:
+Problem Setup.
+Given a graph $G=(V,E,\mathbf{X})$ with node set $V$, edge set $E$, and node attribute matrix $\mathbf{X}$.
+To construct spiking-compatible inputs for SGNs, we sample the binary features $S$ from the Bernoulli distribution with probability of $\mathbf{X}$ (i.e., $S\sim Ber(\mathbf{X})$) as input of SGNs (Zhao et al., 2024).
+In this paper, we focus on the problem of spiking GDA for graph classification.
+The source domain $\mathcal{D}^{s}=\{(G_{i}^{s},y_{i}^{s})\}_{i=1}^{N_{s}}$ is labeled, where $N^{s}$ is the number of source-domain graphs $\{G_{i}^{s}\}$ and $y_{i}^{s}$ is the label of $G_{i}^{s}$.
+The target domain $\mathcal{D}^{t}=\{G_{j}^{t}\}_{j=1}^{N_{t}}$ is unlabeled and contains $N^{t}$ graphs.
+Both domains share the same label space $\mathcal{Y}$ but can have different graph topologies or attribute distributions.
+Domain Adaptation with Optimal Transport (OT).
+Following You et al.
+(2023), we factorize a trained model $h$ as $g\circ f$, where $f:\mathcal{D}\mapsto\mathbb{R}^{d}$ is the feature extractor ($Z=f(\mathcal{D})$) and $g:\mathbb{R}^{d}\mapsto\mathcal{Y}$ is the discriminator ($Y=g(Z)$).
+For simplicity, we focus on binary classification where $\mathcal{Y}=[0,1]$.
+Denote the classifier that predicts labels from the feature representation as $\hat{g}:\mathbb{R}^{d}\mapsto\mathcal{Y}$.
+The (empirical) source and target risks are given by $\hat{\epsilon}_{S}(g,\hat{g})=\frac{1}{N_{s}}\sum_{n=1}^{N_{s}}|g(Z_{n})-\hat{g}(Z_{n})|$ and $\epsilon_{T}(g,\hat{g})=\mathbb{E}_{\mathbb{P}_{T}(Z)}|g(Z)-\hat{g}(Z)|$, respectively.
+DA with optimal transport (OT) (Villani et al., 2008) addresses covariate shift by finding the optimal way to transport masses between the two source and target distributions while minimizing the transport cost.
+The following Theorem shows that the generalization gap depends on both the domain divergence $2C_{g}W_{1}(\mathbb{P}_{S}(Z),\mathbb{P}_{T}(Z))$ and model discriminability $\omega$.
 Theorem 1
+(Redko et al., 2017; Shen et al., 2018) Assume that the discriminator $g$ is $C_{g}$-Lipschitz.
+[^1] The Lipschitz norm $\|g\|_{Lip}=\max_{Z_{1},Z_{2}}\frac{|g(Z_{1})-g(Z_{2})|}{\rho(Z_{1},Z_{2})}=C_{g}$ for some given distance function $\rho$.
+Let :=$\mathcal{H}:=\{g:\mathcal{Z}\rightarrow\mathcal{Y}\}$ (where $\mathcal{Z}$ is the feature space) be the set of bounded real-valued functions with pseudo-dimension $Pdim(\mathcal{H})=d$.
+For any $g\in\mathcal{H}$, the following holds with probability at least $1-\delta$:
+$\epsilon_{T}(g,\hat{g})\leq$ $\hat{\epsilon}_{S}(g,\hat{g})+\sqrt{\frac{4d}{N_{S}}\log\left(\frac{eN_{S}}{d}\right)+\frac{1}{N_{S}}\log\left(\frac{1}{\delta}\right)}+2C_{g}W_{1}(\mathbb{P}_{S}(Z),\mathbb{P}_{T}(Z))+\omega,$
+where $\omega=\min_{\|g\|_{Lip}\leq C_{g}}\{\epsilon_{S}(g,\hat{g})+\epsilon_{T}(g,\hat{g})\}$ is the discriminative ability to capture source and target data, and $W_{1}(\mathbb{P},\mathbb{Q})$ is the distribution divergence defined in (Redko et al., 2017; Shen et al., 2018).
+Applying OT-based DA methods to graph data introduces additional challenges due to the non-Euclidean nature of graphs and intricate dependencies between nodes.
 You et al.
-(2023) Assuming that the learned discriminator is $C_{g}$-Lipschitz continuous as described in Redko et al.
-(2017), and the graph feature extractor $f$ (also referred to as GNN) is $C_{f}$-Lipschitz that $||f||_{Lip}=\max_{G_{1},G_{2}}\frac{||f(G_{1})-f(G_{2})||_{2}}{\eta(G_{1},G_{2})}=C_{f}$ for some graph distance measure $\eta$.
-Let :=$\mathcal{H}:=\{h:\mathcal{G}\rightarrow\mathcal{Y}\}$ be the set of bounded real-valued functions with the pseudo-dimension $Pdim(\mathcal{H})=d$ that $h=g\circ f\in\mathcal{H}$, with probability at least $1-\delta$ the following inequality holds:
-$\epsilon_{T}(h,\hat{h})\leq$ $\ \hat{\epsilon}_{S}(h,\hat{h})+\sqrt{\frac{4d}{N_{S}}\log\left(\frac{eN_{S}}{d}\right)+\frac{1}{N_{S}}\log\left(\frac{1}{\delta}\right)}$
-$+2C_{f}C_{g}W_{1}(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G))+\omega,$
-where the (empirical) source and target risks are $\hat{\epsilon}_{S}(h,\hat{h})=\frac{1}{N_{S}}\sum_{n=1}^{N_{S}}|h(G_{n})-\hat{h}(G_{n})|$ and $\epsilon_{T}(h,\hat{h})=\mathbb{E}_{\mathbb{P}_{T}(G}\{|h(G)-\hat{h}(G)|\}$, respectively, where $\hat{h}:\mathcal{G}\rightarrow\mathcal{Y}$ is the labeling function for graphs and $\omega=\min_{||g||_{Lip}\leq C_{g},||f||_{Lip}\leq C_{f}}\{\epsilon_{S}(h,\hat{h})+\epsilon_{T}(h,\hat{h})\}$.
-The first Wasserstein distance is defined as Villani et al.
-(2009): $W_{1}(\mathbb{P},\mathbb{Q})=\sup_{||g||_{Lip}\leq 1}\left\{\mathbb{E}_{\mathbb{P}_{S}(Z)}g(Z)-\mathbb{E}_{\mathbb{P}_{T}(Z)}g(Z)\right\}$.
-The comprehensive justification of the OT-based graph domain adaptation bound demonstrates that the generalization gap relies on both the domain divergence $2C_{f}C_{g}W_{1}(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G))$ and model discriminability $\omega$.
-Spiking Graph Networks.
-In contrast to traditional artificial neural networks, SGNs Xu et al.
-(2021a); Zhu et al.
-(2022) convert input data into binary spikes over time, with each neuron in the SGNs maintaining a membrane potential that accumulates input spikes.
-A spike is produced as an output when the membrane potential exceeds a threshold, which is formulated as:
-$u_{\tau+1,i}=\lambda(u_{\tau,i}-V_{th}s_{\tau,i})+\sum_{j}w_{ij}\mathcal{A}\left(A,s_{\tau,j}\right)+b,\vspace{-4pt}$
-$s_{\tau+1,i}=\mathbb{H}(u_{\tau+1,i}-V_{th}),$ (1)
-where $\mathbb{H}(x)$ is the Heaviside function, which is the non-differentiable spiking function, $\mathcal{A}$ is the graph aggregation operation, and $A$ is the adjacency matrix of a graph.
-$s_{\tau,i}$ denotes the binary spike train of neuron $i$, and $\lambda$ is the constant.
-$w_{ij}$ and $b$ are the weights and bias of each neuron.
-4 Methodology
-This work studies the spiking graph domain adaptation problem and proposes a new approach DeSGDA.
-DeSGDA consists of three parts: Degree-aware personalized spiking representation utilizes different thresholds for different degrees, effectively addressing the inflexible architecture challenge; Adversarial distribution alignment uses the adversarial training on membrane potential against a domain discriminatory to align distribution between different domains, and Pseudo-label distillation further applies the pseudo-label to enhance model performance.
-We provide the theoretical guarantee of DeSGDA to ensure the effectiveness.
-The overview of DeSGDA is shown in Figure 1.
-Problem Setup.
-Given a graph $G=(V,E,\mathbf{X})$ with the node set $V$, the edge set $E$, and the node attribute matrix $\mathbf{X}$.
-Denote $S$ as the binary input sampled from Bernoulli distribution with probability of $\mathbf{X}$.
-The labeled source domain is denoted as $\mathcal{D}^{s}=\{(G_{i}^{s},y_{i}^{s})\}_{i=1}^{N_{s}}$, where $y_{i}^{s}$ denotes the labels of $G_{i}^{s}$; the unlabeled target domain is $\mathcal{D}^{t}=\{G_{j}^{t}\}_{j=1}^{N_{t}}$, where $N^{s}$ and $N^{t}$ denote the number of source graphs and target graphs.
-Both domains share the same label space $\mathcal{Y}$ but have different distributions in the graph space.
-We aim to train a spiking graph model using labeled source graphs and unlabeled target graphs to achieve superior performance in the target domain.
-4.1 Degree-aware Personalized Spiking Representation
-In this part, we first study the disadvantages of directly applying SNNs to graphs and then propose the degree-aware personalized spiking representation.
-Existing SGNs Li et al.
-(2023); Yao et al.
-(2023a); Duan et al.
-(2024) usually employ a global threshold for membrane potential firing.
-However, the global threshold can lead to the inflexible architecture issue since nodes with higher degrees are more likely to trigger spikes than those with lower degrees.
-As shown in Eq.
-1, nodes with higher degrees have more neighbors, and the aggregation operation allows for more significant feature accumulation, making it easier for these nodes to trigger spikes compared to those with fewer neighbors.
-To alleviate this issue, we propose the degree-aware thresholds and iteratively update their values.
-Specifically, we first $set$ all the degrees of nodes in the source domain graphs, i.e., $D^{s}=set(D_{1}^{s}\cup\cdots\cup D_{N_{s}}^{s})$, where $D_{i}$ denotes the degree set of graph $G_{i}^{s}$, and $set(\cdot)$ operation is an unordered sequence of non-repeating elements.
-Considering that low-degree nodes are more challenging to trigger while high-degree nodes trigger more easily, we propose setting higher thresholds for high-degree nodes and lower for low-degree nodes, which is formulated as:
-$s_{\tau}^{d_{i}^{s}}=\mathbb{H}(u_{\tau}$ $-V_{th}^{d_{i}^{s}}),\quad S^{d_{i}^{s}}=avg(s_{\tau}^{d_{i}^{s}}),$ (2)
-$V_{th}^{d_{i}^{s}}$ $=(1-\alpha)V_{th}^{d_{i}^{s}}+\alpha S^{d_{i}^{s}},$
-where $V_{th}^{d_{i}^{s}}$ is the threshold of degree $d_{i}^{s}\in D^{s}$, initially set to $V_{th}$, and $\alpha$ is a hyper-parameter.
-The $avg(\cdot)$ operation takes the average of spiking representation with degree $d_{i}^{s}$.
-Consequently, high-degree nodes tend to achieve high $S^{d_{i}^{s}}$, which leads to an iterative increase in the threshold corresponding to degree $d_{i}^{s}$ and conversely for lower-degree nodes.
-To further explore the background mechanism of the hypothesis, we have the following analysis.
-Hypothesis 1
-In graph spiking networks, nodes with low-degree are more challenging to trigger while high-degree nodes trigger more easily.
-The details analysis are introduced in Appendix A.
-With different thresholds for different node degrees, we can obtain the personalized node spiking representation $\mathbf{s}_{v\in G_{i}^{s}}^{d_{j}^{s}}$.
-Then, we summarize all node representations with a readout function into the graph-level representation and output the prediction with a multi-layer perception (MLP) classifier:
-$\mathbf{s}_{i}=$ $\operatorname{READOUT}\left(\left\{\mathbf{s}_{v}^{d_{j}^{s}}\right\}_{v\in G_{i}^{s}}\right),$ (3)
-$\hat{y}_{i}^{s}=H(\mathbf{s}_{i}),$
-where $\hat{y}_{i}$ is the predicted result and $H(\cdot)$ is the classifier. After that, the source classification loss is:
-$\mathcal{L}_{S}=\mathbb{E}_{G_{i}^{s}\in\mathcal{D}^{s}}{l}(y_{i}^{s},\hat{y}_{i}^{s}),$ (4)
-where $l(\cdot)$ is the loss function and $y_{i}^{s}$ is the ground truth of the $i$-th graph $G_{i}^{s}$ in the source domain.
-However, in the scenario of domain adaptation, two significant issues remain unresolved.
-The first issue is that degrees in the target domain may be unseen in the source, rendering the thresholds ineffective for these degrees.
-The second is that the thresholds in the target may differ from those in the source, simply applying the source domain’s thresholds could lead to performance degradation.
-To alleviate the first issue, we initialize the threshold $V_{th}^{d_{i}^{t}}$ with the same value, where $d_{i}^{t}\notin D^{s}$.
-Then, with the training process of adversarial alignment, we iteratively update the threshold for degree $d_{i}^{t}$ with Eq.
-2.
-To address the second issue, we incorporate the pseudo-label distillation module in Section 4.3 to guide the update of source degree thresholds on the target domain.
-4.2 Adversarial Distribution Alignment
-To eliminate the discrepancy between the source and target domains, we propose the adversarial distribution alignment module.
-Specifically, for each source graph $G_{i}^{s}$ and target graph $G_{i}^{t}$, we use the degree-aware personalized spiking GNNs-based encoder $F(\cdot)$ and semantic classifier $H(\cdot)$ to produce predicted labels.
-Then, a domain discriminator $Q(\cdot)$ is trained to distinguish features from the source and target domains.
-The encoder and classifier are adversarial trained to align the feature spaces of the source and target domains.
-$\mathcal{L}_{AD}$ $=\mathbb{E}_{G_{i}^{s}\in\mathcal{D}^{s}}\log Q\left(F(G_{i}^{s}),H(G_{i}^{s})|V_{th}^{D^{s}}\right)$
-$+\mathbb{E}_{G_{j}^{t}\in\mathcal{D}^{t}}\log\left(1-Q\left(F\left(G_{j}^{t}\right),H(G_{j}^{t})|V_{th}^{D^{s}}\right)\right).$
-However, the degree in the target domain may be unseen by the source. Thus, we further initialize the threshold with $V_{th}^{d_{j}^{t}}$ and $d_{j}^{t}\notin D^{s}$, which is formulated as:
-$\mathcal{L}_{AD}$ $=\mathbb{E}_{\begin{subarray}{c}G_{j}^{t}\in\mathcal{D}^{t}\\ D_{j}^{t}\subset{D^{s}}\end{subarray}}\log\left(1-Q\left(F\left(G_{j}^{t}\right),H(G_{j}^{t})|V_{th}^{D^{s}}\right)\right)$
-$+\mathbb{E}_{G_{i}^{s}\in\mathcal{D}^{s}}\log Q\left(F(G_{i}^{s}),H(G_{i}^{s})|V_{th}^{D^{s}}\right)$
-$+\mathbb{E}_{\begin{subarray}{c}G_{j}^{t}\in\mathcal{D}^{t}\\ \exists d_{j}^{t}\notin{D^{s}}\end{subarray}}\log\left(1-Q\left(F\left(G_{j}^{t}\right),H(G_{j}^{t})|V_{th}^{D^{s}},V_{th}^{D^{t}}\right)\right),$
-where $D^{t}=\{d_{i}^{t}|d_{i}^{t}\in D^{t},d_{i}^{t}\notin D^{s}\}$.
-We iteratively update $V_{th}^{D^{t}}$ with Eq.
+(2023) extends the OT-based DA framework to graphs, and provides a generalization bound for GDA.
+Details are in Appendix A.
+Spiking Graph Networks (SGNs).
+Spiking Neural Networks (SNNs) are brain-inspired models that communicate through discrete spikes rather than continuous-valued activations (Maass, 1997; Gerstner and Kistler, 2002; Bohte et al., 2000).
+This provides significant advantages in temporal information processing and energy efficiency.
+More details are in Appendix B.
+SGNs are a specialized form of SNNs tailored for graph data (Xu et al., 2021a; Zhu et al., 2022), where each node is modeled as a spiking neuron.
+The membrane of each node evolves based on both the temporal spiking dynamics and the graph’s structural connectivity.
+Let $u_{\tau,i}$ be the membrane potential of node $i$ at latency step $\tau$.
+SGNs first updates the membrane potentials via input aggregation:
+$u_{\tau+1,i}=\lambda(u_{\tau,i}-V_{th}s_{\tau,i})+\sum\nolimits_{j}w_{ij}\mathcal{A}(A,s_{\tau,j})+b_{i},$ (1)
+where $s_{\tau,i}$ is the spiking representation, $\lambda\in(0,1)$ is the leak factor, $V_{th}$ is the firing threshold, $b_{i}$ is the bias, $w_{ij}$ is the synaptic weight from node $j$ to node $i$, $A$ is the adjacency matrix, and $\mathcal{A}$ is the graph-based aggregation operation.
+Next, spikes are triggered through thresholding:
+$s_{\tau+1,i}=\mathbb{H}(u_{\tau+1,i}-V_{th}),\quad u_{\tau+1,i}=(1-s_{\tau+1,i})u_{\tau+1,i}+s_{\tau+1,i}V_{reset},$ (2)
+where $V_{\text{reset }}$ is the reset potential, and $\mathbb{H}(x)$ is the Heaviside step function which serves as the non-differentiable spiking function. Finally, the neuron is reset upon firing.
+4 Proposed Methodology
+Figure 1: Overview of the proposed DeSGraDA.
+Degree-Conscious Spiking Representation generates source and target domain spiking representations by adapting neuron firing threshold based on node degrees.
+To align domain distributions, Temporal Distribution Alignment leverages adversarial training on temporal membrane potential to counter domain discrimination.
+Furthermore, Pseudo-labe Distillation is employed to identify reliable target samples and reinforce overall model performance.
+This paper proposes a novel framework DeSGraDA for the problem of spiking graph domain adaptation.
+DeSGraDA consists of three parts: Degree-Conscious Spiking Representation (Section 4.1), which assigns adaptive firing thresholds based on node degrees to address the limitations of rigid, fixed-threshold architectures; Temporal Distribution Alignment (Section 4.2), which employs adversarial training on temporal membrane potentials against a domain discriminator to align spiking dynamics across domains; and Pseudo-label Distillation further applies the pseudo-label to enhance model performance.
+We also provide a theoretical generalization bound to support the effectiveness of DeSGraDA.
+An overview of the framework is in Figure 1.
+4.1 Degree-Conscious Spiking Representation
+First, we study the disadvantages of directly applying SNNs to graphs and then propose the degree-conscious spiking representation module.
+Existing SGNs (Li et al., 2023; Duan et al., 2024) usually employ a fixed global threshold ($V_{th}$ in Eq.
+1 and 2) for firing.
+Assume that the membrane potential of node $u_{\tau,i}$ follows the normal distribution $\mathcal{N}(\mu,\sigma^{2})$ (Kipf and Welling, 2016).
+The following Proposition shows that the high-degree nodes are more likely to trigger spikes than the low-degree ones. Proof is in Appendix C.
+Proposition 1
+With aggregation operation in SGNs (i.e., $\mathcal{A}$ in Eq.
+1), the expectation of the updated node membrane potential is: $\mathbb{E}(u_{\tau+1,i})\sim\mathcal{N}\left(\left(1+\sum\nolimits_{j\in N(i)}w_{ij}\right)\mu,\left(1+\sum\nolimits_{j\in N(i)}w_{ij}\right)^{2}\sigma^{2}\right)$, where $N(i)$ is the set of node $i$’s neighbors, and $w_{ij}$ is the weight between nodes $i$ and $j$.
+Figure 2: Correlation between spiking frequency and aggregation weights under fixed thresholds on the PROTEINS and DD datasets.
+From Proposition 1, we observe that node $i$ follows a normal distribution with a mean of $(1+\sum_{j\in N(i)}w_{ij})\mu$, determined by the aggregated weights of its neighboring nodes [^2] This does not account for models that constrain the sum of the neighboring weights (e.g., GAT (Veličković et al., 2018))..
+We conduct an experiment to investigate the relationship between the aggregated weight ($\sum\nolimits_{j\in N(i)}w_{ij}$) and spiking frequency (i.e., count of $u_{T}>V_{th}$) for a fixed threshold $V_{th}$.
+As shown in Figure 2, the spiking frequency and aggregation weights under fixed thresholds exhibit a relatively high correlation coefficient, indicating that nodes with higher degrees tend to accumulate more features from neighbors, making it easier to trigger spikes than those with fewer neighbors.
+Fixed thresholds inherently bias spiking activation toward high-degree nodes, leading to under-representation of structurally important yet sparsely connected nodes and undermining both expressiveness and generalization in SGNs.
+To alleviate this problem, we propose the use of degree-conscious thresholds.
+Specifically, let $B_{i}^{s}$ be the set of degrees for the nodes in graph $G_{i}^{s}$.
+We collect all unique node degrees across the source domain graphs with $B^{s}=set(B_{1}^{s}\cup\cdots\cup B_{N_{s}}^{s})$, where $set(\cdot)$ operation is an unordered sequence of distinct elements.
+With the observation from Proposition 1, we propose setting higher thresholds for high-degree nodes and lower for low-degree nodes.
+For node $v$ with degree $d_{v}^{s}$, we have:
+$\;\;s_{\tau,v}^{d_{v}^{s}}=\mathbb{H}(u_{\tau,v}-V_{th}^{d_{v}^{s}}),\quad V_{th,\tau+1}^{d_{v}^{s}}=(1-\alpha)V_{th,\tau}^{d_{v}^{s}}+\alpha\bar{s}^{d_{v}^{s}}_{\tau,v},$ (3)
+$u_{\tau+1,v}=\lambda(u_{\tau,v}-V_{th}s_{\tau,v}^{d_{v}^{s}})+\sum\nolimits_{j\in N(v)}w_{ij}\mathcal{A}\left(A,s_{\tau,v}^{d_{j}^{s}}\right)+b_{i},$
+where $V_{th}^{d_{v}^{s}}$ is the threshold for nodes with degree $d_{v}^{s}\in B^{s}$, initially set to $V_{th}$, and $\alpha$ is a hyper-parameter.
+$\bar{s}^{d_{v}^{s}}_{\tau}$ is the average of spiking representation $s_{\tau}^{d_{v}^{s}}$ with degree $d_{v}^{s}$ on latency step $\tau$.
+In Eq.
+3, we dynamically update the threshold $V^{d^{s}_{v}}_{th,\tau+1}$ with $(1-\alpha)V^{d^{s}_{v}}_{th,\tau}+\alpha\bar{s}^{d^{s}_{v}}_{\tau,v}$ based on the average spiking frequency $\bar{s}^{d^{s}_{v}}_{\tau,v}$ of nodes with degree $d^{s}_{v}$.
+Consequently, high-degree nodes typically exhibit higher $\bar{s}^{d^{s}_{v}}_{\tau,v}$, resulting in increased thresholds $V^{d^{s}_{v}}_{th,\tau+1}$ to suppress over-activation, while low-degree nodes yield lower thresholds to enhance activation, enabling adaptive spiking control across structurally diverse nodes.
+For each node $v$ in graph $G_{i}^{s}$, we calculate the membrane potential $u_{\tau,v}^{s}$, and summarize all the node representations with a readout function (Xu et al., 2018; Wang et al., 2024a) into the graph-level representation:
+$\mathbf{s}_{G_{i}^{s}}=\operatorname{READOUT}\left(\left\{{s}_{T}^{d_{v}^{s}}\right\}_{v\in G_{i}^{s}}\right),$ (4)
+where $T$ is the total number of latency steps.
+Finally, we output the prediction $\hat{y}_{G_{i}^{s}}^{s}=H(\mathbf{s}_{G_{i}^{s}})$ with a classifier $H(\cdot)$ by minimizing the source classification loss $\mathcal{L}_{S}=\mathbb{E}_{G_{i}^{s}\in\mathcal{D}^{s}}{l}(y_{G_{i}^{s}}^{s},\hat{y}_{G_{i}^{s}}^{s})$, where $l(\cdot)$ is the loss function and $y_{i}^{s}$ is the ground truth of graph $G_{i}^{s}$ in the source domain.
+4.2 Temporal Distribution Alignment
+Unlike GNNs, spiking models rely on membrane potential dynamics to generate sparse spike trains, making their spike representations highly dynamic and non-differentiable.
+Existing GDA (Yin et al., 2023; Chen et al., 2025) methods assume continuous, static embeddings and thus fail to align the time-dependent neural dynamics in SGNs.
+To address this, we propose a temporal-based alignment framework that captures and matches the evolution of spiking patterns across domains.
+First, we introduce a temporal attention mechanism that adaptively aggregates time-dependent neuronal states.
+Specifically, given the temporal membrane potential of graphs on each latency step, we have $[\mathbf{u}_{1,G_{i}},\ldots,\mathbf{u}_{\tau,G_{i}},\cdots,\mathbf{u}_{T,G_{i}}]$, where $\mathbf{u}_{\tau,G_{i}}=\operatorname{READOUT}(\left\{{u}_{\tau,v}\right\}_{v\in G_{i}})$ and $G_{i}\in\{\mathcal{D}^{s},\mathcal{D}^{t}\}$.
+We stack the $T$ steps membrane potential into $\mathbf{U}_{G_{i}}\in\mathbb{R}^{T\times d}$, where $d$ is the hidden dimension.
+The goal is to learn an importance-weighted $\alpha_{\tau}$ to summarize of temporal membrane potential representation, which is formulated as follows:
+$\alpha_{\tau}=\operatorname{Atten}(\mathbf{U}_{G_{i}}),\quad\tilde{\mathbf{U}}_{G_{i}}=\sum\nolimits_{\tau=1}^{T}\alpha_{\tau}\mathbf{u}_{\tau,G_{i}},$ (5)
+where $\operatorname{Atten}(\cdot)$ is the self-attention operation (Vaswani et al., 2017).
+Then, we propose the adversarial distribution alignment module to eliminate the discrepancy between the source and target domains.
+Specifically, for each source graph $G_{i}^{s}$ and target graph $G_{j}^{t}$, we denote the semantic classifier as $H(\cdot)$ to produce predicted labels, and a domain discriminator $Q(\cdot)$ to distinguish features from the source and target domains.
+The temporal-based distribution alignment module is adversarially trained to align the feature spaces of the source and target domains, which is formulated as:
+$\mathcal{L}_{AD}=\mathbb{E}_{G_{i}^{s}\in\mathcal{D}^{s}}\log Q\left(H(\tilde{\mathbf{U}}_{G_{i}}^{s})|V_{th}^{B^{s}}\right)+\mathbb{E}_{\begin{subarray}{c}G_{j}^{t}\in\mathcal{D}^{t},\exists d_{j}^{t}\notin{B^{s}}\end{subarray}}\log\left(1-Q\left(H(\tilde{\mathbf{U}}_{G_{i}}^{t})|V_{th}^{B^{s}},V_{th}^{B^{t}}\right)\right),$
+where $B^{t}=\{d_{i}^{t}|d_{i}^{t}\in B^{t},d_{i}^{t}\notin B^{s}\}$.
+We iteratively update $V_{th}^{B^{t}}$ with Eq.
 3 on each latency.
-Furthermore, we present an upper bound for adversarial distribution alignment.
+Furthermore, we present an upper bound for temporal-based distribution alignment.
 Theorem 2
-Assuming that the learned discriminator is $C_{g}$-Lipschitz continuous as described in Theorem 1, the graph feature extractor $f$ (also referred to as GNN) is $C_{f}$-Lipschitz that $||f||_{Lip}=\max_{G_{1},G_{2}}\frac{||f(G_{1})-f(G_{2})||_{2}}{\eta(G_{1},G_{2})}=C_{f}$ for some graph distance measure $\eta$ and the loss function bounded by $C>0$.
-Let :=$\mathcal{H}:=\{h:\mathcal{G}\rightarrow\mathcal{Y}\}$ be the set of bounded real-valued functions with the pseudo-dimension $Pdim(\mathcal{H})=d$ that $h=g\circ f\in\mathcal{H}$, and provided the spike training data set $S_{n}=\{(\mathbf{X}_{i},y_{i})\in\mathcal{X}\times\mathcal{Y}\}_{i\in[n]}$ drawn from $\mathcal{D}^{s}$, with probability at least $1-\delta$ the following inequality :
-$\epsilon_{T}(h,$ $\hat{h}_{T}(\mathbf{X}))\leq\hat{\epsilon}_{S}(h,\hat{h}_{S}(\mathbf{S}))+2\mathbb{E}\left[\sup\frac{1}{N_{S}}\sum_{i=1}^{N_{S}}\epsilon_{i}h(\mathbf{X}_{i},y_{i},p_{i})\right]$
-$+C\sqrt{\frac{ln(2/\delta)}{N_{S}}}+\min\Big{(}|\epsilon_{S}(h,\hat{h}_{S}(\mathbf{X}))-\epsilon_{S}(h,\hat{h}_{T}(\mathbf{X}))|,$
-$|\epsilon_{T}(h,\hat{h}_{S}(\mathbf{X}))-\epsilon_{T}(h,\hat{h}_{T}(\mathbf{X}))|\Big{)}+2C_{f}C_{g}W_{1}\left(\mathbb{P}_{S},\mathbb{P}_{T}\right),$
-where the (empirical) source and target risks are $\hat{\epsilon}_{S}(h,\hat{h}(\mathbf{S}))=\frac{1}{N_{S}}\sum_{n=1}^{N_{S}}|h(\mathbf{S}_{n})-\hat{h}(\mathbf{S}_{n})|$ and $\epsilon_{T}(h,\hat{h}(\mathbf{X}))=\mathbb{E}_{\mathbb{P}_{T}(G}\{|h(G)-\hat{h}(G)|\}$, respectively, where $\hat{h}:\mathcal{G}\rightarrow\mathcal{Y}$ is the labeling function for graphs and $\omega=\min_{||g||_{Lip}\leq C_{g},||f||_{Lip}\leq C_{f}}\{\epsilon_{S}(h,\hat{h}(\mathbf{X}))+\epsilon_{T}(h,\hat{h}(\mathbf{X}))\}$, $\epsilon_{i}$ is the Rademacher variable and $p_{i}$ is the $i^{th}$ row of $\mathbf{P}$, which is the probability matrix with:
-$\mathbf{P}_{kt}=\begin{array}[]{lr}\exp\left(\frac{u_{k}(t)-V_{th}}{\sigma(u_{k}(t)-u_{reset})}\right),&if\quad u_{\theta}\leq u(t)\leq V_{th},\\ 0,&if\quad u_{reset}\leq u_{k}(t)\leq u_{\theta}.\end{array}$ (5)
-Theorem 2 proves the generalization bound of spiking graph domain adaptation. More details can be found in Appendix B.
+Assume that the learned discriminator is $C_{g}$-Lipschitz continuous, the feature extractor $f$ is $C_{f}$-Lipschitz that $||f||_{Lip}=\max_{G_{1},G_{2}}\frac{||f(G_{1})-f(G_{2})||_{2}}{\eta(G_{1},G_{2})}=C_{f}$ for some graph distance measure $\eta$ and the loss function bounded by $C>0$.
+Let :=$\mathcal{H}:=\{h:\mathcal{G}\rightarrow\mathcal{Y}\}$ be the set of bounded functions with the pseudo-dimension $Pdim(\mathcal{H})=d$ that $h=g\circ f\in\mathcal{H}$, and provided the spike training data set $S_{n}=\{(\mathbf{X}_{i},y_{i})\in\mathcal{X}\times\mathcal{Y}\}_{i\in[n]}$ drawn from $\mathcal{D}^{s}$, with probability at least $1-\delta$ the inequality :
+$\epsilon_{T}(h,\hat{h}_{T}(\mathbf{X}))\leq$ $\hat{\epsilon}_{S}(h,\hat{h}_{S}(\mathbf{S}))+2\mathbb{E}\left[\sup\frac{1}{N_{S}}\sum_{i=1}^{N_{S}}\epsilon_{i}h(\mathbf{X}_{i},y_{i},p_{i})\right]+C\sqrt{\frac{ln(2/\delta)}{N_{S}}}$ (6)
+$+\omega+2C_{f}C_{g}W_{1}\left(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G)\right),$
+[…]
+The proof is proposed in Appendix D.
+Theorem 2 justifies that the generalization gap of spiking GDA relies on the domain divergence $2C_{f}C_{g}W_{1}(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G))$ and model discriminability $\omega$, as well as the model’s ability to avoid overfitting to the training data, which is quantified by the Rademacher complexity term $2\mathbb{E}\left[\sup\frac{1}{N_{S}}\sum_{i=1}^{N_{S}}\epsilon_{i}h(\mathbf{X}_{i},y_{i},p_{i})\right]$.
+In the application of spiking GDA, this term captures how the model’s sensitivity to random fluctuations in the node feature aggregation (especially for higher-degree nodes) can lead to overfitting, thus affecting the model’s ability to generalize to the target domain.
+This overfitting risk is particularly relevant when the model is too flexible in fitting the training data, exacerbating the generalization gap, especially under domain shifts.
 4.3 Pseudo-label Distillation for Discrimination Learning
-To further address the variance in thresholds between the target and source domains, we incorporate the pseudo-label distillation module into the DeSGDA framework.
-With reliable pseudo-labels, we can effectively update the source degree thresholds in the target domain.
-The goal of the pseudo-label distilling procedure is to keep those examples and their corresponding pseudo-labels from the deep feature space that aligns with the shallow feature space.
-Specifically, we denote $\mathbf{s^{\prime}}_{i}^{t}$ as the shallow spiking graph representation on the $L^{\prime}$-th layer, where $L^{\prime}<L$, and $\hat{y}_{i}^{t}$ as the prediction of graph $G_{i}^{t}$ on the $L$-th layer.
-Then, to enhance alignment between the shallow and deep feature spaces and facilitate the generation of more accurate pseudo-labels, we cluster the shallow features $\mathbf{s^{\prime}}^{t}$ into $C$ clusters and each cluster $\mathcal{E}_{j}$ includes graphs $\{G_{j}^{t}\}$.
-After that, we find the dominating labels $e_{r}$ in the cluster, i.e., $\max_{r}|\{\mathcal{E}_{r}:e_{r}=\hat{y}_{j}^{t}\}|$, and remove other instances with the same pseudo-label but in different clusters.
-Formally, the pseudo-labels are signed with:
-$\mathcal{P}=\left\{\left({G}_{j}^{t},\hat{y}_{j}^{t}\right):e_{j}=\max_{r}\left|\left\{\mathcal{E}_{r}:e_{r}=\hat{y}_{j}^{t}\right\}\right|\right\}.$ (6)
+To further tiny the generalization gap between the target and source domains, we incorporate the pseudo-label distillation module into the DeSGraDA framework.
+The goal of the module is to ensure consistent prediction between the shallow and deep layers.
+Specifically, let $\mathbf{s^{\prime}}_{\tau,G_{i}}^{t}$ be the shallow spiking graph representation of $G_{i}$ on the latency step $\tau$ ($\tau<T$) in the target domain, and $\hat{y}_{G_{i}}^{t}=H(\mathbf{s}_{G_{i}})$ be the prediction of graph $G_{i}$.
+Then, to enhance consistency between the shallow and deep feature spaces and facilitate the generation of more accurate predictions, we cluster the shallow graphs features in the target domain into $C$ clusters, and each cluster $\mathcal{E}_{j}$ includes graphs $\{G_{j}^{t}\}$.
+After that, we find the dominating labels $e_{r}$ in the cluster, i.e., $\max_{r}|\{\mathcal{E}_{r}:e_{r}=\hat{y}_{G_{j}}^{t}\}|$, and remove other samples with the same prediction but in different clusters.
+Formally, the pseudo-labels are signed as:
+$\mathcal{P}=\left\{\left({G}_{j}^{t},\hat{y}_{G_{j}}^{t}\right):e_{j}=\max_{r}\left|\left\{\mathcal{E}_{r}:e_{r}=\hat{y}_{G_{j}}^{t}\right\}\right|\right\}.$ (7)
 Finally, we utilize the distilled pseudo-labels to guide the update of source degree thresholds on the target domain with Eq. 3, and to direct classification in the target domain:
-$\mathcal{L}_{T}=\mathbb{E}_{{G}_{j}^{t}\in\mathcal{P}}l\left({H(\mathbf{s}}_{j}^{t}),\hat{y}_{j}^{t}\right),$ (7)
-where $H(\cdot)$ and $\mathbf{s}_{j}^{t}$ are the classifier and spiking graph representation, which are defined in Eq. 3. $l(\cdot)$ is the loss function, and we implement it with cross-entropy loss.
-Theorem 3
-Under the assumption of Theorem 1, we further assume that there exists a small amount of i.i.d.
-samples with pseudo labels $\{(G_{n},Y_{n})\}_{n=1}^{N^{\prime}_{T}}$ from the target distribution $\mathbb{P}_{T}(G,Y)$ $(N^{\prime}_{T}\ll N_{S})$ and bring in the conditional shift assumption that domains have different labeling function $\hat{h}_{S}\neq\hat{h}_{T}$ and $\max_{G_{1},G_{2}}\frac{|\hat{h}_{D}(G_{1})-\hat{h}_{D}(G_{2})|}{\eta(G_{1},G_{2})}=C_{h}\leq C_{f}C_{g}(D\in\{S,T\})$ for some constant $C_{h}$ and distance measure $\eta$, and the loss function bounded by $C>0$.
-Let :=$\mathcal{H}:=\{h:\mathcal{G}\rightarrow\mathcal{Y}\}$ be the set of bounded real-valued functions with the pseudo-dimension $Pdim(\mathcal{H})=d$, and provided the spike training data set $S_{n}=\{(\mathbf{X}_{i}^{s},y_{i}^{s})\}_{i\in[n]}$, with probability at least $1-\delta$ the following inequality holds:
-$\epsilon_{T}(h,\hat{h}_{T}($ $\mathbf{X}))\leq\frac{N^{\prime}_{T}}{N_{S}+N^{\prime}_{T}}\hat{\epsilon}_{T}(h,\hat{h}_{T}(S))+\frac{N_{S}}{N_{S}+N^{\prime}_{T}}\cdot\Bigg{(}\hat{\epsilon}_{S}(h,\hat{h}_{S}(S))+2C_{f}C_{g}W_{1}\left(\mathbb{P}_{S},\mathbb{P}_{T}\right)$
-$+2\mathbb{E}\left[\sup\frac{1}{N_{S}}\sum_{i=1}^{N_{S}}\epsilon_{i}h(\mathbf{X}_{i},y_{i},p_{i})\right]+C\sqrt{\frac{ln(2/\delta)}{N_{S}}}$
-$+\min\Big{(}|\epsilon_{S}(h,\hat{h}_{S}(\mathbf{X})))-\epsilon_{S}(h,\hat{h}_{T}(\mathbf{X})))|,$
-$|\epsilon_{T}(h,\hat{h}_{S}(\mathbf{X})))-\epsilon_{T}(h,\hat{h}_{T}(\mathbf{X})))|\Big{)}\Bigg{)}\hat{\epsilon}_{S}(h,\hat{h}_{S}(S))+2\mathbb{E}\left[\sup\frac{1}{N_{S}}\sum_{i=1}^{N_{S}}\epsilon_{i}h(\mathbf{X}_{i},y_{i},p_{i})\right]$
-$+C\sqrt{\frac{ln(2/\delta)}{N_{S}}}+2C_{f}C_{g}W_{1}\left(\mathbb{P}_{S},\mathbb{P}_{T}\right)+\omega^{\prime},$
-where the (empirical) source and target risks are $\hat{\epsilon}_{S}(h,\hat{h})=\frac{1}{N_{S}}\sum_{n=1}^{N_{S}}|h(G_{n})-\hat{h}(G_{n})|$ and $\epsilon_{T}(h,\hat{h})=\mathbb{E}_{\mathbb{P}_{T}(G}\{|h(G)-\hat{h}(G)|\}$, respectively, where $\hat{h}:\mathcal{G}\rightarrow\mathcal{Y}$ is the labeling function for graphs and $\omega^{\prime}=\min_{||g||_{Lip}\leq C_{g},||f||_{Lip}\leq C_{f}}\{\epsilon_{S}(h,\hat{h})+\epsilon_{T}(h,\hat{h})\}$, $\epsilon_{i}$ is the Rademacher variable and $p_{i}$ is the $i^{th}$ row of $\mathbf{P}$, which is defined in Theorem 2.
-The proof is detailed in Appendix C.
-From Theorem 3, we observe that the bound of DeSGDA is lower than simply aligning the distributions by incorporating the highly reliable pseudo-labels, demonstrating the effectiveness of pseudo labels for spiking graph domain adaptation.
+$\mathcal{L}_{T}=\mathbb{E}_{{G}_{j}^{t}\in\mathcal{P}}l\left({H(\mathbf{s}}_{G_{j}^{t}}^{t}),\hat{y}_{G_{j}^{t}}^{t}\right),$ (8)
+where $\mathbf{s}_{G_{j}^{t}}^{t}$ is the spiking graph representation of $G_{j}^{t}$ in the target domain.
+$l(\cdot)$ is the loss function, and we implement it with cross-entropy loss.
+We further analyze the generalization bound by applying the pseudo-label distillation module, and the proof is detailed in Appendix E.
+From the proof, we observe that the bound is lower than simply aligning the distributions by incorporating the highly reliable pseudo-labels, demonstrating the effectiveness of pseudo-labels for spiking graph domain adaptation.
 4.4 Learning Framework
-Finally, the overall training objective of DeSGDA integrates classification loss $\mathcal{L}_{S}$, adversarial training loss $\mathcal{L}_{AD}$, and pseudo-label distillation loss $\mathcal{L}_{T}$, which is formulated as:
+Overall, the training objective of DeSGraDA integrates classification loss $\mathcal{L}_{S}$, temporal-based distribution alignment loss $\mathcal{L}_{AD}$, and pseudo-label distillation loss $\mathcal{L}_{T}$, which is formulated as:
 $\mathcal{L}=\mathcal{L}_{S}+\mathcal{L}_{T}-\lambda\mathcal{L}_{AD},$ (9)
-where $\lambda$ is a hyper-parameter to balance the adversarial training loss and classification loss. The learning procedure is illustrated in Algorithm D, and the complexity is shown in Appendix E .
+where $\lambda$ is a hyper-parameter to balance the distribution alignment loss and classification loss.
+The learning procedure is illustrated in Algorithm F, and the complexity is shown in Appendix G.
+[…]
+Methods SEED BCI
+[…]
+WL subtree 40.7 36.8 42.6 35.0 35.6 38.3 47.9 47.7 46.0 46.7 47.7 47.5
+GCN 46.5$\pm$0.6 47.4$\pm$0.9 46.6$\pm$1.2 47.7$\pm$1.4 45.8$\pm$1.2 47.1$\pm$1.6 49.6$\pm$2.5 48.7$\pm$2.8 51.1$\pm$1.0 51.5$\pm$2.0 49.6$\pm$2.4 49.1$\pm$1.7
+GIN 47.4$\pm$1.7 48.0$\pm$1.6 47.3$\pm$1.4 47.5$\pm$1.8 41.6$\pm$2.0 46.1$\pm$1.3 49.4$\pm$2.5 48.4$\pm$2.1 51.8$\pm$1.4 51.2$\pm$2.5 50.0$\pm$1.7 48.7$\pm$2.1
+GMT 46.5$\pm$0.5 47.8$\pm$1.3 47.2$\pm$0.7 47.2$\pm$1.3 46.4$\pm$0.9 46.4$\pm$1.2 48.8$\pm$1.3 47.8$\pm$1.1 49.4$\pm$1.0 48.5$\pm$1.5 50.7$\pm$0.9 51.5$\pm$1.5
+CIN 46.9$\pm$0.5 48.4$\pm$1.1 47.0$\pm$1.5 47.3$\pm$0.7 47.0$\pm$1.6 47.0$\pm$0.9 50.3$\pm$1.6 48.8$\pm$1.5 50.4$\pm$1.3 50.4$\pm$1.2 50.1$\pm$1.5 50.9$\pm$1.7
+SpikeGCN 46.3$\pm$1.0 47.4$\pm$0.8 45.8$\pm$1.2 47.7$\pm$1.1 45.8$\pm$1.5 46.4$\pm$1.2 52.5$\pm$1.6 52.8$\pm$1.3 54.1$\pm$1.9 52.1$\pm$1.0 51.7$\pm$1.8 50.5$\pm$2.3
+DRSGNN 47.1$\pm$1.0 48.5$\pm$0.9 46.5$\pm$1.2 48.1$\pm$1.3 46.9$\pm$0.8 47.6$\pm$1.4 52.7$\pm$1.3 52.8$\pm$1.9 53.8$\pm$1.5 52.7$\pm$1.6 53.0$\pm$2.1 51.3$\pm$1.8
+CDAN 52.6$\pm$1.2 54.5$\pm$0.7 53.9$\pm$0.7 55.9$\pm$1.3 51.6$\pm$1.1 53.6$\pm$0.8 51.9$\pm$1.3 52.6$\pm$1.4 51.8$\pm$1.1 55.4$\pm$1.8 52.5$\pm$1.5 53.1$\pm$1.4
+ToAlign 51.2$\pm$1.3 52.3$\pm$0.8 51.5$\pm$0.9 49.7$\pm$1.5 49.6$\pm$1.1 49.4$\pm$1.3 52.5$\pm$1.7 53.7$\pm$1.5 52.2$\pm$1.4 54.4$\pm$1.2 52.7$\pm$1.0 51.8$\pm$1.3
+MetaAlign 51.2$\pm$1.4 53.7$\pm$0.9 52.2$\pm$1.1 53.8$\pm$0.8 51.2$\pm$1.4 52.0$\pm$1.2 51.1$\pm$1.5 51.8$\pm$1.2 50.4$\pm$1.7 52.5$\pm$1.5 51.7$\pm$1.5 51.3$\pm$1.1
+DEAL 57.4$\pm$1.1 57.5$\pm$1.4 56.6$\pm$0.7 58.1$\pm$1.2 53.9$\pm$0.7 57.8$\pm$1.3 53.7$\pm$1.4 52.5$\pm$2.2 52.6$\pm$1.6 54.5$\pm$1.4 52.7$\pm$1.7 52.8$\pm$1.2
+CoCo 55.5$\pm$1.5 56.7$\pm$0.7 56.3$\pm$1.3 58.8$\pm$0.8 54.2$\pm$1.2 57.5$\pm$1.3 54.0$\pm$1.3 55.2$\pm$2.5 52.7$\pm$2.1 52.7$\pm$1.9 51.7$\pm$2.8 51.0$\pm$2.4
+SGDA 47.1$\pm$0.6 41.6$\pm$1.4 43.8$\pm$0.7 45.9$\pm$1.2 49.4$\pm$1.1 50.1$\pm$1.5 49.7$\pm$1.6 48.4$\pm$1.5 50.6$\pm$1.0 50.4$\pm$1.3 50.5$\pm$1.2 50.7$\pm$1.4
+StruRW 47.1$\pm$0.9 45.9$\pm$0.7 46.5$\pm$1.3 48.2$\pm$1.2 46.9$\pm$1.2 47.3$\pm$1.4 48.7$\pm$1.1 47.3$\pm$1.7 49.5$\pm$1.1 49.7$\pm$1.5 50.0$\pm$1.8 50.2$\pm$1.6
+A2GNN 47.6$\pm$1.2 47.6$\pm$0.9 46.2$\pm$0.8 48.3$\pm$1.1 46.2$\pm$1.0 47.9$\pm$0.6 52.0$\pm$1.7 53.0$\pm$1.4 52.0$\pm$1.0 53.7$\pm$1.1 52.2$\pm$1.3 51.8$\pm$1.7
+PA-BOTH 48.2$\pm$1.4 48.2$\pm$0.8 47.3$\pm$1.2 48.3$\pm$1.0 48.5$\pm$1.2 45.2$\pm$0.6 49.2$\pm$1.6 50.0$\pm$1.2 51.1$\pm$1.3 51.3$\pm$1.5 50.5$\pm$1.6 48.8$\pm$1.4
+DeSGraDA 58.0$\pm$1.5 58.2$\pm$1.4 57.0$\pm$1.8 58.3$\pm$1.6 55.9$\pm$2.1 58.1$\pm$1.6 54.1$\pm$1.5 53.6$\pm$1.6 54.9$\pm$1.1 56.2$\pm$1.8 55.0$\pm$1.3 54.6$\pm$1.2
 5 Experiment
@@
 Dataset.
-To demonstrate the effectiveness of DeSGDA, we conduct extensive experiments on four widely-used graph classification datasets from TUDataset  [^1] https://chrsmrrs.github.io/datasets/, including PROTEINS Dobson and Doig (2003), NCI1 Wale et al.
-(2008), FRANKENSTEIN Orsini et al.
-(2015), and MUTAGENICITY Kazius et al.
-(2005).
-To better address the variation in domain distributions within each dataset, we divided them into source and target domains based on the edge density, node density, and graph flux (i.e., the ratio of the number of nodes to the number of edges).
-The specific statistics, distribution visualization, and details introduction of experimental datasets are presented in Appendix F.
+To evaluate the effectiveness of DeSGraDA, we conduct extensive experiments across two types of domain shifts: (1) structure-based domain shifts, where the discrepancy between domains arises primarily from differences in graph topology, such as variations in node and edge densities.
+This category includes datasets DD, PROTEINS (Dobson and Doig, 2003), SEED Zheng and Lu (2015); Duan et al.
+(2013), and BCI (Brunner et al., 2008); (2) feature-based domain shifts, where domains differ mainly in semantic information while maintaining similar graph structures.
+This setting involves datasets including DD, PROTEINS, BZR, BZR_MD, COX2, and COX2_MD (Dobson and Doig, 2003; Sutherland et al., 2003).
+The specific statistics, distribution visualization, and detailed introduction of experimental datasets are presented in Appendix H.
 Baselines.
-We compare the proposed DeSGDA with competitive baselines on the aforementioned datasets, including one graph kernel method: WL subtree Shervashidze et al.
-(2011); four general graph neural networks: GCN Kipf and Welling (2017), GIN Xu et al.
-(2018), CIN Bodnar et al.
-(2021) and GMT Baek et al.
-(2021); two spiking graph neural networks: SpikeGCN Zhu et al.
-(2022) and DRSGNN Zhao et al.
-(2024); three domain adaptation methods: CDAN Long et al.
-(2018), ToAlign Wei et al.
-(2021b), and MetaAlign Wei et al.
-(2021a); and six graph domain adaptation methods: DEAL Yin et al.
-(2022b), CoCo Yin et al.
-(2023a), SGDA Qiao et al.
-(2023), DGDA Cai et al.
-(2024), A2GNN Liu et al.
-(2024a) and PA-BOTH Liu et al.
-(2024b).
-More detailed settings about the compared baselines are introduced in Appendix G.
+We compare the proposed DeSGraDA with competitive baselines on the aforementioned datasets, including one graph kernel method: WL subtree (Shervashidze et al., 2011); four graph-based methods: GCN (Kipf and Welling, 2017), GIN (Xu et al., 2018), CIN (Bodnar et al., 2021) and GMT (Baek et al., 2021); two spiking-based graph methods: SpikeGCN (Zhu et al., 2022) and DRSGNN (Zhao et al., 2024); three domain adaptation methods: CDAN (Long et al., 2018), ToAlign (Wei et al., 2021b), and MetaAlign (Wei et al., 2021a); and six graph domain adaptation methods: DEAL (Yin et al., 2022), CoCo (Yin et al., 2023), SGDA (Qiao et al., 2023), StruRW (Liu et al., 2023), A2GNN (Liu et al., 2024a) and PA-BOTH (Liu et al., 2024b).
+More settings about baselines are introduced in Appendix I and J.
+Table 2: Graph classification results (in %) under node and edge density domain shifts on the PROTEINS dataset, and feature domain shifts on DD, PROTEINS, BZR, BZR_MD, COX2, and COX2_MD.
+For convenience, PROTEINS, DD, COX2, COX2_MD, BZR, and BZR_MD are abbreviated as P, D, C, CM, B, and BM, respectively.
+Bold results indicate the best performance.
+Methods Node Shift Edge Shift Feature Shift
 […]
+WL subtree 69.1 61.2 41.6 68.7 50.7 58.1 43.0 42.2 53.1 58.2 51.3 44.0
+GCN 73.7$\pm$0.3 57.6$\pm$0.2 24.4$\pm$0.4 73.4$\pm$0.2 57.6$\pm$0.2 24.0$\pm$0.1 48.9$\pm$2.0 60.9$\pm$2.3 51.2$\pm$1.8 66.9$\pm$1.8 48.7$\pm$2.0 78.8$\pm$1.7
+GIN 71.8$\pm$2.7 58.5$\pm$4.3 74.2$\pm$1.7 62.5$\pm$4.7 53.0$\pm$4.6 73.7$\pm$0.8 57.3$\pm$2.2 61.9$\pm$1.9 53.8$\pm$2.5 55.6$\pm$2.0 49.9$\pm$2.4 79.2$\pm$2.8
+GMT 73.7$\pm$0.2 57.6$\pm$0.3 75.6$\pm$1.4 73.4$\pm$0.3 57.6$\pm$0.1 24.0$\pm$0.1 59.5$\pm$2.5 50.7$\pm$2.2 49.3$\pm$1.8 58.2$\pm$2.0 50.2$\pm$2.3 74.4$\pm$1.8
+CIN 74.1$\pm$0.6 60.1$\pm$2.1 75.6$\pm$0.2 74.5$\pm$0.2 57.8$\pm$0.2 75.6$\pm$0.6 59.1$\pm$2.6 58.0$\pm$2.7 51.2$\pm$2.0 55.6$\pm$1.5 49.2$\pm$1.4 74.2$\pm$1.9
+SpikeGCN 71.8$\pm$0.9 64.9$\pm$1.4 71.1$\pm$1.9 71.8$\pm$0.8 63.8$\pm$1.0 68.6$\pm$1.1 59.6$\pm$2.2 63.3$\pm$1.8 52.6$\pm$2.5 68.6$\pm$1.8 53.3$\pm$1.7 76.1$\pm$2.0
+DRSGNN 73.6$\pm$1.1 64.6$\pm$1.2 70.2$\pm$1.7 72.6$\pm$0.6 63.1$\pm$1.4 70.4$\pm$1.9 60.9$\pm$2.4 65.4$\pm$1.9 52.9$\pm$1.8 66.9$\pm$2.3 52.8$\pm$1.7 76.4$\pm$2.7
+CDAN 75.9$\pm$1.0 60.8$\pm$0.6 75.8$\pm$0.3 72.2$\pm$1.8 59.8$\pm$2.1 69.3$\pm$4.1 59.4$\pm$2.0 63.1$\pm$2.7 51.2$\pm$2.3 68.2$\pm$1.8 50.7$\pm$1.6 75.2$\pm$1.9
+ToAlign 73.7$\pm$0.4 57.6$\pm$0.6 24.4$\pm$0.1 73.4$\pm$0.1 57.6$\pm$0.1 24.0$\pm$0.3 62.1$\pm$2.1 66.5$\pm$2.3 53.2$\pm$2.6 55.8$\pm$2.3 56.2$\pm$3.0 78.8$\pm$2.4
+MetaAlign 74.3$\pm$0.8 60.6$\pm$1.7 76.3$\pm$0.3 75.5$\pm$0.9 64.8$\pm$1.6 69.3$\pm$2.7 63.3$\pm$2.1 66.2$\pm$1.9 51.2$\pm$2.0 69.5$\pm$2.3 48.7$\pm$1.8 76.8$\pm$2.7
+DEAL 75.4$\pm$1.2 68.1$\pm$1.9 73.8$\pm$1.4 76.5$\pm$0.4 67.5$\pm$1.3 76.0$\pm$0.2 70.6$\pm$1.9 66.8$\pm$2.5 50.9$\pm$2.4 67.8$\pm$1.9 51.1$\pm$2.3 79.4$\pm$2.2
+CoCo 74.8$\pm$0.6 65.5$\pm$0.4 72.4$\pm$2.9 75.5$\pm$0.2 59.8$\pm$0.5 73.6$\pm$2.3 66.0$\pm$2.7 61.2$\pm$2.3 53.6$\pm$1.8 78.2$\pm$2.0 57.8$\pm$1.6 79.8$\pm$1.8
+SGDA 64.2$\pm$0.5 66.9$\pm$1.2 65.4$\pm$1.6 63.8$\pm$0.6 66.7$\pm$1.0 60.1$\pm$0.8 48.3$\pm$2.0 55.8$\pm$2.6 49.8$\pm$1.8 66.9$\pm$2.3 50.3$\pm$2.1 78.8$\pm$2.6
+A2GNN 65.7$\pm$0.6 66.3$\pm$0.9 65.2$\pm$1.4 65.4$\pm$1.3 68.2$\pm$1.4 65.4$\pm$0.7 57.8$\pm$2.1 60.3$\pm$1.5 51.5$\pm$1.8 67.7$\pm$2.1 51.6$\pm$2.3 77.5$\pm$1.9
+StruRW 71.9$\pm$2.3 66.7$\pm$1.8 52.8$\pm$1.9 72.6$\pm$2.2 66.2$\pm$2.2 48.9$\pm$2.0 59.1$\pm$2.3 58.8$\pm$2.8 51.2$\pm$2.0 54.8$\pm$2.9 49.2$\pm$1.4 74.7$\pm$2.1
+PA-BOTH 61.0$\pm$0.8 60.3$\pm$0.6 63.7$\pm$1.5 63.1$\pm$0.7 64.3$\pm$0.5 66.3$\pm$0.7 54.2$\pm$3.2 56.7$\pm$2.6 52.9$\pm$2.8 61.8$\pm$2.0 47.5$\pm$3.0 78.8$\pm$1.9
+DeSGraDA 76.3$\pm$1.9 69.2$\pm$2.3 77.5$\pm$2.2 76.8$\pm$1.9 68.6$\pm$1.8 76.5$\pm$2.8 73.6$\pm$1.9 71.2$\pm$1.6 54.7$\pm$1.8 78.6$\pm$2.2 56.3$\pm$1.5 80.3$\pm$1.9
+5.2 Performance Comparison
+Figure 3: Correlation comparisons of spiking frequency and aggregation weights under adaptive thresholds on PROTEINS and DD datasets.
+We present the results of the proposed DeSGraDA with all baselines under two types of domain shifts on different datasets in Tables 1, 2, and 11-14.
+From these tables, we observe that: (1) GDA methods outperform traditional graph-based and spiking-based graph methods in most cases, highlighting the adverse impact of domain distribution shifts on conventional approaches and underscoring the importance of advancing research in spiking graph domain adaptation.
+(2) The spiking-based graph methods (i.e., SpikeGCN and DRSGNN) outperform the models specific for node classification (i.e., SGDA, StruRW, A2GNN, and PA-BOTH) but fall short compared to models for graph classification (i.e., DEAL and CoCo).
+This performance gap is primarily due to the limited exploration of graph classification under domain shift.
+Although spiking-based methods exhibit advantages over adapted node classification models, they remain less effective than specialized graph domain adaptation methods explicitly designed for graph-level tasks.
+(3) DeSGraDA outperforms all baselines in most cases, demonstrating its advantage over other methods.
+The superior performance can be attributed to two key factors.
+First, the degree-conscious spiking representations dynamically adjust node-specific firing thresholds in SNNs, enabling the model to capture more expressive and discriminative graph features.
+Second, the temporal-based distribution alignment aligns source and target domain representations by matching spiking membrane dynamics over time, effectively mitigating distributional discrepancies.
+Moreover, the pseudo-label distillation helps refine degree thresholds in the target domain, further enhancing generalization.
+More results can be found in Appendix K.1.
+We further conduct experiments to examine the correlation between spiking frequency and aggregation weights under adaptive thresholds across different node degrees.
+As shown in Figure 3, the correlation coefficient under adaptive thresholds is significantly lower than that under fixed thresholds, demonstrating that DeSGraDA effectively smooths spiking frequency across node degrees, mitigates over-activation in high-degree nodes, and promotes balanced information aggregation.
+Table 3: The results of ablation studies on the PROTEINS dataset (source → target). Bold results indicate the best performance.
 […]
-WL subtree 73.4 72.7 70.5 73.0 72.8 59.0 66.5 71.6 60.6 58.3 76.3 64.0 68.2
-GCN 57.2$\pm$2.7 62.8$\pm$1.7 67.6$\pm$0.5 58.5$\pm$1.3 67.7$\pm$0.4 61.0$\pm$0.3 65.0$\pm$0.8 51.1$\pm$1.3 65.6$\pm$2.2 55.4$\pm$0.4 68.5$\pm$3.1 67.7$\pm$0.5 62.3
-GIN 69.3$\pm$2.3 65.8$\pm$0.8 69.3$\pm$1.7 69.8$\pm$1.6 71.4$\pm$2.1 52.4$\pm$1.8 64.0$\pm$2.4 65.7$\pm$3.2 53.4$\pm$3.7 58.1$\pm$0.8 72.6$\pm$0.3 64.6$\pm$2.3 64.7
-GMT 67.8$\pm$1.3 69.6$\pm$0.7 74.5$\pm$0.5 67.6$\pm$2.5 69.9$\pm$2.1 55.8$\pm$0.7 74.8$\pm$1.4 60.1$\pm$2.4 71.4$\pm$3.3 51.5$\pm$0.5 69.0$\pm$0.5 63.3$\pm$1.3 66.3
-CIN 62.6$\pm$0.5 59.4$\pm$0.5 64.0$\pm$0.9 58.5$\pm$1.8 71.9$\pm$1.7 60.6$\pm$2.1 63.7$\pm$0.5 61.2$\pm$2.1 73.2$\pm$0.5 57.7$\pm$3.0 68.1$\pm$0.4 58.5$\pm$2.7 63.3
-SpikeGCN 66.1$\pm$0.4 67.8$\pm$0.7 68.5$\pm$0.9 67.2$\pm$0.3 70.1$\pm$1.1 68.6$\pm$1.3 66.1$\pm$0.7 65.8$\pm$1.2 74.5$\pm$0.9 65.4$\pm$0.6 73.8$\pm$1.1 70.1$\pm$1.6 68.6
-DRSGNN 68.3$\pm$0.5 71.9$\pm$0.9 71.5$\pm$1.1 75.1$\pm$1.7 76.7$\pm$1.3 74.4$\pm$0.8 70.5$\pm$0.7 67.8$\pm$0.8 77.0$\pm$1.5 65.9$\pm$1.6 75.2$\pm$1.9 69.2$\pm$2.1 71.8
-CDAN 75.6$\pm$0.5 70.5$\pm$0.6 71.6$\pm$0.5 69.8$\pm$0.5 76.6$\pm$0.8 71.4$\pm$0.3 71.4$\pm$0.3 72.1$\pm$0.3 75.5$\pm$0.7 74.3$\pm$0.8 78.2$\pm$1.1 74.0$\pm$0.8 73.4
-ToAlign 51.1$\pm$0.6 55.8$\pm$0.1 63.3$\pm$0.2 55.8$\pm$0.4 68.1$\pm$0.7 55.8$\pm$0.3 63.3$\pm$0.5 51.1$\pm$0.2 68.1$\pm$1.0 51.1$\pm$0.4 68.1$\pm$0.6 63.3$\pm$0.2 59.6
-MetaAlign 59.4$\pm$1.1 62.2$\pm$1.0 68.9$\pm$0.3 65.3$\pm$0.8 75.1$\pm$0.7 67.5$\pm$2.1 70.9$\pm$1.4 60.6$\pm$2.3 72.4$\pm$1.4 59.4$\pm$0.6 74.6$\pm$0.7 67.8$\pm$1.3 67.0
-DEAL 76.6$\pm$0.4 62.8$\pm$0.8 72.8$\pm$1.3 67.3$\pm$2.2 77.2$\pm$2.3 67.6$\pm$1.9 71.2$\pm$1.6 56.0$\pm$2.5 73.9$\pm$2.1 66.0$\pm$0.3 76.4$\pm$1.1 65.5$\pm$2.1 69.4
-CoCo 73.4$\pm$0.5 73.6$\pm$0.8 73.4$\pm$1.0 71.6$\pm$0.5 75.2$\pm$1.6 74.6$\pm$0.3 70.7$\pm$0.8 68.4$\pm$1.5 75.0$\pm$0.2 72.7$\pm$0.4 76.3$\pm$1.1 75.0$\pm$1.8 73.3
-SGDA 63.8$\pm$0.8 65.2$\pm$0.5 66.7$\pm$0.3 59.1$\pm$1.2 62.3$\pm$0.7 60.6$\pm$0.4 65.2$\pm$0.9 61.8$\pm$1.0 64.5$\pm$1.3 60.9$\pm$0.8 59.4$\pm$1.2 64.9$\pm$1.1 62.9
-DGDA 59.4$\pm$0.7 62.3$\pm$1.1 63.1$\pm$0.5 61.2$\pm$0.9 60.4$\pm$0.6 58.8$\pm$1.0 60.3$\pm$0.8 63.5$\pm$1.2 61.9$\pm$0.8 60.4$\pm$1.6 64.2$\pm$1.3 62.6$\pm$1.4 61.5
-A2GNN 65.4$\pm$0.7 66.4$\pm$1.1 65.7$\pm$1.3 66.0$\pm$0.6 64.9$\pm$1.2 65.8$\pm$1.6 65.5$\pm$1.8 66.0$\pm$1.4 65.8$\pm$2.1 65.6$\pm$1.9 66.1$\pm$1.7 66.0$\pm$2.0 65.8
-PA-BOTH 66.9$\pm$0.5 67.1$\pm$0.8 67.3$\pm$1.1 65.8$\pm$0.7 69.1$\pm$1.0 66.1$\pm$1.4 66.7$\pm$1.3 67.4$\pm$1.4 66.3$\pm$1.8 66.0$\pm$1.2 66.8$\pm$0.8 66.3$\pm$1.5 66.8
-DeSGDA 78.7$\pm$1.3 78.4$\pm$1.1 74.8$\pm$0.6 77.6$\pm$0.9 79.5$\pm$1.2 76.7$\pm$0.8 74.9$\pm$0.7 71.2$\pm$1.7 79.5$\pm$1.4 72.8$\pm$0.8 81.0$\pm$1.5 75.1$\pm$1.0 76.7
-Implementation Details.
-DeSGDA and all baseline models are implemented using PyTorch [^2] https://pytorch.org/ and PyTorch Geometric [^3] https://www.pyg.org/.
-For DeSGDA, we deploy the GIN Xu et al.
-(2018) as the backbone of the degree-aware personalized spiking graph encoder, incorporating a mean-pooling layer for the readout function.
-We conduct experiments for DeSGDA and all baselines on NVIDIA A100 GPUs for a fair comparison, where the learning rate of Adam optimizer set to ${10^{-4}}$, hidden embedding dimension 256, weight decay ${10^{-12}}$, and GNN layers 4.
-Additionally, DeSGDA and all baseline models are trained using all labeled source samples and evaluated on unlabeled target samples Wu et al.
-(2020).
-The performances of all models are measured and averaged on all samples for five different runs.
-Table 2: The results of ablation studies on PROTEINS (source → target). Bold results indicate the best performance.
-[…]
-DeSGDA w/o CA 72.3 72.6 71.9 75.1 74.1 71.3 71.5 70.4 71.6 70.5 78.3 71.5 72.6
-DeSGDA w/o PL 72.1 68.7 67.8 69.4 63.7 55.7 68.3 69.5 70.2 69.6 76.7 66.9 68.3
-DeSGDA w/o CF 56.7 54.2 60.7 62.1 73.4 68.3 55.3 69.5 76.7 65.5 73.7 63.3 65.0
-DeSGDA w/o TL 61.7 43.8 56.1 71.2 49.5 69.0 42.4 68.3 73.7 64.0 70.5 43.5 59.5
-DeSGDA w/ PT 71.5 68.3 66.1 72.4 66.7 70.1 67.9 70.8 66.3 71.8 74.1 72.9 70.0
-DeSGDA w/ CL 74.9 73.3 73.7 75.1 77.0 71.3 73.9 70.0 78.1 77.6 78.8 74.9 74.9
-DeSGDA 78.7 78.4 74.8 77.6 79.5 76.7 74.9 71.2 79.5 72.8 81.0 75.1 76.7
-5.2 Performance Comparision
-We present the results of our proposed DeSGDA with all baseline models under the setting of graph domain adaptation on different datasets in Tables 1 and 13-23.
-From these tables, we observe that: (1) The performance of graph domain adaptation methods surpasses graph-based and spike-based graph methods.
-We attribute this improvement to the fact that domain distribution shifts degrade the performance of traditional graph methods.
-(2) The graph domain adaptation methods (DEAL and CoCo) outperform the spike-based graph methods (SpikeGCN and DRSGNN), underscoring the necessity of the research in spiking graph domain adaptation.
-(3) The WL subtree method outperforms SGDA, DGDA, A2GNN, and PA-BOTH but falls short compared to DEAL and CoCO.
-We attribute this to the relatively limited research specifically addressing the graph domain adaptation problem (e.g., DEAL and CoCo).
-To bridge this gap, we adapted node classification methods for graph classification tasks (e.g., SGDA, DGDA, A2GNN, and PA-BOTH).
-While the WL subtree method demonstrates superior performance over these adapted node classification-based methods, it remains less effective than dedicated graph domain adaptation methods tailored for graph classification tasks.
-(4) Our DeSGDA outperforms all baselines for most cases, which demonstrates its superiority over other methods.
-The remarkable performance of DeSGDA lies in two main reasons: (i) The degree-aware personalized spiking representations can capture more expressive information for graph classification by dynamically adjusting the thresholds of nodes in SNNs.
-(ii) The adversarial distribution alignment effectively addresses domain discrepancies by adversarially training the encoder and domain discriminator to align feature spaces.
-Moreover, the pseudo-label distillation aids in updating the source degree thresholds in the target domain, thereby ensuring optimal performance.
-More results on other datasets can be found in Appendix H.1.
-Figure 2: The performance with different GNN architectures on PROTEINS.
-Additionally, we conduct experiments to explore the flexibility of DeSGDA.
-Specifically, we replace the backbone of the degree-aware personalized spiking graph encoder (GIN) with various GNN methods (i.e., GCN and GAT), and the results are shown in Figure 2.
-From those results, we observe that GIN consistently outperforms other GNN architectures in most cases, demonstrating its powerful representation capability.
-This phenomenon also justifies our choice of using GIN to enhance the performance of the proposed DeSGDA.
-More results are reported in Appendix H.1.
+DeSGraDA w/ CDAN 73.3 82.4 67.8 76.5 74.4 78.7 66.5 71.3 73.7 70.2 74.1 68.8
+DeSGraDA w/o PL 73.7 81.2 67.1 81.2 75.9 79.6 67.8 71.9 74.7 69.0 75.4 68.3
+DeSGraDA w/o CF 65.0 67.4 53.5 64.5 66.6 68.8 56.7 63.4 66.1 60.9 53.9 56.6
+DeSGraDA w/o TL 73.6 80.6 65.6 80.6 73.1 78.4 63.9 69.3 69.6 68.7 72.9 64.5
+DeSGraDA 76.3 84.6 69.2 83.6 77.5 83.7 69.8 74.0 76.2 73.0 77.8 70.5
+5.3 Ablation Study
+We conduct ablation studies to examine the contributions of each component: (1) DeSGraDA w/ CDAN: It utilizes the static distribution alignment instead of temporal-based module; (2) DeSGraDA w/o PL: It removes the pseudo-label distilling module; (3) DeSGraDA w/o CF: It removes the classification loss $\mathcal{L}_{S}$; (4) DeSGraDA w/o TL: It utilizes the fixed global thresholds on all nodes.
+Experimental results are shown in Table 3.
+From the table, we find that: (1) The degree-aware thresholding mechanism substantially improves representational capacity.
+When replaced with fixed global thresholds (DeSGraDA w/o TL), performance consistently declines, suggesting that dynamically adjusting thresholds based on node degree helps the model better handle structural heterogeneity and encode informative spiking representations.
+(2) The temporal-based distribution alignment and pseudo-label distillation modules are crucial for effective domain adaptation.
+Removing the temporal alignment module (DeSGraDA w/ CDAN) leads to consistent performance drops across most tasks, highlighting its role in mitigating domain shifts by aligning membrane dynamics across domains.
+Similarly, eliminating pseudo-label distillation (DeSGraDA w/o PL) degrades performance, demonstrating the importance of leveraging confident target predictions to refine threshold adaptation and support generalization.
+(3) DeSGraDA outperforms all ablated variants across all domain shifts, confirming the complementary strengths of its core components.
+Notably, removing the classification loss (DeSGraDA w/o CF) results in the largest performance degradation, highlighting the necessity of source supervision for learning discriminative features.
+Furthermore, we provide ablation studies to examine the effect of directly replacing SGNs with commonly used GNNs for DeSGraDA, and the results are shown in Tables 7, 8.
+More ablation studies are reported in Appendix K.3.
 5.4 Energy Efficiency Analysis
-Figure 3: Energy consumption of DeSGDA and baselines on different datasets.
-To assess the energy efficiency of DeSGDA, we use the metric from Zhu et al.
-(2022) and quantify the energy consumption for graph classification in the inference stage.
-Specifically, the graph domain adaption methods are evaluated on GPUs (NVIDIA A100), and the spiking-based methods are evaluated on neuromorphic chips (ROLLS Indiveri et al.
-(2015)) following Zhu et al.
-(2022).
-The results are shown in Figure 3, where we find that compared with traditional graph domain adaptation methods, the spike-based methods (DeSGDA and DRSGNN) have significantly lower energy consumption, demonstrating the superior energy efficiency of SGNs.
-Moreover, although the energy consumption of DeSGDA is slightly higher than DRSGNN due to additional computations required for domain adaptation, the performance improvement justifies the deployment of DeSGDA in low-power devices.
-Additionally, we present a comparison of training time and memory usage between DeSGDA and other graph domain adaptation methods.
-The results are detailed in Tables 9 and  10.
-5.4 Ablation Study
-We conduct ablation studies to examine the contributions of each component in the proposed DeSGDA: (1) DeSGDA w/o CA: It removes the adversarial distribution alignment module; (2) DeSGDA w/o PL: It removes the pseudo-label distilling module; (3) DeSGDA w/o CF: It removes the classification loss $\mathcal{L}_{S}$; (4) DeSGDA w/o TL: It utilizes the global thresholds on all nodes; (5) DeSGDA w/ PT: It deploys the adaptive perturbations Yin et al.
-(2022b) on source data for alignment; (6) DeSGDA w/ CL: It replaces the adversarial learning with the cross-domain contrastive learning Yin et al.
-(2023a).
-Experimental results are shown in Table 2.
-From the table, we find that: (1) DeSGDA outperforms DeSGDA w/o CA, DeSGDA w/o PL, and DeSGDA w/o CF, demonstrating that the adversarial distribution alignment module effectively reduces domain discrepancies, ensuring well-aligned feature spaces between source and target domains.
-Additionally, the pseudo-label distillation module addresses variance in thresholds across domains, while the classification loss $\mathcal{L}_{S}$ enables DeSGDA to effectively learn from labeled source data and generalize to the target domain.
-(2) DeSGDA w/o TL shows lower performance compared to DeSGDA, showing that degree-aware thresholds, iteratively updated during training, resolve the issue of inflexible architecture in SGNs.
-By using these thresholds, DeSGDA effectively learns meaningful representations for nodes with various degrees.
-(3) DeSGDA outperforms DeSGDA w/ PT and DeSGDA w/ CL.
-We attribute this to the adaptive perturbations method (DeSGDA w/ PT) fail to help DeSGDA overcome domain discrepancies effectively.
-Additionally, the cross-domain contrastive learning method (DeSGDA w/ CL) is less effective at aligning the source and target distribution compared to adversarial learning.
-Furthermore, we provide ablation studies to examine the effect of directly replacing SGNs with commonly used Graph Neural Networks (GNNs) for generating representations for DeSGDA, and the results are shown in Tables 11, 12.
-More details about ablation results on other datasets are reported in Appendix H.3.
-(a) Time Latency $\tau$
-(b) Initial Threshold $V_{th}^{degree}$
-(c) Balance Ratio $\alpha$
-Figure 4: Hyperparameter sensitivity analysis of time latency $\tau$, initial threshold $V_{th}^{degree}$ in SNNs, and balance ratio $\alpha$ on PROTEINS.
+To assess the energy efficiency of DeSGraDA, we use the metric from (Zhu et al., 2022) and quantify the energy consumption for graph classification in the inference stage.
+Specifically, the graph domain adaptation methods are evaluated on GPUs (NVIDIA A100), and the spiking-based methods are evaluated on neuromorphic chips (ROLLS (Indiveri et al., 2015)) following (Zhu et al., 2022).
+The results are shown in Figure 4(a), where we find that compared with traditional graph domain adaptation methods, the spike-based methods (DeSGraDA and DRSGNN) have significantly lower energy consumption, demonstrating the superior energy efficiency of SGNs.
+Moreover, although the energy consumption of DeSGraDA is slightly higher than DRSGNN due to additional computations required for domain adaptation, the performance improvement justifies deploying DeSGraDA in low-power devices.
+Additionally, we present a comparison of training time and memory usage between DeSGraDA and other GDA methods, and the results are detailed in Tables 5 and  6.
+(a) Energy Consumption
+(b) Latency Step $T$
+(c) Initial Threshold $V_{th}^{degree}$
+Figure 4: (a) Energy efficiency analysis and (b), (c) hyperparameter sensitivity analysis of latency step $T$ and initial threshold $V_{th}^{degree}$ on the PROTEINS dataset.
 5.5 Sensitivity Analysis
-We conduct the sensitivity analysis of DeSGDA with respect to the impact of its hyperparameters: time latency $\tau$, degree threshold value $V_{th}^{degree}$ in SNNs, and balance ratio $\alpha$, which plays a crucial role in the performance of DeSGDA.
-In particular, $\tau$ controls the number of SNNs propagation steps; $V_{th}^{degree}$ determines when a neuron fires; $\alpha$ governs the changing ratio of degree-aware thresholds.
-Figure 4 illustrates how $\tau$, $V_{th}^{degree}$, and $\alpha$ affects the performance of DeSGDA on the PROTEINS dataset.
-More results on other datasets are shown in Appendix H.4.
-We vary $\tau$ within the range of $\{5,6,7,8,9,10\}$, $V_{th}^{degree}$ in $\{0.05,0.1,0.2,0.5,1.0,2.0,5.0\}$, and $\alpha$ in $\{0.1,0.3,0.5,0.7,0.9\}$.
-From the results, we observe that: (1) The performance of DeSGDA in Figure 4(a) generally exhibits an increasing trend at the beginning and then stabilizes when $\tau$ is greater than 8.
-We attribute this to smaller values of $\tau$ potentially losing important information for representation, while larger values significantly increase model complexity.
-To balance performance and complexity, we set $\tau$ to 8 as default.
-(2) Figure 4(b) indicates an initial increase followed by a decreasing trend in performance as $V_{th}^{degree}$ increases.
-This trend occurs because a lower threshold may trigger more spikes for high-degree nodes, leading to a drastic change in threshold, which degrades performance.
-Conversely, a higher threshold for low-degree nodes could result in fewer spikes, affecting model’s ability to process information.
-Thus, we set $V_{th}^{degree}$ to 0.5 as default.
-(3) From Figure 4(c), we find that the performance of DeSGDA initially increases and then decreases as $\alpha$ increases.
-The potential reason is that smaller $\alpha$ may delay the updating of the threshold, leading to performance degradation.
-Contrarily, a larger $\alpha$ tends to introduce more spikes that change dynamically at each step, resulting in instability in the model’s performance.
-Therefore, we set $\alpha$ to 0.5 as default.
+We conduct the sensitivity analysis of DeSGraDA to investigate the impact of key hyperparameters: latency step $T$ and degree threshold $V_{th}^{degree}$ in SGNs.
+Specifically, $T$ controls the number of SGNs propagation steps, and $V_{th}^{degree}$ determines the firing threshold of each neuron based on node degree.
+Figure 4(b) and 4(c) illustrates how $T$ and $V_{th}^{degree}$ affects the performance of DeSGraDA on the PROTEINS dataset.
+More results on other datasets are shown in Appendix K.4.
+We vary $T$ in $\{5,6,7,8,9,10\}$ and $V_{th}^{degree}$ in $\{0.05,0.1,0.2,0.5,1.0,2.0,5.0\}$.
+From the results, we observe that: (1) The performance of DeSGraDA in Figure 4(b) generally exhibits an increasing trend at the beginning and then stabilizes when $T>9$.
+We attribute this to smaller values of $T$ potentially losing important information for representation, while larger values significantly increase model complexity.
+To balance effectiveness and efficiency, we set $T=9$ as default.
+(2) Figure 4(c) indicates an initial increase followed by a decreasing trend in performance as $V_{th}^{degree}$ increases.
+This trend arises because a lower threshold may cause excessive spiking for high-degree nodes, leading to unstable representation, while a higher threshold may suppress spiking for low-degree nodes, reducing information flow.
+Accordingly, we set $V_{th}^{degree}$ to 0.2 as default.
 6 Conclusion
-In this paper, we propose the problem of spiking graph domain adaptation and introduce a novel framework, DeSGDA, for graph classification.
-This framework enhances the adaptability and performance of SGNs through three key aspects: node degree-aware personalized spiking representation, adversarial feature distribution alignment, and pseudo-label distillation.
-Our approach enables expressive information capture through degree-dependent spiking thresholds, aligns feature distributions via adversarial training, and utilizes pseudo-labels to leverage unlabeled data effectively.
-Extensive experimental validation across benchmark datasets demonstrates that DeSGDA surpasses existing methods in accuracy while maintaining efficient energy consumption, making it a promising solution for advancing the domain adaptation capabilities of spiking graph networks.
-In the future, we will apply SGNs in source-free domain adaptation and domain generalization scenarios.
-Appendix A Proof of Hypothesis 1
+In this paper, we propose the problem of spiking graph domain adaptation and introduce a novel framework DeSGraDA for graph classification.
+DeSGraDA enhances the adaptability and performance of SGNs through three key aspects: degree-conscious spiking representation, temporal distribution alignment, and pseudo-label distillation.
+DeSGraDA captures expressive information via degree-dependent spiking thresholds, aligns feature distributions through temporal dynamics, and effectively exploits unlabeled target data via pseudo-label refinement.
+Extensive experiments on benchmark datasets demonstrate that DeSGraDA surpasses existing methods in accuracy while maintaining energy efficiency, showcasing its potential as a strong solution for DA in SGNs.
+In the future, we plan to extend DeSGraDA to source-free and domain-generalization scenarios.
+Appendices
+Appendix A DA Bound for Graphs
+DA Bound for Graphs.
+Due to the DA theory is agnostic to data structures and encoders, You et al.
+[2023] directly rewrite it for graph-structured data ($G$) accompanied with graph feature extractors ($f$) as follows, and the covariate shift assumption is reframed as $\mathbb{P}_{S}(Y|G)=\mathbb{P}_{T}(Y|G)$.
+Theorem 3
+[You et al., 2023] Assume that the learned discriminator is $C_{g}$-Lipschitz continuous as in [Redko et al., 2017], and the graph feature extractor $f$ is $C_{f}$-Lipschitz that $||f||_{Lip}=\max_{G_{1},G_{2}}\frac{||f(G_{1})-f(G_{2})||_{2}}{\eta(G_{1},G_{2})}=C_{f}$ for some graph distance measure $\eta$.
+Let :=$\mathcal{H}:=\{h:\mathcal{G}\rightarrow\mathcal{Y}\}$ be the set of bounded real-valued functions with pseudo-dimension $Pdim(\mathcal{H})=d$ that $h=g\circ f\in\mathcal{H}$, with probability at least $1-\delta$ the following inequality holds:
+$\epsilon_{T}(h,\hat{h})\leq\ \hat{\epsilon}_{S}(h,\hat{h})+\sqrt{\frac{4d}{N_{S}}\log\left(\frac{eN_{S}}{d}\right)+\frac{1}{N_{S}}\log\left(\frac{1}{\delta}\right)}+2C_{f}C_{g}W_{1}(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G))+\omega,$
+where $\epsilon_{T}(h,\hat{h})=\mathbb{E}_{\mathbb{P}_{T}(G}\{|h(G)-\hat{h}(G)|\}$ is the (empirical) target risk, $\hat{\epsilon}_{S}(h,\hat{h})=\frac{1}{N_{S}}\sum_{n=1}^{N_{S}}|h(G_{n})-\hat{h}(G_{n})|$ is the (empirical) source risk, $\hat{h}:\mathcal{G}\rightarrow\mathcal{Y}$ is the labeling function for graphs and $\omega=\min_{||g||_{Lip}\leq C_{g},||f||_{Lip}\leq C_{f}}\{\epsilon_{S}(h,\hat{h})+\epsilon_{T}(h,\hat{h})\}$, and $W_{1}(\mathbb{P},\mathbb{Q})=\sup_{||g||_{Lip}\leq 1}\left\{\mathbb{E}_{\mathbb{P}_{S}(Z)}g(Z)-\mathbb{E}_{\mathbb{P}_{T}(Z)}g(Z)\right\}$ is the first Wasserstein distance [Villani et al., 2008].
+Appendix B Spiking Neural Networks
+Spiking Neural Networks (SNNs) are brain-inspired models that communicate through discrete spike events rather than continuous-valued activations Maass [1997], Gerstner and Kistler [2002], Bohte et al.
+[2000].
+This design provides significant advantages in temporal information processing and energy efficiency.
+Different from traditional neural networks, SNNs emulate biological mechanisms such as membrane potential integration, threshold-triggered firing, and post-spike resetting Cao et al.
+[2015], Zhu et al.
+[2022].
+To capture these biological dynamics, SNNs employ neuron models that mathematically describe the temporal evolution of membrane potentials and the conditions for spike generation Lagani et al.
+[2023].
+A widely used neuron model in SNNs is the Leaky Integrate-and-Fire (LIF) model Tal and Schwartz [1997], Lansky and Ditlevsen [2008], which operates through three fundamental stages:
+(1) Integrate: At each lantency step $t$, the membrane potential $V[t]$ is updated by integrating the input current $I[t]$ and applying a decay to the previous potential $V[t-1]$:
+$V[t]=\lambda V[t-1](1-S[t-1])+I[t]$ (10)
+where $\lambda\in(0,1)$ is the decay factor that controls the leakage rate, and $S[t-1]$ is the binary spike indicator from the previous time step.
+(2) Fire: A spike is emitted when the membrane potential exceeds the threshold $V_{th}$:
+$S[t]=H(V[t]-V_{th}),$ (11)
+where $H(\cdot)$ denotes the Heaviside step function, a non-differentiable function defined as:
+$H(x)=\begin{cases}1,&x\geq 0,\\ 0,&\text{ otherwise. }\end{cases}$ (12)
+(3) Reset: After a spike is emitted, the membrane potential is reset according to the following rule:
+$V[t]=(1-S[t])V[t]+S[t]V_{reset},$ (13)
+where $V_{reset}$ denotes the resting potential, typically set to zero.
+Due to the non-differentiability of $H(\cdot)$, surrogate gradient methods are commonly employed to approximate its derivative during backpropagation, enabling gradient-based optimization in deep SNN architectures Bohte et al.
+[2000], Sun et al.
+[2024].
+Appendix C Proof of Proposition 1
 Assuming that the node feature $h_{i}$ follows a normal distribution $\mathcal{N}(\mu,\sigma^{2})$, then for each node in the graph, we follow the message-passing mechanism and have the information aggregation as:
-$h_{i}=h_{i}+\sum_{j\in N(i)}w_{ij}h_{j}.$ (9)
-Therefore, we have the expectation:
+$h_{i}=h_{i}+\sum_{j\in N(i)}w_{ij}h_{j},$ (14)
+where $N(i)$ is the neighbor set of node $i$. Therefore, we have the expectation:
 $\mathbb{E}(h_{i})=\mathbb{E}(h_{i})+\sum_{j\in N(i)}w_{ij}\mathbb{E}(h_{j}).$ (15)
 Since $\mathbb{E}(h_{j})\sim\mathcal{N}(\mu,\sigma^{2})$, we have:
-$\mathbb{E}(h_{i})\sim\mathcal{N}\left((1+\sum_{j\in N(i)}w_{ij})\mu,(1+\sum_{j\in N(i)}w_{ij})\sigma^{2}\right).$ (11)
-From the results, we observe that node $i$ follows a normal distribution with a mean of $(1+\sum_{j\in N(i)}w_{ij})\mu$, determined by the aggregated weights of its neighboring nodes.
-To provide a more intuitive understanding, we visualize the aggregated neighbor weights of GCN Kipf and Welling [2017] and GIN Xu et al.
-[2018] in Figure 5.
+$\mathbb{E}(h_{i})\sim\mathcal{N}\left((1+\sum_{j\in N(i)}w_{ij})\mu,(1+\sum_{j\in N(i)}w_{ij})^{2}\sigma^{2}\right).$ (16)
+From the results, we observe that node $i$ follows a normal distribution with a mean of $(1+\sum_{j\in N(i)}w_{ij})^{2}\mu$, determined by the aggregated weights of its neighboring nodes.
+To provide a more intuitive understanding, we visualize the aggregated neighbor weights of GCN [Kipf and Welling, 2017] and GIN [Xu et al., 2018] in Figure 2.
 The results show that as the degree increases, the aggregated weights also increase progressively.
@@
 Based on this, we propose assigning higher thresholds to high-degree nodes and lower thresholds to low-degree nodes.
-Another observation is that methods that normalize neighbor weights to 1 (e.g., GAT Veličković et al.
-[2017], GraphSAGE Hamilton et al.
-[2017]) still result in aggregated features following the same normal distribution.
-This normalization diminishes the ability to distinguish between nodes with varying degrees, ultimately degrading performance.
-This explains why, when using GAT as the backbone of DeSGDA, the performance is the weakest.
-Figure 5: Visualization of degree-aware thresholds and aggregation weights on PROTEINS and NCI1.
-Appendix B Proof of Theorem 2
+Appendix D Proof of Theorem 2
 Theorem 2 Assuming that the learned discriminator is $C_{g}$-Lipschitz continuous as described in Theorem 3, the graph feature extractor $f$ (also referred to as GNN) is $C_{f}$-Lipschitz that $||f||_{Lip}=\max_{G_{1},G_{2}}\frac{||f(G_{1})-f(G_{2})||_{2}}{\eta(G_{1},G_{2})}=C_{f}$ for some graph distance measure $\eta$ and the loss function bounded by $C>0$.
 Let :=$\mathcal{H}:=\{h:\mathcal{G}\rightarrow\mathcal{Y}\}$ be the set of bounded real-valued functions with the pseudo-dimension $Pdim(\mathcal{H})=d$ that $h=g\circ f\in\mathcal{H}$, and provided the spike training data set $S_{n}=\{(\mathbf{X}_{i},y_{i})\in\mathcal{X}\times\mathcal{Y}\}_{i\in[n]}$ drawn from $\mathcal{D}^{s}$, with probability at least $1-\delta$ the following inequality:
-$\epsilon_{T}(h,\hat{h}_{T}(\mathbf{X}))\leq$ $\hat{\epsilon}_{S}(h,\hat{h}_{S}(S_{n}))+2\mathbb{E}\left[\sup\frac{1}{N_{S}}\sum_{i=1}^{N_{S}}\epsilon_{i}h(\mathbf{X}_{i},y_{i},p_{i})\right]+C\sqrt{\frac{ln(2/\delta)}{N_{S}}}$ (12)
-$+\min\left(|\epsilon_{S}(h,\hat{h}_{S}(\mathbf{X}))-\epsilon_{S}(h,\hat{h}_{T}(\mathbf{X}))|,|\epsilon_{T}(h,\hat{h}_{S}(\mathbf{X}))-\epsilon_{T}(h,\hat{h}_{T}(\mathbf{X}))|\right)$
-$+2C_{f}C_{g}W_{1}\left(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G)\right),$
-where the (empirical) source and target risks are $\hat{\epsilon}_{S}(h,\hat{h})=\frac{1}{N_{S}}\sum_{n=1}^{N_{S}}|h(G_{n})-\hat{h}(G_{n})|$ and $\epsilon_{T}(h,\hat{h})=\mathbb{E}_{\mathbb{P}_{T}(G}\{|h(G)-\hat{h}(G)|\}$, respectively, where $\hat{h}:\mathcal{G}\rightarrow\mathcal{Y}$ is the labeling function for graphs and $\omega=\min_{||g||_{Lip}\leq C_{g},||f||_{Lip}\leq C_{f}}\{\epsilon_{S}(h,\hat{h})+\epsilon_{T}(h,\hat{h})\}$, $\epsilon_{i}$ is the Rademacher variable and $p_{i}$ is the $i^{th}$ row of $\mathbf{P}$, which is the probability matrix with:
+[…]
+where the (empirical) source and target risks are $\hat{\epsilon}_{S}(h,\hat{h}(\mathbf{S}))=\frac{1}{N_{S}}\sum_{n=1}^{N_{S}}|h(\mathbf{S}_{n})-\hat{h}(\mathbf{S}_{n})|$ and $\epsilon_{T}(h,\hat{h}(\mathbf{X}))=\mathbb{E}_{\mathbb{P}_{T}(G}\{|h(G)-\hat{h}(G)|\}$, respectively, where $\hat{h}:\mathcal{G}\rightarrow\mathcal{Y}$ is the labeling function for graphs and $\omega=\min\left(|\epsilon_{S}(h,\hat{h}_{S}(\mathbf{X}))-\epsilon_{S}(h,\hat{h}_{T}(\mathbf{X}))|,|\epsilon_{T}(h,\hat{h}_{S}(\mathbf{X}))-\epsilon_{T}(h,\hat{h}_{T}(\mathbf{X}))|\right)$, $\epsilon_{i}$ is the Rademacher variable and $p_{i}$ is the $i^{th}$ row of $\mathbf{P}$, which is the probability matrix with:
 $\mathbf{P}_{kt}=\left\{\begin{array}[]{lr}\exp\left(\frac{u_{k}(t)-V_{th}}{\sigma(u_{k}(t)-u_{reset})}\right),&if\quad u_{\theta}\leq u(t)\leq V_{th},\\ 0,&if\quad u_{reset}\leq u_{k}(t)\leq u_{\theta}.\end{array}\right.$ (18)
@@
 $\overset{(a)}{\leq}|\epsilon_{S}(h,\hat{h}_{S})-\epsilon_{S}(h,\hat{h}_{T})|+2C_{f}C_{g}W_{1}\left(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G)\right),$
-where $(a)$ results from Shen et al.
-[2018] Theorem 1 with the assumption $\max(||h||_{Lip},\max_{G_{1},G_{2}}\frac{|\hat{h}_{D}(G_{1})-\hat{h}_{D}(G_{2})|}{\eta(G_{1},G_{2})})\leq C_{f}C_{g}$, $D\in\{S,T\}$.
+where $(a)$ results from [Shen et al., 2018] Theorem 3 with the assumption $\max(||h||_{Lip},\max_{G_{1},G_{2}}\frac{|\hat{h}_{D}(G_{1})-\hat{h}_{D}(G_{2})|}{\eta(G_{1},G_{2})})\leq C_{f}C_{g}$, $D\in\{S,T\}$.
 Similarly, we obtain:
@@
 $\epsilon_{T}(h,\hat{h}_{T})\leq$ $\epsilon_{S}(h,\hat{h}_{S})+2C_{f}C_{g}W_{1}\left(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G)\right)$ (23)
-$+\min\left(|\epsilon_{S}(h,\hat{h}_{S})-\epsilon_{S}(h,\hat{h}_{T})|,|\epsilon_{T}(h,\hat{h}_{S})-\epsilon_{T}(h,\hat{h}_{T})|\right)$
+$+\min\left(|\epsilon_{S}(h,\hat{h}_{S})-\epsilon_{S}(h,\hat{h}_{T})|,|\epsilon_{T}(h,\hat{h}_{S})-\epsilon_{T}(h,\hat{h}_{T})|\right).$
 The $\hat{h}$ above is the abbreviation of $\hat{h}(x)$, which means the input is the continuous feature.
-Provided the spike training data set $S_{n}=\{(\mathbf{X}_{i},y_{i})\in\mathcal{X}\times\mathcal{Y}\}_{i\in[n]}$ drawn from $\mathcal{D}$, and motivated by Yin et al.
-[2024a], we have:
+Provided the spike training data set $S_{n}=\{(\mathbf{X}_{i},y_{i})\in\mathcal{X}\times\mathcal{Y}\}_{i\in[n]}$ drawn from $\mathcal{D}$, and motivated by [Yin et al., 2024], we have:
 $\lim_{\tau\to\infty}P\left(\hat{h}(S_{n})_{\tau,i}>\hat{h}(\mathbf{X}_{\tau,i})+\epsilon\right)\leq e^{-\epsilon^{2}/2(\sigma+\hat{w}_{i}\epsilon/3)},$ (24)
@@
 From Equation 2, we observe that as $\tau\to\infty$, the difference between spike and real-valued features will be with the probability of $p=e^{-\epsilon^{2}/2(\sigma+\hat{w}_{i}\epsilon/3)}$ to exceed the upper and lower bounds.
-Furthermore, motivated by the techniques given by Bartlett and Mendelson [2002], we have:
+Furthermore, motivated by the techniques given by [Bartlett and Mendelson, 2002], we have:
 $\epsilon_{S}(h,\hat{h}_{S}(S_{n}))\leq\hat{\epsilon}_{S}(h,\hat{h}_{S}(S_{n}))+\underbrace{\sup[\epsilon_{S}(h,\hat{h}_{S}(S_{n}))-\hat{\epsilon}_{S}(h,\hat{h}_{S}(S_{n}))]}_{R(S_{n},\mathbf{P})},$ (25)
@@
 $\left\{\begin{array}[]{lr}|R(S_{n},\mathbf{P})-R(S^{\prime}_{n},\mathbf{P})|\leq C/n,&\\ |R(S_{n},\mathbf{P})-R(S_{n},\mathbf{P}^{\prime})|\leq C/n.&\end{array}\right.$ (28)
-From McDiarmid’s inequality McDiarmid et al. [1989], with probability at least $1-\delta$, we have:
+From McDiarmid’s inequality [McDiarmid et al., 1989], with probability at least $1-\delta$, we have:
 $R(S_{n},\mathbf{P})\leq\mathbb{E}_{S_{n}\in\mathcal{D},\mathbf{P}}[R(S_{n},\mathbf{P})]+C\sqrt{\frac{ln(2/\delta)}{N_{S}}}.$ (29)
@@
 $\leq 2\mathbb{E}\left[\sup\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}\hat{h}(\mathbf{X}_{i},y_{i},p_{i})\right],$
-where $\epsilon_{i}$ is the Rademacher variable. Combining Eq. 23 24 26, we have:
+where $\epsilon_{i}$ is the Rademacher variable. Combining Eq. 28 29 31, we have:
 $\epsilon_{S}(h,\hat{h}_{S}(\mathbf{X}_{n}))\leq\hat{\epsilon}_{S}(h,\hat{h}_{S}(S_{n}))+2\mathbb{E}\left[\sup\frac{1}{N_{S}}\sum_{i=1}^{N_{S}}\epsilon_{i}h(\mathbf{X}_{i},y_{i},p_{i})\right]+C\sqrt{\frac{ln(2/\delta)}{N_{S}}}.$ (32)
@@
 $+2C_{f}C_{g}W_{1}\left(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G)\right).$
-Appendix C Proof of Theorem 3
-Theorem 3 Under the assumption of Theorem 1, we further assume that there exists a small amount of i.i.d.
+Appendix E Generalization Bound with Pseudo-label Distillation Module
+Theorem 4
+Under the assumption of Theorem 3, we further assume that there exists a small amount of i.i.d.
 samples with pseudo labels $\{(G_{n},Y_{n})\}_{n=1}^{N^{\prime}_{T}}$ from the target distribution $\mathbb{P}_{T}(G,Y)$ $(N^{\prime}_{T}\ll N_{S})$ and bring in the conditional shift assumption that domains have different labeling function $\hat{h}_{S}\neq\hat{h}_{T}$ and $\max_{G_{1},G_{2}}\frac{|\hat{h}_{D}(G_{1})-\hat{h}_{D}(G_{2})|}{\eta(G_{1},G_{2})}=C_{h}\leq C_{f}C_{g}(D\in\{S,T\})$ for some constant $C_{h}$ and distance measure $\eta$, and the loss function bounded by $C>0$.
 Let :=$\mathcal{H}:=\{h:\mathcal{G}\rightarrow\mathcal{Y}\}$ be the set of bounded real-valued functions with the pseudo-dimension $Pdim(\mathcal{H})=d$, and provided the spike training data set $S_{n}=\{(\mathbf{X}_{i},y_{i})\in\mathcal{X}\times\mathcal{Y}\}_{i\in[n]}$ drawn from $\mathcal{D}^{s}$, with probability at least $1-\delta$ the following inequality holds:
-$\epsilon_{T}(h,\hat{h}_{T}(\mathbf{X}))\leq$ $\frac{N^{\prime}_{T}}{N_{S}+N^{\prime}_{T}}\hat{\epsilon}_{T}(h,\hat{h}_{T}(S))+\frac{N_{S}}{N_{S}+N^{\prime}_{T}}\Bigg{(}\hat{\epsilon}_{S}(h,\hat{h}_{S}(S))+2C_{f}C_{g}W_{1}\left(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G)\right)$ (29)
-$+2\mathbb{E}\left[\sup\frac{1}{N_{S}}\sum_{i=1}^{N_{S}}\epsilon_{i}h(\mathbf{X}_{i},y_{i},p_{i})\right]+C\sqrt{\frac{ln(2/\delta)}{N_{S}}}$
-$+\min\left(|\epsilon_{S}(h,\hat{h}_{S}(\mathbf{X})))-\epsilon_{S}(h,\hat{h}_{T}(\mathbf{X})))|,|\epsilon_{T}(h,\hat{h}_{S}(\mathbf{X})))-\epsilon_{T}(h,\hat{h}_{T}(\mathbf{X})))|\right)\Bigg{)}$
-$\leq$ $\hat{\epsilon}_{S}(h,\hat{h}_{S}(S))+2\mathbb{E}\left[\sup\frac{1}{N_{S}}\sum_{i=1}^{N_{S}}\epsilon_{i}h(\mathbf{X}_{i},y_{i},p_{i})\right]+C\sqrt{\frac{ln(2/\delta)}{N_{S}}}$
-$+2C_{f}C_{g}W_{1}\left(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G)\right)+\omega^{\prime}.$
-where the (empirical) source and target risks are $\hat{\epsilon}_{S}(h,\hat{h})=\frac{1}{N_{S}}\sum_{n=1}^{N_{S}}|h(G_{n})-\hat{h}(G_{n})|$ and $\epsilon_{T}(h,\hat{h})=\mathbb{E}_{\mathbb{P}_{T}(G}\{|h(G)-\hat{h}(G)|\}$, respectively, where $\hat{h}:\mathcal{G}\rightarrow\mathcal{Y}$ is the labeling function for graphs and $\omega^{\prime}=\min_{||g||_{Lip}\leq C_{g},||f||_{Lip}\leq C_{f}}\{\epsilon_{S}(h,\hat{h})+\epsilon_{T}(h,\hat{h})\}$, $\epsilon_{i}$ is the Rademacher variable and $p_{i}$ is the $i^{th}$ row of $\mathbf{P}$, which is the probability matrix with:
-$\mathbf{P}_{kt}=\left\{\begin{array}[]{lr}\exp\left(\frac{u_{k}(t)-V_{th}}{\sigma(u_{k}(t)-u_{reset})}\right),&if\quad u_{\theta}\leq u(t)\leq V_{th},\\ 0,&if\quad u_{reset}\leq u_{k}(t)\leq u_{\theta}.\end{array}\right.$ (30)
+[…]
+where the (empirical) source and target risks are $\hat{\epsilon}_{S}(h,\hat{h})=\frac{1}{N_{S}}\sum_{n=1}^{N_{S}}|h(G_{n})-\hat{h}(G_{n})|$ and $\epsilon_{T}(h,\hat{h})=\mathbb{E}_{\mathbb{P}_{T}(G}\{|h(G)-\hat{h}(G)|\}$, respectively, where $\hat{h}:\mathcal{G}\rightarrow\mathcal{Y}$ is the labeling function for graphs and $\omega=\min\left(|\epsilon_{S}(h,\hat{h}_{S}(\mathbf{X})))-\epsilon_{S}(h,\hat{h}_{T}(\mathbf{X})))|,|\epsilon_{T}(h,\hat{h}_{S}(\mathbf{X})))-\epsilon_{T}(h,\hat{h}_{T}(\mathbf{X})))|\right)$, $\epsilon_{i}$ is the Rademacher variable and $p_{i}$ is the $i^{th}$ row of $\mathbf{P}$, which is defined in Theorem 2.
 Proof.
@@
 $\epsilon_{T}(h,\hat{h}_{T}(\mathbf{X}_{n}))\leq\hat{\epsilon}_{T}(h,\hat{h}_{T}(S_{n}))+2\mathbb{E}\left[\sup\frac{1}{N^{\prime}_{T}}\sum_{i=1}^{N^{\prime}_{T}}\epsilon_{i}h(\mathbf{X}_{i},y_{i},p_{i})\right]+C\sqrt{\frac{ln(2/\delta)}{N^{\prime}_{T}}}.$ (36)
-Combining Eq. 31 and 32, we have:
+Combining Eq. 35 and 36, we have:
 $\epsilon_{T}(h,\hat{h}_{T}(\mathbf{X}))\overset{(a)}{\leq}$ $\frac{N^{\prime}_{T}}{N_{S}+N^{\prime}_{T}}\left(\hat{\epsilon}_{T}(h,\hat{h}_{T}(S))+2\mathbb{E}\left[\sup\frac{1}{N^{\prime}_{T}}\sum_{i=1}^{N^{\prime}_{T}}\epsilon_{i}h(\mathbf{X}_{i},y_{i},p_{i})\right]+C\sqrt{\frac{ln(2/\delta)}{N^{\prime}_{T}}}\right)$
@@
 $+\min\left(|\epsilon_{S}(h,\hat{h}_{S}(\mathbf{X})))-\epsilon_{S}(h,\hat{h}_{T}(\mathbf{X})))|,|\epsilon_{T}(h,\hat{h}_{S}(\mathbf{X})))-\epsilon_{T}(h,\hat{h}_{T}(\mathbf{X})))|\right)\Bigg{)}$
-where (a) is the outcome of applying the union bound with coefficient $\frac{N^{\prime}_{T}}{N_{S}+N^{\prime}_{T}}$, $\frac{N_{S}}{N_{S}+N^{\prime}_{T}}$ respectively; (b) additionally adopt the assumption $N^{\prime}_{T}\ll N_{S}$, following the sleight-of-hand in Li et al.
-[2021a] Theorem 3.2.
+where (a) is the outcome of applying the union bound with coefficient $\frac{N^{\prime}_{T}}{N_{S}+N^{\prime}_{T}}$, $\frac{N_{S}}{N_{S}+N^{\prime}_{T}}$ respectively; (b) additionally adopt the assumption $N^{\prime}_{T}\ll N_{S}$, following the sleight-of-hand in [Li et al., 2021] Theorem 3.2.
 Due to the sampels are selected with high confidence, thus, we have the following assumption:
 $\hat{\epsilon}_{T}\leq\epsilon_{T}\leq$ $\hat{\epsilon}_{S}(h,\hat{h}(\mathbf{X})))+2\mathbb{E}\left[\sup\frac{1}{N_{S}}\sum_{i=1}^{N_{S}}\epsilon_{i}h(\mathbf{X}_{i},y_{i},p_{i})\right]$ (37)
-$+C\sqrt{\frac{ln(2/\delta)}{N_{S}}}+2C_{f}C_{g}W_{1}(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G))+\omega^{\prime},$
-where $\omega^{\prime}=\min_{||g||_{Lip}\leq C_{g},||f||_{Lip}\leq C_{f}}\{\epsilon_{S}(h,\hat{h})+\epsilon_{T}(h,\hat{h})\}$, $\hat{\epsilon}_{T}$ is the empirical risk on the high confidence samples, $\epsilon_{T}$ is the empirical risk on the target domain.
+$+C\sqrt{\frac{ln(2/\delta)}{N_{S}}}+2C_{f}C_{g}W_{1}(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G))+\omega,$
+where $\omega=\min\left(|\epsilon_{S}(h,\hat{h}_{S}(\mathbf{X})))-\epsilon_{S}(h,\hat{h}_{T}(\mathbf{X})))|,|\epsilon_{T}(h,\hat{h}_{S}(\mathbf{X})))-\epsilon_{T}(h,\hat{h}_{T}(\mathbf{X})))|\right)$, $\hat{\epsilon}_{T}$ is the empirical risk on the high confidence samples, $\epsilon_{T}$ is the empirical risk on the target domain.
 Besides, we have:
@@
 $\leq$ $\hat{\epsilon}_{S}(h,\hat{h}_{S}(S))+2\mathbb{E}\left[\sup\frac{1}{N_{S}}\sum_{i=1}^{N_{S}}\epsilon_{i}h(\mathbf{X}_{i},y_{i},p_{i})\right]+C\sqrt{\frac{ln(2/\delta)}{N_{S}}}$
-$+2C_{f}C_{g}W_{1}\left(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G)\right)+\omega^{\prime}.$
-Appendix D Algorithm
-Algorithm 1 Learning Algorithm of DeSGDA
+$+2C_{f}C_{g}W_{1}\left(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G)\right)+\omega.$
+Appendix F Algorithm
+Algorithm 1 Learning Algorithm of DeSGraDA
 Input: Source data $\mathcal{D}^{s}$, Target data $\mathcal{D}^{t}$
-Output: Node degree-aware personalized spiking graph encoder parameters $\theta$, domain discriminator $\gamma$
-1:Initialize model parameters $\theta$ and $\gamma$
+Output: The parameters $\theta$ of Degree-Conscious spiking encoder, parameters $\gamma$ of domain discriminator, and parameter $\eta$ of semantic classifier.
+1:Initialize model parameters $\theta$, $\gamma$, and $\eta$
 2:while not converged do
 3:     Sample mini-batches $\mathcal{B}^{s}$ and $\mathcal{B}^{t}$ from $\mathcal{D}^{s}$ and $\mathcal{D}^{t}$
-4:     Forward propagate $\mathcal{B}^{s}$ and $\mathcal{B}^{t}$ through the node degree-aware personalized spiking graph encoder
+4:     Forward propagate $\mathcal{B}^{s}$ and $\mathcal{B}^{t}$ through the Degree-Conscious spiking encoder
 5:     Perform pseudo-label distillation
 6:     Compute the loss function using Eq. 9
-7:     Update model parameters $\theta$ and $\gamma$ via backpropagation
+7:     Update model parameters $\theta$, $\gamma$, and $\eta$ via backpropagation
 8:end while
-Appendix E Complexity Analysis
-Here we analyze the computational complexity of the proposed DeSGDA.
-The computational complexity primarily relies on node degree-aware personalized spiking representations.
+Appendix G Complexity Analysis
+Here we analyze the computational complexity of the proposed DeSGraDA.
+The computational complexity primarily relies on Degree-Conscious spiking representations.
 For a given graph $G$, $\|A\|_{0}$ denotes the number of nonzeros in the adjacency matrix of $G$.
 $d$ is the feature dimension.
-$L$ denote the layer number of spiking graph encoder.
+$L$ denote the layer number of spiking encoder.
 $|V|$ is the number of nodes.
-$T$ denotes the number of time latency.
-The spiking graph encoder takes $\mathcal{O}\left(T\cdot L\cdot\left(\|A\|_{0}\cdot d+|V|\cdot d^{2}\right)\right)$ computational time for each graph.
-As a result, the complexity of our DeSGDA is proportional to both $|V|$ and $\|A\|_{0}$.
-Figure 6: Visualization of different distributions on PROTEINS.
-Figure 7: Visualization of different distributions on NCI1.
-Figure 8: Visualization of different distributions on FRANKENSTEIN.
-Figure 9: Visualization of different distributions on MUTAGNENICITY.
-Appendix F Dataset
+$T$ denotes the number of latency step.
+The spiking encoder takes $\mathcal{O}\left(T\cdot L\cdot\left(\|A\|_{0}\cdot d+|V|\cdot d^{2}\right)\right)$ computational time for each graph.
+As a result, the complexity of our DeSGraDA is proportional to both $|V|$ and $\|A\|_{0}$.
+Figure 5: Visualization of different distributions on the PROTEINS dataset.
+Figure 6: Visualization of different distributions on the DD dataset.
+Figure 7: Visualization of edge density distributions on the BCI (left) and EEG (right) datasets.
+Appendix H Dataset
 Table 4: Statistics of the experimental datasets.
 Datasets Graphs Avg. Nodes Avg. Edges Classes
-PROTEINS 1,113 39.1 72.8 2
-NCI1 4,110 29.87 32.30 2
-MUTAGENICITY 4,337 30.32 30.77 2
-FRANKENSTEIN 4,337 16.9 17.88 2
-F.1 Dataset Description
-We conduct extensive experiments on four public benchmark graph datasets from TUDataset. The dataset statistics can be found in Table 3, and their details are shown as follows:
-  • PROTEINS.
-The PROTEINS dataset Dobson and Doig [2003] consists of protein graphs.
-Each label indicates whether the protein graph is an enzyme.
-Nodes within these graphs represent amino acids connected by edges if neighbors along the amino acid sequence are spaced less than 6 Angstroms apart.
-We divide the PROTEINS dataset into four parts, i.e., P0, P1, P2, and P3, based on edge density, node density, and graph flux; the sub-datasets exhibit substantial domain disparities among them.
-  • NCI1.
-The NCI1 Wale et al.
-[2008] dataset contains 4,100 chemical compounds with atoms as nodes and bonds as edges.
-Each label indicates the characteristics that hinder the growth of cancer cells.
-Like the PROTEINS dataset, we divide the NCI1 dataset into four parts, i.e., N0, N1, N2, and N3, based on edge density, node density, and graph flux.
-  • FRANKENSTEIN.
-The FRANKENSTEIN Orsini et al.
-[2015] dataset consists of 4,337 molecular graphs, where nodes represent atoms and edges depict chemical bonds.
-Each graph is labeled to classify molecules based on their biological activity.
-Like the PROTEINS dataset, the entire FRANKENSTEIN dataset is divided into four segments (F0, F1, F2, and F3) based on edge density, node density, and graph flux.
-  • MUTAGENICITY.
-The MUTAGENICITY Kazius et al.
-[2005] dataset contains 4,337 chemical compounds, each represented as a graph where nodes represent atoms and edges indicate bonds.
-Each graph can be used to identify mutagenic compounds, aiding studies in toxicology and chemical safety.
-Like the PROTEINS dataset, the entire MUTAGENICITY dataset is divided into four segments (M0, M1, M2, and M3) based on edge density, node density, and graph flux.
-Table 4: The results of ablation studies on NCI1 (source → target). Bold results indicate the best performance per column.
+SEED 3,818 62.00 125.74 3
+BCI 1,440 22.00 119.80 2
+PROTEINS 1,113 39.10 72.80 2
+DD 1,178 284.32 715.66 2
+COX2 467 41.22 43.45 2
+COX2_MD 303 26.28 335.12 2
+BZR 405 35.75 38.36 2
+BZR_MD 306 21.30 225.06 2
+H.1 Dataset Description
+We conduct extensive experiments on different types of datasets. The dataset statistics can be found in Table 4, and their details are shown as follows:
+  • SEED.
+The SJTU Emotion EEG Dataset (SEED) [Zheng and Lu, 2015, Duan et al., 2013] is a widely used Electroencephalography (EEG) benchmark for emotion recognition.
+It contains EEG recordings from 15 participants watching emotional movie clips that evoke positive, neutral, and negative emotions.
+Each recording consists of 62-channel EEG signals sampled at 1000 Hz.
+Based on the edge density, we partition the SEED dataset into three sub-datasets, namely S0, S1, and S2.
+The sub-datasets exhibit substantial domain disparities among them.
+  • BCI.
+The Brain-Computer Interface (BCI) Competition IV-2a dataset Brunner et al.
+[2008] is a widely used benchmark for motor imagery EEG classification.
+It includes EEG recordings from 9 subjects performing four motor imagery tasks: left hand, right hand, feet, and tongue.
+Each subject completed two sessions on different days, with signals recorded from 22 EEG channels at a sampling rate of 250 Hz.
+Based on the edge density, we partition the dataset into three parts: B0, B1, and B2.
+  • PROTEINS and DD.
+The PROTEINS [Dobson and Doig, 2003] and DD datasets comprise protein structure graphs commonly used for graph classification tasks.
+In both datasets, nodes represent amino acids, forming edges between spatially or sequentially adjacent residues.
+In PROTEINS, each graph is labeled to indicate whether the protein is an enzyme, with edges defined between residues less than 6 Angstroms apart.
+The DD dataset, derived from the Protein Data Bank, focuses on classifying proteins by structural class and typically exhibits denser graph connectivity than PROTEINS.
+Additionally, we partition the PROTEINS and DD datasets into four parts based on edge density and node density, P0 to P3 and D0 to D3, respectively.
+  • COX2 and COX2_MD.
+The COX2 dataset Sutherland et al.
+[2003] consists of 467 molecular graphs, while COX2_MD contains 303 structurally modified counterparts.
+In both datasets, nodes represent atoms and edges correspond to chemical bonds.
+Specifically, COX2_MD introduces structural variations to the original COX2 molecules while preserving their semantic labels.
+  • BZR and BZR_MD.
+The BZR dataset Sutherland et al.
+[2003] comprises 405 molecular graphs, while BZR_MD includes 306 structurally modified graphs derived from BZR.
+In both datasets, nodes represent atoms and edges denote chemical bonds.
+BZR_MD introduces structural variations to simulate domain shifts while preserving the original label semantics.
+H.2 Data processing
+In our implementation, we process the above datasets as follows:
+  • For datasets from TUDataset  [^3] https://chrsmrrs.github.io/datasets/, including PROTEINS, DD, BZR, BZR_MD, COX2, and COX2_MD, we utilize the TUDataset module from PyTorch Geometric for loading.
+Self-loops are added during the preprocessing stage to prevent isolated nodes.
+  • For the SEED dataset, we utilize the TorchEEG library [^4] https://torcheeg.readthedocs.io/en/latest/ to transform raw EEG signals into graph-structured data.
+During graph construction, edges are removed following the approach described in Klepl et al.
+[2022].
+  • For the BCI dataset, we focus on a binary classification subset involving left-hand and right-hand motor imagery tasks, a widely adopted evaluation setting in BCI research.
+Following the construction protocols proposed in Altaheri et al.
+[2022, 2023], we randomly remove a portion of edges from each graph during preprocessing.
+Appendix I Baselines
+In this part, we introduce the details of the compared baselines as follows:
+Graph kernel method. We compare DeSGraDA with one graph kernel method:
+  • WL subtree: Weisfeiler-Lehman (WL) subtree [Shervashidze et al., 2011] is a graph kernel method, which calculates the graph similarity by a kernel function, where it encodes local neighborhood structures into subtree patterns, efficiently capturing the topology information contained in graphs.
+Graph-based methods. We compare DeSGraDA with four widely used graph-based methods:
+  • GCN: GCN [Kipf and Welling, 2017] is a spectral-based neural network that iteratively updates node representations by aggregating information from neighboring nodes, effectively capturing both local graph structure and node features.
+  • GIN: GIN [Xu et al., 2018] is a message-passing neural network designed to distinguish graph structures using an injective aggregation function, theoretically achieving the expressive power of the Weisfeiler-Lehman test.
+  • CIN: CIN [Bodnar et al., 2021] extends the Weisfeiler-Lehman framework by integrating cellular complexes into graph neural networks, allowing for the capture of higher-dimensional topological features.
+  • GMT: GMT [Baek et al., 2021] utilizes self-attention mechanisms to dynamically adjust the importance of nodes based on their structural dependencies, thereby enhancing both adaptability and performance.
+Spiking-based graph methods. We compare DeSGraDA with two spiking-based graph methods:
+  • SpikeGCN: SpikeGCN [Zhu et al., 2022] introduces an end-to-end framework designed to integrate the fidelity characteristics of SNNs with graph node representations.
+  • DRSGNN: DRSGNN [Zhao et al., 2024] dynamically adapts to evolving graph structures and relationships through a novel architecture that updates node representations in real-time.
+Domain adaption methods. We compare DeSGraDA with two recent domain adaption methods:
+  • CDAN: CDAN [Long et al., 2018] employs a conditional adversarial learning strategy to reduce domain discrepancy by conditioning adversarial adaptation on discriminative information from multiple domains.
+  • ToAlign: ToAlign [Wei et al., 2021b] uses token-level alignment strategies within Transformer architectures to enhance cross-lingual transfer, optimizing the alignment of semantic representations across languages.
+  • MetaAlign: MetaAlign [Wei et al., 2021a] is a meta-learning framework for domain adaptation that dynamically aligns feature distributions across domains by learning domain-invariant representations.
+Graph domain adaptation methods. We compare DeSGraDA with six graph domain adaption methods:
+  • DEAL: DEAL [Yin et al., 2022] uses domain adversarial learning to align graph representations across different domains without labeled data, overcoming discrepancies between the source and target domains.
+  • CoCo: CoCo [Yin et al., 2023] leverages contrastive learning to align graph representations between source and target domains, enhancing domain adaptation by promoting intra-domain cohesion and inter-domain separation in an unsupervised manner.
+  • SGDA: SGDA [Qiao et al., 2023] utilizes labeled data from the source domain and a limited amount of labeled data from the target domain to learn domain-invariant graph representations.
+  • StruRW: StruRW [Liu et al., 2023] introduces a structural re-weighting mechanism that dynamically adjusts the importance of nodes and edges based on their domain relevance.
+It enhances feature alignment by emphasizing transferable structures while suppressing domain-specific noise.
+  • A2GNN: A2GNN [Liu et al., 2024a] introduces a novel propagation mechanism to enhance feature transferability across domains, improving the alignment of graph structures and node features in an unsupervised setting.
+  • PA-BOTH: PA-BOTH [Liu et al., 2024b] aligns node pairs between source and target graphs, optimizing feature correspondence at a granular level to improve the transferability of structural and feature information across domains.
+Table 5: GPU memory consumption of different graph domain methods in training stage for each training epoch (in GB).
+DeSGraDA DEAL CoCo SGDA StruRW A2GNN PA-BOTH
+PROTEINS 1.0 1.2 1.2 1.0 1.2 22.3 1.7
+DD 5.6 6.4 2.5 3.9 4.5 35.1 16.8
+SEED 1.1 1.4 1.5 2.6 0.8 16.8 2.8
+BCI 0.7 0.8 0.7 0.7 0.7 14.5 1.6
+Table 6: Time consumption of different graph domain methods in training stage for each training epoch (in seconds).
+DeSGraDA DEAL CoCo SGDA StruRW A2GNN PA-BOTH
+PROTEINS 0.195 0.103 22.123 0.088 0.088 1.313 0.949
+DD 0.427 0.400 184.015 0.135 0.140 2.263 0.787
+SEED 0.192 0.137 26.187 0.126 0.075 1.414 0.311
+BCI 0.224 0.211 35.162 0.103 0.086 0.781 0.123
+Appendix J Implementation Details
+DeSGraDA and all baseline models are implemented using PyTorch [^5] https://pytorch.org/ and PyTorch Geometric [^6] https://www.pyg.org/.
+We conduct experiments for DeSGraDA and all baselines on NVIDIA A100 GPUs for a fair comparison, where the learning rate of Adam optimizer set to ${10^{-4}}$, hidden embedding dimension 256, weight decay ${10^{-12}}$, and GNN layers 4.
+Additionally, DeSGraDA and all baseline models are trained using all labeled source samples and evaluated on unlabeled target samples [Wu et al., 2020].
+The performances of all models are measured and averaged on all samples for five different runs.
+Appendix K More experimental results
+K.1 More performance comparison
+In this part, we provide additional results for our proposed method DeSGraDA compared with all baseline models across various datasets, as illustrated in Table 11-14.
+These results consistently show that DeSGraDA outperforms the baselines in most cases, validating the superiority of our proposed method.
+K.2 Training time and memory comparison
+We provide detailed comparisons of GPU memory consumption and training time per epoch for DeSGraDA and other graph domain adaptation methods under identical experimental settings in this part, as shown in Tables 5 and 6.
+It is worth noting that the training phase is typically conducted on more powerful hardware to achieve optimal performance within a reasonable time frame.
+Table 7: The results of DeSGraDA with different widely used graph neural networks (GIN, GCN and SAGE) on the PROTEINS and DD dataset. Bold results indicate the best performance.
+Methods PROTEINS DD
 […]
-DeSGDA w/o CA 58.9 62.6 61.7 60.1 58.2 62.0 63.4 61.7 63.5 60.7 68.3 64.0 62.1
-DeSGDA w/o PL 62.8 65.1 55.0 56.9 55.3 54.4 58.6 62.9 61.2 62.2 59.7 64.3 59.8
-DeSGDA w/o CF 52.7 62.6 58.5 56.5 56.3 63.8 63.1 52.7 65.4 65.1 58.2 63.7 59.9
-DeSGDA w/o TL 64.6 62.8 60.6 54.9 44.9 65.4 60.8 64.8 56.1 64.1 57.8 54.0 59.3
-DeSGDA w PT 64.4 59.2 64.6 62.2 64.5 60.1 61.1 60.8 62.0 64.2 62.7 65.5 62.6
-DeSGDA w CL 65.3 70.1 68.9 65.2 65.6 63.0 67.6 64.6 68.9 66.8 68.6 70.7 67.1
-DeSGDA 68.5 71.4 70.1 69.0 68.9 66.3 69.6 70.2 71.1 69.3 74.4 70.0 69.9
-Table 5: The results of ablation studies on FRANKENSTEIN (source → target). Bold results indicate the best performance.
+DeSGraDA w GCN 71.9 76.8 64.8 68.6 58.2 70.2 57.6 64.1
+DeSGraDA w SAGE 73.9 79.8 65.5 74.5 57.8 71.6 59.1 66.7
+DeSGraDA w GIN 74.1 81.5 66.4 78.7 59.3 73.0 61.1 68.8
+DeSGraDA 76.3 84.6 69.2 83.6 60.1 76.1 63.9 71.7
+Table 8: The results of DeSGraDA with different widely used graph neural networks (GIN, GCN and SAGE) on the SEED and BCI dataset. Bold results indicate the best performance.
+Methods SEED BCI
 […]
-DeSGDA w/o CA 59.6 60.7 60.2 60.7 77.9 58.5 61.3 60.5 76.8 60.0 80.1 63.9 65.0
-DeSGDA w/o PL 58.9 56.6 55.9 56.6 69.8 53.6 55.0 57.7 70.4 53.2 68.8 61.7 59.9
-DeSGDA w/o CF 55.9 59.2 59.5 54.5 74.9 52.9 59.9 54.6 78.8 56.1 74.5 58.5 61.6
-DeSGDA w/o TL 56.2 57.2 60.7 48.2 53.3 45.8 59.5 53.0 48.8 54.8 51.1 54.7 53.6
-DeSGDA w PT 60.9 60.2 61.0 61.5 76.2 60.9 60.3 62.2 75.3 57.2 74.8 63.3 64.5
-DeSGDA w CL 62.2 61.3 61.7 60.7 79.3 59.7 63.6 64.1 79.5 61.1 78.6 62.7 66.2
-DeSGDA 63.5 64.1 62.6 63.0 81.1 62.4 65.5 65.4 81.9 60.8 82.0 65.9 68.2
-Table 6: The results of ablation studies on MUTAGENICITY (source → target). Bold results indicate the best performance.
-[…]
-DeSGDA w/o CA 58.3 63.3 62.1 60.1 81.4 62.7 63.8 59.1 80.3 57.1 78.9 63.0 65.8
-DeSGDA w/o PL 58.8 57.6 60.2 61.4 73.3 54.5 63.7 61.0 61.2 61.9 71.0 60.2 62.1
-DeSGDA w/o CF 54.7 58.3 59.6 58.7 72.1 60.7 57.9 54.9 78.0 57.4 63.2 60.4 61.3
-DeSGDA w/o TL 56.5 60.7 59.1 60.7 43.4 56.3 51.0 56.7 53.3 56.9 63.1 52.5 55.9
-DeSGDA w PT 62.1 64.3 63.9 61.8 77.3 62.4 66.6 63.9 73.5 66.1 82.1 64.9 67.4
-DeSGDA w CL 66.1 62.6 63.3 64.7 81.5 60.7 69.8 67.0 82.2 67.8 82.4 63.6 69.3
-DeSGDA 65.4 65.9 65.5 65.6 82.8 63.6 70.7 68.2 82.9 67.6 83.9 66.5 70.8
-F.2 Data processing
-In our implementation, we process these four TUDatasets by adding a self-loop connection for each node.
-Additionally, we utilize one-hot embeddings to represent node attributes for datasets (FRANKENSTEIN and MUTAGENICITY) where node features are unavailable.
-Appendix G Baselines
-In this part, we introduce the details of the compared baselines as follows:
-Graph kernel method. We compare DeSGDA with one graph kernel method:
-  • WL subtree: Weisfeiler-Lehman (WL) subtree Shervashidze et al.
-[2011] is a graph kernel method, which calculates the graph similarity by a kernel function, where It encodes local neighborhood structures into subtree patterns, efficiently capturing the topology information contained in graphs.
-Graph neural networks. We compare DeSGDA with four widely used general graph neural networks:
-  • GCN: GCN Kipf and Welling [2017] is a spectral-based neural network that iteratively updates node representations by aggregating information from neighboring nodes, effectively capturing both local graph structure and node features.
-  • GIN: GIN Xu et al.
-[2018] is a message-passing neural network designed to distinguish graph structures using an injective aggregation function, theoretically achieving the expressive power of the Weisfeiler-Lehman test.
-  • CIN: CIN Bodnar et al.
-[2021] extends the Weisfeiler-Lehman framework by integrating cellular complexes into graph neural networks, allowing for the capture of higher-dimensional topological features.
-  • GMT: GMT Baek et al.
-[2021] utilizes self-attention mechanisms to dynamically adjust the importance of nodes based on their structural dependencies, thereby enhancing both adaptability and performance.
-Spiking graph neural networks. We compare DeSGDA with two spiking graph neural networks:
-  • SpikeGCN: SpikeGCN Zhu et al. [2022] introduces an end-to-end framework designed to integrate the fidelity characteristics of SNNs with graph node representations.
-  • DRSGNN: DRSGNN Zhao et al. [2024] dynamically adapts to evolving graph structures and relationships through a novel architecture that updates node representations in real-time.
-Domain adaption methods. We compare DeSGDA with two recent domain adaption methods:
-  • CDAN: CDAN Long et al.
-[2018] employs a conditional adversarial learning strategy to reduce domain discrepancy by conditioning adversarial adaptation on discriminative information from multiple domains.
-  • ToAlign: ToAlign Wei et al.
-[2021b] uses token-level alignment strategies within Transformer architectures to enhance cross-lingual transfer, optimizing the alignment of semantic representations across languages.
-  • MetaAlign: MetaAlign Wei et al.
-[2021a] is a meta-learning framework for domain adaptation that dynamically aligns feature distributions across domains by learning domain-invariant representations.
-Graph domain adaptation methods. We compare DeSGDA with six SOTA graph domain adaption methods:
-  • DEAL: DEAL Yin et al.
-[2022b] uses domain adversarial learning to align graph representations across different domains without labeled data, overcoming discrepancies between the source and target domains.
-  • CoCo: CoCo Yin et al.
-[2023a] leverages contrastive learning to align graph representations between source and target domains, enhancing domain adaptation by promoting intra-domain cohesion and inter-domain separation in an unsupervised manner.
-  • SGDA: SGDA Qiao et al. [2023] utilizes labeled data from the source domain along with a limited amount of labeled data from the target domain to learn domain-invariant graph representations.
-  • DGDA: DGDA Cai et al.
-[2024] employs generative models to capture the underlying distribution of graph data across domains, facilitating the transfer of graph structures and features by learning shared latent spaces.
-  • A2GNN: A2GNN Liu et al.
-[2024a] introduces a novel propagation mechanism to enhance feature transferability across domains, improving the alignment of graph structures and node features in an unsupervised setting.
-  • PA-BOTH: PA-BOTH Liu et al.
-[2024b] aligns node pairs between source and target graphs, optimizing feature correspondence at a granular level to improve the transferability of structural and feature information across domains.
-For GCN and GIN, we use the Pytorch Geometric  [^4] https://www.pyg.org/ to implement the model.
-For other baseline methods, we use the source codes provided by the corresponding paper.
-For all baseline methods, we vary the dropout rate in the range of {0.1,0.3,0.5,0.7} and then choose the best one.
-The hidden dimension in these methods is set to 256 for a fair comparison.
-Table 7: Statistics of the SEED dataset.
-Datasets Graphs Avg. Nodes Avg. Edges Classes
-SEED 7,636 62.0 168.2 3
-[…]
-[…]
-GCN 46.0 $\pm$ 0.9 47.8 $\pm$ 1.0 47.7 $\pm$ 1.4 49.7 $\pm$ 0.7 51.3 $\pm$ 0.8 49.8 $\pm$ 1.0
-GIN 48.9 $\pm$ 0.5 50.3 $\pm$ 0.6 49.5 $\pm$ 0.7 50.7 $\pm$ 1.0 52.7 $\pm$ 1.0 52.1 $\pm$ 0.9
-DEAL 53.5 $\pm$ 0.4 56.2 $\pm$ 0.7 53.2 $\pm$ 0.8 53.7 $\pm$ 1.1 55.1 $\pm$ 0.8 56.0 $\pm$ 0.7
-CoCo 53.9 $\pm$ 0.5 53.0 $\pm$ 0.6 54.1 $\pm$ 0.7 54.3 $\pm$ 0.7 55.3 $\pm$ 1.0 55.9 $\pm$ 0.6
-DeSGDA 54.5 $\pm$ 0.6 55.6 $\pm$ 0.7 54.6 $\pm$ 0.5 54.5 $\pm$ 0.7 55.8 $\pm$ 1.1 56.6 $\pm$ 0.8
-Table 9: GPU memory consumption of different graph domain methods in training stage for each training epoch (in GB).
-DeSGDA DEAL CoCo A2GNN PA-BOTH
-PROTEINS 2.3 1.4 1.2 33.1 5.6
-NCI1 5.7 2.9 3.3 34.8 11.9
-MUTAGENICITY 5.0 2.8 3.2 35.0 12.3
-FRANKENSTEIN 2.9 2.4 1.9 33.6 7.3
-Table 10: Time consumption of different graph domain methods in training stage for each training epoch (in seconds).
-DeSGDA DEAL CoCo A2GNN PA-BOTH
-PROTEINS 0.587 0.126 22.123 0.869 0.283
-NCI1 0.855 0.518 58.564 1.483 0.597
-MUTAGENICITY 0.887 0.612 52.740 1.553 0.511
-FRANKENSTEIN 0.663 0.375 26.837 1.016 0.275
-Appendix H More experimental results
-H.1 More performance comparison
-In this part, we provide additional results for our proposed method DeSGDA compared with all baseline models across various datasets, as illustrated in Table 1 and Table 13-23.
-These results consistently show that DeSGDA outperforms the baselines in most cases, validating the superiority of our proposed method.
-Additionally, the performance of DeSGDA with different GNN architectures on the NCI1, FRANKENSTEIN, and MUTAGENICITY datasets is shown in Figure 10.
-It is evident that GIN consistently outperforms other GNN architectures in most cases.
-To verify the efffectiveness of DeSGDA in EEG data and multi-class scenarios, we conducted additional experiments using the SEED dataset Duan et al.
-[2013], Zheng and Lu [2015], which is a well-known EEG dataset for emotion classification.
-For EEG data processing, we utilized the torcheeg  [^5] https://github.com/torcheeg/torcheeg library to convert standard EEG data into graph structures.
-During graph construction, we remove edges from each graph and partitioning the dataset into source and target domains based on edge density Klepl et al.
-[2022].
-The statistics of SEED dataset is shown in Table 7.
-We compared the DeSGDA with two general graph neural networks (GCN Kipf and Welling [2017] and GIN Xu et al.
-[2018]) and two graph domain adaptation methods (DEAL Yin et al.
-[2022b] and CoCo Yin et al.
-[2023a]).
-The results in Table 8 show that DeSGDA still outperforms the other methods in most cases.
-H.2 Training time and memory comparison
-We provide detailed comparisons of GPU memory consumption and training time per epoch for DeSGDA and other graph domain adaptation methods under identical experimental settings in this part, as shown in Tables 9 and 10.
-It is worth noting that the training phase is typically conducted on more powerful hardware to achieve optimal performance within a reasonable time frame.
-Table 11: The results of DeSGDA with different widely used graph neural networks (GIN, GCN and SAGE) on PROTEINS and NCI1 dataset. Bold results indicate the best performance.
-[…]
-DeSGDA w GCN 76.6 70.5 71.8 74.1 66.3 68.3 68.5 67.1
-DeSGDA w SAGE 75.8 73.3 72.4 75.2 67.2 69.5 66.6 68.4
-DeSGDA w GIN 77.3 75.8 73.8 77.1 69.0 69.8 68.8 68.8
-DeSGDA 78.7 78.4 74.8 77.6 68.5 71.4 70.1 69.0
-Table 12: The results of DeSGDA with different widely used graph neural networks (GIN, GCN and SAGE) on MUTAGENICITY and FRANKENSTEIN dataset. Bold results indicate the best performance.
-[…]
-DeSGDA w GCN 60.9 60.7 63.6 60.6 61.3 62.2 60.7 61.1
-DeSGDA w SAGE 61.1 62.3 64.2 61.1 61.9 62.6 61.3 61.6
-DeSGDA w GIN 64.5 65.4 65.0 63.9 62.8 63.6 61.8 62.8
-DeSGDA 65.4 65.9 65.5 65.6 63.5 64.1 62.6 63.0
-H.3 More Ablation study
-To validate the effectiveness of the different components in DeSGDA, we conduct more experiments with sive variants on NCI1, FRANKENSTEIN and MUTAGENICITY datasets, i.e., DeSGDA w/o CA, DeSGDA w/o PL, DeSGDA w/o CF, DeSGDA w/o TL, DeSGDA w PT and DeSGDA w CL.
-The results are shown in Table 4 , 5 and 6.
+DeSGraDA w GCN 55.7 54.5 53.7 54.0 52.8 51.6 52.2 54.2
+DeSGraDA w SAGE 55.3 54.5 54.2 54.9 52.7 52.2 52.8 54.7
+DeSGraDA w GIN 56.6 57.1 55.8 56.9 53.3 52.9 53.8 55.4
+DeSGraDA 58.0 58.2 57.0 58.3 54.1 53.6 54.9 56.2
+K.3 More Ablation study
+To validate the effectiveness of the different components in DeSGraDA, we conduct more experiments with four variants on DD, SEED, and BCI datasets, i.e., DeSGraDA w CDAN, DeSGraDA w/o PL, DeSGraDA w/o CF and DeSGraDA w/o TL.
+The results are shown in Table 9 and 10.
 From the results, we have similar observations as summarized in Section 5.3.
-Additionally, we conduct ablation studies to examine the effect of directly replacing the SGNs with commonly used Graph Neural Networks (GNNs) for generating representations for DeSGDA: (1) DeSGDA w GCN: It replaces SGNs with GCN Kipf and Welling [2017]; (2) DeSGDA w GIN: It replaces SGNs with GIN Xu et al.
-[2018]; (3) DeSGDA w SAGE: It replaces SGNs with GraphSAGE Hamilton et al.
-[2017].
-The experimental results across the PROTEINS, NCI1, MUTAGENICITY, and FRANKENSTEIN datasets are shown in Table 11 and 12.
+Additionally, we conduct ablation studies to examine the effect of directly replacing the SGNs with commonly used Graph Neural Networks (GNNs) for generating representations for DeSGraDA: (1) DeSGraDA w GCN: It replaces SGNs with GCN [Kipf and Welling, 2017]; (2) DeSGraDA w GIN: It replaces SGNs with GIN [Xu et al., 2018]; (3) DeSGraDA w SAGE: It replaces SGNs with GraphSAGE [Hamilton et al., 2017].
+The experimental results across the PROTEINS, DD, SEED, and BCI datasets are shown in Table 7 and 8.
 However, the critical aspect of our work lies in the specific problem we set up, i.e., low-power and distribution shift environments.
 In this context, directly replacing SGNs with commonly used GNNs like GIN or GCN is not feasible, as these models are unsuitable for deployment on low-energy devices.
-As demonstrated in Section 5.3, GNN based methods have much higher energy consumption than the spike based methods.
-H.4 More Sensitivity Analysis
-In this part, we provide additional sensitivity analysis of the proposed DeSGDA with respect to the impact of its hyperparameters: the time latency $\tau$, degree threshold value $V_{th}^{degree}$ in SNNs, and balance ratio $\alpha$ on NCI1, FRANKENSTEIN, and MUTAGENICITY datasets.
-The results are illustrated in Figure 11, 12 and 13, where we observe trends similar to those discussed in Section 5.5.
-(a) NCI1
-(b) FRANKENSTEIN
-(c) MUTAGENICITY
-Figure 10: The performance with different GNN architectures on different datasets.
-(a) Time Latency
-(b) Initial Threshold
-(c) Balance Tatio
-Figure 11: Hyperparameter sensitivity analysis of time latency $\tau$, initial threshold $V_{th}^{degree}$ in SNNs, and balance ratio $\alpha$ on NCI1.
-(a) Time Latency
-(b) Initial Threshold
-(c) Balance Ratio
-Figure 12: Hyperparameter sensitivity analysis of time latency $\tau$, initial threshold $V_{th}^{degree}$ in SNNs, and balance ratio $\alpha$ on FRANKENSTEIN.
-(a) Time Latency
-(b) Initial Threshold
-(c) Balance Ratio
-Figure 13: Hyperparameter sensitivity analysis of time latency $\tau$, initial threshold $V_{th}^{degree}$ in SNNs, and balance ratio $\alpha$ on MUTAGENICITY.
+As demonstrated in Section 5.4, GNN-based methods have much higher energy consumption than the spike-based methods.
+K.4 More Sensitivity Analysis
+In this part, we provide additional sensitivity analysis of the proposed DeSGraDA with respect to the impact of its hyperparameters: the latency step $T$ and initial threshold value $V_{th}^{degree}$ in SNNs on the DD, SEED and BCI datasets.
+The results are illustrated in Figure 8 and 9, where we observe trends similar to those discussed in Section 5.5.
+(a) DD
+(b) SEED
+(c) BCI
+Figure 8: Hyperparameter sensitivity analysis of latency step $\mathit{T}$ in SNNs on the DD, SEED and BCI datasets.
+(a) DD
+(b) SEED
+(c) BCI
+Figure 9: Hyperparameter sensitivity analysis of the initial threshold $V_{th}^{degree}$ in SNNs on the DD, SEED and BCI datasets.
+Table 9: The results of ablation studies on the DD dataset (source → target). Bold results indicate the best performance.
 […]
+DeSGraDA w CDAN 57.3 72.7 59.9 68.8 57.4 58.1 68.1 60.9 75.9 59.3 80.2 77.9
+DeSGraDA w/o PL 58.3 71.8 58.1 69.8 57.9 58.2 67.1 59.3 75.1 60.7 79.1 76.8
+DeSGraDA w/o CF 51.0 62.0 57.1 61.8 54.0 51.5 59.0 54.3 62.3 52.9 64.1 69.3
+DeSGraDA w/o TL 58.0 72.0 60.4 68.2 58.4 58.8 68.7 60.8 77.6 60.5 76.6 77.4
+DeSGraDA 60.1 76.1 63.9 71.7 61.1 62.3 73.6 64.2 80.9 64.7 82.0 80.6
+Table 10: The results of ablation studies on the SEED and BCI datasets (source → target). Bold results indicate the best performance.
+Methods SEED BCI
 […]
-WL subtree 75.9 70.4 64.3 63.9 60.6 64.7 73.2 78.9 66.8 69.2 74.2 72.9 69.6
-GCN 49.2$\pm$1.7 55.8$\pm$1.5 46.8$\pm$0.5 54.6$\pm$2.2 43.4$\pm$0.6 46.7$\pm$0.2 50.0$\pm$1.8 57.2$\pm$2.2 44.2$\pm$0.4 51.6$\pm$0.8 62.7$\pm$2.1 56.8$\pm$1.3 51.6
-GIN 68.8$\pm$2.5 70.6$\pm$1.0 64.2$\pm$1.1 67.2$\pm$2.4 62.2$\pm$1.8 62.5$\pm$1.5 68.7$\pm$2.4 72.5$\pm$0.6 63.3$\pm$1.6 65.2$\pm$0.6 62.4$\pm$0.3 70.9$\pm$0.5 66.6
-GMT 66.7$\pm$0.3 58.2$\pm$0.5 63.9$\pm$0.3 58.4$\pm$0.3 63.8$\pm$0.4 56.7$\pm$0.5 63.9$\pm$0.7 66.3$\pm$1.0 63.8$\pm$1.1 66.6$\pm$0.4 63.8$\pm$0.2 62.6$\pm$0.7 62.9
-CIN 58.7$\pm$2.4 54.9$\pm$0.2 52.0$\pm$0.3 54.8$\pm$0.1 56.6$\pm$0.2 54.9$\pm$0.1 52.9$\pm$1.4 52.8$\pm$0.5 56.5$\pm$0.6 52.8$\pm$2.1 58.5$\pm$0.8 56.6$\pm$1.4 55.1
-SpikeGCN 58.9$\pm$0.9 65.2$\pm$1.2 60.8$\pm$1.3 62.0$\pm$1.5 62.3$\pm$0.8 58.7$\pm$1.6 64.1$\pm$0.6 66.7$\pm$1.3 60.5$\pm$1.7 63.8$\pm$1.4 62.2$\pm$2.1 61.1$\pm$1.5 62.1
-DRSGNN 58.0$\pm$0.6 64.3$\pm$1.1 61.2$\pm$0.8 62.2$\pm$1.0 62.9$\pm$1.5 64.0$\pm$1.3 60.6$\pm$1.6 64.0$\pm$1.4 67.6$\pm$2.1 62.4$\pm$1.9 71.3$\pm$2.3 68.8$\pm$2.0 63.9
-CDAN 64.0$\pm$1.1 68.1$\pm$0.3 60.1$\pm$0.5 64.0$\pm$1.3 60.9$\pm$0.2 57.8$\pm$1.0 64.3$\pm$1.6 61.2$\pm$0.2 66.3$\pm$0.7 59.0$\pm$0.5 68.9$\pm$0.3 63.7$\pm$0.6 63.2
-ToAlign 52.8$\pm$0.5 54.8$\pm$0.2 48.2$\pm$1.1 54.8$\pm$1.5 44.0$\pm$0.8 54.8$\pm$2.0 48.2$\pm$1.7 52.8$\pm$0.6 44.0$\pm$0.2 52.8$\pm$0.3 44.0$\pm$1.0 48.2$\pm$1.2 50.0
-MetaAlign 63.1$\pm$0.3 63.8$\pm$1.3 58.9$\pm$2.4 58.5$\pm$0.4 59.1$\pm$2.1 59.2$\pm$1.6 70.1$\pm$0.8 63.3$\pm$1.4 66.5$\pm$2.7 60.9$\pm$1.1 71.4$\pm$0.2 67.5$\pm$0.8 63.5
-DEAL 70.7$\pm$0.9 72.3$\pm$0.2 69.9$\pm$0.8 68.9$\pm$0.7 64.1$\pm$0.6 65.6$\pm$0.9 71.9$\pm$0.4 69.9$\pm$1.7 70.6$\pm$0.4 66.5$\pm$0.3 71.6$\pm$0.7 69.9$\pm$0.5 69.3
-CoCo 64.0$\pm$1.3 63.9$\pm$0.6 65.8$\pm$1.8 59.9$\pm$1.7 62.2$\pm$2.1 60.6$\pm$1.6 65.0$\pm$2.1 64.8$\pm$1.4 60.0$\pm$0.8 61.3$\pm$0.5 68.5$\pm$0.4 67.1$\pm$0.6 63.6
-SGDA OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
-DGDA OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
-A2GNN 58.9$\pm$0.9 60.1$\pm$0.7 59.8$\pm$1.2 59.4$\pm$1.0 62.3$\pm$1.5 60.9$\pm$1.6 61.6$\pm$1.3 59.9$\pm$1.9 64.9$\pm$1.6 62.9$\pm$2.1 65.4$\pm$1.5 63.3$\pm$2.3 61.7
-PA-BOTH 61.1$\pm$0.5 60.9$\pm$0.4 61.6$\pm$0.6 61.2$\pm$0.8 60.8$\pm$0.6 61.5$\pm$0.5 62.2$\pm$1.0 61.9$\pm$0.7 61.8$\pm$1.1 61.1$\pm$0.9 60.9$\pm$0.8 61.3$\pm$1.2 61.6
-DeSGDA 68.5$\pm$1.2 71.4$\pm$1.3 70.1$\pm$0.7 69.0$\pm$1.1 68.9$\pm$1.0 66.3$\pm$1.4 69.6$\pm$1.3 70.2$\pm$1.7 71.1$\pm$1.6 69.3$\pm$1.8 74.4$\pm$1.6 70.0$\pm$1.9 69.9
-[…]
-[…]
-WL subtree 58.4 51.8 58.7 51.3 64.3 48.9 64.9 58.9 78.5 54.6 57.1 61.3 59.1
-GCN 56.2$\pm$0.2 59.0$\pm$1.3 41.4$\pm$0.4 45.8$\pm$0.5 21.2$\pm$0.7 41.4$\pm$1.7 42.5$\pm$1.6 49.0$\pm$0.4 24.1$\pm$1.6 44.8$\pm$0.7 81.4$\pm$0.3 58.8$\pm$0.2 47.1
-GIN 60.7$\pm$0.6 58.0$\pm$1.0 61.0$\pm$2.3 58.9$\pm$2.3 77.5$\pm$2.2 45.3$\pm$2.5 62.5$\pm$0.2 59.2$\pm$3.0 71.4$\pm$2.8 49.8$\pm$1.7 77.9$\pm$1.4 59.9$\pm$0.5 61.8
-GMT 56.2$\pm$0.4 59.8$\pm$0.2 41.4$\pm$0.3 59.8$\pm$0.7 21.2$\pm$1.1 59.8$\pm$0.5 41.4$\pm$0.2 56.2$\pm$0.2 21.1$\pm$1.1 56.2$\pm$1.4 78.8$\pm$0.6 58.6$\pm$0.8 50.9
-CIN 57.8$\pm$1.1 60.1$\pm$0.7 58.6$\pm$0.2 59.8$\pm$0.2 78.9$\pm$0.1 59.9$\pm$0.4 58.8$\pm$0.3 57.4$\pm$0.5 78.8$\pm$0.6 57.7$\pm$1.2 78.8$\pm$0.7 60.1$\pm$1.1 63.9
-SpikeGCN 56.1$\pm$0.7 59.7$\pm$1.0 58.8$\pm$0.6 57.8$\pm$0.2 77.1$\pm$1.3 53.2$\pm$1.6 41.4$\pm$1.9 56.1$\pm$1.5 70.1$\pm$0.9 59.9$\pm$1.5 76.8$\pm$1.8 58.5$\pm$1.4 60.4
-DRSGNN 60.2$\pm$0.9 59.9$\pm$0.8 57.3$\pm$1.2 59.0$\pm$1.0 74.2$\pm$1.9 54.6$\pm$1.7 58.5$\pm$1.5 58.9$\pm$1.8 77.7$\pm$2.3 56.9$\pm$2.0 78.9$\pm$2.4 58.8$\pm$1.6 62.9
-CDAN 60.9$\pm$0.7 59.8$\pm$0.5 61.1$\pm$1.3 61.0$\pm$0.2 80.5$\pm$1.2 59.8$\pm$0.3 64.0$\pm$0.4 61.4$\pm$0.1 81.8$\pm$0.1 58.0$\pm$1.2 81.8$\pm$0.3 63.8$\pm$0.7 66.1
-ToAlign 56.2$\pm$0.2 59.8$\pm$0.2 41.4$\pm$0.1 59.8$\pm$0.2 21.1$\pm$0.3 59.8$\pm$0.7 41.4$\pm$1.1 56.2$\pm$1.2 21.1$\pm$0.4 56.2$\pm$0.6 21.1$\pm$1.3 41.4$\pm$0.5 44.6
-MetaAlign 57.3$\pm$2.4 59.1$\pm$1.1 60.9$\pm$1.5 60.2$\pm$0.4 80.3$\pm$2.1 60.4$\pm$0.6 64.0$\pm$1.1 64.9$\pm$0.6 81.4$\pm$1.2 58.5$\pm$2.3 80.8$\pm$0.5 63.4$\pm$1.8 65.9
-DEAL 65.3$\pm$0.6 64.0$\pm$0.2 61.3$\pm$0.6 61.0$\pm$0.9 78.3$\pm$2.1 55.5$\pm$1.8 64.9$\pm$1.2 64.8$\pm$1.1 80.1$\pm$1.3 60.1$\pm$2.1 81.8$\pm$0.4 65.7$\pm$0.7 66.9
-CoCo 63.5$\pm$2.4 61.5$\pm$1.0 64.4$\pm$1.0 61.2$\pm$0.7 81.7$\pm$0.4 55.0$\pm$1.6 64.5$\pm$0.6 64.6$\pm$1.1 80.4$\pm$1.5 60.6$\pm$1.5 81.5$\pm$0.6 62.2$\pm$1.7 66.8
-SGDA 55.7$\pm$0.5 55.4$\pm$0.9 54.8$\pm$0.3 55.3$\pm$0.7 56.1$\pm$0.5 55.4$\pm$0.8 53.2$\pm$1.1 55.1$\pm$0.6 58.4$\pm$0.4 55.3$\pm$0.5 57.7$\pm$1.0 54.9$\pm$0.6 55.7
-DGDA OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
-A2GNN 56.0$\pm$0.3 56.3$\pm$0.6 55.6$\pm$0.4 57.3$\pm$0.7 58.6$\pm$0.6 55.9$\pm$0.9 55.5$\pm$0.5 55.3$\pm$0.2 61.2$\pm$1.3 56.6$\pm$0.9 65.5$\pm$0.8 56.0$\pm$1.0 57.5
-PA-BOTH 62.2$\pm$0.5 60.7$\pm$0.7 61.5$\pm$0.6 61.2$\pm$1.0 61.9$\pm$1.3 61.1$\pm$0.8 62.3$\pm$0.4 61.7$\pm$0.8 62.0$\pm$0.9 61.1$\pm$1.2 61.2$\pm$1.5 60.9$\pm$0.6 61.5
-DeSGDA 63.5$\pm$1.1 64.1$\pm$0.9 62.6$\pm$1.3 63.0$\pm$0.8 81.1$\pm$1.2 62.4$\pm$1.5 65.5$\pm$0.6 65.4$\pm$1.7 81.9$\pm$1.0 60.8$\pm$1.4 82.0$\pm$2.1 65.9$\pm$1.8 68.2
-[…]
-[…]
-WL subtree 74.4 72.9 64.9 68.9 49.1 59.8 70.0 70.5 76.9 60.7 82.6 70.5 68.5
-GCN 63.1$\pm$1.0 68.1$\pm$0.3 48.8$\pm$0.4 62.6$\pm$0.3 29.1$\pm$2.1 38.8$\pm$0.3 54.3$\pm$0.1 61.8$\pm$0.5 30.4$\pm$0.2 43.6$\pm$0.3 67.8$\pm$0.1 57.9$\pm$1.3 52.2
-GIN 68.1$\pm$1.6 74.2$\pm$0.6 59.6$\pm$2.3 65.2$\pm$1.4 40.3$\pm$2.7 54.6$\pm$1.8 61.3$\pm$1.1 63.1$\pm$3.2 71.6$\pm$3.0 60.0$\pm$1.4 79.7$\pm$1.3 69.2$\pm$0.7 63.9
-GMT 56.5$\pm$0.3 60.7$\pm$0.4 57.9$\pm$0.2 40.2$\pm$1.2 80.6$\pm$0.4 39.3$\pm$0.6 57.9$\pm$1.1 45.0$\pm$2.1 80.6$\pm$0.5 43.5$\pm$1.1 80.6$\pm$1.4 57.9$\pm$2.2 58.4
-CIN 64.1$\pm$3.0 61.3$\pm$0.5 63.5$\pm$2.3 63.6$\pm$1.5 78.2$\pm$0.5 63.9$\pm$2.7 60.6$\pm$1.5 57.0$\pm$0.4 73.7$\pm$3.2 61.4$\pm$1.0 79.1$\pm$2.1 61.1$\pm$1.9 65.6
-SpikeGCN 56.4$\pm$1.2 60.7$\pm$0.9 59.5$\pm$1.3 57.7$\pm$1.6 59.0$\pm$1.1 60.1$\pm$1.8 54.2$\pm$1.0 59.9$\pm$2.1 50.2$\pm$2.6 55.1$\pm$1.7 80.1$\pm$2.5 57.9$\pm$2.2 59.3
-DRSGNN 56.7$\pm$0.7 61.0$\pm$1.1 57.2$\pm$1.0 57.7$\pm$1.6 52.1$\pm$1.4 55.2$\pm$1.5 59.4$\pm$1.8 56.3$\pm$1.7 75.9$\pm$2.3 60.7$\pm$1.9 80.6$\pm$0.8 58.0$\pm$1.3 60.9
-CDAN 62.8$\pm$0.3 68.2$\pm$0.6 63.6$\pm$0.6 66.9$\pm$1.7 81.2$\pm$0.5 65.0$\pm$2.1 65.8$\pm$0.2 64.7$\pm$1.2 80.7$\pm$0.1 62.5$\pm$2.3 82.4$\pm$0.4 66.0$\pm$0.5 69.1
-ToAlign 43.5$\pm$0.4 39.3$\pm$0.7 57.9$\pm$1.0 39.3$\pm$1.4 80.6$\pm$1.1 39.3$\pm$0.7 57.9$\pm$0.3 43.5$\pm$2.1 80.6$\pm$1.8 43.5$\pm$0.4 80.6$\pm$0.9 57.9$\pm$1.0 55.3
-MetaAlign 63.1$\pm$2.5 68.8$\pm$2.6 63.3$\pm$0.6 65.2$\pm$2.2 81.9$\pm$0.1 64.5$\pm$1.4 65.0$\pm$0.6 68.3$\pm$0.6 81.0$\pm$0.3 65.2$\pm$0.2 82.5$\pm$0.4 68.3$\pm$0.6 69.7
-DEAL 64.6$\pm$0.5 65.5$\pm$0.8 64.2$\pm$1.0 63.1$\pm$2.1 82.7$\pm$0.8 62.8$\pm$0.7 70.2$\pm$0.4 67.3$\pm$0.4 79.6$\pm$0.1 63.9$\pm$1.4 75.7$\pm$0.3 67.0$\pm$0.2 68.9
-CoCo 65.7$\pm$1.8 74.1$\pm$0.7 65.1$\pm$0.2 67.6$\pm$0.9 80.5$\pm$1.3 56.5$\pm$1.7 68.4$\pm$1.3 70.7$\pm$0.4 78.9$\pm$1.2 67.3$\pm$0.3 83.7$\pm$0.1 71.5$\pm$0.9 70.8
-SGDA OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
-DGDA OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
-A2GNN 55.4$\pm$0.3 55.7$\pm$0.7 55.6$\pm$0.5 54.7$\pm$0.8 63.3$\pm$1.0 56.6$\pm$0.9 55.3$\pm$0.6 55.7$\pm$0.5 65.5$\pm$0.8 56.6$\pm$1.2 69.9$\pm$1.4 55.0$\pm$0.5 58.3
-PA-BOTH 61.2$\pm$0.9 62.0$\pm$0.4 60.7$\pm$0.8 61.7$\pm$0.5 60.9$\pm$1.2 61.1$\pm$0.7 61.5$\pm$0.9 60.2$\pm$1.2 61.3$\pm$1.5 61.8$\pm$0.8 62.2$\pm$0.9 62.0$\pm$1.0 61.4
-DeSGDA 65.4$\pm$1.3 65.9$\pm$0.9 65.5$\pm$1.4 65.6$\pm$1.1 82.8$\pm$1.6 63.6$\pm$1.5 70.7$\pm$1.8 68.2$\pm$0.7 82.9$\pm$1.2 67.6$\pm$0.8 83.9$\pm$1.0 66.5$\pm$1.2 70.8
+DeSGraDA w CDAN 54.1 55.2 53.5 56.8 53.9 55.7 52.0 52.3 51.5 53.6 52.7 52.4
+DeSGraDA w/o PL 55.6 55.7 54.3 56.1 53.7 54.5 53.3 52.9 52.6 52.8 52.1 52.2
+DeSGraDA w/o CF 29.7 47.2 45.8 47.5 45.6 32.8 43.0 44.3 44.2 44.7 33.1 34.2
+DeSGraDA w/o TL 53.8 54.6 53.2 55.4 52.0 53.4 51.1 50.9 52.2 53.9 50.5 51.3
+DeSGraDA 58.0 58.2 57.0 58.3 55.9 58.1 54.1 53.6 54.9 56.2 55.0 54.6
 […]
@@
 SGDA 64.2$\pm$0.5 61.0$\pm$0.7 66.9$\pm$1.2 61.9$\pm$0.9 65.4$\pm$1.6 66.5$\pm$1.0 64.6$\pm$1.1 60.1$\pm$0.5 66.3$\pm$1.3 59.3$\pm$0.8 66.0$\pm$1.6 66.2$\pm$1.3 64.1
-DGDA 58.1$\pm$0.4 58.6$\pm$0.6 58.9$\pm$1.0 61.0$\pm$0.9 59.6$\pm$0.7 60.2$\pm$1.5 56.7$\pm$0.6 56.8$\pm$0.8 58.1$\pm$0.4 58.8$\pm$1.1 57.0$\pm$1.2 62.2$\pm$1.6 58.9
+StruRW 71.9$\pm$2.3 82.6$\pm$1.9 66.7$\pm$1.8 74.5$\pm$2.8 52.8$\pm$1.9 57.3$\pm$2.0 62.2$\pm$2.4 63.3$\pm$2.1 59.5$\pm$1.6 56.3$\pm$2.0 66.6$\pm$2.3 52.4$\pm$2.0 63.8
 A2GNN 65.7$\pm$0.6 65.9$\pm$0.8 66.3$\pm$0.9 65.6$\pm$1.1 65.2$\pm$1.4 65.6$\pm$1.3 65.9$\pm$1.7 65.8$\pm$1.6 65.0$\pm$1.5 66.1$\pm$1.2 65.2$\pm$1.9 65.9$\pm$1.8 65.7
 PA-BOTH 61.0$\pm$0.8 61.2$\pm$1.3 60.3$\pm$0.6 66.7$\pm$2.1 63.7$\pm$1.5 61.9$\pm$2.0 66.2$\pm$1.4 69.9$\pm$2.3 68.0$\pm$0.7 69.4$\pm$1.8 61.5$\pm$0.4 67.6$\pm$1.0 64.9
-DeSGDA 77.6$\pm$0.9 84.3$\pm$1.1 70.5$\pm$0.6 84.8$\pm$1.4 76.6$\pm$0.7 83.9$\pm$0.9 71.9$\pm$0.6 76.9$\pm$1.1 76.1$\pm$0.8 73.7$\pm$1.0 76.0$\pm$1.2 70.4$\pm$0.7 76.9
-[…]
-[…]
-WL subtree 73.5 79.5 64.8 75.9 58.9 68.4 72.5 72.0 69.7 63.6 76.1 74.0 70.7
-GCN 51.2$\pm$0.1 71.1$\pm$0.4 42.7$\pm$0.4 27.8$\pm$0.3 32.1$\pm$1.1 27.0$\pm$0.2 55.2$\pm$0.6 50.5$\pm$0.7 50.9$\pm$1.1 49.1$\pm$0.3 67.1$\pm$0.6 57.3$\pm$0.6 48.5
-GIN 66.9$\pm$2.2 78.9$\pm$2.3 60.3$\pm$3.1 72.8$\pm$0.3 51.1$\pm$0.6 68.6$\pm$1.8 63.5$\pm$2.1 67.8$\pm$3.7 65.9$\pm$1.7 60.3$\pm$1.8 71.1$\pm$1.1 67.2$\pm$1.3 66.2
-GMT 50.9$\pm$0.5 73.0$\pm$0.1 57.3$\pm$0.3 73.0$\pm$0.4 66.5$\pm$0.2 73.0$\pm$0.3 72.4$\pm$0.6 50.9$\pm$0.1 66.5$\pm$0.4 58.3$\pm$0.2 66.5$\pm$0.5 72.8$\pm$0.3 65.1
-CIN 60.1$\pm$0.7 73.1$\pm$1.1 57.5$\pm$0.2 73.0$\pm$0.4 66.5$\pm$1.1 73.1$\pm$0.7 58.5$\pm$2.1 52.9$\pm$1.4 66.5$\pm$1.3 56.1$\pm$0.1 66.5$\pm$0.4 57.4$\pm$0.7 63.4
-SpikeGCN 63.3$\pm$0.4 72.6$\pm$0.7 60.6$\pm$0.2 73.1$\pm$0.8 65.4$\pm$0.6 66.6$\pm$1.2 64.3$\pm$1.4 64.7$\pm$1.0 65.6$\pm$1.2 59.8$\pm$0.8 70.1$\pm$1.7 60.9$\pm$1.4 65.6
-DRSGNN 63.9$\pm$0.8 73.3$\pm$1.2 56.9$\pm$1.5 72.5$\pm$1.3 66.7$\pm$2.1 64.4$\pm$1.8 65.6$\pm$1.3 64.8$\pm$1.9 66.7$\pm$1.4 58.4$\pm$1.6 68.6$\pm$2.0 61.4$\pm$1.3 65.3
-CDAN 57.1$\pm$0.4 75.0$\pm$0.7 61.2$\pm$0.4 73.7$\pm$0.1 68.2$\pm$0.4 73.3$\pm$0.3 60.2$\pm$0.1 56.5$\pm$1.4 68.2$\pm$0.2 53.9$\pm$1.4 68.4$\pm$0.2 59.6$\pm$0.5 64.6
-ToAlign 49.1$\pm$0.3 27.0$\pm$0.6 57.3$\pm$0.5 27.0$\pm$0.4 66.5$\pm$0.5 27.0$\pm$0.2 57.3$\pm$0.3 49.1$\pm$0.4 66.5$\pm$0.2 49.1$\pm$0.3 66.5$\pm$0.1 57.3$\pm$0.4 50.0
-MetaAlign 65.6$\pm$1.8 77.7$\pm$0.2 63.5$\pm$1.4 75.7$\pm$0.7 66.4$\pm$0.3 74.0$\pm$0.3 66.3$\pm$1.1 64.6$\pm$1.2 66.7$\pm$0.2 59.5$\pm$2.6 66.7$\pm$0.3 66.7$\pm$2.7 67.8
-DEAL 64.0$\pm$0.9 71.9$\pm$1.2 61.4$\pm$0.3 73.3$\pm$0.3 64.9$\pm$1.4 71.9$\pm$1.9 62.5$\pm$2.1 66.2$\pm$0.5 54.2$\pm$1.4 55.6$\pm$0.8 64.6$\pm$0.4 58.8$\pm$0.4 64.1
-CoCo 69.7$\pm$0.1 80.4$\pm$0.4 64.7$\pm$1.2 76.5$\pm$0.4 65.0$\pm$1.7 73.9$\pm$0.3 68.9$\pm$1.3 70.7$\pm$0.9 68.2$\pm$1.2 61.4$\pm$1.7 73.0$\pm$0.1 65.2$\pm$0.9 69.8
-SGDA OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
-DGDA OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
-A2GNN 59.0$\pm$0.6 58.3$\pm$1.1 58.5$\pm$0.8 58.6$\pm$1.3 58.7$\pm$1.0 59.0$\pm$0.7 58.5$\pm$1.1 58.7$\pm$1.5 59.1$\pm$0.6 58.3$\pm$1.2 58.6$\pm$0.7 59.0$\pm$0.5 58.7
-PA-BOTH 57.7$\pm$0.4 58.0$\pm$0.6 57.9$\pm$0.5 56.9$\pm$0.8 57.4$\pm$0.6 58.3$\pm$0.5 57.1$\pm$1.2 58.8$\pm$0.9 58.1$\pm$0.7 58.0$\pm$0.9 57.9$\pm$0.5 58.3$\pm$0.8 57.9
-DeSGDA 64.4$\pm$1.2 76.9$\pm$1.5 64.9$\pm$0.9 76.6$\pm$1.2 68.6$\pm$1.8 74.1$\pm$1.3 66.9$\pm$0.8 65.1$\pm$1.2 69.9$\pm$1.5 63.9$\pm$2.0 70.9$\pm$1.6 64.2$\pm$1.4 68.9
-[…]
-[…]
-WL subtree 65.7 71.8 57.9 71.1 47.4 43.4 65.5 75.1 45.3 34.9 52.7 49.8 56.7
-GCN 70.6$\pm$2.1 60.3$\pm$1.5 60.5$\pm$3.4 62.3$\pm$1.1 58.4$\pm$0.5 43.2$\pm$0.2 63.8$\pm$1.0 70.3$\pm$0.3 50.6$\pm$1.0 32.8$\pm$0.3 50.1$\pm$0.4 42.2$\pm$0.2 55.4
-GIN 66.7$\pm$2.1 73.7$\pm$2.4 57.3$\pm$3.1 69.4$\pm$2.3 58.6$\pm$0.4 43.1$\pm$0.3 66.4$\pm$2.7 74.8$\pm$1.8 42.2$\pm$1.6 33.5$\pm$1.0 57.4$\pm$0.8 43.9$\pm$2.3 57.2
-GMT 67.3$\pm$0.3 56.8$\pm$0.4 58.0$\pm$0.2 56.8$\pm$0.2 60.6$\pm$0.3 56.8$\pm$0.5 57.8$\pm$0.1 67.3$\pm$0.1 39.5$\pm$0.3 67.3$\pm$0.2 39.5$\pm$0.5 57.8$\pm$0.4 57.1
-CIN 67.6$\pm$0.4 63.7$\pm$2.1 58.9$\pm$1.0 56.8$\pm$0.4 63.6$\pm$0.4 59.5$\pm$2.7 58.7$\pm$1.2 67.0$\pm$0.5 61.7$\pm$1.6 67.8$\pm$0.7 62.2$\pm$2.1 56.0$\pm$1.3 61.9
-SpikeGCN 67.2$\pm$0.5 57.2$\pm$1.2 57.9$\pm$0.8 57.1$\pm$0.6 61.1$\pm$1.3 58.9$\pm$1.6 60.0$\pm$1.2 67.2$\pm$0.9 53.9$\pm$2.1 64.4$\pm$0.8 57.8$\pm$1.0 59.9$\pm$1.2 60.2
-DRSGNN 67.4$\pm$0.4 58.4$\pm$1.0 59.0$\pm$1.2 57.4$\pm$0.5 62.3$\pm$1.1 60.4$\pm$1.3 61.1$\pm$1.6 67.9$\pm$1.5 56.2$\pm$1.8 66.2$\pm$2.1 60.9$\pm$1.4 58.6$\pm$2.5 61.3
-CDAN 72.9$\pm$0.4 72.7$\pm$0.4 65.4$\pm$0.3 72.9$\pm$0.1 61.2$\pm$0.3 70.3$\pm$0.2 65.7$\pm$0.4 72.7$\pm$0.1 61.0$\pm$0.1 72.1$\pm$1.2 60.7$\pm$0.2 65.3$\pm$0.6 67.7
-ToAlign 32.7$\pm$2.0 43.2$\pm$0.1 42.2$\pm$1.3 43.2$\pm$0.9 60.5$\pm$0.7 43.2$\pm$1.2 42.2$\pm$0.4 32.7$\pm$1.2 60.5$\pm$0.9 32.7$\pm$0.3 60.5$\pm$0.7 42.2$\pm$0.4 44.7
-MetaAlign 67.3$\pm$0.7 56.8$\pm$0.2 57.8$\pm$0.6 56.8$\pm$0.4 60.5$\pm$1.3 56.8$\pm$0.8 57.8$\pm$1.1 67.3$\pm$1.2 60.5$\pm$0.4 67.3$\pm$0.6 60.5$\pm$0.7 57.8$\pm$0.6 60.6
-DEAL 75.0$\pm$0.9 76.3$\pm$2.4 65.9$\pm$1.8 77.5$\pm$2.7 60.3$\pm$4.5 69.7$\pm$3.2 67.2$\pm$1.5 75.3$\pm$1.7 57.4$\pm$4.1 71.1$\pm$2.2 65.7$\pm$2.7 66.4$\pm$1.6 69.0
-CoCo 74.2$\pm$1.7 74.3$\pm$0.6 65.9$\pm$1.2 72.7$\pm$2.1 61.1$\pm$0.2 71.0$\pm$1.7 68.6$\pm$0.3 75.9$\pm$0.2 60.7$\pm$0.2 73.9$\pm$0.4 59.7$\pm$1.1 67.3$\pm$0.8 68.8
-SGDA 55.9$\pm$0.6 57.1$\pm$0.5 56.1$\pm$0.4 54.6$\pm$0.8 55.8$\pm$1.1 57.7$\pm$0.6 54.3$\pm$0.7 53.6$\pm$1.3 59.1$\pm$0.8 56.7$\pm$0.6 55.4$\pm$1.2 53.8$\pm$0.5 55.9
-DGDA OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
-A2GNN 55.9$\pm$0.7 55.7$\pm$0.4 56.6$\pm$0.6 57.1$\pm$1.0 56.1$\pm$1.2 55.8$\pm$0.5 56.5$\pm$0.7 55.5$\pm$0.4 55.9$\pm$0.8 56.2$\pm$0.6 56.5$\pm$1.5 56.0$\pm$0.5 56.2
-PA-BOTH 56.4$\pm$0.5 55.9$\pm$0.6 56.0$\pm$0.5 56.4$\pm$0.4 56.3$\pm$0.6 57.7$\pm$0.7 56.6$\pm$0.2 58.8$\pm$0.9 56.9$\pm$0.7 57.2$\pm$0.3 56.5$\pm$0.5 58.3$\pm$0.8 56.9
-DeSGDA 75.2$\pm$0.8 74.3$\pm$1.2 68.0$\pm$1.5 73.6$\pm$2.0 62.3$\pm$1.7 71.2$\pm$2.5 68.8$\pm$1.3 76.0$\pm$1.2 61.9$\pm$2.2 71.4$\pm$1.6 62.8$\pm$1.2 65.7$\pm$0.9 69.3
-[…]
-[…]
-WL subtree 78.0 68.7 70.1 70.5 59.0 61.2 71.7 78.0 49.9 56.3 69.4 71.9 67.1
-GCN 74.5$\pm$0.2 60.8$\pm$2.1 69.7$\pm$0.4 68.5$\pm$1.7 54.1$\pm$0.9 55.2$\pm$0.9 68.6$\pm$1.6 75.5$\pm$0.5 51.5$\pm$1.3 46.4$\pm$1.7 58.6$\pm$0.4 60.2$\pm$0.2 61.9
-GIN 77.9$\pm$3.1 70.7$\pm$2.4 70.9$\pm$0.8 69.2$\pm$1.2 64.1$\pm$1.0 61.9$\pm$2.4 78.5$\pm$0.2 79.8$\pm$3.3 65.5$\pm$2.7 71.5$\pm$0.9 69.5$\pm$1.8 73.5$\pm$2.6 71.1
-GMT 67.3$\pm$0.2 52.5$\pm$0.1 59.9$\pm$0.3 47.5$\pm$0.2 53.5$\pm$0.2 52.5$\pm$0.4 59.9$\pm$0.1 67.3$\pm$0.2 46.7$\pm$0.5 67.3$\pm$0.3 53.3$\pm$0.1 59.9$\pm$0.4 57.1
-CIN 70.8$\pm$1.1 66.9$\pm$3.4 61.7$\pm$0.6 62.6$\pm$2.4 56.3$\pm$3.1 62.9$\pm$1.3 65.1$\pm$1.0 68.8$\pm$1.7 56.6$\pm$1.4 66.9$\pm$1.0 58.1$\pm$1.3 62.5$\pm$0.9 63.3
-SpikeGCN 63.2$\pm$0.6 55.9$\pm$1.7 59.8$\pm$1.2 56.5$\pm$1.6 54.1$\pm$2.2 60.3$\pm$1.2 60.1$\pm$0.8 68.8$\pm$1.4 54.4$\pm$2.0 64.2$\pm$1.6 56.6$\pm$0.8 58.5$\pm$1.5 59.4
-DRSGNN 56.9$\pm$0.9 53.4$\pm$1.2 58.9$\pm$1.0 57.7$\pm$1.4 53.2$\pm$1.6 61.2$\pm$1.5 60.6$\pm$2.1 67.9$\pm$1.2 57.1$\pm$1.4 67.3$\pm$2.0 59.9$\pm$1.9 57.1$\pm$0.8 59.3
-CDAN 75.5$\pm$0.1 71.3$\pm$0.4 70.7$\pm$0.3 70.3$\pm$0.1 58.7$\pm$0.6 58.4$\pm$0.6 70.2$\pm$0.5 76.1$\pm$0.5 58.5$\pm$0.6 69.4$\pm$1.5 59.0$\pm$0.1 63.7$\pm$1.4 66.8
-ToAlign 67.3$\pm$0.2 47.5$\pm$0.4 59.9$\pm$0.6 47.5$\pm$0.5 46.7$\pm$0.4 47.5$\pm$0.2 59.9$\pm$0.7 67.3$\pm$0.3 46.7$\pm$0.1 67.3$\pm$0.4 46.7$\pm$0.5 59.9$\pm$0.3 55.4
-MetaAlign 76.5$\pm$0.4 71.8$\pm$1.1 71.8$\pm$0.8 71.4$\pm$0.9 59.3$\pm$0.8 63.0$\pm$1.0 74.2$\pm$1.6 78.0$\pm$0.2 61.7$\pm$1.2 69.9$\pm$1.6 62.2$\pm$0.4 68.3$\pm$1.5 69.0
-DEAL 76.6$\pm$0.8 68.8$\pm$1.0 69.9$\pm$0.4 66.4$\pm$0.8 59.3$\pm$2.1 64.2$\pm$2.2 79.1$\pm$0.1 81.9$\pm$0.6 64.5$\pm$1.1 75.3$\pm$0.6 69.8$\pm$1.6 76.5$\pm$0.2 71.0
-CoCo 75.5$\pm$0.4 71.7$\pm$0.7 68.7$\pm$1.1 69.2$\pm$2.0 60.8$\pm$1.1 65.7$\pm$0.3 79.2$\pm$1.2 76.8$\pm$0.6 63.8$\pm$0.5 73.8$\pm$0.4 64.6$\pm$0.8 70.1$\pm$1.1 70.0
-SGDA OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
-DGDA OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
-A2GNN 55.4$\pm$0.6 56.3$\pm$0.2 55.6$\pm$0.8 55.1$\pm$0.5 55.3$\pm$1.1 55.9$\pm$0.4 56.1$\pm$0.7 55.7$\pm$0.6 57.1$\pm$0.3 56.6$\pm$1.2 55.2$\pm$0.7 56.8$\pm$1.0 55.9
-PA-BOTH 55.9$\pm$1.0 56.0$\pm$0.5 56.1$\pm$0.7 56.6$\pm$1.2 55.9$\pm$0.6 56.0$\pm$0.7 57.3$\pm$0.8 56.8$\pm$1.3 55.9$\pm$1.2 56.3$\pm$1.0 56.4$\pm$0.9 57.1$\pm$1.3 56.4
-DeSGDA 76.7$\pm$1.0 72.5$\pm$0.6 72.0$\pm$2.3 71.5$\pm$1.5 61.0$\pm$1.2 67.6$\pm$0.8 75.5$\pm$1.0 79.8$\pm$1.2 62.8$\pm$1.6 75.9$\pm$1.3 65.5$\pm$2.1 72.3$\pm$1.7 71.1
+DeSGraDA 76.3$\pm$1.9 84.6$\pm$2.5 69.2$\pm$2.3 83.6$\pm$2.6 77.5$\pm$2.2 83.7$\pm$1.9 69.8$\pm$2.4 74.0$\pm$1.6 76.2$\pm$2.0 73.0$\pm$2.1 77.8$\pm$2.3 70.5$\pm$1.7 76.4
 […]
@@
 SGDA 63.8$\pm$0.6 65.2$\pm$1.3 66.7$\pm$1.0 59.1$\pm$1.5 60.1$\pm$0.8 64.4$\pm$1.2 65.2$\pm$0.7 63.9$\pm$0.9 64.5$\pm$0.6 61.1$\pm$1.3 58.9$\pm$1.4 64.9$\pm$1.2 63.2
-DGDA 58.7$\pm$0.8 59.9$\pm$1.2 57.1$\pm$0.6 57.9$\pm$0.8 59.2$\pm$1.3 58.9$\pm$0.4 61.1$\pm$1.2 60.3$\pm$1.6 58.6$\pm$0.9 57.5$\pm$1.2 58.4$\pm$0.5 62.3$\pm$1.5 59.2
+StruRW 72.6$\pm$2.2 84.5$\pm$1.7 66.2$\pm$2.2 72.5$\pm$2.4 48.9$\pm$2.0 56.5$\pm$2.3 63.1$\pm$1.8 64.4$\pm$2.4 55.8$\pm$2.0 56.6$\pm$2.4 67.0$\pm$2.6 42.4$\pm$2.0 62.5
 A2GNN 65.4$\pm$1.3 66.3$\pm$1.1 68.2$\pm$1.4 66.3$\pm$1.2 65.4$\pm$0.7 65.9$\pm$0.9 66.9$\pm$1.3 65.4$\pm$1.2 65.6$\pm$0.9 65.5$\pm$1.2 66.1$\pm$2.0 66.0$\pm$1.8 66.1
 PA-BOTH 63.1$\pm$0.7 67.2$\pm$1.1 64.3$\pm$0.5 72.1$\pm$1.8 66.3$\pm$0.7 64.1$\pm$1.2 69.7$\pm$2.1 67.5$\pm$1.8 61.2$\pm$1.4 67.7$\pm$2.3 61.2$\pm$1.6 65.5$\pm$0.6 65.9
-DeSGDA 76.7$\pm$0.8 84.6$\pm$0.9 69.4$\pm$0.6 85.2$\pm$1.5 76.2$\pm$1.1 83.9$\pm$1.2 69.9$\pm$0.6 76.3$\pm$1.4 75.9$\pm$1.0 73.5$\pm$1.3 76.3$\pm$1.6 68.3$\pm$0.7 76.4
+DeSGraDA 76.8$\pm$1.9 87.0$\pm$2.1 68.6$\pm$1.8 83.7$\pm$2.5 76.5$\pm$2.8 83.9$\pm$2.3 70.3$\pm$1.8 75.4$\pm$2.2 76.7$\pm$1.9 73.7$\pm$2.7 79.9$\pm$3.2 67.9$\pm$1.3 76.7
 […]
 […]
-WL subtree 72.6 80.3 62.7 75.5 52.0 63.6 69.1 69.8 70.7 59.4 80.0 70.6 68.9
-GCN 49.5$\pm$0.4 71.1$\pm$0.4 46.8$\pm$0.5 33.7$\pm$2.8 32.7$\pm$0.4 27.4$\pm$0.1 56.2$\pm$1.5 55.3$\pm$0.4 58.2$\pm$1.7 51.0$\pm$0.2 60.7$\pm$3.7 53.2$\pm$0.2 49.6
-GIN 67.3$\pm$2.7 67.9$\pm$4.8 61.5$\pm$4.2 65.4$\pm$3.7 58.9$\pm$4.1 61.0$\pm$3.4 62.5$\pm$3.2 66.2$\pm$2.1 69.7$\pm$0.9 56.8$\pm$0.7 72.4$\pm$2.8 64.0$\pm$1.6 64.5
-GMT 50.3$\pm$1.2 42.5$\pm$3.4 51.1$\pm$3.7 42.5$\pm$4.5 56.1$\pm$4.7 42.5$\pm$4.1 53.2$\pm$4.9 51.0$\pm$0.2 68.2$\pm$0.4 51.0$\pm$0.3 68.2$\pm$0.5 53.2$\pm$0.4 52.5
-CIN 51.1$\pm$0.2 72.6$\pm$0.1 54.0$\pm$0.9 72.6$\pm$0.2 68.2$\pm$0.3 71.5$\pm$1.3 55.0$\pm$2.1 53.5$\pm$1.8 68.2$\pm$0.3 52.0$\pm$0.3 68.3$\pm$0.1 53.6$\pm$0.6 61.7
-SpikeGCN 62.8$\pm$0.8 73.1$\pm$1.2 61.4$\pm$0.8 70.9$\pm$0.6 57.7$\pm$1.6 66.2$\pm$1.1 61.2$\pm$1.6 64.5$\pm$1.0 62.3$\pm$1.4 57.3$\pm$0.9 68.9$\pm$1.2 60.1$\pm$1.0 63.6
-DRSGNN 64.3$\pm$0.6 76.3$\pm$0.9 56.7$\pm$1.1 73.2$\pm$0.8 58.6$\pm$1.4 63.9$\pm$1.8 63.0$\pm$2.1 65.1$\pm$1.6 64.1$\pm$1.9 59.2$\pm$2.2 70.8$\pm$2.5 56.6$\pm$1.4 64.3
-CDAN 59.6$\pm$0.3 73.8$\pm$0.5 56.7$\pm$1.4 73.7$\pm$0.3 71.2$\pm$0.4 73.2$\pm$0.3 55.5$\pm$0.2 57.3$\pm$1.1 69.9$\pm$0.2 54.6$\pm$2.0 69.8$\pm$1.4 56.6$\pm$0.3 64.3
-ToAlign 51.0$\pm$0.2 27.4$\pm$0.1 53.2$\pm$0.4 27.4$\pm$0.2 68.2$\pm$0.3 27.4$\pm$0.3 53.2$\pm$0.1 51.0$\pm$0.2 68.2$\pm$0.2 51.0$\pm$0.4 68.2$\pm$0.3 53.2$\pm$0.2 50.0
-MetaAlign 65.0$\pm$0.7 77.6$\pm$1.6 62.0$\pm$0.6 77.1$\pm$0.9 68.2$\pm$0.8 74.5$\pm$2.0 64.2$\pm$0.9 65.4$\pm$0.3 68.0$\pm$0.3 56.1$\pm$2.3 68.2$\pm$0.1 66.2$\pm$1.1 67.7
-DEAL 65.6$\pm$0.6 73.0$\pm$0.9 58.0$\pm$0.3 71.6$\pm$1.6 60.1$\pm$2.8 73.1$\pm$0.5 62.8$\pm$1.0 65.0$\pm$2.4 65.8$\pm$0.8 53.9$\pm$2.6 57.6$\pm$2.8 56.7$\pm$3.1 63.6
-CoCo 70.4$\pm$0.7 80.4$\pm$0.9 62.4$\pm$0.8 75.8$\pm$1.2 65.7$\pm$2.0 73.7$\pm$0.3 67.0$\pm$0.8 70.4$\pm$0.7 69.7$\pm$0.4 62.7$\pm$0.9 74.4$\pm$0.5 63.7$\pm$0.9 69.7
-SGDA OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
-DGDA OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
-A2GNN 59.2$\pm$0.8 58.7$\pm$0.5 59.0$\pm$1.1 58.7$\pm$0.8 58.9$\pm$0.6 59.2$\pm$1.2 58.7$\pm$0.6 58.6$\pm$1.2 59.0$\pm$1.0 59.5$\pm$0.6 58.7$\pm$0.5 58.5$\pm$1.1 58.9
-PA-BOTH 57.6$\pm$0.5 58.4$\pm$0.4 58.9$\pm$0.6 57.4$\pm$0.6 57.1$\pm$1.0 58.4$\pm$0.5 58.0$\pm$1.0 58.1$\pm$0.5 58.4$\pm$0.6 57.7$\pm$1.1 57.5$\pm$0.6 58.0$\pm$0.4 58.0
-DeSGDA 66.0$\pm$0.5 76.2$\pm$0.4 62.8$\pm$0.6 77.6$\pm$0.6 68.5$\pm$1.0 74.6$\pm$0.5 65.1$\pm$1.0 65.4$\pm$0.5 70.9$\pm$0.6 64.8$\pm$1.1 72.6$\pm$0.6 66.3$\pm$0.4 69.3
+WL subtree 49.2 56.8 29.6 20.1 21.0 18.4 59.5 50.5 57.3 48.1 63.9 66.9 49.3
+GCN 48.9$\pm$2.8 59.0$\pm$1.7 20.7$\pm$2.0 27.3$\pm$2.3 15.1$\pm$1.8 26.9$\pm$2.2 61.6$\pm$1.9 53.6$\pm$1.5 68.1$\pm$1.6 52.9$\pm$2.6 64.9$\pm$2.1 69.7$\pm$2.3 47.4
+GIN 48.8$\pm$1.9 24.7$\pm$2.1 44.1$\pm$1.8 22.4$\pm$2.3 57.0$\pm$2.1 18.4$\pm$2.0 73.0$\pm$1.8 52.5$\pm$2.3 63.2$\pm$1.6 53.6$\pm$1.5 70.3$\pm$2.6 69.4$\pm$1.8 49.6
+GMT 49.1$\pm$1.9 32.9$\pm$2.2 31.8$\pm$1.8 27.3$\pm$2.3 52.5$\pm$1.5 27.6$\pm$1.8 75.4$\pm$1.9 53.2$\pm$2.1 74.1$\pm$2.6 57.9$\pm$2.4 70.9$\pm$1.8 71.1$\pm$2.7 52.0
+CIN 50.4$\pm$1.8 18.4$\pm$2.0 21.2$\pm$2.1 36.8$\pm$1.8 43.0$\pm$2.1 22.9$\pm$1.9 53.4$\pm$1.7 56.5$\pm$1.5 62.3$\pm$1.6 53.3$\pm$1.9 75.0$\pm$2.1 69.3$\pm$2.0 51.4
+SpikeGCN 53.4$\pm$1.3 61.5$\pm$1.1 28.4$\pm$2.1 48.0$\pm$1.4 20.8$\pm$1.4 51.2$\pm$1.4 69.3$\pm$2.0 62.4$\pm$1.6 77.9$\pm$1.1 64.1$\pm$1.5 76.8$\pm$2.1 71.9$\pm$1.8 57.1
+DRSGNN 52.1$\pm$1.5 69.7$\pm$1.8 28.7$\pm$1.7 42.4$\pm$2.1 18.6$\pm$2.0 48.3$\pm$2.8 79.5$\pm$2.0 59.5$\pm$1.5 77.7$\pm$1.9 63.1$\pm$1.5 76.4$\pm$2.1 72.1$\pm$2.3 57.4
+CDAN 49.6$\pm$2.2 69.8$\pm$1.6 44.1$\pm$1.8 33.9$\pm$1.9 43.1$\pm$2.4 42.3$\pm$2.0 70.5$\pm$1.8 60.3$\pm$2.2 76.6$\pm$2.1 60.1$\pm$1.4 75.8$\pm$2.5 70.5$\pm$2.3 58.1
+ToAlign 54.0$\pm$2.4 71.0$\pm$2.1 34.8$\pm$1.8 46.9$\pm$1.9 29.6$\pm$2.3 45.7$\pm$1.8 71.9$\pm$2.2 61.6$\pm$2.1 76.7$\pm$1.9 62.8$\pm$2.3 76.4$\pm$2.0 71.0$\pm$1.3 58.5
+MetaAlign 48.1$\pm$2.0 70.0$\pm$1.9 30.7$\pm$1.4 18.4$\pm$1.8 24.9$\pm$2.3 18.4$\pm$1.9 70.1$\pm$2.3 51.9$\pm$1.5 74.6$\pm$2.4 51.9$\pm$2.2 75.1$\pm$1.8 69.3$\pm$1.7 52.7
+DEAL 57.9$\pm$2.3 71.6$\pm$2.0 57.2$\pm$1.8 59.3$\pm$1.5 62.2$\pm$2.2 59.4$\pm$1.9 72.1$\pm$2.3 63.9$\pm$1.7 78.2$\pm$2.2 62.6$\pm$1.8 78.3$\pm$2.1 77.3$\pm$1.9 66.5
+CoCo 59.5$\pm$2.1 70.4$\pm$1.7 56.6$\pm$2.6 58.3$\pm$2.2 59.4$\pm$1.9 53.9$\pm$2.5 74.7$\pm$1.4 62.7$\pm$2.0 70.6$\pm$1.6 63.1$\pm$2.0 77.2$\pm$1.7 76.4$\pm$2.4 65.2
+SGDA 57.7$\pm$1.5 63.8$\pm$2.1 49.8$\pm$2.3 54.1$\pm$1.7 42.6$\pm$2.3 54.9$\pm$1.6 74.1$\pm$2.8 63.0$\pm$2.3 78.7$\pm$2.7 64.5$\pm$2.1 76.8$\pm$2.2 74.2$\pm$1.6 62.9
+StruRW 50.0$\pm$2.3 53.1$\pm$1.9 32.4$\pm$2.4 40.6$\pm$2.1 26.0$\pm$2.4 38.4$\pm$2.0 73.3$\pm$1.6 61.7$\pm$1.8 71.2$\pm$1.9 53.6$\pm$2.3 75.2$\pm$2.1 71.0$\pm$2.2 53.9
+A2GNN 56.1$\pm$2.0 68.5$\pm$1.6 48.7$\pm$2.1 52.5$\pm$1.8 42.9$\pm$1.4 48.4$\pm$1.8 70.8$\pm$1.7 51.9$\pm$2.0 76.3$\pm$2.2 51.9$\pm$1.8 75.1$\pm$1.6 69.3$\pm$2.4 59.4
+PA-BOTH 51.4$\pm$1.8 62.7$\pm$2.1 31.8$\pm$2.0 40.5$\pm$2.3 28.5$\pm$1.9 45.0$\pm$2.2 69.5$\pm$1.6 61.0$\pm$1.5 68.4$\pm$1.7 57.8$\pm$2.2 73.0$\pm$2.4 73.3$\pm$2.5 55.3
+DeSGraDA 60.1$\pm$2.2 76.1$\pm$1.8 63.9$\pm$1.9 71.7$\pm$2.0 61.1$\pm$1.9 62.3$\pm$1.6 73.6$\pm$2.3 64.2$\pm$2.0 80.9$\pm$2.2 64.7$\pm$1.8 82.0$\pm$2.5 80.6$\pm$2.1 70.1
 […]
 […]
-WL subtree 71.6 72.1 62.1 71.2 57.8 67.7 64.0 75.3 41.1 59.2 55.9 55.4 62.8
-GCN 66.5$\pm$0.4 60.0$\pm$0.8 55.4$\pm$0.3 60.0$\pm$0.1 39.6$\pm$0.3 40.0$\pm$0.4 55.4$\pm$0.2 66.5$\pm$0.1 39.6$\pm$0.6 33.5$\pm$0.3 39.6$\pm$0.1 44.7$\pm$0.2 50.1
-GIN 71.4$\pm$4.7 73.4$\pm$3.4 60.8$\pm$2.7 66.0$\pm$3.4 50.5$\pm$3.7 51.6$\pm$1.8 64.8$\pm$1.0 71.3$\pm$3.5 48.3$\pm$4.2 57.4$\pm$3.8 55.1$\pm$3.4 52.6$\pm$4.3 60.3
-GMT 67.4$\pm$1.0 61.7$\pm$2.1 55.8$\pm$0.7 57.0$\pm$2.4 60.2$\pm$0.5 58.2$\pm$2.0 57.8$\pm$2.1 65.7$\pm$1.3 60.2$\pm$0.3 57.3$\pm$2.3 60.7$\pm$0.6 57.1$\pm$1.2 59.9
-CIN 70.4$\pm$2.8 66.5$\pm$4.3 58.5$\pm$2.6 64.2$\pm$2.7 60.6$\pm$3.0 64.2$\pm$3.2 58.7$\pm$2.4 69.1$\pm$2.7 57.5$\pm$3.4 67.7$\pm$2.1 59.5$\pm$2.3 56.1$\pm$1.2 62.7
-SpikeGCN 66.5$\pm$0.9 60.1$\pm$1.5 55.9$\pm$0.8 60.2$\pm$0.6 54.8$\pm$2.1 59.9$\pm$1.6 55.8$\pm$0.8 62.9$\pm$1.6 58.4$\pm$1.2 61.1$\pm$1.3 58.8$\pm$1.9 62.1$\pm$1.5 59.8
-DRSGNN 67.5$\pm$1.2 61.2$\pm$0.7 55.6$\pm$1.4 61.1$\pm$0.9 52.4$\pm$2.3 61.0$\pm$1.2 56.9$\pm$0.7 66.7$\pm$1.5 60.3$\pm$0.5 62.0$\pm$2.0 59.8$\pm$1.4 59.2$\pm$1.1 60.3
-CDAN 72.9$\pm$0.2 74.0$\pm$0.3 62.7$\pm$0.3 73.8$\pm$0.5 61.2$\pm$1.0 70.0$\pm$1.2 62.8$\pm$0.1 73.0$\pm$0.3 60.6$\pm$0.2 71.6$\pm$1.5 60.5$\pm$0.2 61.1$\pm$1.4 67.0
-ToAlign 68.0$\pm$3.8 73.4$\pm$2.7 64.5$\pm$1.1 63.7$\pm$2.4 60.6$\pm$1.2 61.9$\pm$1.3 64.8$\pm$1.3 74.0$\pm$1.3 60.0$\pm$0.6 65.7$\pm$3.1 61.0$\pm$1.4 56.2$\pm$2.3 64.5
-MetaAlign 73.6$\pm$0.2 72.7$\pm$1.9 63.9$\pm$1.0 67.9$\pm$4.3 60.4$\pm$0.7 65.4$\pm$1.8 65.2$\pm$0.8 73.2$\pm$2.3 60.0$\pm$0.6 66.7$\pm$2.4 61.2$\pm$1.1 56.8$\pm$2.1 65.6
-DEAL 75.4$\pm$0.3 74.6$\pm$1.1 66.1$\pm$0.6 74.6$\pm$0.8 53.8$\pm$1.0 69.6$\pm$1.8 66.4$\pm$0.3 73.9$\pm$0.6 61.6$\pm$1.4 69.8$\pm$0.2 60.7$\pm$1.0 58.3$\pm$0.9 67.1
-CoCo 74.6$\pm$0.9 77.2$\pm$0.6 64.1$\pm$3.4 73.8$\pm$1.1 60.5$\pm$0.2 71.5$\pm$0.7 65.9$\pm$0.5 76.0$\pm$0.5 61.4$\pm$0.4 72.6$\pm$0.6 59.6$\pm$1.0 64.7$\pm$1.0 68.5
-SGDA 56.6$\pm$0.6 56.9$\pm$0.8 55.3$\pm$1.2 54.6$\pm$0.5 57.9$\pm$1.3 58.3$\pm$0.4 56.1$\pm$0.9 55.9$\pm$0.6 54.6$\pm$1.3 56.7$\pm$0.5 53.3$\pm$0.7 56.8$\pm$1.1 56.1
-DGDA OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
-A2GNN 55.4$\pm$0.8 56.1$\pm$0.6 56.7$\pm$1.0 55.3$\pm$0.5 54.9$\pm$0.7 57.2$\pm$0.9 55.7$\pm$0.5 56.5$\pm$1.3 54.5$\pm$0.6 56.8$\pm$0.5 56.2$\pm$1.0 58.8$\pm$0.8 56.1
-PA-BOTH 56.1$\pm$0.5 56.0$\pm$0.4 56.3$\pm$0.7 56.4$\pm$0.4 56.0$\pm$0.6 57.1$\pm$0.7 56.2$\pm$1.1 58.3$\pm$0.9 56.5$\pm$0.6 57.2$\pm$0.9 56.9$\pm$0.4 57.7$\pm$0.8 56.8
-DeSGDA 75.9$\pm$1.0 74.7$\pm$1.2 66.8$\pm$1.5 71.3$\pm$1.9 63.2$\pm$1.5 69.4$\pm$1.7 66.5$\pm$1.3 74.3$\pm$2.4 62.6$\pm$1.5 73.0$\pm$1.8 61.6$\pm$2.2 63.6$\pm$1.5 68.6
-[…]
-[…]
-WL subtree 74.9 74.8 67.3 69.9 57.8 57.9 73.7 80.2 60.0 57.9 70.2 73.1 68.1
-GCN 73.0$\pm$1.7 68.7$\pm$1.5 66.8$\pm$3.5 69.2$\pm$0.9 53.9$\pm$3.4 53.4$\pm$2.7 69.3$\pm$0.8 74.0$\pm$1.1 55.1$\pm$1.3 42.6$\pm$1.9 55.5$\pm$3.5 57.9$\pm$2.9 61.6
-GIN 74.1$\pm$1.8 73.4$\pm$3.4 65.4$\pm$1.5 70.4$\pm$2.9 58.9$\pm$2.7 61.2$\pm$1.1 73.2$\pm$3.8 77.7$\pm$3.0 63.1$\pm$3.7 63.9$\pm$2.4 67.4$\pm$2.3 73.2$\pm$1.9 68.5
-GMT 69.0$\pm$4.0 67.4$\pm$3.8 60.3$\pm$4.2 66.5$\pm$3.8 54.9$\pm$1.6 54.8$\pm$3.6 65.6$\pm$4.2 70.4$\pm$3.2 64.0$\pm$2.3 56.8$\pm$4.3 64.7$\pm$1.5 61.1$\pm$3.5 63.0
-CIN 68.5$\pm$2.1 65.1$\pm$2.6 65.4$\pm$1.3 63.6$\pm$2.8 57.3$\pm$3.4 59.0$\pm$3.1 59.3$\pm$1.5 68.3$\pm$1.3 58.1$\pm$2.4 71.1$\pm$3.1 60.7$\pm$1.7 61.7$\pm$2.4 63.2
-SpikeGCN 66.7$\pm$1.5 65.5$\pm$2.0 57.9$\pm$0.4 60.2$\pm$1.6 53.2$\pm$1.4 60.1$\pm$1.5 57.7$\pm$1.2 67.3$\pm$1.5 57.7$\pm$2.1 60.1$\pm$1.9 59.9$\pm$2.4 63.3$\pm$1.8 60.1
-DRSGNN 66.9$\pm$1.2 62.1$\pm$0.7 57.1$\pm$1.2 63.3$\pm$2.1 56.6$\pm$0.9 62.1$\pm$1.3 56.9$\pm$1.0 67.2$\pm$1.8 58.1$\pm$0.6 61.3$\pm$2.5 58.8$\pm$1.0 64.7$\pm$1.7 61.3
-CDAN 74.2$\pm$0.3 73.7$\pm$0.5 68.8$\pm$0.2 71.8$\pm$0.4 59.9$\pm$2.0 58.6$\pm$1.9 70.7$\pm$1.4 74.3$\pm$0.3 59.2$\pm$1.2 69.0$\pm$1.6 60.0$\pm$1.2 62.7$\pm$1.3 66.9
-ToAlign 75.5$\pm$1.9 67.1$\pm$3.8 68.1$\pm$1.5 63.3$\pm$2.7 55.6$\pm$1.2 67.3$\pm$4.3 69.4$\pm$3.3 77.0$\pm$1.2 57.6$\pm$1.6 74.9$\pm$2.4 59.0$\pm$3.3 64.6$\pm$3.4 66.6
-MetaAlign 74.5$\pm$0.9 73.8$\pm$0.6 69.4$\pm$1.2 72.6$\pm$1.3 59.8$\pm$1.8 70.7$\pm$2.7 72.0$\pm$0.5 75.6$\pm$0.6 62.4$\pm$2.1 72.3$\pm$1.9 62.2$\pm$1.1 72.0$\pm$1.2 69.7
-DEAL 76.3$\pm$0.2 72.4$\pm$0.7 68.8$\pm$1.0 72.5$\pm$0.7 57.6$\pm$0.6 67.6$\pm$1.9 77.4$\pm$0.6 80.0$\pm$0.7 64.9$\pm$0.7 72.8$\pm$1.4 70.3$\pm$0.3 76.2$\pm$1.3 71.4
-CoCo 77.5$\pm$0.4 75.7$\pm$1.3 68.3$\pm$3.7 74.9$\pm$0.5 65.1$\pm$2.1 74.0$\pm$0.4 76.9$\pm$0.6 77.4$\pm$3.4 66.4$\pm$1.5 71.2$\pm$2.7 62.8$\pm$4.2 77.1$\pm$0.6 72.2
-SGDA OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
-DGDA OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
-A2GNN 55.3$\pm$0.3 54.9$\pm$0.6 55.8$\pm$0.4 55.1$\pm$0.8 54.2$\pm$1.0 57.1$\pm$1.2 56.1$\pm$0.5 55.2$\pm$0.7 57.9$\pm$1.5 56.3$\pm$0.6 54.4$\pm$0.5 58.1$\pm$1.5 55.8
-PA-BOTH 56.3$\pm$0.5 57.7$\pm$0.9 56.9$\pm$0.6 56.2$\pm$1.0 55.7$\pm$0.8 56.5$\pm$0.9 57.8$\pm$1.2 56.9$\pm$2.1 56.5$\pm$1.5 56.2$\pm$1.8 56.8$\pm$1.4 57.4$\pm$0.7 56.8
-DeSGDA 75.8$\pm$1.4 74.5$\pm$1.7 69.5$\pm$1.3 75.0$\pm$2.0 61.0$\pm$1.5 69.2$\pm$1.3 69.5$\pm$1.6 76.1$\pm$1.5 65.0$\pm$1.4 75.5$\pm$2.2 63.4$\pm$1.8 68.3$\pm$1.3 70.3
+WL subtree 51.5 59.5 28.6 23.8 23.4 19.4 56.8 51.2 54.9 50.2 61.5 57.3 44.8
+GCN 49.6$\pm$2.2 62.7$\pm$2.3 22.8$\pm$2.0 26.9$\pm$1.4 13.9$\pm$2.0 22.6$\pm$1.9 74.6$\pm$1.3 58.7$\pm$2.4 75.1$\pm$1.1 52.2$\pm$1.6 76.6$\pm$1.3 67.5$\pm$2.1 50.3
+GIN 48.9$\pm$2.8 25.9$\pm$1.8 44.6$\pm$1.5 23.0$\pm$2.0 57.2$\pm$1.8 19.4$\pm$2.0 71.8$\pm$1.2 54.8$\pm$2.4 62.2$\pm$1.5 52.7$\pm$1.9 71.3$\pm$1.7 67.5$\pm$2.4 50.0
+GMT 50.8$\pm$2.2 42.7$\pm$2.5 34.9$\pm$2.7 34.8$\pm$1.8 48.2$\pm$2.0 29.6$\pm$2.5 68.9$\pm$1.5 52.6$\pm$1.6 71.2$\pm$1.5 57.1$\pm$2.1 75.9$\pm$1.4 67.8$\pm$1.6 52.9
+CIN 50.4$\pm$2.2 29.4$\pm$2.0 23.1$\pm$1.4 31.6$\pm$1.7 42.8$\pm$2.0 24.6$\pm$1.6 54.2$\pm$1.2 57.5$\pm$1.6 73.5$\pm$2.2 52.7$\pm$1.3 75.6$\pm$1.6 67.1$\pm$2.1 48.6
+SpikeGCN 56.4$\pm$1.9 70.5$\pm$2.1 34.1$\pm$2.6 53.2$\pm$2.9 20.7$\pm$1.6 49.1$\pm$1.7 79.7$\pm$2.4 66.5$\pm$1.2 77.3$\pm$2.1 61.7$\pm$1.6 78.7$\pm$2.0 71.0$\pm$1.5 59.9
+DRSGNN 55.3$\pm$2.4 69.9$\pm$2.2 27.4$\pm$2.0 47.6$\pm$2.7 17.9$\pm$1.6 47.4$\pm$2.1 70.7$\pm$2.0 65.9$\pm$1.7 76.9$\pm$2.1 62.2$\pm$1.4 78.5$\pm$1.8 71.4$\pm$1.6 57.6
+CDAN 49.7$\pm$1.9 65.3$\pm$2.3 45.4$\pm$1.8 43.1$\pm$2.1 42.8$\pm$1.7 51.8$\pm$1.6 71.5$\pm$2.0 64.9$\pm$1.6 74.5$\pm$2.5 59.2$\pm$2.2 77.9$\pm$2.1 69.0$\pm$1.5 59.5
+ToAlign 52.3$\pm$2.5 66.5$\pm$2.0 47.1$\pm$1.6 45.6$\pm$1.8 41.2$\pm$2.2 51.2$\pm$1.8 73.9$\pm$1.9 65.9$\pm$2.3 77.6$\pm$2.0 60.8$\pm$1.6 78.1$\pm$2.4 70.2$\pm$2.1 60.9
+MetaAlign 48.1$\pm$2.0 67.3$\pm$1.7 32.8$\pm$2.0 19.4$\pm$1.8 23.9$\pm$2.5 19.4$\pm$1.7 70.1$\pm$1.8 51.9$\pm$2.1 77.3$\pm$3.2 51.9$\pm$1.6 76.1$\pm$1.8 70.5$\pm$2.0 50.7
+DEAL 58.4$\pm$1.5 70.6$\pm$2.0 63.9$\pm$1.6 54.1$\pm$2.1 66.9$\pm$2.4 51.8$\pm$1.6 75.1$\pm$2.5 67.4$\pm$1.8 77.8$\pm$1.9 60.3$\pm$2.1 80.5$\pm$1.8 75.0$\pm$2.0 66.8
+CoCo 60.9$\pm$2.3 69.6$\pm$1.2 62.2$\pm$2.2 66.2$\pm$2.0 66.0$\pm$1.8 52.5$\pm$2.3 71.1$\pm$2.4 65.3$\pm$1.5 78.9$\pm$1.3 60.3$\pm$1.4 79.6$\pm$2.1 73.5$\pm$1.8 67.3
+SGDA 57.2$\pm$1.5 68.8$\pm$1.8 42.3$\pm$2.0 61.4$\pm$1.7 39.8$\pm$2.2 52.0$\pm$1.8 66.7$\pm$1.9 66.4$\pm$2.3 78.1$\pm$2.1 63.6$\pm$2.6 73.6$\pm$1.6 70.8$\pm$1.9 61.7
+StruRW 52.5$\pm$2.5 56.7$\pm$1.3 39.0$\pm$2.3 40.1$\pm$2.0 24.4$\pm$2.1 29.7$\pm$2.4 75.4$\pm$1.7 63.3$\pm$2.0 74.8$\pm$1.6 53.4$\pm$1.5 75.4$\pm$1.4 68.7$\pm$1.7 54.6
+A2GNN 53.1$\pm$2.0 65.3$\pm$1.7 42.8$\pm$1.9 40.5$\pm$2.1 33.9$\pm$2.5 39.4$\pm$1.8 69.8$\pm$2.2 61.9$\pm$1.9 77.3$\pm$2.1 61.9$\pm$2.0 76.1$\pm$2.3 67.2$\pm$1.8 57.4
+PA-BOTH 51.9$\pm$1.8 50.6$\pm$2.0 35.8$\pm$1.5 37.7$\pm$1.7 27.6$\pm$2.3 43.7$\pm$1.9 62.1$\pm$1.6 61.2$\pm$1.9 65.7$\pm$2.0 58.2$\pm$1.5 73.3$\pm$1.8 69.5$\pm$2.5 53.1
+DeSGraDA 62.1$\pm$2.0 72.0$\pm$2.4 69.1$\pm$1.7 71.7$\pm$2.2 68.7$\pm$1.6 58.6$\pm$2.1 76.1$\pm$1.9 66.4$\pm$1.7 80.7$\pm$2.2 63.6$\pm$1.6 81.9$\pm$1.5 78.6$\pm$2.3 70.8
+Appendix L Limitation
+The proposed DeSGraDA framework, while demonstrating significant improvements in spiking graph domain adaptation, does have some limitations.
+It assumes a substantial domain shift between the source and target domains, which may not always be applicable in real-world scenarios where domain shifts are minimal.
+Additionally, the computational complexity introduced by adversarial feature distribution alignment and pseudo-label distillation could become a bottleneck, especially for large-scale datasets.
+The framework’s sensitivity to hyperparameters, such as time latency and threshold values, also requires careful tuning for different datasets, which may hinder its practical application.
+Furthermore, while the method provides a generalization bound, its robustness in diverse real-world settings and its ability to address privacy or fairness concerns in sensitive domains remain underexplored.
+These aspects highlight opportunities for further refinement and broader applicability of DeSGraDA.
+Appendix M Broader Impacts
+The work on spiking graph domain adaptation offers significant positive societal impacts, particularly in enhancing domain adaptation for graph-based learning tasks in healthcare, biology, and environmental sciences, where it could improve predictions, decision-making, and personalized outcomes.
+Additionally, its energy-efficient SNNs could contribute to more sustainable AI models.
+However, potential negative impacts include the risk of reinforcing biases in sensitive applications such as criminal justice or hiring, particularly if the training data is not diverse or representative.
+The computational complexity may also hinder adoption in resource-limited settings, potentially widening the digital divide.
+Moreover, privacy concerns in healthcare and finance must be carefully addressed to avoid using sensitive data.
+In conclusion, while DeSGraDA holds promise, careful attention to fairness, accessibility, and privacy is necessary for its responsible implementation.